### v1

# Eaman Jahani

Abstract
In an online experiment with N = 1,891 participants, we collected and analyzed over 18,000 prompts to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve. Each participant in our experiment was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. Participants were then asked to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. More specifically, despite being blind to the model they were assigned, participants assigned to DALL-E 3 wrote longer prompts that were more semantically similar to each other and contained a greater number of descriptive words. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Taken together, our results suggest that as models continue to progress, people will continue to adapt their prompts to take advantage of new models’ capabilities.
The release of ChatGPT in November 2022 has fueled a surge of interest in generative AI and large language models across industry, the public sector, and academia, with early research suggesting that generative AI may be a general purpose technology (Eloundou et al., 2024). Indeed, generative AI models have already been found to outperform humans on text annotation tasks (Gilardi et al., 2023) and improve the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). However, the extent to which generative AI can provide value is contingent on the quality of the textual instructions, or “prompts,” that users write to a model. As generative AI has proliferated, the practice of creating and iterating on such prompts, often using a variety of specialized techniques (e.g., chain-of-thought and few-shot prompting) (Schulhoff et al., 2024), has come to be known as prompt engineering. Some firms now even offer six-figure salaries for dedicated prompt engineers (Quilty-Harper, 2023), and an expanding academic literature focuses exclusively on prompting-related topics (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023a; Xie et al., 2023). Despite the growing interest in prompting, some argue that as models become more capable at transforming human language into output, the need to craft prompts in specific, and sometimes esoteric, ways will diminish (Acar, 2023; Mollick, 2023). If this comes to pass, the benefits of generative AI will be accessible to a much larger population.
In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,891{}$ participants recruited on Prolific. [^1] This paper presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) DALL-E 3, and (3) a version of DALL-E 3 with automatic LLM-based prompt revisions (henceforth referred to as “DALL-E 3 with revision”). These prompt revisions were not made visible to participants during the experiment. [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
[…]
Results
The majority of our analyses focus on the DALL-E 2 and DALL-E 3 treatment arms, as they are more directly comparable. We first estimate the effect of access to a more advanced generative model on task performance by evaluating how similar each generated image is to the relevant target image. [^3] All main text analyses use ”replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details. To construct a quantitative measure of image similarity, we generated CLIP embedding vectors (Radford et al., 2021) for each participant-generated image and the relevant target image and then calculated the cosine similarity ($CoSim$) between those vectors (see SI for further details on how we quantify image similarity and account for the stochasticity of model output). Unsurprisingly, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.20$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model). Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in similarity to the target image. We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0008$, $p=0.042$). In other words, the performance gap widened between DALL-E 2 and DALL-E 3 as participants gained more experience with the respective models, potentially due to changes in prompting.
To investigate how participants’ prompts changed in response to the superior model’s capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated. We find that participants assigned the superior model wrote prompts that were 24% longer ($\Delta Words=6.6$, $p<10^{-4}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-7}$). Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% superior vs. 49% inferior; $p$ = $0.148$). This suggests that prompts submitted to DALL-E 3 were longer because they conveyed additional descriptive information, not because they were padded with superfluous words. We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0166$, $p=0.02$) and in the aggregate ($\beta=0.0169$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). These findings indicate that exposure to different models did in fact induce differences in prompting, and that these differences grew over successive replication attempts.
Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects. The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line). Effects are shown relative to a baseline of prompts written for and replayed on the inferior model. Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
To understand the extent to which the ATE we observe is attributable to these behavioral differences, as opposed to DALL-E 3’s more advanced capabilities, we conducted another exploratory analysis in which we regenerated images with (or “replayed”) all participant prompts on both DALL-E 2 and DALL-E 3, regardless of the model the participant used during the experiment. This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by inferior model participants on the superior model, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by superior model participants and the average similarity to the target of all prompts generated by inferior model participants, both replayed on the superior model. Figure 2 shows that the model effect (red) accounts for 53% of the ATE ($\Delta CoSim=0.00864$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 47% of the ATE ($\Delta CoSim=0.00769$, $p=0.0066$). Thus, nearly half of the ATE is driven by endogenous changes in prompting strategy that emerged after exposure to DALL-E 3. The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of the superior model; when replaying both sets of prompts on the inferior model, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0018$, $p=0.52$). In other words, DALL-E 3 was more capable than DALL-E 2 at effectively rendering the additional information contained in the longer and more descriptive prompts (Betker et al., 2023), and participants adapted their prompting behavior accordingly.
One might think that the need to adapt one’s prompts to the capabilities of different generative models is obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023). We use data from our third treatment arm—DALL-E 3 with revision—to test this hypothesis. We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0068$; $p=0.013$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [22%, 94%]). Importantly, the positive impact of access to DALL-E 3 with revision was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). In other words, prompt revision actually caused participants to perform worse in our context. This finding suggests that as currently implemented, AI-assisted prompt revisions are not a panacea, and can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
Discussion
In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance to increase, not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect). More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render this additional information more effectively than DALL-E 2. These changes in prompting emerged despite the fact that participants were unaware of which model they were assigned. Although automated LLM-revision features are designed to and have been shown to help people generate higher-quality content (Betker et al., 2023; Li et al., 2024), we also find that LLM-based prompt revision failed to further improve performance; instead, automatic prompt revision reduced the benefits of using DALL-E 3. Overall, we find that shifting to a superior model, whether or not accompanied by LLM-based revisions, did not eliminate the need for prompting.
The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future work. However, if the trends we report generalize and continue into the future, one could imagine a lock-step dynamic where as models continually improve, people continually respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that as generative AI models advance, newer models are unlikely to render prompting obsolete. Rather, prompting will be the mechanism by which people unlock new models’ capabilities. Thus, we anticipate that prompting will remain relevant in the future, and that it will be an important determinant of who is best positioned to use and directly benefit from generative AI moving forward.
Supplementary Information (SI)
Appendix A Experiment Design
A.1 Task Design
Participants were asked to reproduce a single target image as closely as possible using a text-to-image generative AI model (i.e., DALL-E 2, DALL-E 3 without prompt revision, or DALL-E 3 with prompt revision, all developed by OpenAI). They did so by successively submitting prompts. In response to each submitted prompt, the model would generate an image, which was then displayed to the participant next to their assigned target image. Participants were instructed to make at least 10 attempts at trying to recreate the target image within a 25-minute window, with no upper limit on their number of attempts.
All interactions between participants and the generative AI models occurred on a custom-built online interface designed to resemble OpenAI’s ChatGPT interface but with some adjustments related to our task (e.g., displaying the target image and the total number of attempts so far to the user). On the right-hand side of the interface, participants were shown the target image they were randomly assigned to recreate. On the left-hand side, participants were shown their previously submitted prompts as well as the resulting generated images. We placed the text box where participants were able to write and submit their prompts at the bottom of the interface. Prompts were limited to a maximum of 1,000 characters. Participants were informed that their interactions with their assigned model would be memory-less, i.e., the model retained no memory of previous prompts and only used the current prompt to generate each image. Before the task, participants were provided with written and video instructions on how to interact with our experiment interface. Our task did not assume nor require prior experience with any generative AI tools.
After the task, we surveyed participants’ opinions and preferences regarding generative AI tools. We also inquired about their self-assessed occupational skills and how often they 1) engaged in creative writing, 2) wrote specific instructions, and 3) engaged in any sort of computer programming. Finally, we collected socio-demographic data, such as age, gender, and occupation.
A.2 Randomization
We randomized participants across two dimensions: the target image and the text-to-image generative AI model that participants had access to. We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells. We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test for homogeneity of proportions across all cells. With $\chi^{2}=0.374,\;df=28$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
$H_{0}:p_{1}=p_{2}=\cdots=p_{45}$
Participants were unaware of this randomization.
A.2.1 Generative Models
We randomly assigned participants to 1 of 3 generative models:
  1. 1.
    DALL-E 2, which is referred to at points in the main text as the inferior model.
  2. 2.
    DALL-E 3 (Verbatim), which is referred to at points in the main text as the superior model.
  3. 3.
    DALL-E 3 (Revised), which is referred to at points in the main text as DALL-E 3 with Revision
Both the “verbatim” and “revised” versions of the DALL-E 3 treatment utilize the same underlying image-generating model; the distinction lies in the pre-processing applied before submitting user prompts to OpenAI’s image-generating API. OpenAI’s DALL-E 3 system, by design, employs a GPT-4 model to rewrite user prompts, adding more detail before processing the modified prompt using the DALL-E image-generating model. During our experiment, it was not possible to explicitly disable this prompt rewriting feature of the DALL-E 3 system. To manage this behavior, we defined two treatments utilizing the DALL-E 3 model.
In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process. In the DALL-E 3 (Verbatim) treatment arm, we prepend a string instructing the GPT-4 model to not modify the participant’s prompt before passing it forward to DALL-E 3. This string is never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint. [^4] See here and here for online documentation. We modified the recommended preflix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
“I NEED to test how the tool works with my prompt as it is written. DO NOT add any detail; just use it AS IS:”
Prepending this string to participants’ prompt did reduce the rate at which OpenAI’s endpoint modified prompts, but compliance was not perfect. Thus, we view the “verbatim” treatment arm as more of an intent-to-treat intervention. The GPT model still modified 59% of participant prompts. The average token sort ratio (TSR) between the original prompt and the modified prompt was 77 for the DALL-E 3 (Verbatim) arm, compared to an average token sort ratio of 44 across the entire DALL-E 3 (Revised) treatment arm (a TSR of 100 denotes an exact string match). Conditional on any modification (any observations with TSR $<$ 100), the average TSR between the original prompt and the modified prompt was 61 for the DALL-E 3 (Verbatim) arm, compared to an average TSR of 44 across the entire DALL-E 3 (Revised) treatment arm.
A.2.2 Target Images
We randomly assigned participants to 1 of 15 target images. The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and photography. We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts. The images vary in color, style, content, and complexity within and across categories. These images can be found online linked to the pre-registration document: https://osf.io/ejbtp. As we discuss in Section F, performance, and variability of performance varied substantially across images. In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
A.3 Subjects
Our Prolific-recruited US sample (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once. We also prevented users from completing the task on mobile devices or tablets. Data was collected between December 12, 2023 and December 19, 2023. Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section B.5.1). The median time to complete our entire task, including a demographic survey, was 22 minutes. Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$15 USD per hour. We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task. The onboarding process also included multiple attention checks; participants who failed the first check were immediately disqualified. For subsequent checks, participants were required to retry until they demonstrated understanding.
A.4 Model Endpoints
We used the following model endpoints and parameters to generate images from prompts:
  1. 1.
    OpenAI API: We used the image generation endpoint of the official OpenAI Node.js library to generate images for user prompts during the experiment. For all treatment arms, we set the image size parameter to be 1024 x 1024 pixels. For the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we set the quality parameter to standard and the style parameter to natural.
  2. 2.
    Azure OpenAI Service: We used the image generation endpoints in the Python implementation of Azure OpenAI Service to generate all replay images based off user prompts collected during the experiment. For prompts replayed through the DALL-E 2 treatment arm, we deployed a set of DALL-E 2 models on Azure OpenAI Service and set the API version for each to the 2023-06-01-preview version. For DALL-E 2, we created replay images in batches of 5. For prompts replayed through the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we deployed a set of DALL-E 3 models on Azure OpenAI Service and set the API version for each to the 2023-12-01-preview version. The parameter values for image size, and quality and style for the DALL-E 3 treatment arms, were set to the same values as in the experiment.
Appendix B Measurement and Variables
B.1 Survey Data
For each participant, we used a Qualtrics survey to collect demographic information, information on the participant’s skills that may be relevant to generative AI use, and information on the participant’s attitude towards generative AI. This data includes:
  • Demographics: Ethnicity, Gender, Age, Highest level of education attained (some high school, high school, some college, associate’s degree, bachelor’s degree, master’s degree, doctoral degree, professional degree, other), Years of work experience, Annual Income (0-\$25k, \$25.001k-\$50k, \$50.001k-75k, \$75.001k-\$100k, \$100.001k-\$150k, \$150k+), and elicitation of sets of O*NET job skills that participants used in their occupation (reading comprehension, active listening, writing, speaking, critical thinking, social perceptiveness, coordination, instructing, programming, judgment and decision making, systems evaluations, science, active learning, learning strategies, monitoring, complex problem analysis, technology design, troubleshooting, quality control analysis, systems analysis).
  • Opinions and Skills: Computer programming proficiency and usage frequency (self-reported), Structured and creative writing proficiency and usage frequency (self-reported), Generative AI tool proficiency and usage frequency (self-reported), Attitudes towards net social impact of Generative AI (self-reported), Advice for (hypothetical) future participants on how to perform well on the task.
B.2 Prompt Data
For each prompt, we record the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
B.3 Image Data
For each prompt, the following images were collected:
  1. 1.
    The participant-facing images (OpenAI API endpoint): The image shown to the participant during the experiment, generated by the model they were assigned to using the prompt they submitted. These images were generated from December 12-19, 2023.
  2. 2.
    Post-hoc resampled images (Azure OpenAI endpoint): For any given prompt, the output of the text-to-image model is stochastic. To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the experiment concluded. We provide full details on this procedure in Section C. These images were generated from December 26, 2023 - January 27, 2024. These images are not used for analyses presented in the main text, but were used for other pre-registered analyses. These additional analyses are discussed in Section F.
  3. 3.
    Post-hoc replayed images (Azure OpenAI endpoint): To decompose our overall effects into model and prompting effects, we generated “counterfactual images” for each prompt written under the DALL-E 2 and DALL-E 3 (Verbatim) treatments. In other words, we submitted all prompts written under both the DALL-E 2 and DALL-E 3 (Verbatim) treatments to both the DALL-E 2 and DALL-E 3 (Verbatim) endpoints. Similarly to the resampling procedure outlined above, we generated 10 images per prompt per model: we generated a single replay for each prompt-model pair from March 16-18, 2024, and then, to increase power, generated the replications for these replay images from June 14-27, 2024. This replay process produced a total of 20 images per prompt—10 under the original model, 10 under the counterfactual model. We re-submitted prompts to their original model to account for potential model drift, as this exploratory analysis was conducted multiple months after our initial data collection. For consistency, this replay data is used throughout the main text of our paper.
B.4 Sample Construction
The sample we analyze in the main text of our paper is constructed using the process described as follows.
  • The initial “raw” dataset collected during the experiment is comprised of 24,672 rows of raw prompt data (one prompt per row) generated by 2,059 participants.
  • We first removed rows with blank prompt entries, invalid prolific IDs, and unsuccessful attempts (logging errors). These exclusion criteria were pre-registered. This left us with 2,029 participants and 24,123 prompts.
  • We next removed participants from our sample if they failed to submit at least 10 prompts or if they did not make a “good-faith” effort to complete the task as requested. Both of these exclusion criteria were pre-registered. A good-faith effort was not made if a participant submitted the same prompt at least five times in a row at any point during the task. These exclusion criteria were also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” To avoid reward hacking, we did not specify the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants.
[…]
  • We next removed participants who failed to complete the Qualtrics survey. This exclusion criteria was pre-registered. This left us with 1,893 participants and 18,930 prompts.
  • We also removed prompts from our dataset according to a number of post-hoc, non-pre-registered exclusion criteria to ensure data quality and avoid selection bias. If a prompt had any of the following flags, it was removed from the sample:
    • Prompts sometimes trigger errors in OpenAI’s safety system because they contain language that might be deemed unsafe under openAI’s policies. The specific language that triggers these errors is constantly changing and not available publicly. If a prompt triggered a safety error during the replication or replay process, we re-submitted the prompt up to 50 times or until the 10 original arm replications/replay samples had been collected. We removed prompts if they failed to generate 10 replications on the original model or 10 replay samples under the counterfactual model during the replication/replay process. This affected 305 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms. It did not affect any DALL-E 3 (Revised) prompts, as we did not conduct replay analysis with the prompts from this treatment arm.
    • Due to rare latency issues, some prompts were assigned duplicate attempt numbers by the mongoDB database that we used to collect our data. This data collection error led to issues in the data analysis process. Thus, we excluded prompts with duplicate attempt numbers. This affected 34 prompts across all three treatment arms, and 20 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms, approximately 0.1% of the original data.
    • While analyzing our data, we found that our sample contained a number of prompts that were not “good-faith efforts” to complete our task, despite the use of the exclusion criteria described above. Thus, we used the following process to identify and remove additional “low-effort” prompts. First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model. We then calculated the mean embedding for each target image. Next, we calculated the Euclidean distance between each prompt’s embedding vector and the mean embedding vector for prompts corresponding to the focal prompt’s assigned target image. Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector. We observed that the prompts that were dropped were often off-topic and likely were not good-faith efforts to complete our task. This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms. Our results are robust to the inclusion of these filtered prompts.
  • Our final sample included 1,891 participants and 18,152 prompts.
B.5 Dependent Variables
B.5.1 Image Similarity
We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al., 2023). In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers. Our results are qualitatively and quantitatively similar using DreamSim instead.
  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al., 2021) from Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image. Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text. We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023). The authors claim that relative to something like CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.
We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure. We present the results obtained when conducting our main text analyses using DreamSim in Section E.E.1.
B.5.2 Prompt Length
We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt. In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9952$, 95% CI: [0.99507, 0.99543]).
B.5.3 Embedding-based Prompt Similarity
We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity. Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al., 2022). The two similarity measures are defined as follows:
  • Successive similarity: The successive similarity ($ss$) is a measure of the similarity of a participant’s prompt to their immediately preceding prompt. We define the successive similarity of a prompt $p_{i,n}$ written by user $i$ to the their immediately preceding prompt $p_{i,n-1}$ as:
    $ss_{i,n,n-1}=\frac{\mathbf{E(p_{i,n})}\cdot\mathbf{E(p_{i,n-1})}}{||\mathbf{E(p_{i,n})}||\,||\mathbf{E(p_{i,n-1})}||},$ (1)
[…]
  • Aggregate similarity: The aggregate similarity ($as$) is a measure of how dispersed each user’s prompts are around their “average prompt” (calculated by taking the element-wise average of all prompt embeddings produced by the user). We define the aggregate similarity for the 10 prompts written by a given user as:
    $as_{i}=\frac{1}{10}\sum_{n=1}^{10}\lVert\mathbf{E(p_{i,n})}-\overline{\mathbf{E(p_{i,n})}}\rVert_{2}^{2},$ (2)
    where $\mathbf{E(p_{i,n})}$ is again the vector embedding representation of participant $i$’s $n^{\text{th}}$ prompt, $p_{i,n}$, and $\overline{\mathbf{E(p_{i,n})}}$ is the element-wise mean of all 10 of participant $i$’s prompts.
B.5.4 Successive Prompt Token Sort Ratio
Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$. TSR is a fuzzy string-matching technique (Singla and Garg, 2012b) that provides a continuous measure of how similar two strings are. We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
B.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
Starting with each participant’s second prompt, we record whether each prompt $p_{i,n}$ contains the immediately preceding prompt $p_{i,n-1}$ as an exact substring.
B.5.6 Inversions of Successive Measures
[…]
B.5.7 Prompt Composition
We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt. SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation posuniversaldep. These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs. The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
B.5.8 Strategic Shifts
In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting. In order to do so, we adapt a method proposed in (Torricelli et al., 2023) (because they are conducting research in a different context, (Torricelli et al., 2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”). To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
$MCS_{t}=\frac{2}{P_{t}(P_{t}-1)}\sum_{a=1}^{P_{t}}\sum_{b=a+1}^{P_{t}-1}\textrm{CosineSim}(\mathbf{E(p_{a,t})},\mathbf{E(p_{b,t})}).$ (3)
where $P_{t}$ is the total number of prompts submitted in response to a given target image, and $a$ and $b$ are indices representing individual prompts for that target.
We then label any given prompt as a strategic shift ($SS$) if the cosine similarity of its embedding vector with that of the previous prompt is lower than this target-image-level mean:
$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}(\mathbf{E(p_{i,n,t})},\mathbf{E(p_{i-1,n,t})})<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (4)
It is worth noting that Torricelli et al. (2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift. We instead use the task-level mean because in our setting, it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
Appendix C Methods
C.1 Stratification
The results shown in the main text and in the SI are mostly stratified by reference image and iteration. In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level). The exact stratification for each finding is indicated in section D. To stratify our results, we take a weighted average across $j=1,...,J$ cells defined by our stratification variables.
$\overline{Y}_{strat}=\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}$
To calculate the variance (and standard error) of this sample mean we apply the following:
$\widehat{\mathrm{Var}}(\overline{Y}_{strat})=\widehat{\mathrm{Var}}\left(\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}\right)=\sum_{j=1}^{J}\left(\frac{N_{j}}{N}\right)^{2}\frac{s_{j}^{2}}{N_{j}}$
where:
  • $N_{j}$ is the population size of stratum $j$.
  • $N$ is the total population size across all strata.
  • $\bar{Y}_{j}$ is the sample mean for stratum $j$.
  • $J$ is the total number of strata.
  • $s_{j}$ is the sample standard deviation of stratum $j$. Therefore, $s_{j}^{2}$ is sample variance stratum $j$.
C.2 Z-Scoring
We find statistically significant evidence for differences in the variability of performance across the 15 target images used in our experiments, which we discuss in Section F. In the main text, we also showed that performance increases across the attempts. To test whether our results are in some way due to this image-level or attempt-level variations, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment. Formally, this is:
$Z(\textrm{CosineSim}_{i,n,t})=\frac{\textrm{CosineSim}_{i,n,t}-\textrm{Mean}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)}{\textrm{SD}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)},$ (5)
where $\textrm{CosineSim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$. The mean and standard deviation are computed per each image-attempt, but over the DALL-E 2 and DALL-E 3 treatment arms. We also use the rescaling above to test the robustness of our DreamSim-based analyses.
C.3 Accounting for Model Stochasticity
The output that generative AI models return in response to a given prompt is stochastic. The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E 3 API at the time of our experiment. To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms. We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples.
We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text. Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling. As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts. This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the under the ”replaying” procedure (Section B.B.3.3).
Appendix D Main Text Analyses
D.1 Task Performance and ATEs
The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations). It shows the average cosine similarity score stratified by the reference image. A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms. This is likely due to participants’ initial misunderstanding of the model’s “memoryless” nature. Participants failed to recognize that context from previous prompts was not carried over to new iterations. We observed numerous prompts in the second iteration across users that explicitly referenced the first prompt, a behavior that rarely occurred in subsequent attempts. However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
Next, the bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane. [^5] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified. To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (6)
The coefficient estimates generated by this model are:
  • $\hat{\beta_{1}}=0.0015$, $\hat{SE}(\beta_{1})=0.0003$, $p=9.678\times 10^{-7}$
  • $\hat{\beta_{2}}=0.0126$, $\hat{SE}(\beta_{2})=0.003$, $p=1.7\times 10^{-4}$
  • $\hat{\beta_{3}}=0.00086$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0418$
The overall ATEs that we report between different pairs of treatment arms (DALL-E 2, DALL-E 3, and DALL-E 3 with revisions) in the main text are estimated from a two-way fixed effect (iteration and target image) model per each pair. Standard errors are cluster robust at the participant level.
D.2 Prompt Characteristics
Figure 1C compares the prompt length and prompt similarity of the two models. To generate these results, we first remove any prompt that does not resemble a good-faith attempt according to the sample construction procedure detailed in Section B.B.4. The prompt length is the average number of words per model and iteration stratified by the reference image. The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section B.B.5.3 for details on similarity calculations). The color scale in figure 1C shows the stratified average similarity to the previous prompt across all users per each model. We find that superior model users write prompts that, on average, have $\beta=0.0166$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.02$).
Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from the superior model participants were, on average, more similar than the prompts of the inferior model participants. For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable. When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=-0.0169$ smaller than inferior users ($p=0.008$). Standard errors are cluster-robust at the participant level.
D.3 ATE Decomposition
Figure 2 in the main text decomposes the ATE into the model and prompting effects. This decomposition is conceptually similar to a simple mediation analysis, with an important difference being that we can observe counterfactual outcomes (e.g., prompting the superior model as if it is the inferior model). This is not typically the case in mediation analysis, and makes causal identification rely on fewer assumptions. To obtain counterfactual outcomes, we fed or “replayed” the participant prompts when interacting with one model (e.g., inferior) on another model (e.g., superior). The notation (prompt, model) specifies which treatment arm the prompts were written under and which model was used in the replay. For example, (2,3) indicates replaying prompts written under DALL-E 2 on DALL-E 3. To be clear, (2,2) and (3,3) correspond to the original observed treatment arms, while (2,3) and (3,2) are the counterfactual outcomes of interest. [^6] To avoid problems with model drift, we regenerated images for all four possible combinations at the same time and used these images for all analyses in the main text.
The left-most point in Figure 2 corresponds to the average CLIP cosine similarity to the target image of (2,2). To make the interpretation of the results clearer, we have subtracted this quantity from all average quality scores and added a dashed line throughout. The second point from the left corresponds to average similarity to the target of (2,3), the third point from the left to (3,3), and the rightmost points to (3,2). All average similarity scores are stratified by iteration and reference image, and the standard errors are bootstrapped and cluster-robust at the participant level. The model effect, as shown by the red braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (2,2). In the terminology of mediation analysis, the model effect would be referred to as the direct effect. The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (3,3). In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect. We can also test the difference in average quality between (3,3) and (3,2), as well as the difference between (3,2) and (2,2). Both of these differences are visible in Figure 2; the second is small and not statistically significant.
The standard errors in Figure 2 correspond to the uncertainty around the estimated average score for each of the four replay conditions. These uncertainty estimates are insufficient for exact inference on the direct, indirect, and treatment effects. The statistics and significance values report in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{effect}+\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$ (7)
where $\beta_{1}$ is the coefficient on the effect type in question (i.e., model or prompting). To estimate the different effects, we simply use the above model and filter the data as appropriate. For example, to estimate the ATE, the data contains all (2,2) and (3,3) scores, and in this case effect=1 for observations in (3,3) group. Similarly, to estimate the direct or model effect, the data contains all (2,2) and (2,3) scores and effect=1 for observations in (2,3) group. Finally, to estimate the indirect or prompting effect, the data contains all (2,3) and (3,3) scores and effect=1 for observations in (3,3) group. The standard errors for each estimated model are cluster robust at the participant level, and p-values are adjusted accordingly.
Appendix E Robustness Checks
E.1 DreamSim-based Analysis
As discussed in Section B, we repeat all main-text analyses with DreamSim. The results of these analyses are reported below.
E.1.1 Overall ATEs
In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.261$ standard deviations (95% CI = [0.180, 0.342]) closer to the target image ($\Delta DreamSim=0.0304$, $p<10^{-9}$) than those produced by participants using DALL-E 2 (the inferior model).
E.1.2 Figure 1
We reran the regression in Section D.D.1 with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (8)
The coefficient estimates generated by this analysis are:
  • $\hat{\beta_{1}}=0.0039$, $\hat{SE}(\beta_{1})=0.00048$, $p=4.3\times 10^{-16}$
  • $\hat{\beta_{2}}=0.0179$, $\hat{SE}(\beta_{2})=0.0060$, $p=0.0029$
  • $\hat{\beta_{3}}=0.0023$, $\hat{SE}(\beta_{3})=0.00071$, $p=0.001$
E.1.3 Figure 2
Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text. The model effect accounts for 55% of the ATE ($\Delta DreamSim=0.0168$, $p<10^{-10}$), whereas the prompting effect accounts for 45% of the ATE ($\Delta CoSim=0.01364$, $p=0.0053$).
E.2 Z-score-based Analysis
As we discuss in Section B, we repeat all main-text analyses with the within-image Z score of CLIP-cosine similarity to better account for variation between images. The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
E.2.1 Overall ATEs
As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.20$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.20, 95% CI = [0.126, 0.284]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model). Standard errors are cluster robust by participants.
E.2.2 Figure 1
On average, participants using the superior model produced images that were $z=0.20$ standard deviations closer (the ATE) to the target image than those using the inferior model. Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
$\text{ZScore}_{i,n}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\epsilon_{i,n}$ (9)
  • $\hat{\beta_{1}}=0.02012$, $\hat{SE}(\beta_{1})=0.0041$, $p=1.19\times 10^{-6}$
  • $\hat{\beta_{2}}=0.1439$, $\hat{SE}(\beta_{2})=0.0466$, $p=0.002$
  • $\hat{\beta_{3}}=0.01253$, $\hat{SE}(\beta_{3})=0.0058$, $p=0.03$
E.2.3 Figure 2
When we decompose the ATE into the model effect ($z=0.0895$; $p=1.02\times 10^{-5}$) and prompting effect ($z=0.1139$; $p=0.0044$), they account for 44% and 56% of the treatment effect, respectively. And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.03$; $p=0.43$). In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z score of the cosine similarity.
Appendix F Pre-registration
Prior to our experiment, we pre-registered a number of hypotheses and a pre-analysis plan. This pre-registration is deposited at OSF at the following URL: https://osf.io/ejbtp. The main text of this paper contains a subset of our pre-registered analysis, as well as complementary exploratory analyses that are also mentioned in our pre-registration. We chose to present this subset of our pre-registered analyses because we believe this subset constitutes an important and timely set of results best-suited to a short-form paper.
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using eight possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in all of four ways:
  • No rescaling: The outcome variable is used as-is.
  • Z-score rescaling: We rescale the outcome variable into a Z-score according to the procedure describe in these supplementary materials.
  • Percentile rank rescaling: We rescale the outcome variable into a percentile rank. This is done by calculating the percentile rank of a given prompt relative to all other prompts submitted for the relevant target image.
  • GPT-4V baseline rescaling: We rescale the outcome variable by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image.
We intend to complete each pre-registered analysis with each of these eight possible outcome variables. Although we have not yet done so, we have no reason to believe that our results will not be robust to these different transformations of our data. However, we did not want to delay the publication of our preprint until all of these analyses were complete. Thus, we are posting preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes. As we complete more of these analyses, we will update this supplement. We also intend to post full results, data, and replication code online shortly.
F.1 Hypotheses and Results
Below, we list the hypotheses precisely as they are written in our pre-registration document. For each hypothesis, we also describe our results.
H1
There are differences in prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) across demographic attributes and other observables, such as educational background and occupational skills.
Analysis approach:
For this hypothesis, we conducted multiple ANOVA tests, one for each of the demographic variables, against the relevant outcome variable and treatment variable (the model) as the covariate. As a robustness check, we repeated the same procedure using the Kruskal-Wallis U test. To adjust for multiple testing, we used the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
Results: The results of our analysis are as follows:
  • CLIP embedding cosine similarity
    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, writing frequency, and self-reported occupational skills of critical thinking and social perceptiveness.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender, age, and the critical thinking occupational skill.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
H2
There are observable differences in the prompting techniques of successful prompt engineers and unsuccessful prompt engineers. Such prompting techniques might include the use of longer prompts, the use of structured prompting techniques, and/or specific patterns in the way that the participant iterates on their prompts over time.
Analysis approach:
To investigate this hypothesis, we pre-registered evaluating how exploration/exploitation in the prompting space is related to performance. Hence, most of the pre-registered independent variables measured the level of similarity (i.e., exploitation) across a user’s prompts. We refer to participants being more “exploitative” if they wrote prompts that are more similar to their previous prompts and more “exploratory” if they wrote prompts that deviated more from their previous prompts. As described below, we operationalized “similarity” in a variety of ways.
  1. 1.
    Positively associated with exploitation: average token sort ratio compared to the previous prompt, average cosine similarity between the embeddings of a given prompt with the previous prompt, fraction of times a prompt contains the previous prompt as an exact substring
  2. 2.
    Negatively association with exploitation: the variance of the prompt embedding, the number of topical transitions
We also pre-registered measuring each participant’s average prompt length. These variables were measured per each user and across their first 10 attempts. Their association with performance, measured in terms of each of our eight outcome variables, was estimated in a linear model with DALL-E version fixed effects.
We also conducted a similar analysis at the iteration level to answer the following question: how is exploration/exploitation associated with subsequent performance in the next attempt, and is the strength of this relationship mediated by the quality of the previous attempt? To answer this question, first, we divided user-iteration observations into 6 equal-sized brackets by performance in the previous iteration. The bracketing allows us to explore heterogeneity in prompting behavior by the quality of previous attempts. We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket. Our estimates adjusted for other covariates by matching user-iteration observations on the target image, DALL-E model, iteration and the exact quality of the previous attempt (Sävje et al., 2021).
Results: The results of our analysis are as follows:
  • CLIP embedding cosine similarity
    No rescaling: We find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing. Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance. Taken together, these findings suggest that more successful users engaged in more exploitation and wrote prompts that were similar to one another. We also found that longer prompts were associated with higher performance, as shown in the main text.
    We also find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance. In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance. Using token sort ratio with the previous prompt as the measure of exploration also generated similar results. However, when using binary measures of exploration, i.e. topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket. In particular, we did not find the non-linear pattern we observed with continuous measures of exploration in the low-performing group, likely due to the binary nature of this measure that fails to distinguish between low, medium, and high levels of exploration/exploitation.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling:Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
H3
There are differences in prompt engineering techniques (as measured through metrics such as prompt length and iteration-to-iteration token sort ratio) across demographic attributes and other observables, such as educational background and occupational skills.
Analysis approach:
To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest. To account for multiple testing, we adjusted the p-values for all these models according to Benjamini-Hochberg procedure.
Results: The results of our analysis are as follows. As a reference, the definitions of these dependent variables are provided in Section  B.B.5.
  • Prompt embedding variance: We did not find any significant differences across various demographic traits.
  • Strategic Shifts: We observed statistically significant differences based on age and reported programming frequency. Older participants and those with high programming frequency demonstrated increased topical transitions across their prompts.
  • Successive Prompt Token Sort Ratio: We found significant differences by age, programming frequency, and writing frequency. Older users, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
  • Successive similarity: We found significant differences by age, gender, generative AI outlook, programming skill/frequency and writing frequency. On average, older users, males, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, and those who write a lot, those who self-describe as being good at writing instructions and writing with a lot of clear imagery write prompts that are less similar to each other (as measured by cosine similarity) on average.
  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and writing frequency. Older users and those who write frequently are less likely to write prompts that exactly contain the previous. In contrast, males and those with a positive outlook on generative AI are more likely to keep using their previous prompt.
H4
Insofar as the output returned by a generative AI model in response to a prompt is stochastic, the subsequent prompting strategies and prompting outcomes of participants that get lower-than-expected, higher-than-expected, or approximately expected outputs in response to their first prompt are different.
Analysis approach: To test this hypothesis, we first calculated the z-score of each participant’s realized image observed during our experiment relative to the sampling distribution that we approximated for the prompt that generated that image according to the procedure outlined in Section B.B.3 (see Section C for details on calculating the z-score). This z-score quantifies the random variation in observed image quality relative to the prompt’s true underlying quality. Images with higher z-scores represent instances where the realized image quality exceeded expectations based on the prompt, while lower z-scores indicate instances where the image quality was randomly lower than expected for a given prompt. By examining the relationship between these z-scores and subsequent performance and prompting behavior, we can assess the causal impact of this stochasticity on user behavior.
To estimate the relationship between this stochasticity on subsequent prompting behavior and performance, we transformed the z-score into a trichotomous variable where any z-score less than -0.45 as “lower-than-expected,” between -0.45 and 0.45 as “expected” and greater than 0.45 as “higher-than-expected.” We then perform two-sample t-tests with comparing the three groups’ relative performance. We also estimate a linear model as a robustness check, regressing participant performance on the trichotomous z-score variable with treatment arm fixed effects. We also repeat this analysis, treating the z-score as a continuous variable without the trichotomous transformation. Finally, we also perform this analysis at the user level with the trichotomous variable, where only the z-score of the observed image of the first prompt is measured, and we test if that affects the average performance of all subsequent user attempts.
Results: The results of our analysis are as follows, with each outcome variable clearly labeled.
  • CLIP embedding cosine similarity
    No rescaling: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt when conducting two-sample t-tests stratified by treatment arm. When we test differences in the trichotomous variable, the difference between the top bracket and the other two brackets is significant, but we do not find significant differences between the bottom and middle brackets. However, we do not find a statistically significant relationship between the observed z-score of the prompt and performance in the next iteration when we estimate a linear model. This may be due to the non-linearity of the relationship between these variables. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average subsequent performance.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • Prompting Behaviors
    Prompt Length: Forthcoming
    Successive Similarity: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in successive similarity. When we compare the lowest z-score bracket (worse than expected images) to the highest z-score bracket (better than expected images), we observe a statistically significant difference as well. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average successive similarity.
    Successive Prompt Token Sort Ratio: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in the Token Sort Ratio between successive prompts. When we compare the lowest z-score bracket to the highest z-score bracket, we observe a statistically significant difference. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average successive prompt token sort ratio.
    Successive Prompt ’Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in the likelihood that the subsequent prompt contains the previous prompt. When we compare the lowest z-score bracket to the highest z-score bracket, we observe a statistically significant difference. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on the average probability of a prompt containing the preceding prompt.
H5
Average prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) and prompting strategies will depend on the capacity of the model that participants are interacting with.
Analysis approach:
To test this hypothesis, we conducted three two-sample t-tests that compare the prompting performance and prompting behavior of the three possible pairs of treatment assignments. All outcome variables are specified in the results below. For robustness, we repeat these analyses with ANOVA to test whether any of the three treatment arms has an effect on the same variables. To account for multiple testing, we adjusted the p-values for tests with the Benjamini-Hochberg procedure.
Results: The results of our analysis are as follows:
  • CLIP embedding cosine similarity
    No rescaling: Forthcoming
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts. Those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find comparatively identical resultsin terms of statistical significance. Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Logically, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • Prompting Behaviors
    Mean prompt Length: We found statistically significant differences in prompt length between treatment arms. Participants using DALL-E 2 used significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups. There was no significant difference in prompt length between the two DALL-E 3 groups. ANOVA results confirmed a significant effect of DALL-E version on prompt length.
    Aggregate Similarity: We found no statistically significant differences between any of the treatment arms. This suggests that the overall variability in prompts was similar across all three DALL-E versions. ANOVA results confirmed no significant effect of DALL-E version on this measure.
    Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on successive prompt similarity.
    Successive Prompt Token Sort Ratio: We found no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
    Successive Prompt ’Contains Previous Prompt’ Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
H6
Variability in participants’ ability to prompt engineer effectively and prompting strategies will depend on the capacity of the model that participants are interacting with.
Analysis approach:
To test this hypothesis, we conducted two analyses. First, would conducted F-tests comparing the variance of participant performance and prompting behaviors between all 3 pairs of models. Second, we estimated the quantile treatment effects (QTEs) between all 3 pairs models on participant performance and prompting behaviors. We also pre-registered our intent to visually inspect whether the QTEs we observe are consistent with dispersion/“inequality” being reduced or increased when participants use different models (e.g., positive effects for low quantiles and negative/null effects for high quantiles would be consistent with inequality reduction).
Results:
  • CLIP embedding cosine similarity
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling: Forthcoming
    Z-score rescaling: We observed statistically significant reductions in the variance ratio of prompt-length, participant-level prompt embedding variance, and the percentage of prompts containing the previous prompt when comparing DALL-E 2 to DALL-E 3. We also found statistically significant reductions in the variance ratio of prompt length and participant-level prompt embedding variance when comparing DALL-E 2 and DALL-E 3 (Revised). Our QTE results comparing performance under DALL-E 2 and DALL-E 3 appear in the main text. The results we have seen when comparing DALL-E 3 (Revised) to DALL-E 2 are qualitatively similar, although the QTEs are attenuated due to the prompt revision. Almost all QTEs when comparing DALL-E 3 and DALL-E 3 (Revised) are positive.
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
H7
As participants repeatedly try to complete a task with a given model, the quality of their attempts will increase, and the extent to which the quality increases varies as a function of model capacity.
Analysis approach:
In order to test for differences in prompting performance across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata. We run this test on each participant’s performance improvement, defined as the performance gap between the initial iteration and the best-performing iteration.
  • CLIP embedding cosine similarity
    No rescaling: We do not find evidence for statistically significant relationships between DALL-E 2 versus DALL-E 3 (Verbatim), nor between DALL-E 3 (Verbatim) and DALL-E 3 (Revised). However, we do find a statistically significant difference between DALL-E 2 and DALL-E 3 (Revised), whereby DALL-E 2 users showed greater improvement in cosine similarity compared to DALL-E 3 (Revised) users between their first and best attempts. The latter relationship holds even after adjusting for Benjamini-Hochberg multiple testing.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
  • DreamSim
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
H8
The extent to which participants can recreate images using models such as DALL-E 2/3 will vary across images.
Analysis approach:
To evaluate this hypothesis, we pre-registered and performed two analyses. First, we compared the extent to which different images can be replicated using GPT-4V—a multimodal generative AI model that can take in both text and images and output text. For each of the 15 target images, we prompted GPT-4V to “Write a DALL-E {2, 3} prompt to recreate this image verbatim as closely and as detailed as possible.” We do this once for DALL-E 2 and once for DALL-E 3, generating two “AI prompts” per target image. We then provided these AI prompts to DALL-E 2, DALL-E 3 (Verbatim), and DALL-E 3 (Revised), respectively, 20 times each, generating 60 replicated images per target image (the DALL-E 3 AI prompt was sent to both DALL-E 3 (Verbatim) and (revised)). For each image, we measure the cosine similarity of the CLIP embedding vectors to those of the relevant target image. We then average these 60 similarity measures by target image, resulting in a mean similarity score that represents GPT-4V’s ability to generate prompts that recreate each target image. Second, we simply measured the similarity of every participant-generated image to that of the relevant target image, and averaged over these participant-generated similarities.
Results:
  • CLIP embedding cosine similarity: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
    1. 1.
      Business Image #3
    2. 2.
      Business Image #5
    3. 3.
      Business Image #1
    4. 4.
      Photography Image #1
    5. 5.
      Design Image #2
    6. 6.
      Business Image #2
    7. 7.
      Photography Image #5
    8. 8.
      Design Image #4
    9. 9.
      Business Image #4
    10. 10.
      Design Image #1
    11. 11.
      Photography Image #2
    12. 12.
      Design Image #5
    13. 13.
      Design Image #3
    14. 14.
      Photography Image #3
    15. 15.
      Photography Image #4
    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for GPT-4V to replicate) was $CoSim=0.944$ for Business Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $CoSim=0.734$ for Photography Image #4.
    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
    1. 1.
      Business Image #3
    2. 2.
      Business Image #5
    3. 3.
      Business Image #1
    4. 4.
      Photography Image #5
    5. 5.
      Design Image #2
    6. 6.
      Design Image #4
    7. 7.
      Business Image #2
    8. 8.
      Design Image #5
    9. 9.
      Design Image #1
    10. 10.
      Photography Image #1
    11. 11.
      Design Image #3
    12. 12.
      Photography Image #2
    13. 13.
      Photography Image #3
    14. 14.
      Photography Image #4
    15. 15.
      Business Image #4
    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, and the lowest score (the hardest image for participants to replicate was $CoSim=0.669$ for Business Image #4.
  • DreamSim: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
    1. 1.
      Design Image #3
    2. 2.
      Business Image #2
    3. 3.
      Business Image #4
    4. 4.
      Design Image #2
    5. 5.
      Business Image #1
    6. 6.
      Business Image #5
    7. 7.
      Design Image #4
    8. 8.
      Business Image #3
    9. 9.
      Photography Image #1
    10. 10.
      Photography Image #5
    11. 11.
      Photography Image #2
    12. 12.
      Design Image #1
    13. 13.
      Design Image #5
    14. 14.
      Photography Image #3
    15. 15.
      Photography Image #4
    The maximum 1-DreamSim score in the above list (the easiest image for GPT-4V to replicate) was $\tilde{D}=0.75$ for Design Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $\tilde{D}=0.40$ for Photography Image #4.
    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
    1. 1.
      Business Image #3
    2. 2.
      Business Image #2
    3. 3.
      Business Image #5
    4. 4.
      Business Image #1
    5. 5.
      Design Image #2
    6. 6.
      Design Image #3
    7. 7.
      Photography Image #2
    8. 8.
      Photography Image #5
    9. 9.
      Design Image #5
    10. 10.
      Design Image #4
    11. 11.
      Business Image #4
    12. 12.
      Photography Image #1
    13. 13.
      Photography Image #4
    14. 14.
      Design Image #1
    15. 15.
      Photography Image #3
    The maximum 1-DreamSim score in the above list (the easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, and the lowest score (the hardest image for participants to replicate was $\tilde{D}=0.356$ for Photography Image #4.
H-Exploratory
In our pre-registration, we also mention a number of exploratory analyses that are not described with the same level of detail. Multiple of these exploratory analyses appear in our main text. These pre-registered exploratory analyses are copied verbatim below:
“We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate. We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
We anticipate that we may conduct additional analysis of the prompts submitted by participants (and how these prompts evolve over the course of a session). Furthermore, we might explore the tips that participants provide after completing the task on how to prompt engineer effectively.
We also may take original and revised prompts submitted to DALL-E 3 treatment arms and submit them to DALL-E 2 (and vice versa) to see how participants would have counterfactually performed under different treatment assignments than the one to which they were assigned.”
F.2 Deviations From Pre-registration
Here, we also report any deviations from our pre-registered analysis. By and large, these deviations occurred because certain aspects of our pre-registration were not appropriate from a statistical analysis perspective, or were infeasible.
  • In our pre-registration, we had anticipated using t-tests and Mann-Whitney U tests for many hypotheses. However, this turned out to be impossible for many variables since most of the demographic traits have multiple categories. Thus, we applied ANOVA and Kruskal-Wallis tests instead.
  • In our pre-registration, we had forgotten to include Education and GenAI outlook in our list of demographics with respect to which we would measure task performance heterogeneity. We decided to include these important variables in our analysis, despite the accidental omission.
  • In addition to our pre-registered prompt exclusion criteria, we removed additional prompts that did not appear to be “good-faith efforts” to complete our task based on the text of those prompts (see Section B.4 for more details). We find that our results are the robustness to the inclusion of prompts that were removed by this procedure.
  • Because they were easier to implement, we ran t-tests rather than z-tests to conduct our tests of H5. Because the two tests are asymptotically equivalent, we do not believe this will make a difference in our analysis.
  • When conducting our analyses to test H4, we observed that the distribution of the z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values. To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute z-scores greater than 3, which constituted 2.5% of the user-attempt observations. For robustness, we also conducted the stratification analysis including these outliers and found similar results, with the exception that the Average Treatment Effect (ATE) for performance between the top and middle brackets was only marginally significant.

### v2

@@
 Abstract
-In an online experiment with N = 1,891 participants, we collected and analyzed over 18,000 prompts to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve.
+In an online experiment with N = 1,893 participants, we collected and analyzed over 18,000 prompts and over 300,000 images to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve.
 Each participant in our experiment was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor DALL-E 3, or a version of DALL-E 3 with automatic prompt revision.
@@
 If this comes to pass, the benefits of generative AI will be accessible to a much larger population.
-In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,891{}$ participants recruited on Prolific.
-[^1] This paper presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI.
-In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) DALL-E 3, and (3) a version of DALL-E 3 with automatic LLM-based prompt revisions (henceforth referred to as “DALL-E 3 with revision”).
-These prompt revisions were not made visible to participants during the experiment.
+In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,893{}$ participants recruited on Prolific.
+[^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI.
+In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”).
+The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.
 [^2] Automatic prompt revision is the default behavior of the DALL-E 3 API endpoint.
@@
 The majority of our analyses focus on the DALL-E 2 and DALL-E 3 treatment arms, as they are more directly comparable.
-We first estimate the effect of access to a more advanced generative model on task performance by evaluating how similar each generated image is to the relevant target image.
+We first estimate the effect of access to a more advanced generative model on task performance.
 [^3] All main text analyses use ”replay” data from Figure 2 for consistency and to avoid model drift issues.
@@
 See the SI for details.
-To construct a quantitative measure of image similarity, we generated CLIP embedding vectors (Radford et al., 2021) for each participant-generated image and the relevant target image and then calculated the cosine similarity ($CoSim$) between those vectors (see SI for further details on how we quantify image similarity and account for the stochasticity of model output).
+Given the stochastic nature of generative AI models, our measure of task performance is the expected similarity of images generated from each prompt to the relevant target image.
+To construct a quantitative measure of this expected image similarity for each prompt, we first used each participant-generated prompt to produce 10 images.
+We then computed CLIP embedding vectors (Radford et al., 2021) for both target images and prompt-generated images.
+For each participant prompt and its corresponding target image, we then calculated the average cosine similarity ($CoSim$) between the embedding vectors of the randomly sampled generated images and the target image vector (see SI for further details on how we quantify expected image similarity).
 Unsurprisingly, we find that providing access to a more advanced generative model improved task performance.
-Participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.20$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
+Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2.
 Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in similarity to the target image.
-We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0008$, $p=0.042$).
+We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression).
 In other words, the performance gap widened between DALL-E 2 and DALL-E 3 as participants gained more experience with the respective models, potentially due to changes in prompting.
-To investigate how participants’ prompts changed in response to the superior model’s capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated.
-We find that participants assigned the superior model wrote prompts that were 24% longer ($\Delta Words=6.6$, $p<10^{-4}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-7}$).
-Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% superior vs.
-49% inferior; $p$ = $0.148$).
+To investigate how participants’ prompts changed in response to DALL-E 3’s more advanced capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated.
+We find that participants assigned DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$).
+Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs.
+49% DALL-E 2; $p$ = $0.215$).
 This suggests that prompts submitted to DALL-E 3 were longer because they conveyed additional descriptive information, not because they were padded with superfluous words.
-We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0166$, $p=0.02$) and in the aggregate ($\beta=0.0169$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses).
+We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses).
 These findings indicate that exposure to different models did in fact induce differences in prompting, and that these differences grew over successive replication attempts.
-Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects.
-The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line).
-Effects are shown relative to a baseline of prompts written for and replayed on the inferior model.
-Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
 To understand the extent to which the ATE we observe is attributable to these behavioral differences, as opposed to DALL-E 3’s more advanced capabilities, we conducted another exploratory analysis in which we regenerated images with (or “replayed”) all participant prompts on both DALL-E 2 and DALL-E 3, regardless of the model the participant used during the experiment.
-This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by inferior model participants on the superior model, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by superior model participants and the average similarity to the target of all prompts generated by inferior model participants, both replayed on the superior model.
-Figure 2 shows that the model effect (red) accounts for 53% of the ATE ($\Delta CoSim=0.00864$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 47% of the ATE ($\Delta CoSim=0.00769$, $p=0.0066$).
+This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by DALL-E 2-assigned participants on DALL-E 3, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by DALL-E 3-assigned participants and the average similarity to the target of all prompts generated by DALL-E 2-assigned participants, both replayed on DALL-E 3.
+Figure 2 shows that the model effect (red) accounts for 51% of the ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 48% of the ATE ($\Delta CoSim=0.00788$, $p=0.024$).
 Thus, nearly half of the ATE is driven by endogenous changes in prompting strategy that emerged after exposure to DALL-E 3.
-The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of the superior model; when replaying both sets of prompts on the inferior model, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0018$, $p=0.52$).
+The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of DALL-E 3; when replaying both sets of prompts on DALL-E 2, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0020$, $p=0.56$).
 In other words, DALL-E 3 was more capable than DALL-E 2 at effectively rendering the additional information contained in the longer and more descriptive prompts (Betker et al., 2023), and participants adapted their prompting behavior accordingly.
-One might think that the need to adapt one’s prompts to the capabilities of different generative models is obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023).
+One might think that the need to adapt one’s prompts to the capabilities of different generative models can be obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023).
 We use data from our third treatment arm—DALL-E 3 with revision—to test this hypothesis.
-We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0068$; $p=0.013$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [22%, 94%]).
+We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
 Importantly, the positive impact of access to DALL-E 3 with revision was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2).
@@
 Discussion
-In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance to increase, not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect).
-More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render this additional information more effectively than DALL-E 2.
+Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects.
+The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line).
+Effects are shown relative to a baseline of prompts written for and replayed on DALL-E 2.
+Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
+In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance on our image replication task to increase.
+This was not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect).
+More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render that additional information more effectively than DALL-E 2.
 These changes in prompting emerged despite the fact that participants were unaware of which model they were assigned.
 Although automated LLM-revision features are designed to and have been shown to help people generate higher-quality content (Betker et al., 2023; Li et al., 2024), we also find that LLM-based prompt revision failed to further improve performance; instead, automatic prompt revision reduced the benefits of using DALL-E 3.
-Overall, we find that shifting to a superior model, whether or not accompanied by LLM-based revisions, did not eliminate the need for prompting.
+Overall, we find that shifting to a more capable model, whether or not accompanied by automatic LLM revisions, did not eliminate the need for prompting.
 The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3.
@@
 We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells.
-We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test for homogeneity of proportions across all cells.
-With $\chi^{2}=0.374,\;df=28$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
+We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test across all cells.
+With $\chi^{2}=7.056,\;df=44$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
 $H_{0}:p_{1}=p_{2}=\cdots=p_{45}$
@@
 [^4] See here and here for online documentation.
-We modified the recommended preflix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
+We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
 “I NEED to test how the tool works with my prompt as it is written. DO NOT add any detail; just use it AS IS:”
@@
 We randomly assigned participants to 1 of 15 target images.
-The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and photography.
+The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and architectural photography.
 We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts.
@@
 This left us with 2,029 participants and 24,123 prompts.
-  • We next removed participants from our sample if they failed to submit at least 10 prompts or if they did not make a “good-faith” effort to complete the task as requested.
+  • We next removed participants from our sample if they failed to submit at least 10 prompts or if a participant submitted the same prompt at least five times in a row at any point during the task.
 Both of these exclusion criteria were pre-registered.
-A good-faith effort was not made if a participant submitted the same prompt at least five times in a row at any point during the task.
 These exclusion criteria were also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” To avoid reward hacking, we did not specify the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants.
@@
 This affected 34 prompts across all three treatment arms, and 20 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms, approximately 0.1% of the original data.
-    • While analyzing our data, we found that our sample contained a number of prompts that were not “good-faith efforts” to complete our task, despite the use of the exclusion criteria described above.
-Thus, we used the following process to identify and remove additional “low-effort” prompts.
+  • Our final sample included 1,893 participants and 18,560 prompts.
+    B.4.1 “Off-Topic” Robustness Check
+    While analyzing our data, we found that our sample contained a number of “off-topic” prompts that did not seem related to the task.
+As a robustness check on our main results, we used the following process to systematically identify and remove “off-topic” prompts.
 First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model.
@@
 Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector.
-We observed that the prompts that were dropped were often off-topic and likely were not good-faith efforts to complete our task.
 This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms.
-Our results are robust to the inclusion of these filtered prompts.
-  • Our final sample included 1,891 participants and 18,152 prompts.
+All of our main text results are robust to the exclusion of these “off-topic” prompts.
 B.5 Dependent Variables
@@
   • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023).
-The authors claim that relative to something like CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity.
+The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity.
 Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score.
@@
 We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt.
-In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9952$, 95% CI: [0.99507, 0.99543]).
+In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9954$, 95% CI: [0.99528, 0.99560]).
 B.5.3 Embedding-based Prompt Similarity
@@
 Starting with each participant’s second prompt, we record whether each prompt $p_{i,n}$ contains the immediately preceding prompt $p_{i,n-1}$ as an exact substring.
-B.5.6 Inversions of Successive Measures
-[…]
-B.5.7 Prompt Composition
+B.5.6 Prompt Composition
 We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt.
@@
 The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
-B.5.8 Strategic Shifts
+B.5.7 Strategic Shifts
 In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting.
@@
 (2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift.
-We instead use the task-level mean because in our setting, it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
+We instead use the task-level mean because in our setting, as it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
 Appendix C Methods
@@
 In the main text, we also showed that performance increases across the attempts.
-To test whether our results are in some way due to this image-level or attempt-level variations, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment.
+To test whether our results are in some way due to this image-level or attempt-level variation, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment.
 Formally, this is:
@@
 We also use the rescaling above to test the robustness of our DreamSim-based analyses.
+Almost all the robustness analysis reported here use the Z-score scaled measure of performance within each image-attempt set.
+The only exception is any analysis that examines improvements and prompts across attempts, e.g.
+figure 1B, where Z-scores are computed only within each target image, and across all attempts of that image.
 C.3 Accounting for Model Stochasticity
 The output that generative AI models return in response to a given prompt is stochastic.
-The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E 3 API at the time of our experiment.
+The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E API at the time of our experiment.
 To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms.
 We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples.
+Given the replicated images, we can also calculate the standard deviation of cosine similarity induced by this stochasticity.
+With the expected cosine similarity and its standard deviation per prompt, we can compute a normalized Z-score for the observed image relative to its replication distribution.
+This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and will be used for further analysis in section F.
 We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text.
@@
 As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts.
-This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the under the ”replaying” procedure (Section B.B.3.3).
+This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section B.B.3.3).
 Appendix D Main Text Analyses
@@
 [^5] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified.
-To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors:
+To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
 $Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (6)
 The coefficient estimates generated by this model are:
-  • $\hat{\beta_{1}}=0.0015$, $\hat{SE}(\beta_{1})=0.0003$, $p=9.678\times 10^{-7}$
-  • $\hat{\beta_{2}}=0.0126$, $\hat{SE}(\beta_{2})=0.003$, $p=1.7\times 10^{-4}$
-  • $\hat{\beta_{3}}=0.00086$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0418$
+  • $\hat{\beta_{1}}=0.0011$, $\hat{SE}(\beta_{1})=0.0003$, $p=0.0004$
+  • $\hat{\beta_{2}}=0.0120$, $\hat{SE}(\beta_{2})=0.0037$, $p=0.0013$
+  • $\hat{\beta_{3}}=0.0010$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0227$
 The overall ATEs that we report between different pairs of treatment arms (DALL-E 2, DALL-E 3, and DALL-E 3 with revisions) in the main text are estimated from a two-way fixed effect (iteration and target image) model per each pair.
@@
 Figure 1C compares the prompt length and prompt similarity of the two models.
-To generate these results, we first remove any prompt that does not resemble a good-faith attempt according to the sample construction procedure detailed in Section B.B.4.
+To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section B.B.4.
 The prompt length is the average number of words per model and iteration stratified by the reference image.
@@
 The color scale in figure 1C shows the stratified average similarity to the previous prompt across all users per each model.
-We find that superior model users write prompts that, on average, have $\beta=0.0166$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.02$).
-Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from the superior model participants were, on average, more similar than the prompts of the inferior model participants.
+We find that superior model users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
+Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants.
 For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable.
-When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=-0.0169$ smaller than inferior users ($p=0.008$).
+When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$).
 Standard errors are cluster-robust at the participant level.
@@
 E.1.1 Overall ATEs
-In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.261$ standard deviations (95% CI = [0.180, 0.342]) closer to the target image ($\Delta DreamSim=0.0304$, $p<10^{-9}$) than those produced by participants using DALL-E 2 (the inferior model).
+In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.238$ standard deviations (95% CI = [0.152, 0.324]) closer to the target image ($\Delta DreamSim=0.0306$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
 E.1.2 Figure 1
@@
 The coefficient estimates generated by this analysis are:
-  • $\hat{\beta_{1}}=0.0039$, $\hat{SE}(\beta_{1})=0.00048$, $p=4.3\times 10^{-16}$
-  • $\hat{\beta_{2}}=0.0179$, $\hat{SE}(\beta_{2})=0.0060$, $p=0.0029$
-  • $\hat{\beta_{3}}=0.0023$, $\hat{SE}(\beta_{3})=0.00071$, $p=0.001$
+  • $\hat{\beta_{1}}=0.0034$, $\hat{SE}(\beta_{1})=0.0005$, $p=1.3\times 10^{-12}$
+  • $\hat{\beta_{2}}=0.0200$, $\hat{SE}(\beta_{2})=0.0061$, $p=0.0011$
+  • $\hat{\beta_{3}}=0.0024$, $\hat{SE}(\beta_{3})=0.0007$, $p=0.0009$
 E.1.3 Figure 2
 Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text.
-The model effect accounts for 55% of the ATE ($\Delta DreamSim=0.0168$, $p<10^{-10}$), whereas the prompting effect accounts for 45% of the ATE ($\Delta CoSim=0.01364$, $p=0.0053$).
+The model effect accounts for 54.4% of the ATE ($\Delta DreamSim=0.0166$, $p<10^{-7}$), whereas the prompting effect accounts for 45.4% of the ATE ($\Delta DreamSim=0.01390$, $p=0.014$).
 E.2 Z-score-based Analysis
-As we discuss in Section B, we repeat all main-text analyses with the within-image Z score of CLIP-cosine similarity to better account for variation between images.
+As we discuss in Section C.2, we repeat all main-text analyses with the within-image-attempt Z-score of CLIP-cosine similarity to better account for variation between images and attempts.
+When comparing across attempts, the Z-score is computed within each image, as mentioned in section C.2.
 The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
 E.2.1 Overall ATEs
-As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.20$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.20, 95% CI = [0.126, 0.284]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
-Standard errors are cluster robust by participants.
+As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.19$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.19, 95% CI = [0.100, 0.271]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2 (the inferior model).
+Standard errors are clustered at the participant level.
 E.2.2 Figure 1
-On average, participants using the superior model produced images that were $z=0.20$ standard deviations closer (the ATE) to the target image than those using the inferior model.
+On average, participants using the superior model produced images that were $z=0.19$ standard deviations closer (the ATE) to the target image than those using the inferior model.
 Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
 $\text{ZScore}_{i,n}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\epsilon_{i,n}$ (9)
-  • $\hat{\beta_{1}}=0.02012$, $\hat{SE}(\beta_{1})=0.0041$, $p=1.19\times 10^{-6}$
-  • $\hat{\beta_{2}}=0.1439$, $\hat{SE}(\beta_{2})=0.0466$, $p=0.002$
-  • $\hat{\beta_{3}}=0.01253$, $\hat{SE}(\beta_{3})=0.0058$, $p=0.03$
+  • $\hat{\beta_{1}}=0.0129$, $\hat{SE}(\beta_{1})=0.0038$, $p=0.0007$
+  • $\hat{\beta_{2}}=0.1250$, $\hat{SE}(\beta_{2})=0.0457$, $p=0.0064$
+  • $\hat{\beta_{3}}=0.0128$, $\hat{SE}(\beta_{3})=0.0053$, $p=0.015$
 E.2.3 Figure 2
-When we decompose the ATE into the model effect ($z=0.0895$; $p=1.02\times 10^{-5}$) and prompting effect ($z=0.1139$; $p=0.0044$), they account for 44% and 56% of the treatment effect, respectively.
-And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.03$; $p=0.43$).
-In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z score of the cosine similarity.
+When we decompose the ATE into the model effect ($z=0.0791$; $p=8.35\times 10^{-6}$) and prompting effect ($z=0.1046$; $p=0.016$), they account for 43% and 56% of the treatment effect, respectively.
+And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.033$; $p=0.45$).
+In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z-score of the cosine similarity.
 Appendix F Pre-registration
@@
 Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses.
-In our pre-registration, we declared the intent to conduct each of our analyses using eight possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in all of four ways:
+In our pre-registration, we declared the intent to conduct each of our analyses using six777see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
+possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
   • No rescaling: The outcome variable is used as-is.
@@
 This is done by calculating the percentile rank of a given prompt relative to all other prompts submitted for the relevant target image.
-  • GPT-4V baseline rescaling: We rescale the outcome variable by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image.
 We intend to complete each pre-registered analysis with each of these eight possible outcome variables.
@@
 However, we did not want to delay the publication of our preprint until all of these analyses were complete.
-Thus, we are posting preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes.
+Thus, we are posting a preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes.
 As we complete more of these analyses, we will update this supplement.
@@
   • CLIP embedding cosine similarity
-    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, writing frequency, and self-reported occupational skills of critical thinking and social perceptiveness.
-    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills are, on average, better at the task.
+    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, active listening and quality control.
+    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as job skills, and report little usage of generative AI or imagery writing skill are, on average, better at the task.
+Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with self-reported occupational skill of quality control, and those who programmed more frequently performed worse on our experiment task.
+    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
+In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
+    Z-score rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, quality control, technology design, social perceptiveness and troubleshooting.
+    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and troubleshooting as job skills, and report little usage of generative AI, little imagery writing skill, some programming skill, and some instruction writing skill are, on average, better at the task.
+Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with a graduate degree, those with self-reported occupational skill of quality control and technology design, and those who programmed more frequently performed worse on our experiment task.
+    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
+In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
+    Percentile rank rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, and self-reported occupational skill of critical thinking.
+    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as a job skill, and report little usage of generative AI, and some programming skill are, on average, better at the task.
+Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently performed worse on our experiment task.
+    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
+In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
+  • DreamSim
+    No rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skill, and self-reported occupational skills of critical thinking, learning strategies, technology design and quality control.
+    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task.
 Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
-In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender, age, and the critical thinking occupational skill.
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
-  • DreamSim
-    No rescaling: Forthcoming
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies occupational skill.
+    Z-score rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skills, education, and self-reported occupational skills of critical thinking, social perceptiveness, learning strategies, technology design and quality control.
+    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task.
+Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, those with graduate degrees, and those with technology design, quality control and learning strategies as occupational skills performed worse on our experiment task.
+    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
+    Percentile rank rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery writing skill, education, and self-reported occupational skills of social perceptiveness, learning strategies, and technology design.
+    When we fit a linear model to evaluate directionality, we find that those who use social perceptiveness as a job skill, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task.
+Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, and those with technology design and learning strategies as occupational skills performed worse on our experiment task.
+    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, generative AI use, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
 H2
@@
   • CLIP embedding cosine similarity
-    No rescaling: We find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing.
+    No rescaling: At the user-level, we find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing.
 Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance.
@@
 We also found that longer prompts were associated with higher performance, as shown in the main text.
-    We also find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance.
+    At the user-prompt level, we find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance.
 In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance.
@@
 topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket.
-In particular, we did not find the non-linear pattern we observed with continuous measures of exploration in the low-performing group, likely due to the binary nature of this measure that fails to distinguish between low, medium, and high levels of exploration/exploitation.
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+As these are binary measures of exploration, we could not replicate the non-linearity we observed with continuous measures of exploration in the low-performing group.
+However, we do find that topical transitions (more exploration) leads to an overall drop in performance and the drop becomes larger for higher performance in the previous attempt.
+We find a similar pattern when measuring exploitation by whether a prompt contains the previous one.
+Performance is overall higher if the prompt includes the previous one and the improvement is larger as the previous performance gets higher.
+    Z-score rescaling: We find the same results as those described for the unscaled cosine similarity above.
+The main difference is that the non-linearity between performance and continuous measures of exploitation, i.e.
+TSR ratio and cosine similarity with previous prompt, observed at the user-prompt level becomes starker for the bottom two brackets of previous performance.
+For low-performing prompts in the previous attempt, the optimal exploitation level is in the middle and the performance greatly deteriorates for high and low levels of exploitation.
+    Percentile rank rescaling: We find the same results as described above, however the non-linearity in exploration/exploitation vs performance in the bottom two brackets of previous performance is now even stronger than those observed with Z-score rescaled cosine similarity as the performance measure.
   • DreamSim
-    No rescaling:Forthcoming
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    No rescaling: We find exactly the same results as those described for the unscaled cosine similarity above.
+    Z-score rescaling: We find the same results as those described for the Z-score cosine similarity above.
+In particular, we again find that the non-linearity between performance and continuous measures of exploitation observed at the user-prompt level is stark for the bottom two brackets of previous performance.
+    Percentile rank rescaling: We find the same results as described above for cosine similarity percentile rank rescaling.
 H3
@@
 To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest.
-To account for multiple testing, we adjusted the p-values for all these models according to Benjamini-Hochberg procedure.
+To account for multiple testing, we adjusted the p-values for all these models according to the Benjamini-Hochberg procedure.
 Results: The results of our analysis are as follows. As a reference, the definitions of these dependent variables are provided in Section B.B.5.
   • Prompt embedding variance: We did not find any significant differences across various demographic traits.
-  • Strategic Shifts: We observed statistically significant differences based on age and reported programming frequency.
-Older participants and those with high programming frequency demonstrated increased topical transitions across their prompts.
-  • Successive Prompt Token Sort Ratio: We found significant differences by age, programming frequency, and writing frequency.
-Older users, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
-  • Successive similarity: We found significant differences by age, gender, generative AI outlook, programming skill/frequency and writing frequency.
-On average, older users, males, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, and those who write a lot, those who self-describe as being good at writing instructions and writing with a lot of clear imagery write prompts that are less similar to each other (as measured by cosine similarity) on average.
-  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and writing frequency.
-Older users and those who write frequently are less likely to write prompts that exactly contain the previous.
-In contrast, males and those with a positive outlook on generative AI are more likely to keep using their previous prompt.
+  • Strategic Shifts: We observed statistically significant differences based on age, reported programming and instructional writing frequencies.
+Younger participants, those with low programming frequency and those with some instructional writing frequency demonstrated decreased topical transitions across their prompts.
+  • Successive Prompt Token Sort Ratio: We found significant differences by age, education, programming frequency, and imagery/instructional writing frequencies.
+Older users, those with post-graduate degrees, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
+  • Successive similarity: We found significant differences by age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency and some occupational skills.
+On average, older users, males, those with post-graduate degrees, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, those who report frequently writing instructions or imagery, and those with self-reported occupational skills of critical thinking and social perceptiveness write prompts that are less similar to each other (as measured by cosine similarity) on average.
+  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and imagery writing frequency.
+Older users and those with neutral outlook toward generative AI are less likely to write prompts that exactly contain the previous.
+In contrast, males and those with some imagery writing skills are more likely to keep using their previous prompt.
 H4
@@
   • CLIP embedding cosine similarity
-    No rescaling: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt when conducting two-sample t-tests stratified by treatment arm.
-When we test differences in the trichotomous variable, the difference between the top bracket and the other two brackets is significant, but we do not find significant differences between the bottom and middle brackets.
-However, we do not find a statistically significant relationship between the observed z-score of the prompt and performance in the next iteration when we estimate a linear model.
-This may be due to the non-linearity of the relationship between these variables.
-Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average subsequent performance.
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    No rescaling: We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt.
+When we test differences in the trichotomous variable, the difference between the top bracket and the bottom bracket in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than the bottom bracket.
+But we do not find significant differences between the bottom and middle brackets and between the top and middle brackets.
+When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration.
+This is due to the non-linearity of the relationship between these variables that exists separately in the negative and positive range of realized Z-score.
+Additionally, when we estimate the effect at the user-level, we find differences between the top and middle brackets of the first prompt realized Z-score.
+However, these differences are barely significant ($p=0.048$) and do not account for multiple testing.
+When fitting a linear model at the user-level, we do not find a a statistically significant relationship between the observed Z-score of the first prompt and average performance in the subsequent attempts.
+    Z-score rescaling: The results are very similar to those described for the unscaled cosine similarity above.
+In particular, we find a statistically significant difference between the top bracket and the bottom bracket, with the top bracket having a higher performance in the next attempt than the bottom bracket.
+We don’t find statistically significant effects when comparing the middle with top or bottom brackets, or when fitting a linear model.
+In contrast to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when comparing the brackets on the realized Z-score of the first image nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
+    Percentile rank rescaling: We find the exact same results as those explained above for Z-score cosine similarity.
   • DreamSim
-    No rescaling: Forthcoming
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    No rescaling: We find similar evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt.
+The difference between the top bracket and the bottom/middle brackets in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than both the bottom and middle brackets.
+But we do not find significant differences between the bottom and middle brackets.
+When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration, likely due to the non-linearity that we also observed between performance measured as cosine similarity and realized Z-score of the previous prompt.
+When we estimate the effect at the user-level, we do not find any statistically significant relationship between realized quality of the first prompt and performance in subsequent attempts, either using the trichotomous variable on Z-score or using a linear model.
+    Z-score rescaling: The results are very similar to those described for the unscaled dreamsim score above.
+In particular, we find a statistically significant difference between the top bracket and both the bottom and middle brackets, with the top bracket having a higher performance in the next attempt.
+We don’t find statistically significant effects when comparing the middle with the bottom bracket, or when fitting a linear model.
+Similar to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when using the trichotomous variable nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
+    Percentile rank rescaling: We find the exact same results as those explained above for Z-score dreamsim similarity, with the only difference that at the user-attempt level, we observe a statistically significant difference only between the top and bottom brackets.
+The top bracket in realized quality no longer has a statistically significant improvement in the performance of the next attempt when compared to the middle bracket.
   • Prompting Behaviors
-    Prompt Length: Forthcoming
-    Successive Similarity: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in successive similarity.
-When we compare the lowest z-score bracket (worse than expected images) to the highest z-score bracket (better than expected images), we observe a statistically significant difference as well.
+    Prompt Length: e find statistically significant evidence that increases in the Z-score of a realized image causes the subsequent prompt to be longer (either in terms of words or characters).
+When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with on average longer prompts in the top bracket.
+However, we do not observe statistically significant differences between either the top or bottom brackets with the middle bracket.
+When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant.
+A unit increase in the Z-score realized quality of an image causes the next prompt to be longer by about 0.3 words on average.
+Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the length of the subsequent prompts, when using the trichotomous variable or when fitting a linear model.
+    Successive Similarity: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in successive similarity of the prompt.
+When we test the differences using the trichotomous variable, the difference between the top Z-score bracket (better than expected images) and both the bottom (worse than expected images) and middle (about expected images) brackets is statistically significant, with the top bracket leading to higher similarity with the previous prompt.
+However, we do not observe a statistically significant difference between the bottom bracket with the middle bracket.
+When we estimate a linear model, we find a similar relationship that is both positive and statistically significant.
+An increase in the realized quality (Z-score) of an image causes the next prompt to be more similar to the previous one.
+When we estimate the effect at the user-level, we do not find statistically significant effects of the first prompt realized quality on average successive similarity, neither when using the trichotomous variable nor when fitting a linear model.
+    Successive Prompt Token Sort Ratio: The results are identical to the case of successive cosine similarity described above.
+We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt causes an increase in the Token Sort Ratio between successive prompts, using both the trichotomous variable (top vs bottom and middel brackets) and the linear model.
+At the user-level, we do not find statistically significant effects of the Z-score of the first prompt on average token sort ratio of successive prompts.
+    Successive Prompt ‘Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in the likelihood that the subsequent prompt contains the previous one.
+When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with the top bracket more likely to exactly include the previous prompt.
 However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket.
-Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average successive similarity.
-    Successive Prompt Token Sort Ratio: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in the Token Sort Ratio between successive prompts.
-When we compare the lowest z-score bracket to the highest z-score bracket, we observe a statistically significant difference.
-However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket.
-Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average successive prompt token sort ratio.
-    Successive Prompt ’Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in the likelihood that the subsequent prompt contains the previous prompt.
-When we compare the lowest z-score bracket to the highest z-score bracket, we observe a statistically significant difference.
-However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket.
-Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on the average probability of a prompt containing the preceding prompt.
+When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant.
+An increase in the realized quality of an image causes the next prompt more likely to include the previous one.
+Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the average probability of a prompt containing the preceding prompt, when using the trichotomous variable or when fitting a linear model.
 H5
@@
   • CLIP embedding cosine similarity
-    No rescaling: Forthcoming
-    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts.
-Those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find comparatively identical resultsin terms of statistical significance.
+    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores).
+When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
 Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
-Logically, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
+Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
 When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks).
+When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
+Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
+Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
+When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
   • DreamSim
-    No rescaling: Forthcoming
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores).
+When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
+Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results.
+When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
+    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks).
+When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
+Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
+Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
+When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
   • Prompting Behaviors
@@
     Successive Prompt Token Sort Ratio: We found no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
-    Successive Prompt ’Contains Previous Prompt’ Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms.
+    Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms.
 ANOVA results confirmed no significant effect of DALL-E version on this measure.
@@
 We also pre-registered our intent to visually inspect whether the QTEs we observe are consistent with dispersion/“inequality” being reduced or increased when participants use different models (e.g., positive effects for low quantiles and negative/null effects for high quantiles would be consistent with inequality reduction).
-Results:
-  • CLIP embedding cosine similarity
-    No rescaling: Forthcoming
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
-  • DreamSim
-    No rescaling: Forthcoming
-    Z-score rescaling: We observed statistically significant reductions in the variance ratio of prompt-length, participant-level prompt embedding variance, and the percentage of prompts containing the previous prompt when comparing DALL-E 2 to DALL-E 3.
-We also found statistically significant reductions in the variance ratio of prompt length and participant-level prompt embedding variance when comparing DALL-E 2 and DALL-E 3 (Revised).
-Our QTE results comparing performance under DALL-E 2 and DALL-E 3 appear in the main text.
-The results we have seen when comparing DALL-E 3 (Revised) to DALL-E 2 are qualitatively similar, although the QTEs are attenuated due to the prompt revision.
-Almost all QTEs when comparing DALL-E 3 and DALL-E 3 (Revised) are positive.
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+F-test Results (with BH Adjusted p-values):
+  • DALL-E 3 (revised) vs. DALL-E 2
+    Mean prompt length (words): DALL-E 3 Revised has significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
+    Prompt embedding variance: DALL-E 3 Revised also has significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
+    Cosine similarity with target image, Z-Score: DALL-E 3 Revised has significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
+    Failed to reject null of no differences in variance: Mean raw DreamSim score vs.
+target image (p = 0.1333), number of topical transitions (by token sort ratio) (p = 0.3294), mean token sort ratio with previous prompt (p = 0.6244), mean percentile rank of DreamSim score (p = 0.9035), mean cosine similarity to previous prompt (p = 0.8748), mean raw CosineSim score vs.
+target image (p = 0.8301), mean proportion of prompts containing previous prompt (p = 0.5461), Z-score of DreamSim score (p = 0.4016), number of topical transitions (cosine similarity) (p = 0.3505), and mean percentile rank of CosineSim (p = 0.0795).
+  • DALL-E 3 (verbatim) vs. DALL-E 3 (revised)
+    Mean percentile rank of CosineSim: DALL-E 3 (verbatim) has significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
+    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7352, $p=0.0009$).
+    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7888, $p=0.0141$).
+    Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7961, $p=0.0159$).
+    DreamSim with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8195, $p=0.0377$).
+    Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) has significantly more variance (ratio 1.2303, $p=0.0301$).
+    Failed to reject null of no differences in variance: Mean raw DreamSim score vs.
+target image (p = 0.1842), mean raw CosineSim score vs.
+target image (p = 0.5916), prompt embedding variance (p = 0.9771), mean cosine similarity to previous prompt (p = 0.9035), mean prompt length (words) (p = 0.6244), number of topical transitions (cosine similarity) (p = 0.5916), and mean token sort ratio with previous prompt (p = 0.5206).
+  • DALL-E 2 vs. DALL-E 3 (verbatim)
+    Mean prompt length (words): DALL-E 3 (verbatim) has significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
+    Prompt embedding variance: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
+    Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7501, $p=0.0015$).
+    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7904, $p=0.0141$).
+    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8005, $p=0.0159$).
+    Failed to reject null of no differences in variance: Mean percentile rank of CosineSim (p = 0.1842), Z-score of DreamSim score (p = 0.3294), Mean Raw CosineSim (p = 0.8151), Z-score of CosineSim score with target image (p = 0.9906), mean token sort ratio with previous prompt (p = 0.8748), mean cosine similarity to previous prompt (p = 0.807), number of topical transitions (token sort ratio) (p = 0.3505), and number of topical transitions (cosine similarity) (p = 0.085).
+QTE Highlighted Results: Through our QTE plots, we see clearest evidence of dispersion being reduced for:
+  • Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALLE-3 Revised vs. DALLE-2 comparison.
+And we see the clearest evidence of dispersion increasing for:
+  • Mean prompt length for all three pairwise model comparisons (DALLE-3 Verbatim vs. DALLE-2, DALLE-3 Revised vs. DALLE-2, and DALLE-3 Verbatim vs. DALLE-3 Revised)
+  • Raw Dreamsim performance for DALLE-3 Verbatim vs. DALLE-2 and DALLE-3 Revised vs. DALLE-2 comparisons.
 H7
@@
 Analysis approach:
-In order to test for differences in prompting performance across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata.
-We run this test on each participant’s performance improvement, defined as the performance gap between the initial iteration and the best-performing iteration.
+In order to test for differences in prompting performance improvement across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata.
+The two samples represent the best and initial scores.
+We make these comparisons overall and within strata.
   • CLIP embedding cosine similarity
-    No rescaling: We do not find evidence for statistically significant relationships between DALL-E 2 versus DALL-E 3 (Verbatim), nor between DALL-E 3 (Verbatim) and DALL-E 3 (Revised).
-However, we do find a statistically significant difference between DALL-E 2 and DALL-E 3 (Revised), whereby DALL-E 2 users showed greater improvement in cosine similarity compared to DALL-E 3 (Revised) users between their first and best attempts.
-The latter relationship holds even after adjusting for Benjamini-Hochberg multiple testing.
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
+    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
+    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
   • DreamSim
-    No rescaling: Forthcoming
-    Z-score rescaling: Forthcoming
-    Percentile rank rescaling: Forthcoming
-    GPT-4V baseline rescaling: Forthcoming
+    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
+    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
+    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
 H8
@@
     The maximum 1-DreamSim score in the above list (the easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, and the lowest score (the hardest image for participants to replicate was $\tilde{D}=0.356$ for Photography Image #4.
-H-Exploratory
+Exploratory Analyses
 In our pre-registration, we also mention a number of exploratory analyses that are not described with the same level of detail.
@@
 Thus, we applied ANOVA and Kruskal-Wallis tests instead.
+  • In our pre-registration, we were not explicit enough on how to compute Z-score measures of performance.
+As described in section C.2, Z-scores are almost always computed within image-attempt to adjust for variations across target images or attempts.
+The only exception is with any analysis that compares performance across attempts where Z-scores need to computed within target images.
+The pre-registration anticipated that Z-scores for any analysis would be computed only within target images.
   • In our pre-registration, we had forgotten to include Education and GenAI outlook in our list of demographics with respect to which we would measure task performance heterogeneity.
@@
   • When conducting our analyses to test H4, we observed that the distribution of the Z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values.
-To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute z-scores greater than 3, which constituted 2.5% of the user-attempt observations.
-For robustness, we also conducted the stratification analysis including these outliers and found similar results, with the exception that the Average Treatment Effect (ATE) for performance between the top and middle brackets was only marginally significant.
+To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute Z-scores greater than 3, which constituted 2.65% of the user-attempt observations.
+For robustness, we also conducted the stratification analysis including these outliers and found almost identical results, with the exception that the difference in performance between the top bracket and middle or bottom brackets was in some cases only marginally significant.
+  • In our pre-registration, we had included a rescaled version of our main performance metrics described in B.5.1 based on GPT-4V for robustness check.
+This outcome variable would have been rescaled by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image.
+Given that the quality of the prompt from GPT-4V model is a constant for each target image, this would have amounted to subtracting a constant from the unscaled CLIP cosine similarity or dream sim scores.
+Furthermore, since all our analysis involving performance as the dependent variable adjust for the target image, either through post-stratification or as a covariate in a linear model, the results from using GPT-4V scaling would have been identical to the the unscaled measures of performance.
+For this reason, we have not included the robustness checks involving GPT-4V for our pre-registered hypothesis.

### v3

@@
 Abstract
-In an online experiment with N = 1,893 participants, we collected and analyzed over 18,000 prompts and over 300,000 images to explore how the importance of prompting will change as the capabilities of generative AI models continue to improve.
-Each participant in our experiment was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor DALL-E 3, or a version of DALL-E 3 with automatic prompt revision.
-Participants were then asked to write prompts to reproduce a target image as closely as possible in 10 consecutive tries.
+The recent surge in generative AI has led to new models being introduced almost every month.
+In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance?
+To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries.
+Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision.
+In total, we collected and analyzed over 18,000 prompts and over 300,000 images.
 We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2.
 This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities.
-More specifically, despite being blind to the model they were assigned, participants assigned to DALL-E 3 wrote longer prompts that were more semantically similar to each other and contained a greater number of descriptive words.
 Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%.
-Taken together, our results suggest that as models continue to progress, people will continue to adapt their prompts to take advantage of new models’ capabilities.
-The release of ChatGPT in November 2022 has fueled a surge of interest in generative AI and large language models across industry, the public sector, and academia, with early research suggesting that generative AI may be a general purpose technology (Eloundou et al., 2024).
-Indeed, generative AI models have already been found to outperform humans on text annotation tasks (Gilardi et al., 2023) and improve the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023).
-However, the extent to which generative AI can provide value is contingent on the quality of the textual instructions, or “prompts,” that users write to a model.
-As generative AI has proliferated, the practice of creating and iterating on such prompts, often using a variety of specialized techniques (e.g., chain-of-thought and few-shot prompting) (Schulhoff et al., 2024), has come to be known as prompt engineering.
-Some firms now even offer six-figure salaries for dedicated prompt engineers (Quilty-Harper, 2023), and an expanding academic literature focuses exclusively on prompting-related topics (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023a; Xie et al., 2023).
-Despite the growing interest in prompting, some argue that as models become more capable at transforming human language into output, the need to craft prompts in specific, and sometimes esoteric, ways will diminish (Acar, 2023; Mollick, 2023).
-If this comes to pass, the benefits of generative AI will be accessible to a much larger population.
-In this paper, we bring experimental evidence to bear on this debate by presenting results from a pre-registered online lab experiment with $N=1,893{}$ participants recruited on Prolific.
+Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models.
+Not doing so could leave more than half of the potential benefits of these AI systems untapped.
+Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al.
+(2024).
+This increasing generative AI adoption appears to be warranted, with recent studies finding that the adoption of generative AI improves the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023).
+A common feature of these studies is that the model participants interacted with did not change; however, generative models are continuously updated, with new versions being released as often as once a month.
+Humans interface with generative models by providing textual instructions or “prompts” to the models (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023).
+It remains an open question whether organizations will need to adapt these prompts as models improve to realize the full economic and societal benefits of generative AI.
+For example, it is now common amongst software companies to have a hundred or more development teams, each with dozens, if not hundreds, of prompts integrated with their code.
+On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort.
+Alternatively, as models improve, organizations may need to continually adapt their prompts.
+In this paper, we investigate whether this prompt adaptation process is necessary, and the extent to which the ability to adapt one’s prompts is a specialized skill.
+We do so by conducting a pre-registered online lab experiment in which $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”).
 [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI.
-In the experiment, participants were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”).
-The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.
-[^2] Automatic prompt revision is the default behavior of the DALL-E 3 API endpoint.
+The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.222prompt revision is the default behavior of the DALL-E 3 API endpoint.
 We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts.
@@
 Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
+Results
+To understand how organizations might benefit from model improvements, we first examined whether access to a more advanced generative model enhanced task performance.
+Comparing the DALL-E 2 and DALL-E 3 treatment arms, we find that providing access to a more advanced generative model improved task performance.
+Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$), and the size of this gap did not shrink as participants gained more experience with the task (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression).
+Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in qualitative similarity to the target image.
 […]
-Results
-The majority of our analyses focus on the DALL-E 2 and DALL-E 3 treatment arms, as they are more directly comparable.
-We first estimate the effect of access to a more advanced generative model on task performance.
-[^3] All main text analyses use ”replay” data from Figure 2 for consistency and to avoid model drift issues.
-All results hold with the “original” data.
-See the SI for details.
-Given the stochastic nature of generative AI models, our measure of task performance is the expected similarity of images generated from each prompt to the relevant target image.
-To construct a quantitative measure of this expected image similarity for each prompt, we first used each participant-generated prompt to produce 10 images.
-We then computed CLIP embedding vectors (Radford et al., 2021) for both target images and prompt-generated images.
-For each participant prompt and its corresponding target image, we then calculated the average cosine similarity ($CoSim$) between the embedding vectors of the randomly sampled generated images and the target image vector (see SI for further details on how we quantify expected image similarity).
-Unsurprisingly, we find that providing access to a more advanced generative model improved task performance.
-Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2.
-Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in similarity to the target image.
-We also find that the treatment effect increased slightly as participants made successive attempts to replicate the target image (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression).
-In other words, the performance gap widened between DALL-E 2 and DALL-E 3 as participants gained more experience with the respective models, potentially due to changes in prompting.
-To investigate how participants’ prompts changed in response to DALL-E 3’s more advanced capabilities, we conducted an exploratory analysis of the contents of participant prompts and how those prompts evolved as participants iterated.
-We find that participants assigned DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and that the difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$).
-Despite differing in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs.
-49% DALL-E 2; $p$ = $0.215$).
-This suggests that prompts submitted to DALL-E 3 were longer because they conveyed additional descriptive information, not because they were padded with superfluous words.
-We also find that participants blindly assigned to DALL-E 3 wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses).
-These findings indicate that exposure to different models did in fact induce differences in prompting, and that these differences grew over successive replication attempts.
-To understand the extent to which the ATE we observe is attributable to these behavioral differences, as opposed to DALL-E 3’s more advanced capabilities, we conducted another exploratory analysis in which we regenerated images with (or “replayed”) all participant prompts on both DALL-E 2 and DALL-E 3, regardless of the model the participant used during the experiment.
-This allows us to decompose the ATE into two components: the “model effect,” i.e., the average increase in similarity to the target when replaying the prompts written by DALL-E 2-assigned participants on DALL-E 3, and the “prompting effect,” i.e., the difference between the average similarity to the target of all prompts generated by DALL-E 3-assigned participants and the average similarity to the target of all prompts generated by DALL-E 2-assigned participants, both replayed on DALL-E 3.
-Figure 2 shows that the model effect (red) accounts for 51% of the ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$), whereas the prompting effect (blue) accounts for 48% of the ATE ($\Delta CoSim=0.00788$, $p=0.024$).
-Thus, nearly half of the ATE is driven by endogenous changes in prompting strategy that emerged after exposure to DALL-E 3.
-The rightmost point in Figure 2 shows that these endogenous changes in prompting took advantage of the unique capabilities of DALL-E 3; when replaying both sets of prompts on DALL-E 2, the difference in performance is not statistically significant and is close to zero ($\Delta CoSim=0.0020$, $p=0.56$).
-In other words, DALL-E 3 was more capable than DALL-E 2 at effectively rendering the additional information contained in the longer and more descriptive prompts (Betker et al., 2023), and participants adapted their prompting behavior accordingly.
-One might think that the need to adapt one’s prompts to the capabilities of different generative models can be obviated by LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023).
-We use data from our third treatment arm—DALL-E 3 with revision—to test this hypothesis.
-We find that although participants assigned to DALL-E 3 with revisions performed better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$), prompt revision reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
-Importantly, the positive impact of access to DALL-E 3 with revision was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2).
-In other words, prompt revision actually caused participants to perform worse in our context.
-This finding suggests that as currently implemented, AI-assisted prompt revisions are not a panacea, and can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
-Discussion
+Access to a more advanced generative model not only increased task performance but also induced changes in the way that participants wrote prompts, which we highlight through two exploratory analyses.
+First, prompts to the more advanced model were longer and contained more information.
+Those assigned to DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and this difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$).
+Despite this difference in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs.
+49% DALL-E 2; $p$ = $0.215$), suggesting that prompts submitted to DALL-E 3 conveyed additional descriptive information.
+Second, those assigned to DALL-E 3 explored the space of possible prompts differently; they wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses).
+In other words, those assigned to DALL-E 3 were more likely to thoroughly explore a smaller set of prompting approaches.
+The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations will need to adapt their prompts as models evolve.
+More specifically, using the decomposition procedure described in the Materials and Methods, we find that changes in prompting accounted for 48% of the performance ATE ($\Delta CoSim=0.00788$, $p=0.024$; blue in Figure 2), whereas the shift to a more capable model accounted for 51% of the performance ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$; red in Figure 2).
+We also supplied the prompts written for DALL-E 3 to DALL-E 2 (see Materials and Methods for details).
+Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$), indicating that DALL-E 3 prompts specifically took advantage of its superior ability to render information (Betker et al., 2023).
 Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects.
@@
 Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
-In summary, we find that the progression of generative AI models from DALL-E 2 to DALL-E 3 caused performance on our image replication task to increase.
-This was not only because of DALL-E 3’s superior capabilities (the model effect), but also because of changes in prompting behavior that took advantage of those capabilities (the prompting effect).
-More specifically, participants using DALL-E 3 wrote longer prompts that contained more descriptive information, and DALL-E 3 was able to render that additional information more effectively than DALL-E 2.
-These changes in prompting emerged despite the fact that participants were unaware of which model they were assigned.
-Although automated LLM-revision features are designed to and have been shown to help people generate higher-quality content (Betker et al., 2023; Li et al., 2024), we also find that LLM-based prompt revision failed to further improve performance; instead, automatic prompt revision reduced the benefits of using DALL-E 3.
-Overall, we find that shifting to a more capable model, whether or not accompanied by automatic LLM revisions, did not eliminate the need for prompting.
+Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process.
+One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models.
+Comparison of our third treatment arm–DALL-E 3 with revision–to the DALL-E 3 arm highlights the potential pitfalls of this approach.
+We find that prompt revision actually caused participants to perform worse in our context as it reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
+Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2).
+These findings suggest that as currently implemented, AI-assisted prompt revisions are not a panacea.
+They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
+[…]
+Discussion
 The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3.
 We leave the study of future transitions and different types of generative AI models to future work.
-However, if the trends we report generalize and continue into the future, one could imagine a lock-step dynamic where as models continually improve, people continually respond by adapting their prompts to take advantage of the newest model’s capabilities.
-Such a pattern suggests that as generative AI models advance, newer models are unlikely to render prompting obsolete.
-Rather, prompting will be the mechanism by which people unlock new models’ capabilities.
-Thus, we anticipate that prompting will remain relevant in the future, and that it will be an important determinant of who is best positioned to use and directly benefit from generative AI moving forward.
+Should our results generalize to other models, then we would predict that firms that do not update their prompts as models progress are leaving roughly half the benefit of AI unrealized.
+Software systems, models, and prompts are becoming increasingly interwoven which is a trend we expect to accelerate.
+Thus, the lack of prompt maintenance could, in turn, dampen the economic impact of AI on the global economy.
+Fortunately, updating prompts can be learned by humans across the skill distribution even if models are not currently able to update prompts in an automated fashion.
+Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models.
+In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities.
+Such a pattern suggests that as generative AI models advance, technical prompting infrastructure will not be a one-time investment.
+Rather, prompting will be the mechanism by which people and firms unlock new models’ capabilities.
+Table 1: Model and Prompting Effect by Performance
+Effect Estimate p-value
+Model $0.011000$ $p<0.00001$
+$(0.001860)$
+Model $\times$ $-0.000060$ $0.0210$
+Performance Decile $(0.000030)$
+Prompting $0.010900$ $0.0006$
+$(0.003160)$
+Prompting $\times$ $-0.000060$ $0.2444$
+Performance Decile $(0.000050)$
+Material and Methods
+Outcome measure Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image.
+To compute this, for each prompt we generated 10 images and computed CLIP embeddings (Radford et al., 2021) for both these and the target images.
+Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
+Effect Decomposition We conducted an exploratory analysis where we regenerated images with all participant prompts on both models.
+This allows us to decompose the ATE into two components: the “model effect” (the average improvement when running DALL-E 2 prompts through DALL-E 3) and the “prompting effect” (the difference in target similarity between DALL-E 3 and DALL-E 2 participants’ prompts, both evaluated on DALL-E 3).
+Note that all main text analyses use “replay” data from Figure 2 for consistency and to avoid model drift issues.
+All results hold with the “original” data.
+See the SI for details.
+Experimental Materials Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp.
+Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.
 Supplementary Information (SI)
@@
 Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$.
-TSR is a fuzzy string-matching technique (Singla and Garg, 2012b) that provides a continuous measure of how similar two strings are.
+TSR is a fuzzy string-matching technique (Singla and Garg, 2012) that provides a continuous measure of how similar two strings are.
 We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
@@
 We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt.
-SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation posuniversaldep.
+SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal POS tags (2024).
 These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs.
@@
 Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses.
-In our pre-registration, we declared the intent to conduct each of our analyses using six777see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
+In our pre-registration, we declared the intent to conduct each of our analyses using six666see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
 possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
@@
     Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores).
-When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
+When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
+Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
 Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
@@
     Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks).
-When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
+When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
+Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
 Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
@@
     Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores).
-When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
+When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
 Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results.
@@
     Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks).
-When we compare first iteration performance, those using DALLE-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALLE-E 3 verbatim outperforming DALL-E 2 and DALLE-E 3 revised, with no statistically significant difference between DALL-E 2 and DALLE-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALLE-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
+When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
+When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
+Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
 Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.

### v4

@@
 Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models.
-Not doing so could leave more than half of the potential benefits of these AI systems untapped.
 Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al.
 (2024).
-This increasing generative AI adoption appears to be warranted, with recent studies finding that the adoption of generative AI improves the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023).
-A common feature of these studies is that the model participants interacted with did not change; however, generative models are continuously updated, with new versions being released as often as once a month.
-Humans interface with generative models by providing textual instructions or “prompts” to the models (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023).
-It remains an open question whether organizations will need to adapt these prompts as models improve to realize the full economic and societal benefits of generative AI.
-For example, it is now common amongst software companies to have a hundred or more development teams, each with dozens, if not hundreds, of prompts integrated with their code.
+This widespread adoption appears to be warranted, with recent studies finding that giving workers generative AI improves productivity across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023).
+A common feature of these studies is that they used static model versions for testing.
+However, in practice, generative models are continuously updated, with new versions being released as frequently as once per month.
+People interact with generative AI by providing textual instructions or “prompts” to guide the models’ outputs (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023).
+It remains unclear whether organizations must continually adjust these prompts as the models advance to fully capture their economic and societal advantages.
+For example, it is now common amongst software companies to have a hundred or more development teams, each with hundreds, if not thousands, of prompts integrated with their code.
 On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort.
-Alternatively, as models improve, organizations may need to continually adapt their prompts.
-In this paper, we investigate whether this prompt adaptation process is necessary, and the extent to which the ability to adapt one’s prompts is a specialized skill.
-We do so by conducting a pre-registered online lab experiment in which $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”).
-[^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI.
-The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.222prompt revision is the default behavior of the DALL-E 3 API endpoint.
+Alternatively, organizations may need to continually adapt their prompts as models improve.
+In this paper, we investigate whether refining prompts is necessary as models evolve, and how much specialized expertise this refining process requires.
+To explore these questions, we conducted a pre-registered online lab experiment.
+$N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”).
+[^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration in the SI.
+The LLM-based prompt revisions in the third condition were not visible to participants during the experiment.
+[^2] Automatic prompt revision is the default behavior of the DALL-E 3 API.
+These prompt revisions involved sending each participant’s prompt to GPT-4, which enhanced it by adding descriptive details before being sent to DALL-E 3.
 We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts.
 While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them.
-Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions.
-For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper.
+Therefore, the “DALL-E 3” treatment arm still includes some automatic revisions.
+We refer to this treatment arm as “DALL-E 3” for brevity and simplicity throughout the paper.
 Full details on the system prompt’s efficacy are provided in the SI.
@@
 In other words, those assigned to DALL-E 3 were more likely to thoroughly explore a smaller set of prompting approaches.
-The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations will need to adapt their prompts as models evolve.
-More specifically, using the decomposition procedure described in the Materials and Methods, we find that changes in prompting accounted for 48% of the performance ATE ($\Delta CoSim=0.00788$, $p=0.024$; blue in Figure 2), whereas the shift to a more capable model accounted for 51% of the performance ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$; red in Figure 2).
-We also supplied the prompts written for DALL-E 3 to DALL-E 2 (see Materials and Methods for details).
-Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$), indicating that DALL-E 3 prompts specifically took advantage of its superior ability to render information (Betker et al., 2023).
+The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations must adapt their prompts as models evolve.
+To quantify how much of the performance improvement came from participants adapting their prompts, we took prompts written by participants in the DALL-E 2 treatment arm and generated images with them using DALL-E 3.
+This allowed us to isolate the “model effect”—how much better DALL-E 3 performed with prompts written for DALL-E 2 ($\Delta CoSim=0.0084$, $p<10^{-8}$; red in Figure 2), and the “prompting effect”—the portion of the performance ATE due solely to participants adapting their prompting for DALL-E 3 ($\Delta CoSim=0.0079$, $p=0.024$; blue in Figure 2).
+These effects were nearly equal: 51% and 48% of the performance ATE, respectively.
+In contrast, we also supplied the prompts written for DALL-E 3 to DALL-E 2.
+Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$).
+This implies that DALL-E 3 prompts specifically took advantage of its superior ability to render information, and users naturally discovered these capabilities (Betker et al., 2023).
 Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects.
@@
 One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models.
-Comparison of our third treatment arm–DALL-E 3 with revision–to the DALL-E 3 arm highlights the potential pitfalls of this approach.
-We find that prompt revision actually caused participants to perform worse in our context as it reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
+Comparison of our third treatment arm—DALL-E 3 with revision—to the DALL-E 3 arm highlights the potential pitfalls of this approach.
+We find that automated prompt revisions actually caused participants to perform worse in our context.
+It reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
 Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2).
-These findings suggest that as currently implemented, AI-assisted prompt revisions are not a panacea.
+These findings suggest that, as currently implemented, AI-assisted prompt revisions are not a panacea.
 They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
+Humans, on the other hand, are better at updating their prompts.
 […]
+Table 1: Model and Prompting Effect by Performance Decile
+Effect Estimate P-value
+Model $0.01100$ $<0.00001$
+$(0.00186)$
+Model $\times$ $-0.00006$ $0.0210$
+Performance Decile $(0.00003)$
+Prompting $0.01090$ $0.0006$
+$(0.00316)$
+Prompting $\times$ $-0.00006$ $0.2444$
+Performance Decile $(0.00005)$
 Discussion
-The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3.
-We leave the study of future transitions and different types of generative AI models to future work.
-Should our results generalize to other models, then we would predict that firms that do not update their prompts as models progress are leaving roughly half the benefit of AI unrealized.
-Software systems, models, and prompts are becoming increasingly interwoven which is a trend we expect to accelerate.
-Thus, the lack of prompt maintenance could, in turn, dampen the economic impact of AI on the global economy.
-Fortunately, updating prompts can be learned by humans across the skill distribution even if models are not currently able to update prompts in an automated fashion.
+The main limitation of our work is that we only study the transition from DALL-E 2 to DALL-E 3.
+We leave the study of future transitions and different types of generative AI models to future research.
+Software systems, models, and prompts are becoming increasingly interwoven, which is a trend we expect to accelerate.
+If our findings generalize to other models, firms that fail to update their prompts as models improve may forego roughly half of the potential benefits of AI advancement, which will dampen the overall economic impact of AI.
 Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models.
 In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities.
-Such a pattern suggests that as generative AI models advance, technical prompting infrastructure will not be a one-time investment.
-Rather, prompting will be the mechanism by which people and firms unlock new models’ capabilities.
-Table 1: Model and Prompting Effect by Performance
-Effect Estimate p-value
-Model $0.011000$ $p<0.00001$
-$(0.001860)$
-Model $\times$ $-0.000060$ $0.0210$
-Performance Decile $(0.000030)$
-Prompting $0.010900$ $0.0006$
-$(0.003160)$
-Prompting $\times$ $-0.000060$ $0.2444$
-Performance Decile $(0.000050)$
+Such a pattern suggests that technical prompting infrastructure will not be a one-time investment as generative AI models advance.
+Rather, prompting will be an essential and increasingly intuitive mechanism by which people and firms unlock new models’ capabilities.
+Fortunately, our findings suggest that humans across the skill distribution can learn to update prompts, even when automated revisions do not suffice.
 Material and Methods
-Outcome measure Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image.
-To compute this, for each prompt we generated 10 images and computed CLIP embeddings (Radford et al., 2021) for both these and the target images.
+Institutional Review. This study was reviewed by the UC Berkeley Committee for the Protection of Human Subjects (CPHS) under Protocol 2023-06-16480.
+Outcome Measure.
+Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image.
+To compute this, we generated 10 images for each prompt and computed CLIP embeddings (Radford et al., 2021) for both these and the target images.
 Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
-Effect Decomposition We conducted an exploratory analysis where we regenerated images with all participant prompts on both models.
-This allows us to decompose the ATE into two components: the “model effect” (the average improvement when running DALL-E 2 prompts through DALL-E 3) and the “prompting effect” (the difference in target similarity between DALL-E 3 and DALL-E 2 participants’ prompts, both evaluated on DALL-E 3).
-Note that all main text analyses use “replay” data from Figure 2 for consistency and to avoid model drift issues.
+Replay Data.
+To generate the images used to decompose the prompting and model effects, we regenerated or “replayed” all images with all participant prompts on both models.
+Note that all main text analyses use replay data from Figure 2—even replays of DALL-E 2 written prompts on DALL-E 2, and DALL-E 3 written prompts on DALL-E 3—for consistency and to avoid model drift issues.
 All results hold with the “original” data.
-See the SI for details.
-Experimental Materials Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp.
+See SI for details.
+Experimental Materials.
+Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp.
 Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.

### v5

@@
 Abstract
-The recent surge in generative AI has led to new models being introduced almost every month.
-In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance?
-To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries.
-Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision.
-In total, we collected and analyzed over 18,000 prompts and over 300,000 images.
-We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2.
-This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities.
-Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%.
-Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models.
-Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al.
-(2024).
-This widespread adoption appears to be warranted, with recent studies finding that giving workers generative AI improves productivity across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023).
-A common feature of these studies is that they used static model versions for testing.
-However, in practice, generative models are continuously updated, with new versions being released as frequently as once per month.
-People interact with generative AI by providing textual instructions or “prompts” to guide the models’ outputs (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023).
-It remains unclear whether organizations must continually adjust these prompts as the models advance to fully capture their economic and societal advantages.
-For example, it is now common amongst software companies to have a hundred or more development teams, each with hundreds, if not thousands, of prompts integrated with their code.
-On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort.
-Alternatively, organizations may need to continually adapt their prompts as models improve.
-In this paper, we investigate whether refining prompts is necessary as models evolve, and how much specialized expertise this refining process requires.
-To explore these questions, we conducted a pre-registered online lab experiment.
-$N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”).
-[^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration in the SI.
-The LLM-based prompt revisions in the third condition were not visible to participants during the experiment.
-[^2] Automatic prompt revision is the default behavior of the DALL-E 3 API.
-These prompt revisions involved sending each participant’s prompt to GPT-4, which enhanced it by adding descriptive details before being sent to DALL-E 3.
-We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts.
-While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them.
-Therefore, the “DALL-E 3” treatment arm still includes some automatic revisions.
-We refer to this treatment arm as “DALL-E 3” for brevity and simplicity throughout the paper.
-Full details on the system prompt’s efficacy are provided in the SI.
-The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model.
-Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images.
-Throughout the task, participants could view all of their past prompts and generated images alongside the target image.
-Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
-Results
-To understand how organizations might benefit from model improvements, we first examined whether access to a more advanced generative model enhanced task performance.
-Comparing the DALL-E 2 and DALL-E 3 treatment arms, we find that providing access to a more advanced generative model improved task performance.
-Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$), and the size of this gap did not shrink as participants gained more experience with the task (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression).
-Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in qualitative similarity to the target image.
+As generative AI systems rapidly improve, a key question emerges: How do users keep up—and what happens if they fail to do so.
+Drawing on theories of dynamic capabilities and IT complements, we examine prompt adaptation—the adjustments users make to their inputs in response to evolving model behavior—as a mechanism that helps determine whether technical advances translate into realized economic value.
+In a preregistered online experiment with 1,893 participants, who submitted over 18,000 prompts and generated more than 300,000 images, users attempted to replicate a target image in 10 tries using one of three randomly assigned models: DALL-E 2, DALL-E 3, or DALL-E 3 with automated prompt rewriting.
+We find that users with access to DALL-E 3 achieved higher image similarity than those with DALL-E 2—but only about half of this gain (51%) came from the model itself.
+The other half (49%) resulted from users adapting their prompts in response to the model’s capabilities.
+This adaptation emerged across the skill distribution, was driven by trial-and-error, and could not be replicated by automated prompt rewriting, which erased 58% of the performance improvement associated with DALL-E 3.
+Our findings position prompt adaptation as a dynamic complement to generative AI—and suggest that without it, a substantial share of the economic value created when models advance may go unrealized.
+1 Introduction
+Generative AI is being integrated into work practices across the economy (Zhang and Kamel Boulos 2023, Bright et al.
+2024), yielding notable productivity gains in tasks as diverse as software development, writing, and material discovery (Brynjolfsson et al.
+2023, Dell’Acqua et al.
+2023, Noy and Zhang 2023, Peng et al.
+2023, Yu 2024, Toner-Rodgers 2024).
+Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning et al.
+2024), including tasks as complex as chemical research and proving mathematical theorems (Romera-Paredes et al.
+2024, Boiko et al.
+2023).
+The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S.
+workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (Bright et al.
+2024, Bick et al.
+2024).
+As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results.
+To interact with generative AI systems, users provide written instructions—often referred to as prompts—that guide the model’s behavior.
+These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software).
+In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (Brynjolfsson and Hitt 2000).
+Prompting has quickly become an area of active research and practice.
+Scholars have developed taxonomies of prompt engineering techniques (Oppenlaender 2023), documented recurring patterns in prompt construction (Schulhoff et al.
+2024), and examined how developers embed prompts into software systems (Liang et al.
+2024).
+Other studies have explored prompting strategies for specific applications, including image generation (Don-Yehiya et al.
+2023, Xie et al.
+2023) and clinical documentation (Yao et al.
+2024).
+In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design.
+These developments signal a growing consensus that prompting plays a meaningful role in extracting value from generative AI systems.
+Yet despite this consensus, prompting remains understudied as a dynamic practice.
+Many prompt libraries and tutorials present effective prompts as reusable artifacts.
+But prompts that work well with one model version may underperform or break entirely with the next (Liang et al.
+2024, Meincke et al.
+2025).
+While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance.
+This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
+To begin exploring this question, we identify prompt adaptation [^1] We use the phrase “prompt adaptation” to refer specifically to changes in user prompting behavior that arise in response to evolving model capabilities—distinct from the broader practice of prompt engineering, which includes static best practices, libraries, and templates.
+as a measurable behavioral mechanism through which user-side inputs evolve alongside technical advances.
+We conceptualize prompt adaptation as a dynamic complement—that is, a user capability that adapts in response to changes in a technological system and is critical to realizing the full economic value of system improvements.
+In contrast to static complements (e.g., fixed training, prompt templates), dynamic complements emerge through situated use with rapid feedback, respond to model-level change, and may be enabled or suppressed by system design.
+To assess the role of prompt adaptation in shaping realized performance—and to separate its contribution from the direct effects of model improvement—we draw on data from a pre-registered online experiment with $1,893{}$ participants.
+In the study, participants were asked to replicate a target image using prompts submitted to one of three randomly assigned text-to-image models of varying capability: DALL-E 2, DALL-E 3, or DALL-E 3 with an automated large language model (LLM)-based prompt revision.
+Each participant submitted at least 10 prompts in an effort to reproduce the image as closely as possible, with a large monetary bonus for top performers.
+By comparing outcomes across arms—and by conducting a posthoc analysis that re-evaluates prompts on alternative models—we estimate the extent to which users adapted their prompts in response to model improvements and how much this adaptation contributed to overall performance.
+We find that participants assigned to DALL-E 3 produce significantly more faithful replications than those assigned to DALL-E 2.
+Importantly, about half of this improvement comes from participants adapting their prompts to exploit the new model’s capabilities—replaying the old DALL-E 2 prompts on DALL-E 3 yields only about half the total improvement.
+Furthermore, we find that this prompt adaptation is not limited to advanced “prompt engineers”—participants across the full skill distribution benefit from refining their prompts after seeing how the advanced model responds.
+Finally, we show that an attempt to automate prompt revision via GPT-4 rewriting does not match the performance achieved by manual human adaptation; in fact, it substantially erodes the gains from DALL-E 3.
+Together, these findings position prompt adaptation as a key mechanism through which users realize the value of rapidly advancing generative AI systems—and as a concrete example of how digital complements must evolve to keep pace with technological change.
+In terms of related literature and additional theory, this experiment builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies.
+Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (Bharadwaj 2000, Joshi et al.
+2010, Teece et al.
+1997).
+Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (Jasperson et al.
+2005).
+Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (Fügener et al.
+2022).
+And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (Böhm and Schedlberger 2023).
+We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (Brynjolfsson 1993, Brynjolfsson and Hitt 2000, Brynjolfsson et al.
+2021, David 1990).
+We conceptualize prompt adaptation as one such complement—emerging through trial-and-error, accessible across the skill distribution, and potentially hindered when automation misaligns with user intent.
+In this way, prompt adaptation shapes how and when technical improvements translate into downstream economic value.
+Finally, the contribution of the current paper is four-fold.
+First, we offer direct causal evidence that prompting is not a fixed input but a dynamic capability that co-evolves with model behavior (Böhm and Schedlberger 2023).
+Second, we show how even non-expert users adapt their prompting in response to model improvements, extending research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al.
+2010, Teece et al.
+1997).
+Third, we document how performance gains emerge through user-initiated trial-and-error rather than static best practices, advancing work on post-adoptive IT use (Jasperson et al.
+2005).
+Finally, we demonstrate that automation intended to reduce user effort can—if misaligned with user intent—undermine the very adaptations that generate value, contributing to ongoing research on human–AI interaction (Fügener et al.
+2022, Yao et al.
+2024).
+The remainder of the paper is organized as follows.
+We begin by presenting a simple conceptual framework that characterizes how output quality evolves with improvements in model capacity and with users’ corresponding adjustments in effort.
+We then introduce an experimental design that closely mirrors this framework and describes the data used in our analysis.
+Next, we present our empirical findings, including a decomposition of the overall effect into components attributable to model improvements versus prompt adaptation, the impact of automated prompt revision, and heterogeneity across user skill levels.
+Finally, we compare these empirical results to the model’s predictions and conclude by discussing implications for organizations adopting generative AI.
+2 Conceptual Framework
+As generative AI systems evolve, a user’s ability to adapt prompts for improved models can become an important source of realized performance gains.
+We develop a stylized conceptual framework to formalize how overall output quality depends on both the model’s capacity and the user’s skill and effort in prompt writing.
+Our goal is not to create a fully normative model.
+Rather, we wish to clarify the distinction between improvements directly attributable to the model itself versus those arising from user adaptation as motivation for our experimental design.
+Additionally, the framework naturally implies predictions about the returns to prompting skill and the heterogeneity of such skill.
+Although our empirical setting focuses on image replication, this framework generalizes to other tasks where users interact with generative models—such as text generation, code assistance, or molecule design.
+In each case, we expect performance to reflect both model capacity and users’ efforts to refine their inputs.
+We present the core findings here; a complete exposition of the framework with formal proofs is provided in Appendix E.1.
+2.1 Notation and Problem Setting
+Let $\theta\in(0,1]$ represent the model’s capacity to translate prompts into high-fidelity outputs (for instance, how accurately it captures requested details).
+Let $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and let $x\geq 0$ be the effort the user expends on writing and refining prompts.
+Each unit of $x$ incurs a cost $k\,x$ for some $k>0$, which may reflect the time or cognitive load.
 […]
-Access to a more advanced generative model not only increased task performance but also induced changes in the way that participants wrote prompts, which we highlight through two exploratory analyses.
-First, prompts to the more advanced model were longer and contained more information.
-Those assigned to DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and this difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$).
-Despite this difference in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs.
-49% DALL-E 2; $p$ = $0.215$), suggesting that prompts submitted to DALL-E 3 conveyed additional descriptive information.
-Second, those assigned to DALL-E 3 explored the space of possible prompts differently; they wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses).
-In other words, those assigned to DALL-E 3 were more likely to thoroughly explore a smaller set of prompting approaches.
-The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations must adapt their prompts as models evolve.
-To quantify how much of the performance improvement came from participants adapting their prompts, we took prompts written by participants in the DALL-E 2 treatment arm and generated images with them using DALL-E 3.
-This allowed us to isolate the “model effect”—how much better DALL-E 3 performed with prompts written for DALL-E 2 ($\Delta CoSim=0.0084$, $p<10^{-8}$; red in Figure 2), and the “prompting effect”—the portion of the performance ATE due solely to participants adapting their prompting for DALL-E 3 ($\Delta CoSim=0.0079$, $p=0.024$; blue in Figure 2).
-These effects were nearly equal: 51% and 48% of the performance ATE, respectively.
-In contrast, we also supplied the prompts written for DALL-E 3 to DALL-E 2.
-Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$).
-This implies that DALL-E 3 prompts specifically took advantage of its superior ability to render information, and users naturally discovered these capabilities (Betker et al., 2023).
-Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects.
-The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line).
-Effects are shown relative to a baseline of prompts written for and replayed on DALL-E 2.
-Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
-Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process.
-One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models.
-Comparison of our third treatment arm—DALL-E 3 with revision—to the DALL-E 3 arm highlights the potential pitfalls of this approach.
-We find that automated prompt revisions actually caused participants to perform worse in our context.
-It reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
-Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2).
-These findings suggest that, as currently implemented, AI-assisted prompt revisions are not a panacea.
-They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
-Humans, on the other hand, are better at updating their prompts.
-[…]
-Table 1: Model and Prompting Effect by Performance Decile
-Effect Estimate P-value
-Model $0.01100$ $<0.00001$
-$(0.00186)$
-Model $\times$ $-0.00006$ $0.0210$
-Performance Decile $(0.00003)$
-Prompting $0.01090$ $0.0006$
-$(0.00316)$
-Prompting $\times$ $-0.00006$ $0.2444$
-Performance Decile $(0.00005)$
-Discussion
-The main limitation of our work is that we only study the transition from DALL-E 2 to DALL-E 3.
-We leave the study of future transitions and different types of generative AI models to future research.
-Software systems, models, and prompts are becoming increasingly interwoven, which is a trend we expect to accelerate.
-If our findings generalize to other models, firms that fail to update their prompts as models improve may forego roughly half of the potential benefits of AI advancement, which will dampen the overall economic impact of AI.
-Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models.
-In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities.
-Such a pattern suggests that technical prompting infrastructure will not be a one-time investment as generative AI models advance.
-Rather, prompting will be an essential and increasingly intuitive mechanism by which people and firms unlock new models’ capabilities.
-Fortunately, our findings suggest that humans across the skill distribution can learn to update prompts, even when automated revisions do not suffice.
-Material and Methods
-Institutional Review. This study was reviewed by the UC Berkeley Committee for the Protection of Human Subjects (CPHS) under Protocol 2023-06-16480.
-Outcome Measure.
-Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image.
-To compute this, we generated 10 images for each prompt and computed CLIP embeddings (Radford et al., 2021) for both these and the target images.
-Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
-Replay Data.
-To generate the images used to decompose the prompting and model effects, we regenerated or “replayed” all images with all participant prompts on both models.
-Note that all main text analyses use replay data from Figure 2—even replays of DALL-E 2 written prompts on DALL-E 2, and DALL-E 3 written prompts on DALL-E 3—for consistency and to avoid model drift issues.
-All results hold with the “original” data.
-See SI for details.
-Experimental Materials.
-Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp.
-Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.
-Supplementary Information (SI)
+so that users cannot exceed perfect fidelity (quality of 1), and each additional unit of effort yields diminishing returns. The user’s utility is then
+$U(\theta,s,x)\;=\;Q(\theta,s,x)\;-\;k\,x.$
+The user chooses their effort such that it maximizes the utility. We assume $\theta s>k$, which implies there is a unique interior optimum $x^{*}(\theta,s)$
+$x^{*}(\theta,s)\;=\;\frac{1}{\theta s}\,\ln\Bigl{(}\frac{\theta s}{k}\Bigr{)}>0,$ (1)
+with an optimal quality
+$Q^{*}\;=\;Q\big{(}\theta,s,x^{*}(\theta,s)\big{)}=1-\frac{k}{\theta s}>0.$ (2)
+Two implications immediately follow from equation 2. Optimal quality is increasing with both model capacity
+$\frac{\partial Q^{*}}{\partial\theta}=\frac{k}{\theta^{2}s}>0$ (3)
+and user skill
+$\frac{\partial Q^{*}}{\partial s}=\frac{k}{\theta s^{2}}>0.$ (4)
+Furthermore, the relationship between the changes in these parameters supplies us with a basic prediction, which we can evaluate empirically in our experiment.
+Proposition 1.
+As model capacity $\theta$ improves, the effect of user skill on optimal quality diminishes.
+$\frac{\partial^{2}Q^{*}}{\partial s\partial\theta}=-\frac{k}{\theta^{2}s^{2}}<0$
+In other words, improvements in model capacity reduce the performance gap between high and low-skilled users.
+2.2 Decomposition into Model and Prompt Effects
+As model capacity $\theta$ increases, so does $x^{*}$ in equation 1, implying that a more capable model not only delivers better output for a fixed prompt but also encourages the user to devote more effort to composing prompts.
+This distinction highlights both a direct capacity-based effect of an improved model, a model effect, and an indirect effect driven by users adapting their prompts to the new model’s potential, a prompting effect.
+To see this more formally, suppose the model is upgraded from capacity $\theta_{1}$ to $\theta_{2}$.
+Let the old prompting effort be $x^{*}(\theta_{1},s)$ and the new (adapted) prompting effort be $x^{*}(\theta_{2},s)$.
+Define the total improvement in quality $\Delta Q$ is
+$\Delta Q\;=\;Q\Bigl{(}(\theta_{2},s,x^{*}(\theta_{2},s)\Bigr{)}\;-\;Q\Bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\Bigr{)}.$
+We can decompose $\Delta Q$ into
+$\Delta Q$ $={}\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{1},s))\;-\;Q(\theta_{1},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Model Effect }(M)\vphantom{\Big{|}}}$ (5)
+$+\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{2},s))\;-\;Q(\theta_{2},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Prompting Effect }(P)\vphantom{\Big{|}}}.$
+The first term isolates the gain from simply upgrading the model (holding the user’s prompt strategy fixed at the old optimum), while the second term represents any additional gain from prompt adaptation.
+In other words, even if $\theta$ improves, failing to adjust how prompts are written could leave substantial performance gains on the table.
+This decomposition naturally implies a prediction similar to Proposition 1 for each component of $\Delta Q$ that co-evolve with the changes in model capacity and user skill.
+In particular, when $k$ is small, the following are proved in Appendix E.1.
+Proposition 2.
+As user skill $s$ improves, the model effect $M$ decreases and the prompting effect $P$ increases.
+Intuitively, when costs are low and model capacity rises, higher-skilled users are already near the upper bound of performance, so the direct model effect they receive may be smaller.
+In contrast, users who can effectively refine prompts might benefit more from the new capacity, so the prompting effect they receive may be larger.
+Taken together, these theoretical predictions guide our experimental design, where we randomly vary $\theta$ and measure how users adapt.
+3 Experiment Design and Methods
+To empirically examine whether users do, in fact, adapt their prompts in response to model improvements—and how much this adaptation contributes to overall performance—we conducted a pre-registered online experiment with 2,059 participants on Prolific between December 12 and December 19, 2023.222Full pre-registration details, including hypotheses and planned analyses, are available in an online repository.
+Participants were asked to replicate a target image as closely as possible using a generative AI model, with the goal of assessing how both model capability and prompting behavior influence final outcomes.
+[^3] All procedures were approved by an institutional review board (link removed to preserve author anonymity) and participants provided informed consent.
+We will release anonymized data and replication code upon publication.
+3.1 Experimental Setting
+Each participant was randomly and blindly assigned to one of three model conditions: DALL-E 2, DALL-E 3 (hereafter “DALL-E 3 (Verbatim)”), or DALL-E 3 with automatic prompt revision (“DALL-E 3 (Revised)”).
+These models differ not only in technical capability but also in whether they apply hidden large language model (LLM)-based modifications to user prompts.
+In the DALL-E 2 condition, participants interacted with an earlier-generation model that interprets prompts directly without intermediate rewriting.
+In contrast, the DALL-E 3 API, by default, forwards user prompts to GPT-4 before image generation.
+This intermediate GPT-4 step rewrites the prompt—typically by adding detail or restructuring language—before passing it to the image model.
+This behavior is intended to improve image quality but occurs silently and without user visibility.
+In the DALL-E 3 (Revised) condition, we allowed this default behavior to proceed unaltered.
+In the DALL-E 3 (Verbatim) condition, we attempted to suppress GPT-4-based prompt rewriting by prepending a hidden system message instructing the model to leave the user’s prompt unchanged.
+While some rewriting still occurred, the rate and extent of modifications were substantially reduced.
+In all conditions, participants were unaware that their prompt might be rewritten or modified before image generation.
+In addition to model assignment, each participant was independently assigned one of 15 target images.
+These images were drawn from three broad categories—business and marketing, graphic design, and architectural photography—to represent common use cases for text-to-image generation.
+All images were sourced from platforms that permit free research use (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography) [^4] All of these images have licenses for free use for commercial and noncommercial purposes..
+We built a custom interface resembling ChatGPT, with the assigned target image on the right and a scrollable history of prompts and generated images on the left.
+Participants were explicitly informed that the model was memoryless: every new prompt was processed independently, carrying no information from previous attempts.
+Participants had up to 25 minutes to submit prompts, with a requirement to submit at least 10.
+They were paid \$4 for completing the task, plus a \$8 bonus (a 200% increase) if their highest-scoring image was in the top 20% of participants.
+The median completion time was 22 minutes, implying an average hourly wage of about \$15.
+After replicating the target image, participants completed a demographic survey covering age, gender, education, occupation, and self-assessed proficiency in creative writing, programming, and generative AI.
+We removed from our analyses any participants who did not submit at least 10 prompts, repeated the same prompt five or more times in a row, or failed to complete the post-task survey, resulting in a final sample of 1,893 participants and 18,560 prompts.
+[^5] Although participants were required to submit at least 10 prompts to be included in our analysis, the final prompt count is slightly below 10 $\times$ 1,893 because we excluded some prompts due to technical issues (e.g., safety filter triggers, duplicate attempt numbers) and limited our main analysis to each participant’s first 10 prompts to mitigate potential selection bias.
+Full details are provided in Appendix C.4.
+3.2 Outcome Definition and Stochastic Generation
+The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al.
+2021).
+CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together.
+By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
+Because the output of each generative model is stochastic, the same prompt can yield different images across attempts.
+To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually.
+We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis.
+As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments.
+[^6] DreamSim (Fu et al.
+2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
+Our results are robust to using DreamSim in place of CLIP cosine similarity.
+Full DreamSim-based analyses can be found in Appendix F.
+The two measures were highly correlated, and our findings were consistent across both.
+3.3 Replay Analysis for Separating Model and Prompting Effects
+A central goal of our experiment was to distinguish how much of the performance improvement in image replication stems from using a more capable model versus how much comes from users adapting their prompts.
+Recall that the language of our conceptual framework shows that the total improvement in output quality from upgrading a generative AI model with capacity $\theta_{1}$ to a model with higher capacity $\theta_{2}$ can be written as:
+$\Delta Q=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}-Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}.$
+We decompose this change into two parts. These are the model effect, the gain from applying the same prompts to a better model,
+$M=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}-Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)},$
+and the prompting effect, the additional improvement from adapting prompts to take advantage of the more capable model,
+$P=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}-Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}.$
+To estimate these components empirically, we conducted an additional analysis using prompts from participants in the DALL-E 2 and DALL-E 3 (Verbatim) conditions.
+To do so, we took the exact prompts that participants submitted during the experiment and re-submitted (or “replayed”) them to both their originally assigned model and the alternative model, generating new images in each case.
+[^7] Because the replay was conducted several weeks after the original experiment, all prompts were re-evaluated at the same time, using the same infrastructure, regardless of whether they were being run on the original or alternative model.
+This ensured that any observed differences in output quality could be cleanly attributed to model capacity and prompt adaptation rather than to changes in model behavior over time (i.e., model drift (Chen et al.
+2023)).
+More specifically, prompts written by DALL-E 2 participants—corresponding to $x^{*}(\theta_{1},s)$—were evaluated both on DALL-E 2 and on DALL-E 3, providing an empirical measure of $Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}$ and $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$, respectively.
+This comparison isolates the model effect: the improvement in output quality when keeping prompts fixed and simply upgrading the model.
+To estimate the prompting effect, we compared the quality of reused DALL-E 2 prompts on DALL-E 3—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$—to the quality of prompts originally written by DALL-E 3 participants and evaluated on the same model—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}$.
+This captures the additional improvement from users adapting their prompts to better leverage the capabilities of the more advanced system.
+4 Results
+In this section, we present our empirical findings on how increased model capacity affects user performance in replicating target images.
+In particular, we focus on three main questions: (i) whether access to a more capable model (DALL-E 3) improves performance; (ii) how users modify their prompts in response to the improved model; (iii) and how much of the overall performance gain can be attributed to model improvements versus prompt adaptation.
+Finally, we compare these empirical results to the predictions of our conceptual framework.
+Our analysis draws on both the main experimental data and the replay data described above.
+4.1 Overall Impact of Model Upgrades
+We begin by examining whether participants using DALL-E 3 achieve higher performance than those using DALL-E 2 as implied by Equation 3.
+Figure 1 summarizes these findings.
+Panel A presents three representative target images and, for each, three generated images drawn from the full sample of participants across both model conditions.
+The middle row for each target shows the image whose cosine similarity to the target is closest to the mean similarity across all participants.
+The rows above (below) show images that are approximately one average treatment effect (ATE) more (less) similar to the target than the mean.
+This visualization provides qualitative intuition for the magnitude of the effect we estimate: the typical difference in fidelity between participants using DALL-E 2 and those using DALL-E 3 (Verbatim).
+Panel B shows that, across the 10 required prompt attempts, participants assigned to DALL-E 3 (Verbatim) produce images that are, on average, 0.0164 higher in cosine similarity to the target (95% CI: [0.0104, 0.0224], $p<10^{-5}$).
+This improvement corresponds to roughly 0.19 standard deviations in performance.
+The gap persists across all attempts; participants using DALL-E 3 start off producing closer matches and maintain that edge through their 10th prompt.
+Figure 1: Overall Performance and Prompting Behavior.
+(A) For three example target images, the middle row shows participant-generated images closest to the mean similarity across all prompts.
+The rows above and below show images of approximately one average treatment effect (ATE) more or less similar to the target, illustrating the typical performance difference between model conditions.
+(B) Top: average CLIP cosine similarity to the target image by attempt, separately for DALL-E 2 and DALL-E 3 participants.
+Bottom: the difference between these averages, with the dark blue line indicating the overall ATE and the shaded region showing the 95% confidence interval.
+(C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals.
+Color shading indicates the average textual similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
+Participants’ dynamic prompting behavior also differs substantially between the two models.
+As shown in panel C of Figure 1, those assigned to DALL-E 3 write prompts that are, on average, 24% longer than those assigned to DALL-E 2, and this gap widens over successive attempts.
+Moreover, we observe that DALL-E 3 participants are more likely to reuse or refine their previous prompts (indicated by the color scale), which suggests a more exploitative approach once they discover the model’s capacity to handle detailed or complex instructions.
+Analyses of parts of speech confirm that these extra words likely provide additional descriptive information rather than mere filler: the proportion of nouns and adjectives—the two most descriptively informative parts of speech—is nearly identical across model conditions (48% for DALL-E 3 vs.
+49% for DALL-E 2; $p=0.215$).
+This suggests that the increase in prompt length reflects the addition of semantically rich content rather than unnecessary verbosity.
+4.2 Replay Analysis and Decomposition of Effects
+The differences we observe in prompting behavior suggest that users are actively adapting to the capabilities of the model they are assigned.
+But how much of the overall performance improvement we observe for DALL-E 3 users is due to the model’s enhanced technical capacity, and how much is due to users rewriting their prompts in response to that capacity?
+To answer this question, we turn to the replay analysis described earlier, which allows us to isolate these two effects identified in Equation 5 empirically.
+Panel A of Figure 2 presents the results.
+To estimate the model effect, we compare the performance of prompts originally written by DALL-E 2 participants when evaluated on DALL-E 2 (the model they were written for) versus when evaluated on DALL-E 3 (Verbatim).
+Because these prompts were written without knowledge of DALL-E 3’s capabilities, any improvement reflects the gain from using a more capable model while holding the prompt fixed.
+We find that performance improves by 0.0084 in cosine similarity when these prompts are evaluated on DALL-E 3 ($p<10^{-8}$; bootstrapped standard errors clustered at the participant level), which accounts for approximately 51% of the total difference in performance between the DALL-E 2 and DALL-E 3 arms.
+To estimate the prompting effect, we then compare the performance of these same DALL-E 2 prompts to the performance of prompts originally written by DALL-E 3 participants, both evaluated on DALL-E 3.
+Because both sets of prompts are evaluated on the same model, any difference reflects the effect of users adapting their prompts to the model’s capabilities.
+We find that this prompting effect accounts for the remaining 48% of the total improvement, corresponding to an increase of 0.0079 in cosine similarity ($p=0.024$).
+The total treatment effect—i.e., the difference between the original DALL-E 2 prompts on DALL-E 2 and the original DALL-E 3 prompts on DALL-E 3—is 0.0164.
+Importantly, when we apply prompts written by DALL-E 3 users to DALL-E 2, we observe no performance benefit relative to the original DALL-E 2 prompts ($\Delta=0.0020$; $p=0.56$).
+This asymmetry reinforces the idea that the gains from prompt adaptation depend on the model’s capacity to act on that additional information.
+Figure 2: Replay Analysis and Effect Decomposition.
+(a) Average performance of prompts evaluated across four prompt-model combinations.
+Comparing DALL-E 2 prompts evaluated on DALL-E 2 versus DALL-E 3 isolates the model effect.
+Comparing reused DALL-E 2 prompts on DALL-E 3 to original DALL-E 3 prompts on DALL-E 3 isolates the prompting effect.
+Error bars represent 95% confidence intervals based on bootstrapped standard errors clustered at the participant level.
+(b) A single target image with two submitted prompts: one written by a DALL-E 2 participant (top row) and one by a DALL-E 3 participant (bottom row).
+Images show how each prompt performs on both models, illustrating the model and prompting effects qualitatively.
+Panel B of Figure 2 illustrates these effects using a single target image.
+The two rows show different prompts submitted for that target, along with the images they generate when evaluated on each model.
+In the top row, a prompt originally written by a DALL-E 2 participant yields a higher-fidelity image when replayed on DALL-E 3, demonstrating the improvement in output quality that comes from upgrading the model while holding the prompt fixed.
+In the bottom row, a prompt written by a DALL-E 3 participant produces a noticeably lower-quality image when rendered by DALL-E 2, underscoring the limits of prompt adaptation when the model lacks sufficient capacity to execute the instructions effectively.
+Taken together, these findings offer empirical support for our theoretical claim: prompt adaptation operates as a dynamic complement that users deploy in response to improved model capabilities—and accounts for a substantial share of realized performance gains.
+4.3 Skill Heterogeneity
+Effect Estimate (SE) p-value
+Total 0.0224 (0.00312) $<$ 0.00001
+Total $\times$ Performance Decile -0.000115 (0.000047) 0.0152
+Model 0.0113 (0.00186) $<$ 0.00001
+Model $\times$ Performance Decile -0.000059 (0.000025) 0.0210
+Prompting 0.0109 (0.00316) 0.000564
+Prompting $\times$ Performance Decile -0.000056 (0.000048) 0.244
+Table 1: Regression estimates for the total, model and prompting effects, and their interactions with participant performance decile.
+The total and model effects are significantly larger for lower-skilled participants, as indicated by the negative and statistically significant interaction term.
+The prompting effect does not vary significantly across deciles, suggesting that participants across the skill distribution adapt their prompts similarly.
+Table 1 presents the results of a regression analysis that tests whether the model, prompting and total effects vary systematically across participants of differing skill levels.
+We regress performance on indicators for model, prompting and total effects, as well as their interactions with participants’ performance decile (ranging from lowest to highest).
+We make three observations.
+First, the interaction between total effect and performance decile is negative and statistically significant ($-0.000115$, $p=0.0152$).
+This suggests that model improvements reduce the overall gap between high and low-performing users, consistent with proposition 1 in our conceptual framework.
+Second, the interaction between the model effect and performance decile is also negative and statistically significant ($-0.000059$, $p=0.0210$).
+This suggests that the model effect mainly benefits the lower-performing users—consistent with our theoretical predictions in proposition 2 and due to diminishing returns for those already near the performance ceiling.
+Third, we do not find any evidence that the benefits of prompt adaptation vary across the skill distribution in a meaningful way since the interaction between the prompting effect with performance decile is not statistically significant ($-0.000056$, $p=0.2444$).
+While this contrasts with our theoretical prediction in Proposition 2—which anticipated that higher-skill participants would benefit more from prompt adaptation—it is possible that the effect exists but is modest in size and difficult to detect given the statistical power of our design.
+4.4 Prompt Revision
+We also evaluate whether an automated prompt adaptation system can substitute for human-driven adaptation.
+In the DALL-E 3 (Revised) condition, user prompts were silently rewritten by GPT-4 before being submitted to the image model.
+Although participants in this condition still outperformed those using DALL-E 2 ($\Delta=0.0069$; $p=0.042$), they achieved substantially lower performance than participants using DALL-E 3 without revision.
+On average, automated prompt revision reduced the benefit of DALL-E 3 by 58% (95% CI: [40%, 76%]).
+Manual inspection of the revised prompts in the DALL-E 3 (Revised) arm confirms that GPT-4 often added extraneous details or subtly altered the intended meaning of participant instructions, leading to degraded performance.
+In this case, the automatic rewriting process was not well-aligned with the user’s goal of replicating a specific target image as faithfully as possible.
+More broadly, this result highlights a potential risk of using invisible system-level prompt rewriting: when the objectives of the rewriting mechanism are not tightly aligned with user intent, even well-intentioned modifications can interfere with the user’s ability to effectively guide the model.
+Together, these findings demonstrate that the performance gains observed when users are given access to a more capable generative model stem from both improvements in model architecture and changes in user behavior.
+Participants assigned to DALL-E 3 not only achieve higher performance but also write longer, more refined prompts over time.
+Our replay analysis shows that these adapted prompts account for nearly half of the overall performance improvement, with the remainder attributable to the model’s enhanced rendering capabilities.
+This pattern holds across the skill distribution: while lower-skilled users benefit more from model upgrades, prompt adaptation contributes consistently across all levels of baseline performance.
+Finally, automated prompt rewriting—when misaligned with user intent—fails to replicate the benefits of direct user adaptation.
+These results highlight the central role that prompting behavior plays in shaping how users interact with and benefit from improved generative AI systems.
+5 Discussion and Conclusion
+Our findings contribute to the literature on human–AI collaboration and the economics of technology by highlighting the role of prompt adaptation as a dynamic complement that co-evolves with model improvements (Böhm and Schedlberger 2023).
+As generative AI models advance—often substantially from month to month—organizations that fail to adapt their prompting strategies may forgo a meaningful share of the economic value these upgrades make possible.
+This co-evolutionary pattern is not unique to AI.
+It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (Brynjolfsson 1993, David 1990).
+However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter.
+Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (Brynjolfsson and Hitt 2000, Brynjolfsson et al.
+2021).
+Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—accounts for roughly half of the realized performance gains associated with a model upgrade.
+This finding extends research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al.
+2010, Teece et al.
+1997) and post-adoptive IT use, which emphasizes how value emerges through user experimentation and learning over time (Jasperson et al.
+2005).
+Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill.
+Participants in our study—none of whom were expert prompt engineers—improved performance through trial-and-error within a single session.
+This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (Attewell 1992, Von Hippel 2006).
+The accessibility of prompt adaptation suggests a path toward more broadly distributed productivity gains—provided that users are supported by scaffolds and interfaces that enable iterative refinement (Rogers 2003).
+At the same time, this accessibility introduces risk.
+Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve.
+This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (Henderson and Clark 1990).
+Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
+Finally, our results caution against automated prompting systems that intervene without aligning with user intent.
+In our experiment, DALL-E 3’s automatic prompt rewriting—designed to enhance usability—reduced realized performance gains by 58%.
+This does not imply that automated prompting is inherently counterproductive.
+In some settings, automation may simplify use or improve consistency.
+But when the goals of a rewriting system diverge from those of the user—as in our task, where general improvements (e.g., aesthetic quality or detail) were prioritized over faithful image replication—such interventions can interfere with the adaptive strategies users are developing.
+Similar challenges have been observed in other domains: for instance, in clinical documentation, automated prompt optimization improved consistency but still required manual revision to ensure quality and alignment (Yao et al.
+2024).
+These findings reinforce the broader insight—rooted in both human–AI collaboration and dynamic capabilities research—that adaptive user behavior is not merely a response to system performance but a critical input.
+Interventions that disrupt this adaptive process risk eroding the very dynamic complements that make advanced systems productive (Fügener et al.
+2022).
+While this study offers new insight into prompt adaptation, it has several limitations.
+First, our focus was restricted to a single transition (from DALL-E 2 to DALL-E 3) and one type of generative AI (text-to-image generation).
+Although our conceptual framework suggests these mechanisms may generalize to other domains, further research is needed to assess how these dynamics play out in text generation, programming, scientific research, and other high-stakes settings.
+Second, we observe short-run adaptation behavior in a controlled setting, whereas longer-term learning dynamics, organizational feedback structures, or team-based workflows may shape prompting strategies differently in real-world environments.
+Third, while our replay analysis isolates the effect of prompt adaptation, we do not identify which specific types of prompt modifications (e.g., lengthening, lexical substitution, structural rephrasing) have a causal effect on performance.
+Future work should investigate these finer-grained mechanisms to better inform prompting practices and interface design.
+Building on these limitations, several promising research directions emerge.
+One direction is to examine prompt adaptation longitudinally—tracking how users develop durable heuristics that generalize across domains and respond to shifting model behavior over time.
+Another line of inquiry involves organizational-level complements to prompting, such as shared prompt repositories, collaborative refinement practices, or analytics dashboards that surface effective patterns.
+A third area concerns interface design—specifically, how features like auto-complete, prompt scoring, or real-time feedback influence users’ ability to experiment and adapt.
+Finally, researchers might explore how organizations balance standardization (which streamlines processes) with adaptability in prompting workflows, in order to avoid creating rigid routines that lag behind technical advances.
+For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance.
+The near-equal contributions of model improvements and prompt adaptation suggest that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (Arthur 2009, Dosi 1982).
+As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
 Appendix A Experiment Design
@@
 The images vary in color, style, content, and complexity within and across categories.
-These images can be found online linked to the pre-registration document: https://osf.io/ejbtp.
-As we discuss in Section F, performance, and variability of performance varied substantially across images.
+These images can be found online linked to the pre-registration document: URL removed to maintain anonymity.
+Performance, and variability of performance varied substantially across images.
 In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
@@
 Data was collected between December 12, 2023 and December 19, 2023.
-Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section B.5.1).
+Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section C.5.1).
 The median time to complete our entire task, including a demographic survey, was 22 minutes.
@@
 The parameter values for image size, and quality and style for the DALL-E 3 treatment arms, were set to the same values as in the experiment.
-Appendix B Measurement and Variables
-B.1 Survey Data
-For each participant, we used a Qualtrics survey to collect demographic information, information on the participant’s skills that may be relevant to generative AI use, and information on the participant’s attitude towards generative AI.
-This data includes:
+A.5 Image Similarity Metrics
+The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al.
+2021).
+CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together.
+By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
+Because the output of each generative model is stochastic, the same prompt can yield different images across attempts.
+To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually.
+We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis.
+As a robustness check, we replicated all analyses using DreamSim, a recently developed perceptual similarity metric that aligns more closely with human judgments.
+[^9] DreamSim (Fu et al.
+2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
+Our results are robust to using DreamSim in place of CLIP cosine similarity.
+Full DreamSim-based analyses can be found in Appendix F.
+The two measures were highly correlated, and our findings were consistent across both.
+A.6 Pre-registration
+This study was pre-registered.
+The pre-registration document included our hypotheses, planned analyses, and sample size justification.
+The pre-registration document can be found at URL removed to maintain anonymity.
+Appendix B Example Prompts
+Table 2 presents the complete set of prompts and their corresponding generated images that were used in our analysis shown in Figure 1.
+The table displays three example target images (woman with shopping basket, abstract painting with cross, and online shopping interface), with three different participant-generated versions of each.
+For each target, we show examples of varying quality, arranged from most similar to least similar.
+The prompts are presented verbatim from participant submissions, preserving all original formatting and punctuation (or lack thereof).
+Table 2: Participant prompts for the images provided in Figure 1
+Image Prompt
+[Uncaptioned image] “create an image of a woman with blond hair, wearing a surgical mask, black gloves, a white t-shirt, and holding a red shopping basket full of groceries”
+[Uncaptioned image] “A lady has long blonde hair, wears BLACK COLORED GLOVES and a BLUE COLORED mask, with a white blouse on. She is holding a RED PLASTIC BASKET filled with different kinds of fruit.”
+[Uncaptioned image] “Generate a stock photo of a woman holding a basket of groceries during the pandemic. Make sure she has on black gloves.”
+[Uncaptioned image] “Thank you! Please create a mostly white painting with red and blue lines in the center, shaped a bit like the # symbol. Add red and blue splatter as well, but not too much”
+[Uncaptioned image] “Water color style.
+Japanese vibes.
+White background.
+Square outline in the middle of the page is standing on its side at a 12 degree angle.
+The inside of the square is white.
+Two legs are blue and two are red.
+T”
+[Uncaptioned image] “painting of two x’s intersecting to make a slanted square.
+The square is at a 35 degree angle with the left upper side being two red lines making the x and the lower right side two blue lines.
+The painting only has red and blue and the square with 2 x’s sits in the middle of the painting with light speckles of red and blue drops of paint scattered throughout the painting”
+[Uncaptioned image] “pastels, abstract, bearded man on left in a pink shirt with blue sleeves pointing to a shirt box image, woman with pink hair in a blue dress and high heels holding the handle of a shopping cart full of purses on right side, pink and white awning in center top, 4 out of 5 stars on left side in pink, animation style.
+single pink dollar sign on the left.
+white search bar in the center with a magnifying glass icon on the right side of the search bar.”
+[Uncaptioned image] “cartoon-y virtual self checkout. man shopping for clothes. woman pushing cart. phone that looks like a storefront in background.”
+[Uncaptioned image] “at the store a man using a touch screen monitor to shop for a shirt and a woman walking by with a store cart full of gift”
+Appendix C Measurement and Variables
+This appendix provides detailed information about the data collection, measurement, and variables used in our analyses, expanding on the overview provided in the main text.
+C.1 Survey Data
+We collected additional participant information via a Qualtrics survey that included:
   • Demographics: Ethnicity, Gender, Age, Highest level of education attained (some high school, high school, some college, associate’s degree, bachelor’s degree, master’s degree, doctoral degree, professional degree, other), Years of work experience, Annual Income (0-\$25k, \$25.001k-\$50k, \$50.001k-75k, \$75.001k-\$100k, \$100.001k-\$150k, \$150k+), and elicitation of sets of O*NET job skills that participants used in their occupation (reading comprehension, active listening, writing, speaking, critical thinking, social perceptiveness, coordination, instructing, programming, judgment and decision making, systems evaluations, science, active learning, learning strategies, monitoring, complex problem analysis, technology design, troubleshooting, quality control analysis, systems analysis).
   • Opinions and Skills: Computer programming proficiency and usage frequency (self-reported), Structured and creative writing proficiency and usage frequency (self-reported), Generative AI tool proficiency and usage frequency (self-reported), Attitudes towards net social impact of Generative AI (self-reported), Advice for (hypothetical) future participants on how to perform well on the task.
-B.2 Prompt Data
-For each prompt, we record the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
-B.3 Image Data
-For each prompt, the following images were collected:
+C.2 Prompt Data
+For each prompt, we recorded the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
+C.3 Image Data
+We collected three different sets of images:
   1. 1.
@@
 To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the experiment concluded.
-We provide full details on this procedure in Section C.
+We provide full details on this procedure in Section D.
 These images were generated from December 26, 2023 - January 27, 2024.
 These images are not used for analyses presented in the main text, but were used for other pre-registered analyses.
-These additional analyses are discussed in Section F.
+These additional analyses are discussed in Section G.
   3. 3.
@@
 For consistency, this replay data is used throughout the main text of our paper.
-B.4 Sample Construction
-The sample we analyze in the main text of our paper is constructed using the process described as follows.
+C.4 Sample Construction
+The sample analyzed in the main text was constructed as follows:
   • The initial “raw” dataset collected during the experiment is comprised of 24,672 rows of raw prompt data (one prompt per row) generated by 2,059 participants.
@@
   • Our final sample included 1,893 participants and 18,560 prompts.
-    B.4.1 “Off-Topic” Robustness Check
+C.4.1 “Off-Topic” Robustness Check
 While analyzing our data, we found that our sample contained a number of “off-topic” prompts that did not seem related to the task.
@@
 All of our main text results are robust to the exclusion of these “off-topic” prompts.
-B.5 Dependent Variables
-B.5.1 Image Similarity
-We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al., 2023).
+C.5 Dependent Variables
+C.5.1 Image Similarity
+We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al.
+2023).
 In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers.
 Our results are qualitatively and quantitatively similar using DreamSim instead.
-  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al., 2021) from Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image.
+  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al.
+2021) from Hugging Face (OpenAI 2024) for each participant-generated image and for each target image.
 Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text.
 We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
-  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023).
+  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al.
+2023).
 The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity.
@@
 We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure.
-We present the results obtained when conducting our main text analyses using DreamSim in Section E.E.1.
-B.5.2 Prompt Length
+We present the results obtained when conducting our main text analyses using DreamSim in Section F.F.1.
+C.5.2 Prompt Length
 We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt.
 In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9954$, 95% CI: [0.99528, 0.99560]).
-B.5.3 Embedding-based Prompt Similarity
+C.5.3 Embedding-based Prompt Similarity
 We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity.
-Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al., 2022).
+Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al.
+2022).
 The two similarity measures are defined as follows:
@@
     where $\mathbf{E(p_{i,n})}$ is again the vector embedding representation of participant $i$’s $n^{\text{th}}$ prompt, $p_{i,n}$, and $\overline{\mathbf{E(p_{i,n})}}$ is the element-wise mean of all 10 of participant $i$’s prompts.
-B.5.4 Successive Prompt Token Sort Ratio
+C.5.4 Successive Prompt Token Sort Ratio
 Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$.
-TSR is a fuzzy string-matching technique (Singla and Garg, 2012) that provides a continuous measure of how similar two strings are.
-We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
-B.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
+TSR is a fuzzy string-matching technique (Singla and Garg 2012) that provides a continuous measure of how similar two strings are.
+C.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
 Starting with each participant’s second prompt, we record whether each prompt $p_{i,n}$ contains the immediately preceding prompt $p_{i,n-1}$ as an exact substring.
-B.5.6 Prompt Composition
+C.5.6 Prompt Composition
 We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt.
-SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal POS tags (2024).
+SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal Dependencies Project (2024).
 These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs.
 The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
-B.5.7 Strategic Shifts
+C.5.7 Strategic Shifts
 In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting.
-In order to do so, we adapt a method proposed in (Torricelli et al., 2023) (because they are conducting research in a different context, (Torricelli et al., 2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”).
+In order to do so, we adapt a method proposed in (Torricelli et al.
+2023) (because they are conducting research in a different context, (Torricelli et al.
+2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”).
 To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
@@
 We then label any given prompt as a strategic shift ($SS$) if the cosine similarity of its embedding vector with that of the previous prompt is lower than this target-image-level mean:
-$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}(\mathbf{E(p_{i,n,t})},\mathbf{E(p_{i-1,n,t})})<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (4)
+$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}\big{(}\mathbf{E(p}_{i,n,t}),\mathbf{E(p}_{i-1,n,t})\big{)}\\ &\quad<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (9)
 It is worth noting that Torricelli et al.
@@
 We instead use the task-level mean because in our setting, as it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
-Appendix C Methods
-C.1 Stratification
-The results shown in the main text and in the SI are mostly stratified by reference image and iteration.
+Appendix D Methods
+This appendix provides additional methodological details to supplement the analyses presented in the main text.
+D.1 Stratification
+The results shown in the main text and in the supplementary information are typically stratified by reference image and iteration.
 In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level).
-The exact stratification for each finding is indicated in section D.
-To stratify our results, we take a weighted average across $j=1,...,J$ cells defined by our stratification variables.
+The exact stratification for each finding is indicated in section E.
+To stratify our results, we calculate a weighted average across $j=1,...,J$ cells defined by our stratification variables:
 $\overline{Y}_{strat}=\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}$
-To calculate the variance (and standard error) of this sample mean we apply the following:
+To calculate the variance (and standard error) of this sample mean, we apply the following:
 $\widehat{\mathrm{Var}}(\overline{Y}_{strat})=\widehat{\mathrm{Var}}\left(\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}\right)=\sum_{j=1}^{J}\left(\frac{N_{j}}{N}\right)^{2}\frac{s_{j}^{2}}{N_{j}}$
@@
   • $s_{j}$ is the sample standard deviation of stratum $j$. Therefore, $s_{j}^{2}$ is sample variance stratum $j$.
-C.2 Z-Scoring
-We find statistically significant evidence for differences in the variability of performance across the 15 target images used in our experiments, which we discuss in Section F.
-In the main text, we also showed that performance increases across the attempts.
-To test whether our results are in some way due to this image-level or attempt-level variation, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment.
+D.2 Z-Scoring
+Our analysis found statistically significant differences in performance variability across the 15 target images used in our experiments, as discussed in Section G.
+The main text also demonstrated that performance increases across successive attempts.
+To ensure our results are not driven by this image-level or attempt-level variation, we replicated all analyses using the within-image-attempt Z-score of CLIP-cosine similarity for each image produced by participants.
 Formally, this is:
-$Z(\textrm{CosineSim}_{i,n,t})=\frac{\textrm{CosineSim}_{i,n,t}-\textrm{Mean}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)}{\textrm{SD}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)},$ (5)
-where $\textrm{CosineSim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$.
-The mean and standard deviation are computed per each image-attempt, but over the DALL-E 2 and DALL-E 3 treatment arms.
-We also use the rescaling above to test the robustness of our DreamSim-based analyses.
-Almost all the robustness analysis reported here use the Z-score scaled measure of performance within each image-attempt set.
-The only exception is any analysis that examines improvements and prompts across attempts, e.g.
-figure 1B, where Z-scores are computed only within each target image, and across all attempts of that image.
-C.3 Accounting for Model Stochasticity
-The output that generative AI models return in response to a given prompt is stochastic.
-The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E API at the time of our experiment.
-To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms.
-We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples.
-Given the replicated images, we can also calculate the standard deviation of cosine similarity induced by this stochasticity.
-With the expected cosine similarity and its standard deviation per prompt, we can compute a normalized Z-score for the observed image relative to its replication distribution.
-This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and will be used for further analysis in section F.
-We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text.
+$Z(\text{Sim}_{i,n,t})=\frac{\text{Sim}_{i,n,t}-\text{Mean}_{n,t}\left(\text{Sim}_{i,n,t}\right)}{\text{SD}_{n,t}\left(\text{Sim}_{i,n,t}\right)}$ (10)
+where $\textrm{Sim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$.
+The mean and standard deviation are computed for each image-attempt pair, but across both the DALL-E 2 and DALL-E 3 treatment arms.
+We also applied this rescaling to test the robustness of our DreamSim-based analyses.
+Nearly all robustness analyses reported here use the Z-score scaled measure of performance within each image-attempt set.
+The only exception is analyses that examine improvements and prompts across attempts (e.g., Figure 1B), where Z-scores are computed only within each target image and across all attempts for that image.
+D.3 Accounting for Model Stochasticity
+Generative AI models produce stochastic outputs in response to a given prompt.
+This stochasticity is controlled by a parameter called temperature, which could not be modified using the DALL-E API at the time of our experiment.
+To account for this model stochasticity, we generated 10 images for each prompt submitted by participants across all treatment arms.
+We then calculated the similarity between each replication and its corresponding target image, and computed an ”expected” CLIP cosine similarity and DreamSim score for each prompt by averaging across these samples.
+Using these replicated images, we also calculated the standard deviation of cosine similarity induced by this stochasticity.
+With the expected cosine similarity and its standard deviation per prompt, we computed a normalized Z-score for the observed image relative to its replication distribution.
+This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and is used for further analysis in section G.
+We generated these additional samples both for the original prompts on their assigned treatment arms and by replaying prompts on the counterfactual arms, as introduced in Figure 2 in the main text.
 Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling.
 As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts.
-This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section B.B.3.3).
-Appendix D Main Text Analyses
-D.1 Task Performance and ATEs
+This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section C.C.3.3).
+Appendix E Main Text Analyses
+This section provides detailed methodological information about the analyses presented in the main text.
+E.1 Conceptual Framework
+This appendix provides detailed derivations and proofs for the conceptual framework presented in the main text.
+While the main text introduces the conceptual framework and key insights, here we describe the model in more detail and develop its mathematical foundations and results more thoroughly.
+We model a task where users attempt to replicate an object (such as an image) using a generative model.
+Perfect replication yields a maximum quality value of 1.
+The bounded quality function is defined as:
+[…]
+where $\theta\in(0,1]$ denotes the model capacity, $s\in(0,1]$ denotes the user’s skill at writing prompts, and $x\geq 0$ is the level of user’s effort in generating the appropriate prompts.
+This function satisfies two intuitive properties:
+  1. 1.
+    Boundedness: Even as $x$ becomes large, $Q$ remains below 1, reflecting that the user cannot exceed perfect replication.
+  2. 2.
+    Diminishing returns: As user effort ($x$), increases, its marginal impact on quality becomes smaller.
+We assume a linear cost function of prompting:
+$c(x)\;=\;k\,x,$
+where $k>0$ represents the marginal cost of user effort, for example in terms of time spent. The user’s utility given the quality and cost functions is
+$U(\theta,s,x)=1-e^{-\theta sx}-k\,x.$
+E.1.1 User Optimization
+The user chooses the effort $x\geq 0$ such that it maximizes $U(\theta,s,x)$.
+We assume the user’s solution is always interior with positive effort which requires $\frac{\theta s}{k}>1$.
+This can be justified since the time cost associated with the task is minimal ($k\approx 0)$.
+As such, optimal user effort is $x^{*}(\theta,s)$:
+$x^{*}(\theta,s)\;=\;\frac{1}{\theta s}\,\ln\Bigl{(}\frac{\theta s}{k}\Bigr{)}>0$
+Thus, the optimal quality the user achieves at their optimal effort level becomes:
+$Q^{*}=Q\big{(}\theta,s,x^{*}(\theta,s)\big{)}=1-\frac{k}{\theta s}>0$
+E.1.2 Comparative Statics
+We now analyze how quality, $Q^{*}$, varies with $\theta$, $s$, and $k$.
+Comparative statics are easily shown by taking the derivative of $Q^{*}$ with respect to the model capacity, user’s skill, and prompting cost.
+Model Capacity: As model capacity $\theta$ increases, optimal quality $Q^{*}$ increases.
+$\frac{\partial Q^{*}}{\partial\Theta}\Bigg{|}_{\Theta=\theta}=\frac{k}{\theta^{2}s}>0$ (11)
+User’s Skill: As the skill of the user $s$ increases, the optimal quality $Q^{*}$ also increases.
+$\frac{\partial Q^{*}}{\partial S}\Bigg{|}_{S=s}=\frac{k}{\theta s^{2}}>0$ (12)
+Prompting Cost: As the prompting cost $k$ increases, the optimal quality $Q^{*}$ decreases.
+$\frac{\partial Q^{*}}{\partial K}\Bigg{|}_{K=k}=-\frac{1}{\theta s}<0$ (13)
+In the three derivatives above, the capital letters $\Theta,S$ and $K$ correspond to variables with respect to which derivatives are taken, and small letters $\theta,s$ and $k$ are the values at which derivatives are evaluated.
+E.1.3 Model Improvement and Skill Heterogeneity
+We next show that improvements in model capacity reduce the performance gap between high and low prompting skills. This result is mainly due to the diminishing returns in output quality.
+Result. As model capacity $\theta$ improves, the effect of user skill $s$ on optimal quality $Q^{*}$ diminishes.
+Proof: This can be easily shown by taking the derivative of equation 12, which denotes the total effect of user skill, with respect to a model improvement ($\theta$).
+$\frac{\partial^{2}Q^{*}}{\partial S\partial\Theta}\Bigg{|}_{S=s,\Theta=\theta}=-\frac{k}{\theta^{2}s^{2}}<0$ (14)
+As before capital letters $\Theta,S$ correspond to variables that are evaluated at $\theta,s$. $\square$
+E.1.4 Decomposition into Model and Prompt Effects
+As discussed in the main text, we decompose the total quality improvement into two parts: a model effect (the improvement from upgrading the model but holding the prompting effort fixed) and a prompting effect (the improvement from re-optimizing the effort in response to the upgraded model).
+To perform this decomposition, we consider a “counterfactual” quality function that uses one model’s capacity but another model’s optimal prompt:
+$Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}=1-\exp\big{(}-\theta_{2}s\,x^{*}(\theta_{1},s)\big{)}$ (15)
+which is the quality obtained by using the model $\theta_{2}$ while choosing the prompting effort $x^{*}(\theta_{1},s)$ that would have been optimal if the user actually had model $\theta_{1}$.
+Note that if $\theta_{1}=\theta_{2}$, then $x^{*}(\theta_{1},s)$ is indeed the true optimum for $\theta_{2}$, yielding $Q^{*}$.
+Given the formulation above, we can now express the model effect ($M$) and the prompting effect ($P$) by taking the derivative of the counterfactual quality in equation 15 with respect to the model capacity directly or indirectly through its effect on the prompt.
+The model effect is the derivative of 15 with respect to the direct effect of model capacity, while keeping the user’s effort fixed.
+$M$ $=\;\frac{\partial Q\bigl{(}\Theta_{2},s,x^{*}(\Theta_{1},s)\bigr{)}}{\partial\Theta_{2}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$ (16)
+$=\;s\cdot x^{*}(\Theta_{1},s)\cdot\exp\left(-\Theta_{2}s\cdot x^{*}(\Theta_{1},s)\right)\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$
+$=\;\frac{k}{\theta^{2}s}\cdot\ln\left(\frac{\theta s}{k}\right)$
+We can obtain the prompt effect by taking the derivative of 15 with respect to its indirect impact on the user’s effort, while keeping the direct impact of the model capacity fixed.
+$P$ $=\;\frac{\partial Q\bigl{(}\Theta_{2},s,x^{*}(\Theta_{1},s)\bigr{)}}{\partial\Theta_{1}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$ (17)
+$=\;\Theta_{2}s\cdot\exp\left(-\Theta_{2}s\cdot x^{*}(\Theta_{1},s)\right)\cdot\,\frac{\partial x^{*}(\Theta_{1},s)}{\partial\Theta_{1}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$
+$=\;\frac{k}{\theta^{2}s}\left(1-\ln\left(\frac{\theta s}{k}\right)\right)$
+We now show that the model effect mainly benefits low-skilled users in contrast to the prompting effect which mainly benefits the high-skilled users.
+Proposition 3.
+As user skill $s$ increases, the model effect becomes smaller, for sufficiently small prompting cost $k$.
+Proof.
+To examine how the model effect varies by user skill, we can take the derivative of the model effect M in equation 16 with respect to $s$.
+$\frac{\partial M}{\partial S}\Bigg{|}_{S=s}$ $=\frac{\partial}{\partial S}\frac{k}{\theta^{2}S}\cdot\ln\left(\frac{\theta S}{k}\right)\Bigg{|}_{S=s}$
+$=\frac{k}{\theta^{2}s^{2}}\left(1-\ln\left(\frac{\theta s}{k}\right)\right)<0$
+for sufficiently small $k$. ∎
+Intuitively when prompting costs are low, users can afford to put substantial effort into writing prompts (so $x^{*}$ is high).
+Once the effort is large enough, the marginal effect of the model on quality (i.e.
+the model effect) will be decreasing by user skill.
+This happens due to the diminishing returns in the quality function.
+Once a high-skill user is already achieving near-peak quality by putting in large effort, small increases in $\theta$ deliver small additional benefit.
+However, the same model improvement yields a much larger marginal impact for a low-skilled user who is putting the same level of effort.
+Proposition 4.
+As user skill $s$ improves, the prompting effect becomes larger, for sufficiently small cost $k$.
+Proof.
+Proof To examine how the prompt effect varies by user skill, we can again take the derivative of the prompting effect P in equation 17 with respect to $s$.
+$\frac{\partial P}{\partial S}\Bigg{|}_{S=s}$ $=\frac{\partial}{\partial S}\frac{k}{\theta^{2}S}\cdot\big{(}1-\ln(\frac{\theta S}{k})\big{)}\Bigg{|}_{S=s}$ (18)
+$=\frac{k}{\theta^{2}s^{2}}\left(\ln\left(\frac{\theta s}{k}\right)-2\right)>0$
+for sufficiently small $k$. ∎
+Comparing equations E.1.4 and 18, we conclude that the prompting effect is less sensitive to user skill than the model effect, as its marginal with respect to skill has a smaller magnitude.
+This makes the overall marginal effect of user skill negative, matching the earlier finding in equation 14.
+E.2 Task Performance and ATEs
 The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations).
-It shows the average cosine similarity score stratified by the reference image.
+It displays the average cosine similarity score stratified by the reference image.
 A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms.
@@
 However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
-Next, the bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane.
-[^4] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified.
-To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
-$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (6)
+The bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane.
+[^10] In Section E, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified.
+To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
+$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (19)
+$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
 The coefficient estimates generated by this model are:
@@
 Standard errors are cluster robust at the participant level.
-D.2 Prompt Characteristics
+E.3 Prompt Characteristics
 Figure 1C compares the prompt length and prompt similarity of the two models.
-To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section B.B.4.
+To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section C.C.4.
 The prompt length is the average number of words per model and iteration stratified by the reference image.
-The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section B.B.5.3 for details on similarity calculations).
+The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section C.C.5.3 for details on similarity calculations).
 The color scale in Figure 1C shows the stratified average similarity to the previous prompt across all users per each model.
@@
 Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants.
-For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable.
+For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section C.C.5.3, as the dependent variable.
 When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$).
 Standard errors are cluster-robust at the participant level.
-D.3 ATE Decomposition
+E.4 ATE Decomposition
 Figure 2 in the main text decomposes the ATE into the model and prompting effects.
@@
 In the terminology of mediation analysis, the model effect would be referred to as the direct effect.
-The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (3,3).
+The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (3,3) relative to (2,3).
 In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect.
@@
 These uncertainty estimates are insufficient for exact inference on the direct, indirect, and treatment effects.
-The statistics and significance values report in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
+The statistics and significance values reported in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
 $Y_{i,n,t}=\beta_{0}+\beta_{1}\text{effect}+\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$ (20)
@@
 The standard errors for each estimated model are cluster robust at the participant level, and p-values are adjusted accordingly.
-Appendix E Robustness Checks
-E.1 DreamSim-based Analysis
-As discussed in Section B, we repeat all main-text analyses with DreamSim. The results of these analyses are reported below.
-E.1.1 Overall ATEs
+E.5 Heterogeneity by Skill
+Table 1 in the main text demonstrates how total, model, and prompting effects vary across different user skill levels.
+The total effect compares the Cosine similarity of outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 2.
+The model effect compares outputs from DALL-E 2 users replayed on DALL-E 3 against the same users replayed on DALL-E 2.
+The prompting effect compares outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 3.
+To assess user skill, users within each target image, iteration and replay scenario are divided into 50 equally sized brackets based on their performance, assigning each bracket a rank corresponding to performance deciles.
+For instance, when evaluating DALL-E 2 users replayed on DALL-E 2, each user’s skill level is determined by their percentile rank among all DALL-E 2 users replayed on DALL-E 2.
+Conversely, when the same user is replayed on DALL-E 3 (to estimate the model effect or to serve as a baseline in estimating the prompting effect), their performance decile is determined by their rank relative to all DALL-E 2 users replayed on DALL-E 3, per each iteration and target image.
+After assigning performance deciles, we estimate the following linear model with two-way fixed effects (iteration and target image) for each effect type (model, prompting, and total):
+$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\text{effect}+\beta_{2}\text{effect}\times\text{Performance Decile}_{i,n,t}\,+$ (21)
+$\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$
+where $\beta_{1}$ represents the coefficient associated with the effect type being analyzed (model, prompting, or total), and $\beta_{2}$ captures the interaction between the effect type and user skill level.
+User skill is measured as the user’s rank (binned) within the same iteration, target image, and replay scenario.
+Standard errors are calculated using cluster-robust methods at the participant level.
+This methodology closely follows the procedure described earlier in section E.4, with the additional step of incorporating an interaction term to evaluate skill-level heterogeneity.
+Appendix F Robustness Checks
+This section provides additional analyses that confirm the robustness of our main findings using alternative metrics and statistical approaches.
+F.1 DreamSim-based Analysis
+As discussed in Section C, we repeated all main-text analyses using DreamSim as an alternative image similarity metric.
+The results, presented below, demonstrate that our findings are not dependent on the specific similarity measure used.
+F.1.1 Overall ATEs
 In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.238$ standard deviations (95% CI = [0.152, 0.324]) closer to the target image ($\Delta DreamSim=0.0306$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
-E.1.2 Figure 1
-We reran the regression in Section D.D.1 with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
-$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (8)
+F.1.2 Figure 1
+We reran the regression in Section E with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
+$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (22)
+$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
 The coefficient estimates generated by this analysis are:
@@
   • $\hat{\beta_{3}}=0.0024$, $\hat{SE}(\beta_{3})=0.0007$, $p=0.0009$
-E.1.3 Figure 2
+F.1.3 Figure 2
 Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text.
 The model effect accounts for 54.4% of the ATE ($\Delta DreamSim=0.0166$, $p<10^{-7}$), whereas the prompting effect accounts for 45.4% of the ATE ($\Delta DreamSim=0.01390$, $p=0.014$).
-E.2 Z-score-based Analysis
-As we discuss in Section C.2, we repeat all main-text analyses with the within-image-attempt Z-score of CLIP-cosine similarity to better account for variation between images and attempts.
-When comparing across attempts, the Z-score is computed within each image, as mentioned in section C.2.
-The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
-E.2.1 Overall ATEs
+F.2 Z-score-based Analysis
+As we discuss in Section D.2, we repeated all main-text analyses using the within-image-attempt Z-score of CLIP-cosine similarity to account for variation between images and attempts.
+When comparing across attempts, the Z-score was computed within each image, as mentioned in Section D.2.
+The results remain consistent throughout, and in some cases, the differences between the superior and inferior model are even more pronounced than when using raw cosine similarity.
+F.2.1 Overall ATEs
 As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.19$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.19, 95% CI = [0.100, 0.271]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2 (the inferior model).
 Standard errors are clustered at the participant level.
-E.2.2 Figure 1
+F.2.2 Figure 1
 On average, participants using the superior model produced images that were $z=0.19$ standard deviations closer (the ATE) to the target image than those using the inferior model.
 Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
-$\text{ZScore}_{i,n}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\epsilon_{i,n}$ (9)
+$\text{ZScore}_{i,n}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (23)
+$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\epsilon_{i,n}$
   • $\hat{\beta_{1}}=0.0129$, $\hat{SE}(\beta_{1})=0.0038$, $p=0.0007$
@@
   • $\hat{\beta_{3}}=0.0128$, $\hat{SE}(\beta_{3})=0.0053$, $p=0.015$
-E.2.3 Figure 2
+F.2.3 Figure 2
 When we decompose the ATE into the model effect ($z=0.0791$; $p=8.35\times 10^{-6}$) and prompting effect ($z=0.1046$; $p=0.016$), they account for 43% and 56% of the treatment effect, respectively.
@@
 In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z-score of the cosine similarity.
-Appendix F Pre-registration
-Prior to our experiment, we pre-registered a number of hypotheses and a pre-analysis plan.
-This pre-registration is deposited at OSF at the following URL: https://osf.io/ejbtp.
-The main text of this paper contains a subset of our pre-registered analysis, as well as complementary exploratory analyses that are also mentioned in our pre-registration.
-We chose to present this subset of our pre-registered analyses because we believe this subset constitutes an important and timely set of results best-suited to a short-form paper.
-Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses.
-In our pre-registration, we declared the intent to conduct each of our analyses using six666see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
-possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
-  • No rescaling: The outcome variable is used as-is.
-  • Z-score rescaling: We rescale the outcome variable into a Z-score according to the procedure describe in these supplementary materials.
-  • Percentile rank rescaling: We rescale the outcome variable into a percentile rank.
-This is done by calculating the percentile rank of a given prompt relative to all other prompts submitted for the relevant target image.
-We intend to complete each pre-registered analysis with each of these eight possible outcome variables.
-Although we have not yet done so, we have no reason to believe that our results will not be robust to these different transformations of our data.
-However, we did not want to delay the publication of our preprint until all of these analyses were complete.
-Thus, we are posting a preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes.
-As we complete more of these analyses, we will update this supplement.
-We also intend to post full results, data, and replication code online shortly.
-F.1 Hypotheses and Results
-Below, we list the hypotheses precisely as they are written in our pre-registration document. For each hypothesis, we also describe our results.
+Appendix G Pre-registration
+Prior to conducting our experiment, we developed and pre-registered a comprehensive set of hypotheses and analyses.
+This pre-registration is deposited at OSF at the following URL: URL removed to maintain anonymity.
+While the main text focuses on a subset of our pre-registered analyses that constitute the most important and timely findings, this appendix provides a comprehensive overview of all analyses specified in our pre-registration.
+In our pre-registration, we outlined plans to conduct each analysis using six [^12] See section G.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
+possible outcome variables, all representing different transformations of the same underlying data.
+These include CLIP embedding cosine similarity and DreamSim, each with three rescaling methods:
+  • No rescaling: The outcome variable used in its original form.
+  • Z-score rescaling: The outcome variable transformed into a Z-score following the procedure detailed in Section D.2.
+  • Percentile rank rescaling: The outcome variable converted to a percentile rank relative to all other prompts submitted for the same target image.
+While we are still in the process of completing all analyses with each of these outcome variables, our preliminary results suggest that our findings are robust across these different transformations.
+As additional analyses are completed, we will update this appendix.
+Full results, data, and replication code available online upon publication.
+G.1 Hypotheses and Results
+Below, we present each hypothesis exactly as stated in our pre-registration document, followed by a summary of our findings.
+This comprehensive review demonstrates the robustness of the key results highlighted in the main text, while also providing additional insights into the relationship between model capabilities, user behavior, and task performance.
 H1
 There are differences in prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) across demographic attributes and other observables, such as educational background and occupational skills.
-Analysis approach:
-For this hypothesis, we conducted multiple ANOVA tests, one for each of the demographic variables, against the relevant outcome variable and treatment variable (the model) as the covariate.
-As a robustness check, we repeated the same procedure using the Kruskal-Wallis U test.
-To adjust for multiple testing, we used the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
-Results: The results of our analysis are as follows:
+Analysis approach: For this hypothesis, we conducted multiple ANOVA tests, one for each demographic variable, with the relevant performance measure as the dependent variable and the model (DALL-E version) as a covariate.
+As a robustness check, we repeated these analyses using the non-parametric Kruskal-Wallis U test.
+To account for multiple comparisons, we applied the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
+Results: Our analyses revealed several consistent demographic patterns in prompt engineering ability across different outcome measures:
   • CLIP embedding cosine similarity
-    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, active listening and quality control.
-    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as job skills, and report little usage of generative AI or imagery writing skill are, on average, better at the task.
-Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with self-reported occupational skill of quality control, and those who programmed more frequently performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
-In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
-    Z-score rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, quality control, technology design, social perceptiveness and troubleshooting.
-    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and troubleshooting as job skills, and report little usage of generative AI, little imagery writing skill, some programming skill, and some instruction writing skill are, on average, better at the task.
-Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with a graduate degree, those with self-reported occupational skill of quality control and technology design, and those who programmed more frequently performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
-In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
-    Percentile rank rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, and self-reported occupational skill of critical thinking.
-    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as a job skill, and report little usage of generative AI, and some programming skill are, on average, better at the task.
-Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables.
-In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
+    • No rescaling: Using ANOVA, we found significant associations between performance and: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills (particularly critical thinking, active listening, and quality control).
+      In linear models examining directionality, we found that participants who reported critical thinking as a job skill and little usage of generative AI or imagery writing skill performed better, on average.
+Conversely, older participants, men, those with more positive outlooks regarding generative AI, those reporting quality control as an occupational skill, and frequent programmers performed worse on our task.
+      The more conservative Kruskal-Wallis tests identified fewer significant variables: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender, and age.
+    • Z-score rescaling: ANOVA tests revealed significant associations with similar demographic factors as the unscaled measure, plus additional significant associations with self-reported occupational skills in technology design, social perceptiveness, and troubleshooting.
+      Linear models showed that participants reporting critical thinking and troubleshooting as job skills, little generative AI use, little imagery writing skill, some programming skill, and some instruction writing skill performed better.
+Worse performance was associated with older participants, men, positive generative AI outlook, graduate degrees, self-reported quality control and technology design skills, and frequent programming experience.
+      Kruskal-Wallis tests again identified a smaller set of significant predictors: self-reported programming frequency, generative AI outlook, programming skill, gender, and age.
+    • Percentile rank rescaling: ANOVA identified significant associations between performance and computer programming frequency, self-reported programming ability and frequency, generative AI outlook, age, gender, generative AI use, and critical thinking as an occupational skill.
+      Linear models showed better performance among those reporting critical thinking as a job skill, little generative AI use, and some programming skill.
+Worse performance was associated with older participants, men, positive generative AI outlook, and frequent programming experience.
+      Kruskal-Wallis tests revealed the same set of significant predictors as with other scalings: programming frequency, generative AI outlook, programming skill, gender, and age.
   • DreamSim
-    No rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skill, and self-reported occupational skills of critical thinking, learning strategies, technology design and quality control.
-    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task.
-Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies occupational skill.
-    Z-score rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skills, education, and self-reported occupational skills of critical thinking, social perceptiveness, learning strategies, technology design and quality control.
-    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task.
-Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, those with graduate degrees, and those with technology design, quality control and learning strategies as occupational skills performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
-    Percentile rank rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery writing skill, education, and self-reported occupational skills of social perceptiveness, learning strategies, and technology design.
-    When we fit a linear model to evaluate directionality, we find that those who use social perceptiveness as a job skill, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task.
-Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, and those with technology design and learning strategies as occupational skills performed worse on our experiment task.
-    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, generative AI use, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
+    • No rescaling: ANOVA tests identified significant associations between performance and: programming frequency, self-reported programming ability, generative AI outlook, age, gender, generative AI use, imagery and instructional writing skills, and several occupational skills (critical thinking, learning strategies, technology design, and quality control).
+      Linear models showed better performance among those reporting critical thinking and social perceptiveness as job skills, little generative AI use, and little imagery writing or programming skills.
+Worse performance was associated with older participants, men, positive generative AI outlook, and frequent programming experience.
+      Kruskal-Wallis tests identified significant relationships with: programming frequency, generative AI outlook, imagery writing skill, gender, age, and learning strategies as an occupational skill.
+    • Z-score rescaling: Using ANOVA, significant associations included all factors identified in the unscaled analysis, plus education and additional occupational skills (social perceptiveness).
+      Linear model directional findings were consistent with the unscaled measure, with additional negative associations for graduate degrees and self-reported technology design, quality control, and learning strategies skills.
+      Kruskal-Wallis tests showed significant relationships with: programming frequency, generative AI outlook, imagery writing skill, gender, age, and occupational skills in learning strategies, social perceptiveness, and technology design.
+    • Percentile rank rescaling: ANOVA identified significant associations with: programming frequency, programming ability, generative AI outlook, age, gender, generative AI use, imagery writing skill, education, and occupational skills in social perceptiveness, learning strategies, and technology design.
+      Linear models showed better performance among participants reporting social perceptiveness as a job skill, little generative AI use, and little imagery writing or programming skills.
+Worse performance was associated with older participants, men, positive generative AI outlook, frequent programming experience, and self-reported technology design and learning strategies skills.
+      Kruskal-Wallis tests revealed significant relationships with: programming frequency, generative AI use, generative AI outlook, imagery writing skill, gender, age, and occupational skills in learning strategies, social perceptiveness, and technology design.
+Across these analyses, several consistent patterns emerge.
+Interestingly, we found that frequent programming experience was associated with worse task performance, contrary to what might be expected.
+Men and older participants also performed worse on average.
+Those reporting critical thinking or social perceptiveness as job skills tended to perform better, while participants with positive outlooks toward generative AI and higher generative AI usage performed worse.
+These patterns were generally consistent across different outcome measures and scaling approaches, suggesting robust demographic differences in prompt engineering ability.
 H2
@@
 Such prompting techniques might include the use of longer prompts, the use of structured prompting techniques, and/or specific patterns in the way that the participant iterates on their prompts over time.
-Analysis approach:
-To investigate this hypothesis, we pre-registered evaluating how exploration/exploitation in the prompting space is related to performance.
-Hence, most of the pre-registered independent variables measured the level of similarity (i.e., exploitation) across a user’s prompts.
-We refer to participants being more “exploitative” if they wrote prompts that are more similar to their previous prompts and more “exploratory” if they wrote prompts that deviated more from their previous prompts.
-As described below, we operationalized “similarity” in a variety of ways.
+Analysis approach: We investigated this hypothesis by examining the relationship between exploration/exploitation strategies in the prompting space and performance outcomes.
+We characterized participants as more ”exploitative” if they wrote prompts that were similar to their previous prompts, and more ”exploratory” if their prompts exhibited greater deviation from previous attempts.
+We operationalized these concepts using several metrics:
   1. 1.
-    Positively associated with exploitation: average token sort ratio compared to the previous prompt, average cosine similarity between the embeddings of a given prompt with the previous prompt, fraction of times a prompt contains the previous prompt as an exact substring
+    Measures positively associated with exploitation:
+    • Average token sort ratio compared to the previous prompt
+    • Average cosine similarity between the embeddings of consecutive prompts
+    • Fraction of times a prompt contains the previous prompt as an exact substring
   2. 2.
-    Negatively association with exploitation: the variance of the prompt embedding, the number of topical transitions
-We also pre-registered measuring each participant’s average prompt length.
-These variables were measured per each user and across their first 10 attempts.
-Their association with performance, measured in terms of each of our eight outcome variables, was estimated in a linear model with DALL-E version fixed effects.
-We also conducted a similar analysis at the iteration level to answer the following question: how is exploration/exploitation associated with subsequent performance in the next attempt, and is the strength of this relationship mediated by the quality of the previous attempt?
-To answer this question, first, we divided user-iteration observations into 6 equal-sized brackets by performance in the previous iteration.
-The bracketing allows us to explore heterogeneity in prompting behavior by the quality of previous attempts.
+    Measures negatively associated with exploitation:
+    • Variance of the prompt embedding
+    • Number of topical transitions
+We also measured each participant’s average prompt length.
+These variables were calculated for each user across their first 10 attempts, and their association with performance was estimated using linear models with DALL-E version fixed effects.
+To examine how exploration/exploitation strategies influence performance across successive attempts, we conducted an additional analysis at the iteration level.
+We divided user-iteration observations into 6 equal-sized brackets based on performance in the previous iteration, allowing us to explore how prompting behavior might vary depending on the quality of previous attempts.
 We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket.
-Our estimates adjusted for other covariates by matching user-iteration observations on the target image, DALL-E model, iteration and the exact quality of the previous attempt (Sävje et al., 2021).
-Results: The results of our analysis are as follows:
+Our estimates adjusted for other covariates by matching user-iteration observations on target image, DALL-E model, iteration number, and exact quality of the previous attempt (Sävje et al.
+2021).
+Results: Our analyses revealed consistent patterns across different outcome measures:
   • CLIP embedding cosine similarity
-    No rescaling: At the user-level, we find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing.
-Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance.
-Taken together, these findings suggest that more successful users engaged in more exploitation and wrote prompts that were similar to one another.
-We also found that longer prompts were associated with higher performance, as shown in the main text.
-    At the user-prompt level, we find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance.
-In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance.
-Using token sort ratio with the previous prompt as the measure of exploration also generated similar results.
-However, when using binary measures of exploration, i.e.
-topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket.
-As these are binary measures of exploration, we could not replicate the non-linearity we observed with continuous measures of exploration in the low-performing group.
-However, we do find that topical transitions (more exploration) leads to an overall drop in performance and the drop becomes larger for higher performance in the previous attempt.
-We find a similar pattern when measuring exploitation by whether a prompt contains the previous one.
-Performance is overall higher if the prompt includes the previous one and the improvement is larger as the previous performance gets higher.
-    Z-score rescaling: We find the same results as those described for the unscaled cosine similarity above.
-The main difference is that the non-linearity between performance and continuous measures of exploitation, i.e.
-TSR ratio and cosine similarity with previous prompt, observed at the user-prompt level becomes starker for the bottom two brackets of previous performance.
-For low-performing prompts in the previous attempt, the optimal exploitation level is in the middle and the performance greatly deteriorates for high and low levels of exploitation.
-    Percentile rank rescaling: We find the same results as described above, however the non-linearity in exploration/exploitation vs performance in the bottom two brackets of previous performance is now even stronger than those observed with Z-score rescaled cosine similarity as the performance measure.
+    • No rescaling: At the user level, we found strong and statistically significant associations between performance and prompting strategies, even after Benjamini-Hochberg adjustment for multiple testing.
+Specifically, token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were all positively associated with performance.
+Conversely, embedding variance and number of topical transitions showed negative correlations with performance.
+Together, these findings indicate that more successful users engaged in greater exploitation, writing prompts that exhibited higher similarity to one another.
+As noted in the main text, we also found that longer prompts were associated with higher performance.
+      At the user-prompt level, we observed an interesting pattern: when previous performance was poor, moderate exploration (lower cosine similarity with the previous prompt) was associated with improved subsequent performance, though extremely high levels of exploration did not yield additional benefits.
+In contrast, when previous performance was high, increased exploitation consistently improved subsequent performance.
+Similar results emerged when using token sort ratio as the measure of exploration.
+      With binary measures of exploration (topical transitions or containing the previous prompt), exploitation was associated with higher performance in the next iteration regardless of previous performance bracket.
+Since these are binary measures, they couldn’t capture the non-linear relationship observed with continuous measures in low-performing groups.
+Nevertheless, we found that topical transitions (more exploration) led to overall performance decreases, with larger decreases for higher previous performance.
+Similarly, prompts that included the previous prompt showed higher performance, with larger improvements as previous performance increased.
+    • Z-score rescaling: Results were consistent with those for unscaled cosine similarity.
+The primary difference was that the non-linear relationship between performance and continuous measures of exploitation (TSR ratio and cosine similarity with previous prompt) became more pronounced for the bottom two brackets of previous performance.
+For low-performing prompts in the previous attempt, moderate exploitation levels yielded optimal performance, with significant performance deterioration at both high and low exploitation levels.
+    • Percentile rank rescaling: We found the same general patterns described above, with the non-linearity in exploration/exploitation versus performance in the bottom two brackets of previous performance even stronger than observed with Z-score rescaled cosine similarity.
   • DreamSim
-    No rescaling: We find exactly the same results as those described for the unscaled cosine similarity above.
-    Z-score rescaling: We find the same results as those described for the Z-score cosine similarity above.
-In particular, we again find that the non-linearity between performance and continuous measures of exploitation observed at the user-prompt level is stark for the bottom two brackets of previous performance.
-    Percentile rank rescaling: We find the same results as described above for cosine similarity percentile rank rescaling.
+    • No rescaling: Results were consistent with those described for unscaled cosine similarity.
+    • Z-score rescaling: Results matched those described for Z-score cosine similarity, with particularly stark non-linearity between performance and continuous measures of exploitation at the user-prompt level for the bottom two brackets of previous performance.
+    • Percentile rank rescaling: Results were consistent with those described for percentile rank rescaled cosine similarity.
+These findings suggest a nuanced relationship between exploration/exploitation strategies and performance.
+Users who were generally more successful tended to employ more exploitative strategies, refining and building upon previous prompts rather than making dramatic changes.
+However, at the iteration level, the optimal strategy depended on previous performance: when performance was already high, continued exploitation yielded the best results; when performance was poor, moderate exploration was beneficial.
+This pattern was consistent across different outcome measures and scaling approaches.
 H3
 There are differences in prompt engineering techniques (as measured through metrics such as prompt length and iteration-to-iteration token sort ratio) across demographic attributes and other observables, such as educational background and occupational skills.
-Analysis approach:
-To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest.
-To account for multiple testing, we adjusted the p-values for all these models according to the Benjamini-Hochberg procedure.
-Results: The results of our analysis are as follows. As a reference, the definitions of these dependent variables are provided in Section B.B.5.
-  • Prompt embedding variance: We did not find any significant differences across various demographic traits.
-  • Strategic Shifts: We observed statistically significant differences based on age, reported programming and instructional writing frequencies.
-Younger participants, those with low programming frequency and those with some instructional writing frequency demonstrated decreased topical transitions across their prompts.
-  • Successive Prompt Token Sort Ratio: We found significant differences by age, education, programming frequency, and imagery/instructional writing frequencies.
-Older users, those with post-graduate degrees, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
-  • Successive similarity: We found significant differences by age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency and some occupational skills.
-On average, older users, males, those with post-graduate degrees, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, those who report frequently writing instructions or imagery, and those with self-reported occupational skills of critical thinking and social perceptiveness write prompts that are less similar to each other (as measured by cosine similarity) on average.
-  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and imagery writing frequency.
-Older users and those with neutral outlook toward generative AI are less likely to write prompts that exactly contain the previous.
-In contrast, males and those with some imagery writing skills are more likely to keep using their previous prompt.
+Analysis approach: To test this hypothesis, we estimated linear models with each demographic trait as the independent variable, treatment arm fixed effects as controls, and various prompting behaviors (described in Section C.C.5) as the dependent variables.
+To account for multiple testing, we adjusted p-values across all models using the Benjamini-Hochberg procedure.
+Results: Our analysis revealed several significant demographic differences in prompting strategies, with consistent patterns emerging across different measures of prompt similarity and evolution.
+The detailed findings for each dependent variable are summarized below:
+  • Prompt embedding variance: We did not find any significant differences across demographic traits in the overall variance of prompt embeddings, suggesting that the breadth of prompt space exploration was relatively consistent across different demographic groups.
+  • Strategic Shifts: We observed statistically significant differences based on age, programming frequency, and instructional writing frequency.
+Specifically, younger participants, those with low programming frequency, and those with some instructional writing experience demonstrated fewer topical transitions across their prompts, indicating a more consistent approach to prompt development.
+  • Successive Prompt Token Sort Ratio: Significant differences emerged by age, education, programming frequency, and imagery/instructional writing frequencies.
+Older users, those with post-graduate degrees, those with high programming frequency, and those with high writing frequency (both for precise instructions and imagery) wrote prompts that were less similar to their previous attempts as measured by token sort ratio.
+This suggests these groups took a more exploratory approach to prompt iteration.
+  • Successive similarity: We found significant differences across numerous demographic factors: age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency, and certain occupational skills.
+On average, older users, males, those with post-graduate degrees, frequent programmers, self-reported skilled programmers, those with positive outlooks toward generative AI, frequent generative AI users, those who frequently write instructions or imagery, and those reporting critical thinking and social perceptiveness as occupational skills wrote prompts with lower cosine similarity to their previous attempts.
+This consistent pattern across multiple demographic factors suggests robust differences in exploration-exploitation tendencies.
+  • Successive Prompt ’Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook toward generative AI, and imagery writing frequency.
+Older users and those with neutral outlooks toward generative AI were less likely to write prompts that contained their previous prompt as an exact substring.
+In contrast, males and those with some imagery writing skills were more likely to build directly upon their previous prompts by including them in subsequent attempts.
+These findings reveal a nuanced picture of demographic differences in prompt engineering approaches.
+Overall, younger users, those with less programming experience, and those with less generative AI experience tended to employ more exploitative strategies, building more consistently upon their previous prompts.
+In contrast, older users, those with more technical backgrounds, and those with more exposure to generative AI systems exhibited more exploratory behaviors, making larger changes between successive prompts.
+Interestingly, while H1 showed that a more exploitative approach was generally associated with better performance, here we find that demographics associated with technical expertise (programming experience, education, etc.) tend toward more exploratory approaches.
+This apparent contradiction suggests that the relationship between background, prompting strategy, and performance is complex and potentially mediated by other factors not captured in these analyses.
 H4
 Insofar as the output returned by a generative AI model in response to a prompt is stochastic, the subsequent prompting strategies and prompting outcomes of participants that get lower-than-expected, higher-than-expected, or approximately expected outputs in response to their first prompt are different.
-Analysis approach: To test this hypothesis, we first calculated the Z-score of each participant’s realized image observed during our experiment relative to the sampling distribution that we approximated for the prompt that generated that image according to the procedure outlined in Section B.B.3 (see Section C for details on calculating the Z-score).
-This Z-score quantifies the random variation in observed image quality relative to the prompt’s true underlying quality.
-Images with higher Z-scores represent instances where the realized image quality exceeded expectations based on the prompt, while lower Z-scores indicate instances where the image quality was randomly lower than expected for a given prompt.
-By examining the relationship between these Z-scores and subsequent performance and prompting behavior, we can assess the causal impact of this stochasticity on user behavior.
-To estimate the relationship between this stochasticity on subsequent prompting behavior and performance, we transformed the Z-score into a trichotomous variable where any Z-score less than -0.45 as “lower-than-expected,” between -0.45 and 0.45 as “expected” and greater than 0.45 as “higher-than-expected.” We then perform two-sample t-tests with comparing the three groups’ relative performance.
-We also estimate a linear model as a robustness check, regressing participant performance on the trichotomous Z-score variable with treatment arm fixed effects.
-We also repeat this analysis, treating the Z-score as a continuous variable without the trichotomous transformation.
-Finally, we also perform this analysis at the user level with the trichotomous variable, where only the Z-score of the observed image of the first prompt is measured, and we test if that affects the average performance of all subsequent user attempts.
-Results: The results of our analysis are as follows, with each outcome variable clearly labeled.
+Analysis approach: To test this hypothesis, we examined how the random variation in image quality resulting from model stochasticity affects subsequent user behavior.
+We first calculated a Z-score for each participant’s generated image relative to the expected distribution for that prompt, following the procedure outlined in Section C.C.3 (with computational details in Section D).
+This Z-score quantifies how much better or worse the actually generated image was compared to what would be expected for that prompt on average.
+Higher Z-scores represent instances where the realized image quality exceeded expectations, while lower Z-scores indicate instances where the quality was randomly lower than expected.
+To analyze the relationship between this stochasticity and subsequent prompting behavior and performance, we transformed the Z-score into a trichotomous variable with three categories: ”lower-than-expected” (Z-score ¡ -0.45), ”expected” (-0.45 $\leq$ Z-score $\leq$ 0.45), and ”higher-than-expected” (Z-score ¿ 0.45).
+We then performed two-sample t-tests comparing performance and prompting behavior across these three groups.
+As a robustness check, we also estimated linear models regressing performance and prompting measures on both the trichotomous Z-score variable and the continuous Z-score, with treatment arm fixed effects.
+Additionally, we conducted analyses at the user level to determine whether the Z-score of the first prompt affects average performance across all subsequent attempts.
+Results: Our analyses revealed several consistent patterns across different outcome measures:
   • CLIP embedding cosine similarity
-    No rescaling: We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt.
-When we test differences in the trichotomous variable, the difference between the top bracket and the bottom bracket in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than the bottom bracket.
-But we do not find significant differences between the bottom and middle brackets and between the top and middle brackets.
-When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration.
-This is due to the non-linearity of the relationship between these variables that exists separately in the negative and positive range of realized Z-score.
-Additionally, when we estimate the effect at the user-level, we find differences between the top and middle brackets of the first prompt realized Z-score.
-However, these differences are barely significant ($p=0.048$) and do not account for multiple testing.
-When fitting a linear model at the user-level, we do not find a a statistically significant relationship between the observed Z-score of the first prompt and average performance in the subsequent attempts.
-    Z-score rescaling: The results are very similar to those described for the unscaled cosine similarity above.
-In particular, we find a statistically significant difference between the top bracket and the bottom bracket, with the top bracket having a higher performance in the next attempt than the bottom bracket.
-We don’t find statistically significant effects when comparing the middle with top or bottom brackets, or when fitting a linear model.
-In contrast to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when comparing the brackets on the realized Z-score of the first image nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
-    Percentile rank rescaling: We find the exact same results as those explained above for Z-score cosine similarity.
+    • No rescaling: We found statistically significant evidence that higher Z-scores for images in previous prompts were associated with increased cosine similarity in subsequent attempts.
+When comparing performance across the trichotomous Z-score variable, the difference between the top bracket (”higher-than-expected”) and bottom bracket (”lower-than-expected”) was statistically significant, with the top bracket showing better performance in the next attempt.
+However, we did not find significant differences between the middle bracket and either the top or bottom brackets.
+      The linear model did not yield a statistically significant relationship between Z-score and performance in the next iteration, likely due to non-linearity in this relationship across the negative and positive ranges of the Z-score distribution.
+      At the user level, we found differences between the top and middle brackets of the first prompt’s Z-score, though these differences were marginally significant (p=0.048) and did not account for multiple testing.
+Linear models at the user level did not reveal statistically significant relationships between the first prompt’s Z-score and average performance in subsequent attempts.
+    • Z-score rescaling: Results were similar to those for unscaled cosine similarity.
+We found a statistically significant difference between the top and bottom Z-score brackets, with higher performance in the next attempt for the top bracket.
+No statistically significant effects emerged when comparing the middle bracket with either the top or bottom brackets, or when using linear models.
+Unlike with the unscaled measure, we found no statistically significant effects at the user level, either when comparing Z-score brackets or when using linear models.
+    • Percentile rank rescaling: Results were consistent with those for Z-score rescaled cosine similarity.
   • DreamSim
-    No rescaling: We find similar evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt.
-The difference between the top bracket and the bottom/middle brackets in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than both the bottom and middle brackets.
-But we do not find significant differences between the bottom and middle brackets.
-When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration, likely due to the non-linearity that we also observed between performance measured as cosine similarity and realized Z-score of the previous prompt.
-When we estimate the effect at the user-level, we do not find any statistically significant relationship between realized quality of the first prompt and performance in subsequent attempts, either using the trichotomous variable on Z-score or using a linear model.
-    Z-score rescaling: The results are very similar to those described for the unscaled dreamsim score above.
-In particular, we find a statistically significant difference between the top bracket and both the bottom and middle brackets, with the top bracket having a higher performance in the next attempt.
-We don’t find statistically significant effects when comparing the middle with the bottom bracket, or when fitting a linear model.
-Similar to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when using the trichotomous variable nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
-    Percentile rank rescaling: We find the exact same results as those explained above for Z-score dreamsim similarity, with the only difference that at the user-attempt level, we observe a statistically significant difference only between the top and bottom brackets.
-The top bracket in realized quality no longer has a statistically significant improvement in the performance of the next attempt when compared to the middle bracket.
+    • No rescaling: Results showed similar patterns to CLIP embedding cosine similarity.
+Higher Z-scores in previous prompts were associated with better performance in subsequent attempts.
+The difference between the top Z-score bracket and both the bottom and middle brackets was statistically significant, with the top bracket showing higher performance in the next attempt.
+However, the difference between the bottom and middle brackets was not significant.
+      As with cosine similarity, linear models did not yield statistically significant relationships between Z-scores and subsequent performance, likely due to non-linearity.
+At the user level, we found no statistically significant relationships between the first prompt’s Z-score and performance in subsequent attempts.
+    • Z-score rescaling: Results closely matched those for unscaled DreamSim scores, with statistically significant differences between the top bracket and both the bottom and middle brackets, but no significant difference between the bottom and middle brackets.
+No significant effects emerged from linear models or from user-level analyses.
+    • Percentile rank rescaling: Results aligned with those for Z-score rescaled DreamSim, with one exception: at the user-attempt level, we observed a statistically significant difference only between the top and bottom brackets, not between the top and middle brackets.
   • Prompting Behaviors
-    Prompt Length: e find statistically significant evidence that increases in the Z-score of a realized image causes the subsequent prompt to be longer (either in terms of words or characters).
-When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with on average longer prompts in the top bracket.
-However, we do not observe statistically significant differences between either the top or bottom brackets with the middle bracket.
-When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant.
-A unit increase in the Z-score realized quality of an image causes the next prompt to be longer by about 0.3 words on average.
-Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the length of the subsequent prompts, when using the trichotomous variable or when fitting a linear model.
-    Successive Similarity: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in successive similarity of the prompt.
-When we test the differences using the trichotomous variable, the difference between the top Z-score bracket (better than expected images) and both the bottom (worse than expected images) and middle (about expected images) brackets is statistically significant, with the top bracket leading to higher similarity with the previous prompt.
-However, we do not observe a statistically significant difference between the bottom bracket with the middle bracket.
-When we estimate a linear model, we find a similar relationship that is both positive and statistically significant.
-An increase in the realized quality (Z-score) of an image causes the next prompt to be more similar to the previous one.
-When we estimate the effect at the user-level, we do not find statistically significant effects of the first prompt realized quality on average successive similarity, neither when using the trichotomous variable nor when fitting a linear model.
-    Successive Prompt Token Sort Ratio: The results are identical to the case of successive cosine similarity described above.
-We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt causes an increase in the Token Sort Ratio between successive prompts, using both the trichotomous variable (top vs bottom and middel brackets) and the linear model.
-At the user-level, we do not find statistically significant effects of the Z-score of the first prompt on average token sort ratio of successive prompts.
-    Successive Prompt ‘Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in the likelihood that the subsequent prompt contains the previous one.
-When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with the top bracket more likely to exactly include the previous prompt.
-However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket.
-When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant.
-An increase in the realized quality of an image causes the next prompt more likely to include the previous one.
-Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the average probability of a prompt containing the preceding prompt, when using the trichotomous variable or when fitting a linear model.
+    • Prompt Length: We found statistically significant evidence that higher Z-scores led to longer subsequent prompts.
+When using the trichotomous variable, the difference between the top and bottom Z-score brackets was significant, with longer prompts on average in the top bracket.
+However, we did not observe significant differences between the middle bracket and either the top or bottom brackets.
+Linear models at the user-attempt level showed a positive and statistically significant relationship, with a one-unit increase in Z-score associated with prompts approximately 0.3 words longer on average.
+At the user level, the Z-score of the first prompt did not significantly affect the length of subsequent prompts.
+    • Successive Similarity: Higher Z-scores were associated with increased similarity between consecutive prompts.
+The top Z-score bracket showed significantly higher prompt similarity compared to both the bottom and middle brackets, though no significant difference emerged between the bottom and middle brackets.
+Linear models confirmed a positive and statistically significant relationship between Z-scores and successive prompt similarity.
+No significant effects were found at the user level.
+    • Successive Prompt Token Sort Ratio: Results mirrored those for successive cosine similarity, with higher Z-scores associated with increased token sort ratios between successive prompts.
+This pattern was consistent across both trichotomous variable comparisons and linear models.
+No significant effects emerged at the user level.
+    • Successive Prompt ’Contains Previous Prompt’ Dummy: Higher Z-scores significantly increased the likelihood that subsequent prompts would contain the previous prompt as an exact substring.
+The difference between top and bottom Z-score brackets was significant, though differences involving the middle bracket were not.
+Linear models at the user-attempt level confirmed a positive and statistically significant relationship.
+No significant effects were found at the user level.
+These findings reveal a consistent pattern: when the stochastic nature of generative AI produces unexpectedly high-quality outputs for a given prompt, users tend to build more directly upon that prompt in their subsequent attempts—writing longer prompts that are more similar to and often directly incorporate the previous prompt.
+This adaptive behavior leads to better performance in subsequent attempts.
+However, this effect appears to be local rather than global; the quality of the first prompt affects the immediate next attempt but does not significantly influence overall performance or prompting behavior across all subsequent attempts.
+This pattern suggests that users are sensitive to and learn from random variation in model outputs, even when they have no way of knowing whether a particularly good or bad result stems from the quality of their prompt or from model stochasticity.
+The tendency to build upon prompts that produce better-than-expected results represents an intuitive but potentially suboptimal learning strategy, as it may attribute too much importance to random variations rather than focusing on the underlying quality of the prompt itself.
 H5
 Average prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) and prompting strategies will depend on the capacity of the model that participants are interacting with.
-Analysis approach:
-To test this hypothesis, we conducted three two-sample t-tests that compare the prompting performance and prompting behavior of the three possible pairs of treatment assignments.
-All outcome variables are specified in the results below.
-For robustness, we repeat these analyses with ANOVA to test whether any of the three treatment arms has an effect on the same variables.
-To account for multiple testing, we adjusted the p-values for tests with the Benjamini-Hochberg procedure.
-Results: The results of our analysis are as follows:
+Analysis approach: To test this hypothesis, we conducted pairwise comparisons across the three treatment arms using two-sample t-tests to evaluate differences in both performance outcomes and prompting behaviors.
+For robustness, we also performed ANOVA tests to examine whether any significant differences existed across all three treatment arms simultaneously.
+P-values were adjusted for multiple testing using the Benjamini-Hochberg procedure.
+Results: Our analyses revealed several key differences in both performance and prompting behavior across the different model conditions:
   • CLIP embedding cosine similarity
-    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores).
-When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
-Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
-When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
-    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks).
-When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
-Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
-When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
+    • Z-score rescaling: We found statistically significant evidence for performance differences across treatment arms when examining participants’ first attempts, average performance across all attempts, and best attempts (all measured using taskwide Z-scores rather than task-iteration Z-scores).
+      For first attempt performance, DALL-E 3 (Verbatim) participants significantly outperformed both DALL-E 3 (Revised) and DALL-E 2 participants, while no statistically significant difference emerged between DALL-E 3 (Revised) and DALL-E 2 participants.
+      Average performance across all attempts showed the same pattern: DALL-E 3 (Verbatim) outperformed both other conditions, with no significant difference between DALL-E 3 (Revised) and DALL-E 2.
+      When comparing participants’ best attempts, we found statistically significant differences between all three treatment arms in a clear hierarchy: DALL-E 3 (Verbatim) produced better results than DALL-E 3 (Revised), which in turn produced better results than DALL-E 2.
+ANOVA tests for all three performance measures (first, average, and best attempts) confirmed significant effects of model assignment.
+    • Percentile rank rescaling: Results using percentile rank rescaling mirrored those found with Z-score rescaling.
+DALL-E 3 (Verbatim) participants outperformed both other conditions on first attempt performance and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2 for these metrics.
+For best attempt performance, all pairwise comparisons revealed significant differences, with DALL-E 3 (Verbatim) outperforming DALL-E 3 (Revised), which outperformed DALL-E 2.
+ANOVA tests confirmed significant effects of model assignment for all three performance measures.
   • DreamSim
-    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores).
-When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results.
-When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
-    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks).
-When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms.
-When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised).
-Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2.
-Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2.
-When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
+    • Z-score rescaling: Results using DreamSim with Z-score rescaling showed the same pattern as CLIP embedding cosine similarity.
+DALL-E 3 (Verbatim) participants significantly outperformed both other conditions on first attempt and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2.
+For best attempts, we again found the same hierarchical pattern of performance differences across all three conditions.
+ANOVA tests confirmed significant effects across all performance measures.
+    • Percentile rank rescaling: Results with percentile rank rescaling of DreamSim scores were consistent with the Z-score findings.
+DALL-E 3 (Verbatim) participants outperformed both other conditions on first and average performance, with no differences between DALL-E 3 (Revised) and DALL-E 2.
+For best attempts, all pairwise comparisons were significant, showing DALL-E 3 (Verbatim) ¿ DALL-E 3 (Revised) ¿ DALL-E 2.
+ANOVA tests confirmed significant effects across all performance measures.
   • Prompting Behaviors
-    Mean prompt Length: We found statistically significant differences in prompt length between treatment arms.
-Participants using DALL-E 2 used significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups.
-There was no significant difference in prompt length between the two DALL-E 3 groups.
-ANOVA results confirmed a significant effect of DALL-E version on prompt length.
-    Aggregate Similarity: We found no statistically significant differences between any of the treatment arms.
-This suggests that the overall variability in prompts was similar across all three DALL-E versions.
-ANOVA results confirmed no significant effect of DALL-E version on this measure.
-    Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences between the treatment arms.
-ANOVA results confirmed no significant effect of DALL-E version on successive prompt similarity.
-    Successive Prompt Token Sort Ratio: We found no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
-    Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms.
-ANOVA results confirmed no significant effect of DALL-E version on this measure.
+    • Mean prompt Length: We found statistically significant differences in prompt length across treatment arms.
+Participants using DALL-E 2 wrote significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups.
+Interestingly, there was no significant difference in prompt length between the two DALL-E 3 variants.
+ANOVA results confirmed a significant effect of model version on prompt length.
+    • Aggregate Similarity: We found no statistically significant differences in the overall variability of prompts across the three treatment arms.
+ANOVA results confirmed no significant effect of model version on this measure, suggesting that participants explored similar breadths of prompt space regardless of the model they were using.
+    • Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences across treatment arms.
+ANOVA results confirmed no significant effect of model version on successive prompt similarity, indicating that the tendency to build upon or deviate from previous prompts was similar across models.
+    • Successive Prompt Token Sort Ratio: No statistically significant differences emerged across treatment arms.
+ANOVA results confirmed no significant effect of model version on this measure of textual similarity between successive prompts.
+    • Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences across treatment arms.
+ANOVA results confirmed no significant effect of model version on this measure.
+These findings paint an interesting picture of how model capacity influences both performance and prompting behavior.
+In terms of performance, DALL-E 3 (Verbatim) consistently outperformed the other conditions across all metrics and scaling approaches, demonstrating the clear advantage of the more advanced model when used without automated prompt revision.
+The DALL-E 3 (Revised) condition generally performed worse than DALL-E 3 (Verbatim) but better than DALL-E 2 on best attempt measures, suggesting that automated prompt revision partially degraded the benefits of the more capable model.
+For prompting behavior, the primary difference observed was in prompt length: participants using either version of DALL-E 3 wrote significantly longer prompts than those using DALL-E 2.
+This suggests that users recognized and adapted to the increased capacity of the more advanced model by providing more detailed instructions.
+However, we found no significant differences across models in measures of prompt similarity or evolution, suggesting that while participants provided more detail when using more capable models, their overall strategies for iterating on prompts remained relatively consistent regardless of model assignment.
 H6
 Variability in participants’ ability to prompt engineer effectively and prompting strategies will depend on the capacity of the model that participants are interacting with.
-Analysis approach:
-To test this hypothesis, we conducted two analyses.
-First, would conducted F-tests comparing the variance of participant performance and prompting behaviors between all 3 pairs of models.
-Second, we estimated the quantile treatment effects (QTEs) between all 3 pairs models on participant performance and prompting behaviors.
-We also pre-registered our intent to visually inspect whether the QTEs we observe are consistent with dispersion/“inequality” being reduced or increased when participants use different models (e.g., positive effects for low quantiles and negative/null effects for high quantiles would be consistent with inequality reduction).
+Analysis approach: To test this hypothesis, we conducted two types of analyses.
+First, we performed F-tests comparing the variance of participant performance and prompting behaviors between all three pairs of model conditions (DALL-E 2 vs.
+DALL-E 3 Revised, DALL-E 3 Revised vs.
+DALL-E 3 Verbatim, and DALL-E 2 vs.
+DALL-E 3 Verbatim).
+Second, we estimated quantile treatment effects (QTEs) between all three pairs of models on participant performance and prompting behaviors.
+We also visually inspected the QTE patterns to determine whether dispersion/”inequality” was being reduced or increased when participants used different models.
+For example, positive effects for low quantiles and negative/null effects for high quantiles would indicate inequality reduction.
 F-test Results (with BH Adjusted p-values):
   • DALL-E 3 (revised) vs. DALL-E 2
-    Mean prompt length (words): DALL-E 3 Revised has significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
-    Prompt embedding variance: DALL-E 3 Revised also has significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
-    Cosine similarity with target image, Z-Score: DALL-E 3 Revised has significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
+    • Mean prompt length (words): DALL-E 3 Revised showed significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
+    • Prompt embedding variance: DALL-E 3 Revised also showed significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
+    • Cosine similarity with target image, Z-Score: DALL-E 3 Revised showed significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
     • Failed to reject null of no differences in variance: Mean raw DreamSim score vs.
@@
   • DALL-E 3 (verbatim) vs. DALL-E 3 (revised)
-    Mean percentile rank of CosineSim: DALL-E 3 (verbatim) has significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
-    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7352, $p=0.0009$).
-    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7888, $p=0.0141$).
-    Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7961, $p=0.0159$).
-    DreamSim with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8195, $p=0.0377$).
-    Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) has significantly more variance (ratio 1.2303, $p=0.0301$).
+    • Mean percentile rank of CosineSim: DALL-E 3 (verbatim) showed significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
+    • Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also showed significantly less variance (ratio 0.7352, $p=0.0009$).
+    • Mean percentile rank of DreamSim: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7888, $p=0.0141$).
+    • Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7961, $p=0.0159$).
+    • DreamSim with target image, Z-Score: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.8195, $p=0.0377$).
+    • Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) showed significantly more variance (ratio 1.2303, $p=0.0301$).
     • Failed to reject null of no differences in variance: Mean raw DreamSim score vs.
@@
   • DALL-E 2 vs. DALL-E 3 (verbatim)
-    Mean prompt length (words): DALL-E 3 (verbatim) has significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
-    Prompt embedding variance: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
-    Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7501, $p=0.0015$).
-    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7904, $p=0.0141$).
-    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8005, $p=0.0159$).
+    • Mean prompt length (words): DALL-E 3 (verbatim) showed significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
+    • Prompt embedding variance: DALL-E 3 (verbatim) also showed significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
+    • Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7501, $p=0.0015$).
+    • Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7904, $p=0.0141$).
+    • Mean percentile rank of DreamSim: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.8005, $p=0.0159$).
     • Failed to reject null of no differences in variance: Mean percentile rank of CosineSim (p = 0.1842), Z-score of DreamSim score (p = 0.3294), Mean Raw CosineSim (p = 0.8151), Z-score of CosineSim score with target image (p = 0.9906), mean token sort ratio with previous prompt (p = 0.8748), mean cosine similarity to previous prompt (p = 0.807), number of topical transitions (token sort ratio) (p = 0.3505), and number of topical transitions (cosine similarity) (p = 0.085).
-QTE Highlighted Results: Through our QTE plots, we see clearest evidence of dispersion being reduced for:
-  • Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALLE-3 Revised vs. DALLE-2 comparison.
-And we see the clearest evidence of dispersion increasing for:
-  • Mean prompt length for all three pairwise model comparisons (DALLE-3 Verbatim vs. DALLE-2, DALLE-3 Revised vs. DALLE-2, and DALLE-3 Verbatim vs. DALLE-3 Revised)
-  • Raw Dreamsim performance for DALLE-3 Verbatim vs. DALLE-2 and DALLE-3 Revised vs. DALLE-2 comparisons.
+QTE Highlighted Results: Through our quantile treatment effects analysis and visual inspection of QTE plots, we found:
+  • Evidence of dispersion reduction: Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALL-E 3 Revised vs.
+DALL-E 2 comparison showed patterns consistent with inequality reduction, with larger positive effects at lower quantiles.
+  • Evidence of dispersion increase: Mean prompt length for all three pairwise model comparisons (DALL-E 3 Verbatim vs.
+DALL-E 2, DALL-E 3 Revised vs.
+DALL-E 2, and DALL-E 3 Verbatim vs.
+DALL-E 3 Revised) showed patterns consistent with increased inequality, with larger positive effects at higher quantiles.
+Similar patterns emerged for raw DreamSim performance in both DALL-E 3 Verbatim vs.
+DALL-E 2 and DALL-E 3 Revised vs.
+DALL-E 2 comparisons.
+These findings present a nuanced picture of how model capacity affects variability in performance and prompting behaviors.
+For prompt length and embedding variance, the more advanced models (both DALL-E 3 variants) showed significantly less variance than DALL-E 2, suggesting more consistent prompting behaviors.
+For performance measures, the results were mixed: DALL-E 3 (Verbatim) generally showed less variance than DALL-E 3 (Revised), indicating more consistent performance across participants, while DALL-E 3 (Revised) showed higher variance in performance than DALL-E 2.
+The QTE analysis suggests that the relationship between model capacity and performance inequality is complex.
+For some measures (e.g., Z-score of cosine similarity), more advanced models appeared to reduce inequality by disproportionately benefiting lower-performing users.
+For other measures (e.g., prompt length and raw DreamSim scores), more advanced models appeared to increase inequality by disproportionately benefiting higher-performing users.
 H7
 As participants repeatedly try to complete a task with a given model, the quality of their attempts will increase, and the extent to which the quality increases varies as a function of model capacity.
-Analysis approach:
-In order to test for differences in prompting performance improvement across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata.
-The two samples represent the best and initial scores.
-We make these comparisons overall and within strata.
+Analysis approach: To evaluate how performance improves across successive attempts and whether this improvement varies by model, we conducted stratified two-sample tests comparing participants’ initial scores with their best scores.
+These comparisons were performed both within each treatment arm (DALL-E 2, DALL-E 3 Verbatim, and DALL-E 3 Revised) and overall across all participants.
+This approach allowed us to determine both whether participants generally improved with practice and whether the rate of improvement differed across models with varying capabilities.
+Results: Our analyses revealed consistent performance improvements across all treatment arms and outcome measures:
   • CLIP embedding cosine similarity
-    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
-    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
-    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
+    • No rescaling: We found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
+This indicates that participants were able to improve their performance through iterative attempts regardless of which model they were using.
+    • Z-score rescaling: Results mirrored those of the unscaled analysis, with statistically significant improvements from initial to best performance observed both within each treatment arm and overall.
+    • Percentile rank rescaling: Consistent with other scaling approaches, we observed statistically significant improvements from initial to best performance both within each treatment arm and overall.
   • DreamSim
-    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
-    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
-    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
+    • No rescaling: DreamSim results aligned with CLIP similarity findings, showing statistically significant improvements from initial to best performance both within each treatment arm and overall.
+    • Z-score rescaling: We again found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
+    • Percentile rank rescaling: Results were consistent with other analyses, showing statistically significant improvements from initial to best performance across all conditions.
+These findings demonstrate that participants consistently improved their performance through iterative attempts across all model conditions and performance measures.
+Users were able to learn from feedback and refine their prompts to achieve better results over time, regardless of which model they were using.
+This learning effect was robust across different outcome measures and scaling approaches, highlighting the importance of iteration and feedback in developing effective prompting strategies.
+Although our original hypothesis also posited that the rate of improvement might vary as a function of model capacity, we did not find consistent evidence for differences in improvement rates across models.
+This suggests that while more capable models yield better absolute performance (as demonstrated in H5), the relative improvement from first to best attempt was similar across model conditions.
+This finding highlights that the benefits of iterative refinement apply broadly across different model capabilities.
 H8
 The extent to which participants can recreate images using models such as DALL-E 2/3 will vary across images.
-Analysis approach:
-To evaluate this hypothesis, we pre-registered and performed two analyses.
-First, we compared the extent to which different images can be replicated using GPT-4V—a multimodal generative AI model that can take in both text and images and output text.
-For each of the 15 target images, we prompted GPT-4V to “Write a DALL-E {2, 3} prompt to recreate this image verbatim as closely and as detailed as possible.” We do this once for DALL-E 2 and once for DALL-E 3, generating two “AI prompts” per target image.
-We then provided these AI prompts to DALL-E 2, DALL-E 3 (Verbatim), and DALL-E 3 (Revised), respectively, 20 times each, generating 60 replicated images per target image (the DALL-E 3 AI prompt was sent to both DALL-E 3 (Verbatim) and (revised)).
-For each image, we measure the cosine similarity of the CLIP embedding vectors to those of the relevant target image.
-We then average these 60 similarity measures by target image, resulting in a mean similarity score that represents GPT-4V’s ability to generate prompts that recreate each target image.
-Second, we simply measured the similarity of every participant-generated image to that of the relevant target image, and averaged over these participant-generated similarities.
+Analysis approach: To evaluate this hypothesis, we employed two complementary approaches.
+First, we used GPT-4V (a multimodal generative AI model capable of processing both text and images) to generate optimal prompts for each target image.
+For each of our 15 target images, we instructed GPT-4V to ”Write a DALL-E 2, 3 prompt to recreate this image verbatim as closely and as detailed as possible.” We generated two such ”AI prompts” per target image—one optimized for DALL-E 2 and one for DALL-E 3.
+We then submitted these AI-generated prompts to the respective models 20 times each, yielding 60 replicated images per target (as the DALL-E 3 prompts were sent to both DALL-E 3 variants).
+We measured the cosine similarity between CLIP embeddings of these generated images and their target images, then averaged these similarity scores to quantify GPT-4V’s ability to generate effective replication prompts for each target.
+Second, we analyzed human performance by measuring the similarity between all participant-generated images and their corresponding target images, averaging these similarities to determine which images were easier or harder for participants to replicate.
+This two-pronged approach allowed us to assess image difficulty from both AI and human perspectives.
 Results:
-  • CLIP embedding cosine similarity: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
+  • CLIP embedding cosine similarity: When ranking target images by GPT-4V’s ability to generate effective prompts, we found substantial variation across images:
     1. 1.
@@
       Photography Image #4
-    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for GPT-4V to replicate) was $CoSim=0.944$ for Business Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $CoSim=0.734$ for Photography Image #4.
-    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
+    The highest average cosine similarity (easiest image for GPT-4V to replicate) was $CoSim=0.944$ for Business Image #3, while the lowest score (hardest image to replicate) was $CoSim=0.734$ for Photography Image #4.
+    When ranking target images by participants’ ability to generate effective prompts, we observed:
     1. 1.
@@
       Business Image #4
-    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, and the lowest score (the hardest image for participants to replicate was $CoSim=0.669$ for Business Image #4.
-  • DreamSim: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
+    The highest average cosine similarity (easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, while the lowest score (hardest image to replicate) was $CoSim=0.669$ for Business Image #4.
+  • DreamSim: When using DreamSim as our similarity metric (inverted as 1-DreamSim to create a similarity rather than distance measure), GPT-4V’s prompt generation performance ranked as follows:
     1. 1.
@@
       Photography Image #4
-    The maximum 1-DreamSim score in the above list (the easiest image for GPT-4V to replicate) was $\tilde{D}=0.75$ for Design Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $\tilde{D}=0.40$ for Photography Image #4.
-    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
+    The highest 1-DreamSim score (easiest image for GPT-4V to replicate) was $\tilde{D}=0.75$ for Design Image #3, while the lowest score (hardest image to replicate) was $\tilde{D}=0.40$ for Photography Image #4.
+    When ranking target images by participants’ replication ability using DreamSim:
     1. 1.
@@
       Photography Image #3
-    The maximum 1-DreamSim score in the above list (the easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, and the lowest score (the hardest image for participants to replicate was $\tilde{D}=0.356$ for Photography Image #4.
+    The highest 1-DreamSim score (easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, while the lowest score (hardest image to replicate) was $\tilde{D}=0.356$ for Photography Image #4.
+These results clearly demonstrate substantial variation in image replicability across our target set.
+Business-related images were generally easier to replicate, while photography images tended to be more challenging.
+While the exact rankings varied somewhat between GPT-4V and human participants, and between similarity metrics, the overall pattern of relative difficulty remained fairly consistent.
+This variation in difficulty confirms the importance of our stratified analysis approach and underscores the need to consider image-specific characteristics when evaluating generative AI performance.
 Exploratory Analyses
-In our pre-registration, we also mention a number of exploratory analyses that are not described with the same level of detail.
-Multiple of these exploratory analyses appear in our main text.
-These pre-registered exploratory analyses are copied verbatim below:
-“We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate.
+Our pre-registration outlined several exploratory analyses that we planned to conduct beyond our primary hypotheses.
+Several of these exploratory analyses appear in our main text, particularly the replay analysis that decomposed performance improvements into model and prompting effects.
+The pre-registered exploratory analyses were described as follows:
+”We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate.
 We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
@@
 We also may take original and revised prompts submitted to DALL-E 3 treatment arms and submit them to DALL-E 2 (and vice versa) to see how participants would have counterfactually performed under different treatment assignments than the one to which they were assigned.”
-F.2 Deviations From Pre-registration
-Here, we also report any deviations from our pre-registered analysis.
-By and large, these deviations occurred because certain aspects of our pre-registration were not appropriate from a statistical analysis perspective, or were infeasible.
-  • In our pre-registration, we had anticipated using t-tests and Mann-Whitney U tests for many hypotheses.
-However, this turned out to be impossible for many variables since most of the demographic traits have multiple categories.
-Thus, we applied ANOVA and Kruskal-Wallis tests instead.
-  • In our pre-registration, we were not explicit enough on how to compute Z-score measures of performance.
-As described in section C.2, Z-scores are almost always computed within image-attempt to adjust for variations across target images or attempts.
-The only exception is with any analysis that compares performance across attempts where Z-scores need to computed within target images.
-The pre-registration anticipated that Z-scores for any analysis would be computed only within target images.
-  • In our pre-registration, we had forgotten to include Education and GenAI outlook in our list of demographics with respect to which we would measure task performance heterogeneity.
-We decided to include these important variables in our analysis, despite the accidental omission.
-  • In addition to our pre-registered prompt exclusion criteria, we removed additional prompts that did not appear to be “good-faith efforts” to complete our task based on the text of those prompts (see Section B.4 for more details).
-We find that our results are the robustness to the inclusion of prompts that were removed by this procedure.
-  • Because they were easier to implement, we ran t-tests rather than z-tests to conduct our tests of H5.
-Because the two tests are asymptotically equivalent, we do not believe this will make a difference in our analysis.
-  • When conducting our analyses to test H4, we observed that the distribution of the Z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values.
-To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute Z-scores greater than 3, which constituted 2.65% of the user-attempt observations.
-For robustness, we also conducted the stratification analysis including these outliers and found almost identical results, with the exception that the difference in performance between the top bracket and middle or bottom brackets was in some cases only marginally significant.
-  • In our pre-registration, we had included a rescaled version of our main performance metrics described in B.5.1 based on GPT-4V for robustness check.
-This outcome variable would have been rescaled by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image.
-Given that the quality of the prompt from GPT-4V model is a constant for each target image, this would have amounted to subtracting a constant from the unscaled CLIP cosine similarity or dream sim scores.
-Furthermore, since all our analysis involving performance as the dependent variable adjust for the target image, either through post-stratification or as a covariate in a linear model, the results from using GPT-4V scaling would have been identical to the the unscaled measures of performance.
-For this reason, we have not included the robustness checks involving GPT-4V for our pre-registered hypothesis.
+As noted in the main text, the third exploratory analysis—submitting prompts from one model condition to another model—became central to our investigation of model versus prompting effects.
+This approach allowed us to isolate how much improvement came from the model’s enhanced capabilities versus users adapting their prompting strategies to take advantage of those capabilities.
+G.2 Deviations From Pre-registration
+We report below all deviations from our pre-registered analysis plan.
+These deviations primarily resulted from statistical or methodological considerations that became apparent during data analysis, rather than from substantive changes to our research questions or hypotheses.
+  • Statistical tests: Our pre-registration specified t-tests and Mann-Whitney U tests for many hypotheses.
+However, this approach proved inappropriate for variables with multiple categories (which characterized most demographic traits).
+We therefore employed ANOVA and Kruskal-Wallis tests instead, which provide equivalent information for multi-category variables.
+  • Z-score computation: Our pre-registration was insufficiently precise about computing Z-score performance measures.
+As detailed in Section D.2, we calculated Z-scores within image-attempt pairs to adjust for variations across target images and attempts, except when comparing performance across attempts, where Z-scores were computed within target images only.
+The pre-registration had anticipated computing Z-scores only within target images for all analyses.
+  • Additional demographic variables: We inadvertently omitted Education and Generative AI outlook from our pre-registered list of demographics for measuring task performance heterogeneity.
+Given their theoretical importance, we included these variables in our analyses despite this accidental omission.
+  • Additional prompt exclusion criteria: Beyond our pre-registered exclusion criteria, we removed additional prompts that did not appear to constitute ”good-faith efforts” based on their text content (see Section C.4 for details).
+Robustness checks confirmed that our results remain consistent when including these prompts.
+  • T-tests vs.
+Z-tests: For testing H5, we used t-tests rather than z-tests because they were easier to implement.
+As these tests are asymptotically equivalent, this change does not meaningfully affect our results.
+  • Z-score outlier handling: When testing H4, we observed that Z-scores did not follow a normal distribution and included extreme values (ranging from -21.9 to 7.1).
+To prevent these outliers from disproportionately influencing our linear models, we excluded observations with absolute Z-scores greater than 3 (2.65
+  • GPT-4V rescaling omission: Our pre-registration included a rescaled version of performance metrics based on GPT-4V prompt quality.
+However, this rescaling would have amounted to subtracting a constant from each unscaled score, and since all our analyses adjusted for target image (either through post-stratification or as a covariate), results would have been identical to those using unscaled measures.
+We therefore omitted these redundant robustness checks.
+G.3 OSF Pre-registration
+A complete copy of the pre-registration deposited on OSF on December 11, 2023 will be included beginning on the next page in the final manuscript, but has been removed here to maintain anonymity.

### v6

@@
 Abstract
-As generative AI systems rapidly improve, a key question emerges: How do users keep up—and what happens if they fail to do so.
-Drawing on theories of dynamic capabilities and IT complements, we examine prompt adaptation—the adjustments users make to their inputs in response to evolving model behavior—as a mechanism that helps determine whether technical advances translate into realized economic value.
-In a preregistered online experiment with 1,893 participants, who submitted over 18,000 prompts and generated more than 300,000 images, users attempted to replicate a target image in 10 tries using one of three randomly assigned models: DALL-E 2, DALL-E 3, or DALL-E 3 with automated prompt rewriting.
-We find that users with access to DALL-E 3 achieved higher image similarity than those with DALL-E 2—but only about half of this gain (51%) came from the model itself.
-The other half (49%) resulted from users adapting their prompts in response to the model’s capabilities.
-This adaptation emerged across the skill distribution, was driven by trial-and-error, and could not be replicated by automated prompt rewriting, which erased 58% of the performance improvement associated with DALL-E 3.
-Our findings position prompt adaptation as a dynamic complement to generative AI—and suggest that without it, a substantial share of the economic value created when models advance may go unrealized.
+As generative AI systems rapidly improve, a key question emerges: how do users adapt to these changes, and when does such adaptation matter for realizing performance gains?
+Drawing on theories of dynamic capabilities and IT complements, we study prompt adaptation—how users adjust their inputs in response to evolving model behavior—using a common experimental design applied to two preregistered tasks with 3,750 total participants who submitted nearly 37,000 prompts.
+We show that the importance of prompt adaptation depends critically on task structure.
+In a task with fixed evaluation criteria and an unambiguous goal, user prompt adaptation accounts for roughly half of the performance gains from a model upgrade.
+In contrast, in an open-ended creative task where the space of acceptable outputs is effectively unbounded and quality is subjective, performance improvements are driven primarily by model capability; prompt adaptation plays a limited role.
+We further show that automated prompt rewriting cannot generally substitute for human adaptation: when aligned with task objectives, it can modestly improve performance, but when misaligned, it can actively undermine the gains from model improvements.
+Together, these findings position prompt adaptation as a dynamic complement whose importance depends on task structure and system design, and suggest that without it, a substantial share of the economic value created by advances in generative models may go unrealized.
 1 Introduction
-Generative AI is being integrated into work practices across the economy (Zhang and Kamel Boulos 2023, Bright et al.
-2024), yielding notable productivity gains in tasks as diverse as software development, writing, and material discovery (Brynjolfsson et al.
-2023, Dell’Acqua et al.
-2023, Noy and Zhang 2023, Peng et al.
-2023, Yu 2024, Toner-Rodgers 2024).
-Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning et al.
-2024), including tasks as complex as chemical research and proving mathematical theorems (Romera-Paredes et al.
-2024, Boiko et al.
-2023).
+Generative AI is being integrated into work practices across the economy (zhang2023generative, bright2024generative), yielding notable productivity gains in tasks as diverse as software development, writing, and scientific research (brynjolfsson2023generative, dell2023navigating, Noy2023, peng2023impact, Yu2024AIImpact).
+Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning2024Automated), including tasks as complex as chemical research and proving mathematical theorems (RomeraParedes2024Math, Boiko2023AutonomousResearch).
 The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S.
-workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (Bright et al.
-2024, Bick et al.
-2024).
+workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (bright2024generative, bick2024rapid).
 As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results.
-To interact with generative AI systems, users provide written instructions—often referred to as prompts—that guide the model’s behavior.
+To interact with generative AI systems, users provide written instructions—or prompts—that guide the model’s behavior.
 These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software).
-In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (Brynjolfsson and Hitt 2000).
+In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (brynjolfsson2000beyond).
 Prompting has quickly become an area of active research and practice.
-Scholars have developed taxonomies of prompt engineering techniques (Oppenlaender 2023), documented recurring patterns in prompt construction (Schulhoff et al.
-2024), and examined how developers embed prompts into software systems (Liang et al.
-2024).
-Other studies have explored prompting strategies for specific applications, including image generation (Don-Yehiya et al.
-2023, Xie et al.
-2023) and clinical documentation (Yao et al.
-2024).
+Scholars have developed taxonomies of prompt engineering techniques (oppenlaender2023), documented recurring patterns in prompt construction (schulhoff2024prompt), and examined how developers embed prompts into software systems (liang2024prompts).
+Other studies have explored prompting strategies for specific applications, including image generation (donyehiya2023, xie2023) and clinical documentation (yao2024clinicians).
 In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design.
@@
 Many prompt libraries and tutorials present effective prompts as reusable artifacts.
-But prompts that work well with one model version may underperform or break entirely with the next (Liang et al.
-2024, Meincke et al.
-2025).
+But prompts that work well with one model version may underperform or break entirely with the next (liang2024prompts, meincke2025prompting).
 While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance.
 This raises a broader question for individuals and organizations investing in prompting capabilities: Are prompt strategies transferable across model versions, or must they be continually revised to match changing model behavior?
-To begin exploring this question, we identify prompt adaptation [^1] We use the phrase “prompt adaptation” to refer specifically to changes in user prompting behavior that arise in response to evolving model capabilities—distinct from the broader practice of prompt engineering, which includes static best practices, libraries, and templates.
+To begin exploring this question, we identify prompt adaptation [^1] We use the phrase “prompt adaptation” in reference to changes in user prompting behavior in response to evolving model capabilities.
+This is distinct from the broader practice of prompt engineering, which includes static best practices, libraries, and templates, among other things.
 as a measurable behavioral mechanism through which user-side inputs evolve alongside technical advances.
@@
 In contrast to static complements (e.g., fixed training, prompt templates), dynamic complements emerge through situated use with rapid feedback, respond to model-level change, and may be enabled or suppressed by system design.
-To assess the role of prompt adaptation in shaping realized performance—and to separate its contribution from the direct effects of model improvement—we draw on data from a pre-registered online experiment with $1,893{}$ participants.
-In the study, participants were asked to replicate a target image using prompts submitted to one of three randomly assigned text-to-image models of varying capability: DALL-E 2, DALL-E 3, or DALL-E 3 with an automated large language model (LLM)-based prompt revision.
-Each participant submitted at least 10 prompts in an effort to reproduce the image as closely as possible, with a large monetary bonus for top performers.
-By comparing outcomes across arms—and by conducting a posthoc analysis that re-evaluates prompts on alternative models—we estimate the extent to which users adapted their prompts in response to model improvements and how much this adaptation contributed to overall performance.
-We find that participants assigned to DALL-E 3 produce significantly more faithful replications than those assigned to DALL-E 2.
-Importantly, about half of this improvement comes from participants adapting their prompts to exploit the new model’s capabilities—replaying the old DALL-E 2 prompts on DALL-E 3 yields only about half the total improvement.
-Furthermore, we find that this prompt adaptation is not limited to advanced “prompt engineers”—participants across the full skill distribution benefit from refining their prompts after seeing how the advanced model responds.
-Finally, we show that an attempt to automate prompt revision via GPT-4 rewriting does not match the performance achieved by manual human adaptation; in fact, it substantially erodes the gains from DALL-E 3.
-Together, these findings position prompt adaptation as a key mechanism through which users realize the value of rapidly advancing generative AI systems—and as a concrete example of how digital complements must evolve to keep pace with technological change.
-In terms of related literature and additional theory, this experiment builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies.
-Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (Bharadwaj 2000, Joshi et al.
-2010, Teece et al.
-1997).
-Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (Jasperson et al.
-2005).
-Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (Fügener et al.
-2022).
-And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (Böhm and Schedlberger 2023).
-We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (Brynjolfsson 1993, Brynjolfsson and Hitt 2000, Brynjolfsson et al.
-2021, David 1990).
-We conceptualize prompt adaptation as one such complement—emerging through trial-and-error, accessible across the skill distribution, and potentially hindered when automation misaligns with user intent.
-In this way, prompt adaptation shapes how and when technical improvements translate into downstream economic value.
-Finally, the contribution of the current paper is four-fold.
-First, we offer direct causal evidence that prompting is not a fixed input but a dynamic capability that co-evolves with model behavior (Böhm and Schedlberger 2023).
-Second, we show how even non-expert users adapt their prompting in response to model improvements, extending research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al.
-2010, Teece et al.
-1997).
-Third, we document how performance gains emerge through user-initiated trial-and-error rather than static best practices, advancing work on post-adoptive IT use (Jasperson et al.
-2005).
-Finally, we demonstrate that automation intended to reduce user effort can—if misaligned with user intent—undermine the very adaptations that generate value, contributing to ongoing research on human–AI interaction (Fügener et al.
-2022, Yao et al.
-2024).
+Generative AI systems support a wide variety of task structures.
+Some tasks are steering-oriented, where users aim to reach a clearly defined goal and success, can be measured against an objective benchmark (e.g., write code to solve a specific problem).
+Others are creative, where goals are open-ended, evaluation is subjective, and improvement depends more on exploration and curation than on precision (e.g., make a beautiful picture).
+Recent work helps to establish these regimes: vafa2025s investigate the steerability of models, i.e., how effectively users can reach a specified target, whereas orwig2024language and zhou2024generative describe generative AI–assisted creation as a co-creative process of idea generation, filtering, and aesthetic evaluation.
+In organizational settings, steering tasks (or bounded tasks, herein) dominate operational workflows such as classification, brand-compliant content generation, or reproducing structured materials where quality can be defined ex ante.
+Creative tasks (or unbounded tasks, herein), by contrast, underpin innovation activities such as logo or campaign design, product ideation, and concept development, where success depends on novelty, appeal, and fit to evolving goals.
+Understanding how prompting and prompt adaptation operate across these two regimes is essential for organizations deploying generative AI at scale.
+To assess the role of prompt adaptation in shaping outcomes across different types of tasks—and to separate its contribution from the direct effects of model improvement—we conducted two pre-registered online experiments with an aggregate sample of 3,750 participants.
+The first, a bounded task, asks participants to iteratively prompt one of three randomly assigned text-to-image models—DALL-E 2, DALL-E 3, or DALL-E 3 with automated large-language-model–based prompt rewriting—to reproduce a reference image as accurately as possible.
+It is explicitly designed to capture steering or target-matching behavior.
+Each participant submitted at least ten prompts and was eligible for a substantial performance-based bonus, incentivizing careful refinement.
+The second, an unbounded task, captures creative or open-ended behavior: participants design a logo for a hypothetical organization based on a short textual vignette without a specified target image or objective benchmark.
+Apart from this difference in goal structure, the two tasks, which were conducted on the same custom interface, were identical from the perspective of participants.
+By comparing outcomes across treatment arms, conducting post-hoc analyses that re-evaluate prompts on alternative models, and contrasting results across the two experiments, we estimate how users adapted their prompts in response to model improvements and how these adaptations contributed to overall performance across both bounded and unbounded settings.
+Across both tasks, we find that participants assigned to DALL-E 3 produced significantly better outputs than those assigned to DALL-E 2, but that the sources of these gains depend on task structure.
+In the bounded image replication task, about half of the gains came from participants adapting their prompts to exploit the new model’s capabilities—replaying DALL-E 2 prompts on DALL-E 3 yields only about half the total improvement.
+By contrast, in the open-ended logo generation task, performance gains were driven primarily by improvements in model capability, with prompt adaptation accounting for a much smaller share of the overall effect (roughly 7%, compared to over 90% attributable to the model).
+Importantly, the prompt adaptation we observe does not appear to be limited to advanced “prompt engineers”: benefits from prompt refinement after interacting with more capable models are observable across the outcome distribution.
+Finally, to assess whether automation can substitute for user-side prompt adaptation, we examine automated prompt revision via GPT-4 rewriting.
+We find that such automation can either erode or modestly enhance performance depending on alignment with the end user’s goals: automated rewriting substantially reduced gains in the replication task, but slightly improved performance in the creative setting.
+Together, these findings position prompt adaptation as a dynamic complement whose importance depends on task structure and system design, shaping how advances in generative AI translate into realized value.
+In terms of related literature and additional theory, our research builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies.
+Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (bharadwaj2000resource, joshi2010changing, teece1997dynamic).
+Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (jasperson2005comprehensive).
+Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (fugener2022cognitive).
+And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (bohm2023co-evolution).
+We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (brynjolfsson1993productivity, brynjolfsson2000beyond, brynjolfsson2021productivity, david1990dynamo).
+We conceptualize prompt adaptation as a dynamic complement that co-evolves with model capability, emerges through use rather than formal training, and whose importance depends on task structure and system design.
+In this way, prompt adaptation shapes how—and under what conditions—technical improvements translate into downstream economic value.
+This paper makes three core contributions.
+First, we conceptualize prompt adaptation as a dynamic complement to improvements in generative AI models; one through which users actively shape how technical advances translate into realized performance.
+Leveraging a replay-based analysis enabled by our experimental design, we provide direct empirical evidence for this distinction by separating performance gains attributable to model capability from those arising through user-side adaptation, and show that prompt adaptation can account for a substantial share of realized performance improvements.
+Second, we identify task structure as a key boundary condition for the importance of prompt adaptation: it is a first-order driver of gains in bounded, steerable tasks with fixed objectives, but plays a much more limited role in open-ended creative tasks where quality is unbounded in nature.
+Finally, we show that automation intended to simplify prompting is not a neutral substitute for user adaptation.
+We find that automated prompt rewriting can either undermine or complement performance, depending on its alignment with user goals.
+Together, these contributions clarify when and how user adaptation amplifies model improvements, positioning prompt adaptation as a dynamic complement whose economic importance depends on task structure and system design.
 The remainder of the paper is organized as follows.
-We begin by presenting a simple conceptual framework that characterizes how output quality evolves with improvements in model capacity and with users’ corresponding adjustments in effort.
-We then introduce an experimental design that closely mirrors this framework and describes the data used in our analysis.
-Next, we present our empirical findings, including a decomposition of the overall effect into components attributable to model improvements versus prompt adaptation, the impact of automated prompt revision, and heterogeneity across user skill levels.
-Finally, we compare these empirical results to the model’s predictions and conclude by discussing implications for organizations adopting generative AI.
+We begin by presenting a simple conceptual framework that characterizes how output quality evolves with improvements in model capacity and with users’ corresponding adjustments in prompting effort.
+We then describe our experimental design and the two task settings we study—a bounded, steerable image replication task and an open-ended logo generation task—along with the data and techniques used in our analyses.
+Next, we present our empirical findings, including a decomposition of the overall effect into components attributable to model improvements versus prompt adaptation, distributional effects across outcome quantiles, and the impact of automated prompt revision.
+We conclude by synthesizing these results and discussing their implications for organizations adopting generative AI.
 2 Conceptual Framework
-As generative AI systems evolve, a user’s ability to adapt prompts for improved models can become an important source of realized performance gains.
-We develop a stylized conceptual framework to formalize how overall output quality depends on both the model’s capacity and the user’s skill and effort in prompt writing.
-Our goal is not to create a fully normative model.
-Rather, we wish to clarify the distinction between improvements directly attributable to the model itself versus those arising from user adaptation as motivation for our experimental design.
-Additionally, the framework naturally implies predictions about the returns to prompting skill and the heterogeneity of such skill.
-Although our empirical setting focuses on image replication, this framework generalizes to other tasks where users interact with generative models—such as text generation, code assistance, or molecule design.
-In each case, we expect performance to reflect both model capacity and users’ efforts to refine their inputs.
-We present the core findings here; a complete exposition of the framework with formal proofs is provided in Appendix E.1.
+We first develop a stylized analytical framework to understand how overall output quality depends jointly on model capacity and on users’ prompting behavior.
+Our goal is not to propose a fully normative model of user–AI interaction, but rather to distinguish improvements directly attributable to the model itself from those arising through user-side prompt adaptation.
+This distinction motivates our experimental design and yields testable predictions about how performance gains and their distributional impacts change as models improve.
+Although our empirical analysis focuses on two tasks—image replication and logo generation—the same conceptual structure applies broadly to settings in which users interact with generative models, including text generation, code assistance, and scientific research.
+In all such environments, output quality reflects a combination of model capability, user skill, and prompting effort, and total performance improvements can be decomposed into direct (model-driven) and behavioral (prompting-driven) components.
+The two tasks we study correspond to different regimes of this framework: image replication is a bounded task with clear ceiling (perfectly replicating the image pixel-for-pixel), whereas logo generation is a more open-ended creative task without an optimal outcome.
+We develop the core logic for the bounded case in the main text and provide complete derivations for both bounded and unbounded formulations in Appendix A.
+Nearly all of the main results carry over to the unbounded case, with minor differences arising from curvature differences across specifications.
+[^2] We note that both the bounded exponential and the unbounded logarithmic forms are illustrative rather than general, and are not intended to be literal representations of the specific tasks we study.
+They adopt simple structures that allow us to decompose total performance gains into a “model effect” (improvements in capacity holding prompts fixed) and a “prompting effect” (improvements from user adaptation holding capacity fixed).
+This decomposition is a simplifying abstraction that isolates these mechanisms analytically, though in practice they may interact.
+The comparative statics of $Q(\theta,s,x)$ also depend on its curvature and on how model capacity and skill interact (e.g., multiplicatively versus additively).
+We use these tractable forms to convey intuition, not to claim functional-form invariance.
 2.1 Notation and Problem Setting
-Let $\theta\in(0,1]$ represent the model’s capacity to translate prompts into high-fidelity outputs (for instance, how accurately it captures requested details).
-Let $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and let $x\geq 0$ be the effort the user expends on writing and refining prompts.
-Each unit of $x$ incurs a cost $k\,x$ for some $k>0$, which may reflect the time or cognitive load.
-[…]
-so that users cannot exceed perfect fidelity (quality of 1), and each additional unit of effort yields diminishing returns. The user’s utility is then
+Let $\theta\in(0,1]$ denote the model’s capacity to translate prompts into high-fidelity outputs, $s\in(0,1]$ denote a user’s baseline skill in prompt engineering, and $x\geq 0$ denote the effort the user expends on writing and refining prompts.
+Each unit of effort incurs a linear cost $c(x)=kx$ for some $k>0$, reflecting the total cognitive and temporal costs of refining one’s prompts.
+We assume $k$ is small ($k\!\approx\!0$), consistent with the low marginal cost of iterative prompting and using generative AI models more generally (shahidi2025coasean).
+We begin with a bounded quality function,
+$Q(\theta,s,x)=1-e^{-\theta sx},$
+which is increasing in model capacity, user skill, and prompting effort, but exhibits diminishing marginal returns and has a natural upper bound on performance. We define the user’s utility as
 $U(\theta,s,x)\;=\;Q(\theta,s,x)\;-\;k\,x.$
-The user chooses their effort such that it maximizes the utility. We assume $\theta s>k$, which implies there is a unique interior optimum $x^{*}(\theta,s)$
+The user chooses their effort to maximize the utility. We assume $\theta s>k$—cost is not prohibitive to effort—which implies there is a unique interior optimum $x^{*}(\theta,s)$:
 $x^{*}(\theta,s)\;=\;\frac{1}{\theta s}\,\ln\Bigl(\frac{\theta s}{k}\Bigr)>0,$ (1)
@@
 $Q^{*}\;=\;Q\big(\theta,s,x^{*}(\theta,s)\big)=1-\frac{k}{\theta s}>0.$ (2)
-Two implications immediately follow from equation 2. Optimal quality is increasing with both model capacity
+It follows immediately from Equation 2 that optimal quality is increasing with both model capacity
 $\frac{\partial Q^{*}}{\partial\theta}=\frac{k}{\theta^{2}s}>0$ (3)
@@
 $\frac{\partial Q^{*}}{\partial s}=\frac{k}{\theta s^{2}}>0.$ (4)
-Furthermore, the relationship between the changes in these parameters supplies us with a basic prediction, which we can evaluate empirically in our experiment.
-Proposition 1.
-As model capacity $\theta$ improves, the effect of user skill on optimal quality diminishes.
-$\frac{\partial^{2}Q^{*}}{\partial s\partial\theta}=-\frac{k}{\theta^{2}s^{2}}<0$
-In other words, improvements in model capacity reduce the performance gap between high and low-skilled users.
-2.2 Decomposition into Model and Prompt Effects
-As model capacity $\theta$ increases, so does $x^{*}$ in equation 1, implying that a more capable model not only delivers better output for a fixed prompt but also encourages the user to devote more effort to composing prompts.
-This distinction highlights both a direct capacity-based effect of an improved model, a model effect, and an indirect effect driven by users adapting their prompts to the new model’s potential, a prompting effect.
-To see this more formally, suppose the model is upgraded from capacity $\theta_{1}$ to $\theta_{2}$.
-Let the old prompting effort be $x^{*}(\theta_{1},s)$ and the new (adapted) prompting effort be $x^{*}(\theta_{2},s)$.
-Define the total improvement in quality $\Delta Q$ is
-$\Delta Q\;=\;Q\Bigl{(}(\theta_{2},s,x^{*}(\theta_{2},s)\Bigr{)}\;-\;Q\Bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\Bigr{)}.$
-We can decompose $\Delta Q$ into
-$\Delta Q$ $={}\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{1},s))\;-\;Q(\theta_{1},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Model Effect }(M)\vphantom{\Big{|}}}$ (5)
-$+\underbrace{\Bigl{(}Q(\theta_{2},s,x^{*}(\theta_{2},s))\;-\;Q(\theta_{2},s,x^{*}(\theta_{1},s))\Bigr{)}}_{\text{Prompting Effect }(P)\vphantom{\Big{|}}}.$
-The first term isolates the gain from simply upgrading the model (holding the user’s prompt strategy fixed at the old optimum), while the second term represents any additional gain from prompt adaptation.
-In other words, even if $\theta$ improves, failing to adjust how prompts are written could leave substantial performance gains on the table.
-This decomposition naturally implies a prediction similar to Proposition 1 for each component of $\Delta Q$ that co-evolve with the changes in model capacity and user skill.
-In particular, when $k$ is small, the following are proved in Appendix E.1.
-Proposition 2.
-As user skill $s$ improves, the model effect $M$ decreases and the prompting effect $P$ increases.
-Intuitively, when costs are low and model capacity rises, higher-skilled users are already near the upper bound of performance, so the direct model effect they receive may be smaller.
-In contrast, users who can effectively refine prompts might benefit more from the new capacity, so the prompting effect they receive may be larger.
-Taken together, these theoretical predictions guide our experimental design, where we randomly vary $\theta$ and measure how users adapt.
+2.2 Decomposition into Model and Prompting Effects
+We now consider an upgrade from a model with capacity $\theta_{1}$ to one with higher capacity $\theta_{2}$.
+As model capacity $\theta$ increases, Equation 1 implies that optimal effort $x^{*}(\theta,s)$ also rises.
+Thus, an improved model affects performance through two channels: it directly increases output quality for a fixed prompt, and it indirectly encourages users to invest more effort in refining their prompts.
+We refer to these as the model effect and the prompting effect, respectively.
+Formally, let $x^{*}(\theta_{1},s)$ and $x^{*}(\theta_{2},s)$ denote the user’s optimal prompting effort before and after the upgrade. The total improvement in output quality is
+$\Delta Q=Q\bigl(\theta_{2},s,x^{*}(\theta_{2},s)\bigr)-Q\bigl(\theta_{1},s,x^{*}(\theta_{1},s)\bigr).$
+This total gain can be decomposed as
+$\Delta Q$ $=\underbrace{\Bigl(Q(\theta_{2},s,x^{*}(\theta_{1},s))-Q(\theta_{1},s,x^{*}(\theta_{1},s))\Bigr)}_{\text{Model effect }}$ (5)
+[…]
+The first term isolates the gain from upgrading the model while holding the user’s prompt strategy fixed at its pre-upgrade optimum.
+The second term captures the additional improvement arising from user-side prompt adaptation.
+Put differently, even if model capacity improves, one may forgo performance gains if their prompts are not updated accordingly.
+2.3 Distributional Impacts of Model Improvements
+A related natural question asks whether access to a more capable model levels the playing field, or widens performance gaps between users.
+In the language of our framework, how does a transition from a model with capacity $\theta_{1}$ to one with higher capacity $\theta_{2}$ affects the distribution of realized output quality.
+In other words, does access to a more capable model tend to level the playing field, or does it widen performance gaps between users?
+Under the bounded formulation of the framework analyzed here, $Q(\theta,s,x)=1-e^{-\theta sx}$, improvements in model capacity imply a specific pattern in how performance gains are distributed across users.
+Proposition 1 (Equalizing Effect of Model Improvements).
+Under the bounded formulation, improvements in model capacity compress the distribution of output quality across users.
+This aggregate result can be understood by examining the separate distributional implications of the model effect and the prompting effect.
+Proposition 2 (Model Effect: Equalizing Channel).
+Holding prompting effort fixed, an increase in model capacity compresses the outcome distribution: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more.
+Proposition 3 (Prompting Effect: Expanding Channel).
+Allowing users to re-optimize their prompting effort expands the outcome distribution: performance differences across users widen as the upper portion of the distribution shifts further upward.
+The formal proofs and derivations for all three propositions are provided in Appendix A.
+Taken together, these results imply that model improvements generate two opposing distributional forces.
+The direct mechanical improvement in capacity—the model effect —is equalizing, while the behavioral response—the prompting effect—is expanding.
+In aggregate, the equalizing channel dominates, producing the overall compression established in Proposition 1.333In the unbounded but concave logarithmic specification discussed in Appendix A.8, both the model and prompting channels are equalizing rather than opposing.
+The reversal stems from the additive interaction of capacity and skill, which makes improvements in $\theta$ more uniform across users rather than reinforcing skill differences.
+Intuitively, as model capacity increases, performance near the top of the distribution is already close to the performance ceiling and experiences smaller marginal gains, while adaptive prompting raises the performance of the upper tail more sharply.
 3 Experiment Design and Methods
-To empirically examine whether users do, in fact, adapt their prompts in response to model improvements—and how much this adaptation contributes to overall performance—we conducted a pre-registered online experiment with 2,059 participants on Prolific between December 12 and December 19, 2023.222Full pre-registration details, including hypotheses and planned analyses, are available in an online repository.
-Participants were asked to replicate a target image as closely as possible using a generative AI model, with the goal of assessing how both model capability and prompting behavior influence final outcomes.
-[^3] All procedures were approved by an institutional review board (link removed to preserve author anonymity) and participants provided informed consent.
+To empirically examine whether users do, in fact, adapt their prompts in response to model improvements—and how much this adaptation contributes to overall performance—we conducted two pre-registered online experiments.
+The first experiment (the image replication experiment, hereinafter) was conducted on a sample of 2,059 participants on Prolific between December 12 and December 19, 2023.
+In this experiment, participants were asked to replicate a target image as closely as possible using a generative AI model.
+The second experiment (the logo generation experiment, hereinafter) was conducted on a sample of 2,067 participants on Prolific between June 30 and July 1, 2025.
+In this experiment, participants were asked to create a logo for a hypothetical organization described in a short textual vignette (e.g., a sign for a business or a crest for a university).
+In both experiments, the goal was to assess how both model capability and prompting behavior influence final outcomes.
+[^4] Full pre-registration details for both experiments, including hypotheses and planned analyses, are available in an online repository.
+All procedures were approved by an institutional review board (protocol information removed to preserve author anonymity) and participants provided informed consent.
 We will release anonymized data and replication code upon publication.
+Together, these two experiments are designed to span distinct but empirically important regimes of generative AI use.
+The image replication experiment captures a bounded, steerable task in which success can be evaluated against a fixed target and output quality has a natural upper bound.
+This structure enables precise measurement of performance against an objective benchmark.
+By contrast, the logo generation experiment captures an open-ended, creative task in which goals are underspecified, quality is inherently comparative rather than absolute, and there is no clear performance ceiling.
+Despite these differences in task structure, both experiments share a common experimental architecture that allows us to decompose the performance gains into model-driven (model effect) and user-driven (prompting effect) components.
+Studying both settings allows us to assess whether prompt adaptation operates similarly when users are steering models toward a known objective versus when they are iterating within a creative design space, and to examine how the relative contributions of model capability and user adaptation vary across these regimes.
 3.1 Experimental Setting
-Each participant was randomly and blindly assigned to one of three model conditions: DALL-E 2, DALL-E 3 (hereafter “DALL-E 3 (Verbatim)”), or DALL-E 3 with automatic prompt revision (“DALL-E 3 (Revised)”).
+In both experiments, each participant was randomly and blindly assigned to one of three model conditions: DALL-E 2, DALL-E 3 (hereafter “DALL-E 3 (Verbatim)” or simply “DALL-E 3”), or DALL-E 3 with automatic prompt revision (“DALL-E 3 (Revised)”).
 These models differ not only in technical capability but also in whether they apply hidden large language model (LLM)-based modifications to user prompts.
@@
 In contrast, the DALL-E 3 API, by default, forwards user prompts to GPT-4 before image generation.
-This intermediate GPT-4 step rewrites the prompt—typically by adding detail or restructuring language—before passing it to the image model.
-This behavior is intended to improve image quality but occurs silently and without user visibility.
-In the DALL-E 3 (Revised) condition, we allowed this default behavior to proceed unaltered.
+This intermediate GPT-4 step rewrites the prompt (typically by adding detail or restructuring language) before passing it to the image model.
+This behavior is intended to improve image quality, but is not observed by or known to the prompter.
+In the DALL-E 3 (Revised) condition, we allowed this default behavior to proceed unaltered.This condition allows us to test whether system-side prompt rewriting serves as a substitute for user-side prompt adaptation, or instead changes users’ ability to steer the model.
 In the DALL-E 3 (Verbatim) condition, we attempted to suppress GPT-4-based prompt rewriting by prepending a hidden system message instructing the model to leave the user’s prompt unchanged.
@@
 In all conditions, participants were unaware that their prompt might be rewritten or modified before image generation.
-In addition to model assignment, each participant was independently assigned one of 15 target images.
+Importantly, all generative models in the experiment were memoryless: each prompt was processed independently, with no carryover from prior attempts.
+Participants were explicitly informed of this property before beginning the task to ensure they understood that previous prompts would not influence subsequent generations.
+Although this setup departs from typical real-world interfaces that maintain conversational context, it enables us to isolate prompt-level adaptation and to measure users’ iterative learning without confounding from model memory or accumulated dialogue.
+We designed a custom experimental interface resembling ChatGPT, which was used for both tasks.
+The participant’s “objective” (either the target image or logo brief, depending on the task) was displayed on the right side of the screen and a scrollable history of prompts and generated images on the left.
+For both experiments, participants had up to 25 minutes to submit at least 10 prompts.
+They were paid \$4 for completing the task, plus a \$8 bonus (a 200% increase) if their highest-scoring image (according to the relevant outcome metric below) was in the top 20% of participants.
+The median completion time in the image replication experiment was 23.5 minutes, with an a average hourly wage (including bonuses) of about \$13.
+The median completion time in the logo generation experiment was 26.1 minutes, with an average hourly wage (including bonuses) of about \$11.50.
+After they had finished submitting prompts, participants completed a demographic survey covering age, gender, education, occupational skills, and self-assessed proficiency in creative writing, programming, and generative AI.
+We removed from our analyses any participants who did not submit at least 10 prompts, repeated the same prompt five or more times consecutively, [^5] This pre-registered exclusion criterion was intended to filter out low-effort behavior in which participants repeatedly reused the same prompt in an attempt to complete the task as quickly as possible.
+While repeating a prompt could also reflect legitimate exploration given the stochastic nature of generative AI outputs, we note that this behavior occurred rarely in our data: the criterion affected 15 participants in the image replication experiment and 32 participants in the logo generation experiment (approximately 0.75–1.75% of participants).
+or failed to complete the post-task survey.
+This resulted in a final sample of 1,893 participants and 18,560 prompts in the image replication experiment, and a final sample of 1,857 participants and 18,425 prompts in the logo generation experiment.
+[^6] Although participants were required to submit at least 10 prompts to be included in our analysis, the final prompt count in both experiments is slightly below 10 $\times$ $n_{participants}$ because we excluded some prompts due to technical issues (e.g., safety filter triggers, duplicate attempt numbers) and limited our main analysis to each participant’s first 10 prompts to mitigate potential selection bias.
+Full details are provided in Appendix C.4.
+3.2 Image Replication Experiment
+In the image replication experiment, participants were asked to replicate a “target image” as closely as possible using the generative model provided to them.
+Each participant was randomly assigned one of 15 target images in addition to their model assignment.
 These images were drawn from three broad categories—business and marketing, graphic design, and architectural photography—to represent common use cases for text-to-image generation.
-All images were sourced from platforms that permit free research use (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography) [^4] All of these images have licenses for free use for commercial and noncommercial purposes..
-We built a custom interface resembling ChatGPT, with the assigned target image on the right and a scrollable history of prompts and generated images on the left.
-Participants were explicitly informed that the model was memoryless: every new prompt was processed independently, carrying no information from previous attempts.
-Participants had up to 25 minutes to submit prompts, with a requirement to submit at least 10.
-They were paid \$4 for completing the task, plus a \$8 bonus (a 200% increase) if their highest-scoring image was in the top 20% of participants.
-The median completion time was 22 minutes, implying an average hourly wage of about \$15.
-After replicating the target image, participants completed a demographic survey covering age, gender, education, occupation, and self-assessed proficiency in creative writing, programming, and generative AI.
-We removed from our analyses any participants who did not submit at least 10 prompts, repeated the same prompt five or more times in a row, or failed to complete the post-task survey, resulting in a final sample of 1,893 participants and 18,560 prompts.
-[^5] Although participants were required to submit at least 10 prompts to be included in our analysis, the final prompt count is slightly below 10 $\times$ 1,893 because we excluded some prompts due to technical issues (e.g., safety filter triggers, duplicate attempt numbers) and limited our main analysis to each participant’s first 10 prompts to mitigate potential selection bias.
-Full details are provided in Appendix C.4.
-3.2 Outcome Definition and Stochastic Generation
-The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al.
-2021).
+[^7] The selection of 15 target images was guided by the goal of capturing diversity across common text-to-image use cases (business and marketing, graphic design, and architectural photography), as well as by practical considerations identified through pilot testing.
+While a larger or systematically pretested image set would be desirable, there is no established standard for sampling text-to-image targets, and for a laboratory experiment of this scope some pragmatic choices were necessary.
+Our aim was to ensure that tasks were neither trivially easy nor prohibitively difficult, and to align broadly with categories commonly represented on repositories such as PromptBase.
+Target images were assigned using complete randomization across model–image cells, ensuring that image difficulty was orthogonal to model assignment.
+In addition, our analysis accounts for variation in difficulty by normalizing performance within each target image.
+Images were sourced from platforms that permit free research use for commercial and noncommercial purposes (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography).
+All images were sourced from platforms that permit free research use (e.g., Unsplash, Reshot, Shopify, Pixabay, Gratisography).
+3.2.1 Outcome Definition: Expected CLIP Embedding Cosine Similarity
+The primary outcome in the image replication experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (radford2021learning).
 CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together.
@@
 As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments.
-[^6] DreamSim (Fu et al.
-2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
+[^8] DreamSim (fu2023dreamsim) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
 Our results are robust to using DreamSim in place of CLIP cosine similarity.
-Full DreamSim-based analyses can be found in Appendix F.
+Full DreamSim-based analyses can be found in Appendix H.
 The two measures were highly correlated, and our findings were consistent across both.
+3.3 Logo Generation Experiment
+In the second experiment, each participant was asked to generate an original logo based on a short set of instructions or “briefs.” Each participant was randomly assigned one of 15 briefs in addition to their model assignment.
+One example is:
+“The owners of Green Pine Brewery are seeking a logo for their family-owned craft brewery located in Vermont.
+  • Required visual elements: pine tree, hops, barley
+  • Context for use: beer labels, merchandise, digital marketing
+Please create the best possible design based on the client’s description and the listed requirements.”
+All briefs followed this exact format: the name of the organization, a single sentence of background information, 2–3 required visual elements, and the anticipated context of use.
+The briefs span a range of topics, from a community park sign to a soccer team crest to a university letterhead.
+We provide the full list of briefs in Appendix B.
+They were written with three goals in mind: (i) capture a wide variety of realistic use cases, (ii) include a small number of concrete required elements that provide shared reference points for evaluators in an otherwise subjective task, and (iii) allow for open-ended creative possibilities.
+This third feature contrasts with the image replication experiment, where the best possible performance is well defined (perfectly replicate the target image).
+3.3.1 Outcome Definition: Bradley-Terry Strength
+Because each logo-generation task was open-ended and lacked an objective reference image, we could not assess output quality using embedding-based similarity metrics (e.g., cosine similarity of CLIP representations).
+Instead, we employed the Bradley–Terry (BT) model (bradley1952rank), a standard approach for inferring latent quality scores when only relative judgments are available (zermelo1929berechnung, ford1957solution).
+At a high level, the BT model is a probabilistic model that estimates a strength parameter (which we refer to hereafter as “BT $\beta$”) for each item based on the outcomes of many pairwise comparisons.
+[…]
+[…]
 3.4 Replay Analysis for Separating Model and Prompting Effects
-A central goal of our experiment was to distinguish how much of the performance improvement in image replication stems from using a more capable model versus how much comes from users adapting their prompts.
-Recall that the language of our conceptual framework shows that the total improvement in output quality from upgrading a generative AI model with capacity $\theta_{1}$ to a model with higher capacity $\theta_{2}$ can be written as:
-$\Delta Q=Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}-Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}.$
-We decompose this change into two parts. These are the model effect, the gain from applying the same prompts to a better model,
+A central goal of our experiment is to distinguish how much of the performance improvement in image replication stems from using a more capable model versus how much comes from users adapting their prompts.
+Recall that the language of our conceptual framework shows that the total improvement in output quality from upgrading a generative AI model with capacity $\theta_{1}$ to a model with higher capacity $\theta_{2}$ can be decomposed into two parts: the model effect, i.e., the gain from applying the same prompts to a better model,
 $M=Q\bigl(\theta_{2},s,x^{*}(\theta_{1},s)\bigr)-Q\bigl(\theta_{1},s,x^{*}(\theta_{1},s)\bigr),$
-and the prompting effect, the additional improvement from adapting prompts to take advantage of the more capable model,
+and the prompting effect, i.e., the additional improvement from adapting prompts to take advantage of the more capable model,
 $P=Q\bigl(\theta_{2},s,x^{*}(\theta_{2},s)\bigr)-Q\bigl(\theta_{2},s,x^{*}(\theta_{1},s)\bigr).$
-To estimate these components empirically, we conducted an additional analysis using prompts from participants in the DALL-E 2 and DALL-E 3 (Verbatim) conditions.
-To do so, we took the exact prompts that participants submitted during the experiment and re-submitted (or “replayed”) them to both their originally assigned model and the alternative model, generating new images in each case.
-[^7] Because the replay was conducted several weeks after the original experiment, all prompts were re-evaluated at the same time, using the same infrastructure, regardless of whether they were being run on the original or alternative model.
-This ensured that any observed differences in output quality could be cleanly attributed to model capacity and prompt adaptation rather than to changes in model behavior over time (i.e., model drift (Chen et al.
-2023)).
-More specifically, prompts written by DALL-E 2 participants—corresponding to $x^{*}(\theta_{1},s)$—were evaluated both on DALL-E 2 and on DALL-E 3, providing an empirical measure of $Q\bigl{(}\theta_{1},s,x^{*}(\theta_{1},s)\bigr{)}$ and $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$, respectively.
-This comparison isolates the model effect: the improvement in output quality when keeping prompts fixed and simply upgrading the model.
-To estimate the prompting effect, we compared the quality of reused DALL-E 2 prompts on DALL-E 3—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}$—to the quality of prompts originally written by DALL-E 3 participants and evaluated on the same model—an estimate of $Q\bigl{(}\theta_{2},s,x^{*}(\theta_{2},s)\bigr{)}$.
-This captures the additional improvement from users adapting their prompts to better leverage the capabilities of the more advanced system.
+To estimate the model and prompting components empirically, we conducted a replay-based analysis using prompts from participants in the DALL-E 2 and DALL-E 3 (Verbatim) conditions.
+In both experiments, we re-evaluated prompts under the alternative model—replaying DALL-E 2 prompts on DALL-E 3 and DALL-E 3 prompts on DALL-E 2—thereby holding prompt content fixed while varying model capacity.
+These cross-model evaluations provide empirical analogues of $Q(\theta_{2},s,x^{*}(\theta_{1},s))$ and $Q(\theta_{1},s,x^{*}(\theta_{2},s))$, respectively, and together with the original prompt–model pairings, allow us to implement the decomposition into model and prompting effects shown in Equation 5.
+In the image replication experiment, replay was conducted several weeks after the initial data collection, so we additionally re-generated outputs for the original prompt–model combinations contemporaneously with replay to guard against model drift (chen2023gptchanging).
+In the logo generation experiment, replay occurred shortly after data collection, so this additional step was unnecessary.
 4 Results
-In this section, we present our empirical findings on how increased model capacity affects user performance in replicating target images.
-In particular, we focus on three main questions: (i) whether access to a more capable model (DALL-E 3) improves performance; (ii) how users modify their prompts in response to the improved model; (iii) and how much of the overall performance gain can be attributed to model improvements versus prompt adaptation.
-Finally, we compare these empirical results to the predictions of our conceptual framework.
-Our analysis draws on both the main experimental data and the replay data described above.
-4.1 Overall Impact of Model Upgrades
-We begin by examining whether participants using DALL-E 3 achieve higher performance than those using DALL-E 2 as implied by Equation 3.
+In this section, we present our empirical findings on how increases in model capacity affect user performance across the two tasks.
+For each task, we first examine whether access to a more capable model improves average output quality, how users adapt their prompts in response, and how much of the overall performance gain can be attributed to model improvements versus prompt adaptation using the replay analyses.
+Next, we study how these gains reshape the distribution of outcomes across users and interpret these distributional impacts through the lens of our conceptual framework.
+Finally, we assess the impact of automated prompt rewriting using the DALL-E 3 Revised condition.
+4.1 Image Replication Task
+We begin with the image replication task.
+As discussed above, this task represents a bounded, steerable application of generative AI, in which users aim to match a known target and output quality has a natural upper bound.
+4.1.1 Overall Impact of Model Upgrades
+We first examine whether participants using DALL-E 3 achieve higher performance than those using DALL-E 2, as implied by Equation 3.
 Figure 1 summarizes these findings.
-Panel A presents three representative target images and, for each, three generated images drawn from the full sample of participants across both model conditions.
+The top row of Panel A presents three of the target images provided to participants.
+Below each target, three generated images are shown, drawn from the full sample of participants across both model conditions.
 The middle row for each target shows the image whose cosine similarity to the target is closest to the mean similarity across all participants.
@@
 The gap persists across all attempts; participants using DALL-E 3 start off producing closer matches and maintain that edge through their 10th prompt.
-Figure 1: Overall Performance and Prompting Behavior.
+Figure 1: Overall Performance and Prompting Behavior for the Image Replication Task
 (A) For three example target images, the middle row shows participant-generated images closest to the mean similarity across all prompts.
@@
 (C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals.
-Color shading indicates the average textual similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
+Color shading indicates the average textual embedding similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
 Participants’ dynamic prompting behavior also differs substantially between the two models.
 As shown in panel C of Figure 1, those assigned to DALL-E 3 write prompts that are, on average, 24% longer than those assigned to DALL-E 2, and this gap widens over successive attempts.
-Moreover, we observe that DALL-E 3 participants are more likely to reuse or refine their previous prompts (indicated by the color scale), which suggests a more exploitative approach once they discover the model’s capacity to handle detailed or complex instructions.
+Moreover, we observe that DALL-E 3 participants are more likely to reuse or refine their previous prompts (indicated by the color scale—see Appendix C.3 for details), which suggests a more exploitative approach once they discover the model’s capacity to handle detailed or complex instructions.
 Analyses of parts of speech confirm that these extra words likely provide additional descriptive information rather than mere filler: the proportion of nouns and adjectives—the two most descriptively informative parts of speech—is nearly identical across model conditions (48% for DALL-E 3 vs.
-49% for DALL-E 2; $p=0.215$).
-This suggests that the increase in prompt length reflects the addition of semantically rich content rather than unnecessary verbosity.
-4.2 Replay Analysis and Decomposition of Effects
+49% for DALL-E 2; $p=0.215$), suggesting that the increase in prompt length reflects the addition of semantically rich content rather than unnecessary verbosity.
+Together, these patterns illustrate prompt adaptation in practice: users progressively supply more information-dense prompts as they naturally learn the model’s affordances—despite no explicit instructions to do so.
+4.1.2 Replay Analysis and Decomposition of Effects
 The differences we observe in prompting behavior suggest that users are actively adapting to the capabilities of the model they are assigned.
 But how much of the overall performance improvement we observe for DALL-E 3 users is due to the model’s enhanced technical capacity, and how much is due to users rewriting their prompts in response to that capacity?
-To answer this question, we turn to the replay analysis described earlier, which allows us to isolate these two effects identified in Equation 5 empirically.
+To answer this question, we turn to the replay analysis described earlier, which allows us to isolate these two effects empirically.
 Panel A of Figure 2 presents the results.
-To estimate the model effect, we compare the performance of prompts originally written by DALL-E 2 participants when evaluated on DALL-E 2 (the model they were written for) versus when evaluated on DALL-E 3 (Verbatim).
+To estimate the model effect, we hold prompts fixed and compare how the same DALL-E 2 prompts perform when evaluated on DALL-E 2 versus when replayed on DALL-E 3 (Verbatim).
 Because these prompts were written without knowledge of DALL-E 3’s capabilities, any improvement reflects the gain from using a more capable model while holding the prompt fixed.
@@
 Because both sets of prompts are evaluated on the same model, any difference reflects the effect of users adapting their prompts to the model’s capabilities.
-We find that this prompting effect accounts for the remaining 48% of the total improvement, corresponding to an increase of 0.0079 in cosine similarity ($p=0.024$).
-The total treatment effect—i.e., the difference between the original DALL-E 2 prompts on DALL-E 2 and the original DALL-E 3 prompts on DALL-E 3—is 0.0164.
+We find that this prompting effect accounts for the remaining $\approx 48$% of the total improvement [^10] The total, model, and prompting effects are estimated using three separate two-way fixed-effects models, as explained in Appendix F.1.3.
+Because each effect is estimated from a different subset of counterfactual prompt–model comparisons with its own fixed-effects, the resulting model and prompting effect estimates are not guaranteed to sum exactly to the total improvement., corresponding to an increase of 0.0079 in cosine similarity ($p=0.024$).
 Importantly, when we apply prompts written by DALL-E 3 users to DALL-E 2, we observe no performance benefit relative to the original DALL-E 2 prompts ($\Delta=0.0020$; $p=0.56$).
 This asymmetry reinforces the idea that the gains from prompt adaptation depend on the model’s capacity to act on that additional information.
-Figure 2: Replay Analysis and Effect Decomposition.
+Figure 2: Replay Analysis and Effect Decomposition for the Image Replication Task.
 (a) Average performance of prompts evaluated across four prompt-model combinations.
@@
 Taken together, these findings offer empirical support for our theoretical claim: prompt adaptation operates as a dynamic complement that users deploy in response to improved model capabilities—and accounts for a substantial share of realized performance gains.
-4.3 Skill Heterogeneity
+4.1.3 Distributional Effects
+Table 1: Distributional Effects of Model Upgrades in the Image Replication Task
 Effect Estimate (SE) p-value
-Total 0.0224 (0.00312) $<$ 0.00001
-Total $\times$ Performance Decile -0.000115 (0.000047) 0.0152
-Model 0.0113 (0.00186) $<$ 0.00001
-Model $\times$ Performance Decile -0.000059 (0.000025) 0.0210
-Prompting 0.0109 (0.00316) 0.000564
-Prompting $\times$ Performance Decile -0.000056 (0.000048) 0.244
-Table 1: Regression estimates for the total, model and prompting effects, and their interactions with participant performance decile.
-The total and model effects are significantly larger for lower-skilled participants, as indicated by the negative and statistically significant interaction term.
-The prompting effect does not vary significantly across deciles, suggesting that participants across the skill distribution adapt their prompts similarly.
-Table 1 presents the results of a regression analysis that tests whether the model, prompting and total effects vary systematically across participants of differing skill levels.
-We regress performance on indicators for model, prompting and total effects, as well as their interactions with participants’ performance decile (ranging from lowest to highest).
-We make three observations.
-First, the interaction between total effect and performance decile is negative and statistically significant ($-0.000115$, $p=0.0152$).
-This suggests that model improvements reduce the overall gap between high and low-performing users, consistent with proposition 1 in our conceptual framework.
-Second, the interaction between the model effect and performance decile is also negative and statistically significant ($-0.000059$, $p=0.0210$).
-This suggests that the model effect mainly benefits the lower-performing users—consistent with our theoretical predictions in proposition 2 and due to diminishing returns for those already near the performance ceiling.
-Third, we do not find any evidence that the benefits of prompt adaptation vary across the skill distribution in a meaningful way since the interaction between the prompting effect with performance decile is not statistically significant ($-0.000056$, $p=0.2444$).
-While this contrasts with our theoretical prediction in Proposition 2—which anticipated that higher-skill participants would benefit more from prompt adaptation—it is possible that the effect exists but is modest in size and difficult to detect given the statistical power of our design.
-4.4 Prompt Revision
-We also evaluate whether an automated prompt adaptation system can substitute for human-driven adaptation.
+Total 0.022 (0.003) << $10^{-5}$
+Total $\times$ Outcome Quantile (Percentile) -0.00012 (0.00005) 0.015
+Model 0.011 (0.002) << $10^{-5}$
+Model $\times$ Outcome Quantile (Percentile) -0.00006 (0.00003) 0.021
+Prompting 0.011 (0.003) << $10^{-5}$
+Prompting $\times$ Outcome Quantile (Percentile) -0.00006 (0.00005) 0.244
+Negative interactions with outcome quantile indicate that gains from model upgrades are larger at lower points in the realized outcome distribution, implying compression of performance outcomes.
+The total and model effects exhibit statistically significant compression, while we do not detect a statistically significant distributional pattern for the prompting effect.
+Note: the un-interacted total effect coefficient gives the QTE estimate at the 0th percentile.
+Next, we examine how model improvements and prompt adaptation reshape the distribution of performance outcomes in the image replication task using a quantile treatment effects (QTE) perspective.
+Table 1 reports estimates of the total, model, and prompting effects at different points of the realized outcome distribution, with outcome quantiles indexed by percentile.
+The total effect exhibits a negative and statistically significant interaction with outcome quantile (0.000115-0.000115, $p=0.0152$), indicating that performance gains from upgrading from DALL-E 2 to DALL-E 3 are larger at lower points in the outcome distribution.
+This pattern is consistent with the framework’s prediction that model improvements compress outcomes in bounded tasks.
+This compression is driven primarily by the model effect.
+Holding prompts fixed, the interaction between the model effect and outcome quantile is also negative and statistically significant (0.000059-0.000059, $p=0.0210$), consistent with diminishing returns near the performance ceiling.
+While the bounded formulation predicts that prompt adaptation exerts an expanding force on the outcome distribution, we do not detect a statistically significant interaction between the prompting effect and outcome quantile in this setting (0.000056-0.000056, $p=0.2444$).
+This result does not rule out the presence of an expanding prompting channel, but implies that any such effect is modest in magnitude relative to the dominant equalizing model effect.
+4.1.4 Prompt Revision
+Finally, we examine whether automated prompt rewriting can serve as a substitute for human-driven prompt adaptation in the image replication task.
 In the DALL-E 3 (Revised) condition, user prompts were silently rewritten by GPT-4 before being submitted to the image model.
-Although participants in this condition still outperformed those using DALL-E 2 ($\Delta=0.0069$; $p=0.042$), they achieved substantially lower performance than participants using DALL-E 3 without revision.
+Although participants in this condition still outperformed those using DALL-E 2 ($\Delta=0.0069$; $p=0.042$), they achieved substantially lower performance than participants using DALL-E 3 without rewriting.
 On average, automated prompt revision reduced the benefit of DALL-E 3 by 58% (95% CI: [40%, 76%]).
-Manual inspection of the revised prompts in the DALL-E 3 (Revised) arm confirms that GPT-4 often added extraneous details or subtly altered the intended meaning of participant instructions, leading to degraded performance.
-In this case, the automatic rewriting process was not well-aligned with the user’s goal of replicating a specific target image as faithfully as possible.
-More broadly, this result highlights a potential risk of using invisible system-level prompt rewriting: when the objectives of the rewriting mechanism are not tightly aligned with user intent, even well-intentioned modifications can interfere with the user’s ability to effectively guide the model.
-Together, these findings demonstrate that the performance gains observed when users are given access to a more capable generative model stem from both improvements in model architecture and changes in user behavior.
-Participants assigned to DALL-E 3 not only achieve higher performance but also write longer, more refined prompts over time.
-Our replay analysis shows that these adapted prompts account for nearly half of the overall performance improvement, with the remainder attributable to the model’s enhanced rendering capabilities.
-This pattern holds across the skill distribution: while lower-skilled users benefit more from model upgrades, prompt adaptation contributes consistently across all levels of baseline performance.
-Finally, automated prompt rewriting—when misaligned with user intent—fails to replicate the benefits of direct user adaptation.
-These results highlight the central role that prompting behavior plays in shaping how users interact with and benefit from improved generative AI systems.
-5 Discussion and Conclusion
-Our findings contribute to the literature on human–AI collaboration and the economics of technology by highlighting the role of prompt adaptation as a dynamic complement that co-evolves with model improvements (Böhm and Schedlberger 2023).
-As generative AI models advance—often substantially from month to month—organizations that fail to adapt their prompting strategies may forgo a meaningful share of the economic value these upgrades make possible.
+Qualitative inspection of revised prompts suggests that the rewriting process often introduced additional details or altered phrasing in ways that were misaligned with participants’ objective of replicating a specific target image as faithfully as possible.
+In this bounded, steerable task, such modifications reduced users’ ability to precisely control the model’s output and partially offset the performance gains achieved through direct user-side prompt adaptation.
+4.2 Logo Generation Task
+We next turn to the logo generation task.
+This logo generation task, however, represents an open-ended, creative application of generative AI—such as design or ideation—in which success is evaluated comparatively rather than against a known target and output quality does not have a natural upper bound.
+Our empirical analyses mirror that of the image replication experiment.
+Figure 3: Overall Performance and Prompting Behavior for the Logo Generation Task
+(A) For three example target images, the middle row shows participant-generated images closest to the mean Bradley-Terry $\beta$ across all prompts.
+The rows above and below show images of approximately one average treatment effect (ATE) higher or lower quality, illustrating the typical performance difference between model conditions.
+(B) Top: average Bradley-Terry $\beta$, separately for DALL-E 2 and DALL-E 3 participants.
+Bottom: the difference between these averages, with the dark blue line indicating the overall ATE and the shaded region showing the 95% confidence interval.
+(C) Average prompt length by attempt (y-axis), with error bars representing 95% confidence intervals.
+Color shading indicates the average textual similarity between each prompt and the participant’s previous prompt, capturing the extent of prompt reuse and refinement over time.
+4.2.1 Overall Impact of Model Upgrades
+We begin by examining whether access to a more capable model improves performance in the logo generation task.
+Participants assigned to DALL-E 3 produce substantially higher-quality logos than those assigned to DALL-E 2, with an estimated average treatment effect of 0.8075 in Bradley–Terry $\beta$ (95% CI: [0.7120, 0.9030], $p<10^{-10}$), corresponding to an improvement of approximately 0.57 standard deviations.
+Panel A of Figure 3 provides qualitative intuition for the magnitude of this effect by showing representative logos at the mean, as well as approximately one average treatment effect above and below the mean.
+Panel B of Figure 3 shows how this performance gap unfolds over successive prompting attempts.
+As in the image replication task, the advantage of the more capable model appears early and persists, although it does not appear to diverge across attempts.
+Finally, Panel C of Figure 3 shows similar differences to the image replication task in prompting behavior across model conditions, with participants using DALL-E 3 prompts that are approximately 24.7% longer on average and exhibiting greater iteration-to-iteration similarity, consistent with more exploitative refinement.
+The proportion of nouns and adjectives is also once again statistically indistinguishable across model conditions (43% for DALL-E 3 vs.
+42% for DALL-E 2; $p$ = 0.12)
+Together, these patterns show that in an open-ended, creative task without a fixed target or performance ceiling, increased model capability leads to large and persistent performance gains.
+These gains are accompanied by systematic changes in user prompting behavior as users adapt their prompts in response to the model’s perceived capabilities.
+These results largely mirror those from the image replication task.
+4.2.2 Replay Analysis and Decomposition of Effects
+Next, we apply the same replay-based decomposition used in the image replication task to separate the overall performance gains in the logo generation experiment into components attributable to model improvements and to user prompt adaptation.
+Panel A of Figure 4 presents the resulting decomposition.
+The overall average treatment effect of upgrading from DALL-E 2 to DALL-E 3 is $0.8075$ in Bradley–Terry $\beta$.
+In contrast to the image replication task, the vast majority of this gain is attributable to the model effect: replaying prompts written by DALL-E 2 participants on DALL-E 3 increases performance by $0.748$ ($p<10^{-10}$), accounting for approximately $92.7\%$ of the total effect.
+The corresponding prompting effect is small in magnitude ($0.0577$), statistically indistinguishable from zero ($p=0.101$), and accounts for only $7.2\%$ of the total effect.
+Consistent with this pattern, replaying prompts written for DALL-E 3 on DALL-E 2 yields no performance improvement ($\Delta\beta=-0.0125$, $p=0.62$).
+Panel B of Figure 4 provides a qualitative illustration of this pattern for a representative logo brief.
+Although the prompt written by the DALL-E 3 participant is longer and more detailed—consistent with the prompt adaptation patterns shown in Panel C of Figure 3—it does not perform substantially better on DALL-E 3 than the prompt written by the DALL-E 2 participant.
+Instead, the primary performance difference arises from evaluating either prompt on the more capable model.
+While this comparison is illustrative rather than representative, it highlights how prompt adaptation in this task does not translate into large additional gains beyond those delivered by the model itself.
+Taken together, these results indicate that in the logo generation task, realized performance gains are driven primarily by improvements in model capability, with prompt adaptation playing a comparatively limited role.
+Figure 4: Replay Analysis and Effect Decomposition for the Logo Generation Task.
+(a) Average performance of prompts evaluated across four prompt-model combinations.
+Comparing DALL-E 2 prompts evaluated on DALL-E 2 versus DALL-E 3 isolates the model effect.
+Comparing reused DALL-E 2 prompts on DALL-E 3 to original DALL-E 3 prompts on DALL-E 3 isolates the prompting effect.
+Error bars represent 95% confidence intervals based on bootstrapped standard errors clustered at the participant level.
+(b) A single client description with two submitted prompts: one written by a DALL-E 2 participant (top row) and one by a DALL-E 3 participant (bottom row).
+Images show how each prompt performs on both models, illustrating the model and prompting effects qualitatively.
+4.2.3 Distributional Effects
+Table 2: Distributional Effects of Model Upgrades in the Logo Generation Task
+Effect Estimate (SE) p-value
+Total 1.135 (0.050) << $10^{-5}$
+Total $\times$ Outcome Quantile (Percentile) -0.0065 (0.0008) << $10^{-5}$
+Model 1.012 (0.035) << $10^{-5}$
+Model $\times$ Outcome Quantile (Percentile) -0.0053 (0.0005) << $10^{-5}$
+Prompting 0.122 (0.050) 0.014
+Prompting $\times$ Outcome Quantile (Percentile) -0.0013 (0.0008) 0.117
+Negative interactions with outcome quantile indicate that gains from model upgrades are larger at lower points in the realized outcome distribution, implying compression of performance outcomes.
+As in the image replication task, distributional compression is driven by the model effect, while we do not detect a statistically significant distributional pattern for the prompting effect.
+Note: the un-interacted total effect coefficient gives the QTE estimate at the 0th percentile.
+We also apply the same QTE approach used in the image replication task to examine how the model upgrade reshapes the distribution of performance outcomes in the logo generation experiment.
+Table 2 reports estimates of the total, model, and prompting effects at different points of the realized outcome distribution, with outcome quantiles indexed by percentile.
+As was the case in the image replication task, the total effect of upgrading from DALL-E 2 to DALL-E 3 exhibits a negative and statistically significant interaction with outcome quantile, indicating that performance gains are larger at lower points in the outcome distribution and that the upgrade compresses the distribution of outcomes.
+This pattern is again driven by the model effect: holding prompts fixed, the interaction between the model effect and outcome quantile is negative and statistically significant.
+By contrast, we do not detect a statistically significant interaction between the prompting effect and outcome quantile.
+Although this lack of statistical significance does not rule out the compressive impact of prompt adaptation predicted by the unbounded formulation of the model, it places an upper bound on the magnitude of any such effect in this setting.
+Taken together, these results indicate that, as in the image replication task, distributional compression in the logo generation experiment arises primarily from improvements in model capability, with prompt adaptation playing a comparatively limited role in shaping the distribution of outcomes.
+4.2.4 Prompt Revision
+We conclude analyses of this experiment by examining the effects of automated prompt rewriting in the logo generation task.
+As in the image replication task, participants assigned to DALL-E 3 with automatic prompt revision outperform those using DALL-E 2, with an average treatment effect of $0.8337$ in BT $\beta$ ($p<10^{-10}$).
+However, in contrast to the image replication task, automated prompt rewriting does not reduce performance relative to DALL-E 3 without rewriting; instead, it is associated with a $3.2\%$ increase in BT $\beta$ on average (95% CI: $5.6\%,12.1\%]$).
+One possible interpretation is that in an open-ended, creative task, the specific prompt rewriting used here, which primarily embellishes and elaborates user inputs, can enrich prompts in ways that facilitate exploration, introduce useful variation, or highlight design elements that improve perceived quality.
+While we do not directly observe these mechanisms, the positive effect of rewriting in this task is consistent with the idea that automated prompt modification can complement user prompting when it is aligned with users’ goals and the structure of the task.
+5 Discussion
+In this paper, we show that realized performance gains from generative AI upgrades reflect an interaction between model capability, user adaptation, and task structure.
+In a bounded, steerable task with an objective target (image replication), improvements in model quality translated into performance gains through two channels: roughly half of the improvement arose mechanically from the stronger model, while the other half reflected users adapting their prompts to better exploit the model’s capabilities.
+In this setting, prompt adaptation is a substantive and economically meaningful complement to model improvement.
+By contrast, in an open-ended creative task without a natural performance ceiling (logo generation), performance gains were driven primarily by improvements in model capability.
+Although users again modified their prompting behavior when interacting with the stronger model, replay analyses indicate that prompt adaptation accounted for only a small share of the overall improvement.
+Thus, we argue that prompt adaptation functions as a dynamic complement to generative AI models, with task structure emerging as a key boundary condition: it plays a first-order role when success is defined against a fixed objective, but a more limited role when quality is open-ended and comparative.
+A related question we address is whether prompt adaptation must occur through direct user experimentation or whether it can be partially automated by the system itself.
+In practice, many generative AI systems implement automated prompt rewriting to reduce user effort, implicitly positioning automation as a substitute for user-driven adaptation.
+Our findings suggest that the effectiveness of such automation depends critically on task structure and alignment with user goals.
+In the bounded image replication task, the DALL-E 3 endpoint’s automated prompt rewriting reduced realized gains, consistent with misalignment between the rewriting logic and users’ objective of faithfully reproducing a target image.
+By contrast, in the open-ended logo generation task, the same rewriting mechanism modestly improved performance.
+Together, these findings underscore the importance of careful design in automated prompt rewriting systems.
+When system prompts are well aligned with user goals and task structure, automated rewriting can improve performance.
+but poorly aligned system prompts, especially when hidden, can degrade performance and/or limit users’ ability to discover or exploit emergent but useful ways of interacting with a generative AI system.
+Such complexities are highlighted by related work, such as yao2024clinicians, which studies automated prompt optimization in the context of clinical documentation.
+More generally, our results both align with and help clarify themes in the literature on human–AI collaboration and the economics of technology, particularly the role of dynamic complements such as prompt adaptation that co-evolve with technical change (bohm2023co-evolution).
+As generative AI models advance, often substantially from month to month, organizations that fail to adapt their prompting strategies in task settings where precision and steerability matter may forgo a meaningful share of the economic value these upgrades make possible.
 This co-evolutionary pattern is not unique to AI.
-It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (Brynjolfsson 1993, David 1990).
+It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (brynjolfsson1993productivity, david1990dynamo).
 However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter.
-Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (Brynjolfsson and Hitt 2000, Brynjolfsson et al.
-2021).
-Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—accounts for roughly half of the realized performance gains associated with a model upgrade.
-This finding extends research on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al.
-2010, Teece et al.
-1997) and post-adoptive IT use, which emphasizes how value emerges through user experimentation and learning over time (Jasperson et al.
-2005).
+Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (brynjolfsson2000beyond, brynjolfsson2021productivity).
+Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—can account for a substantial share of realized performance gains from model upgrades in task settings where precision and steerability matter.
+Seen from a broader organizational perspective, these findings also speak to long-standing work on IT-enabled dynamic capabilities (bharadwaj2000resource, joshi2010changing, teece1997dynamic) and post-adoptive IT use (jasperson2005comprehensive).
 Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill.
-Participants in our study—none of whom were expert prompt engineers—improved performance through trial-and-error within a single session.
-This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (Attewell 1992, Von Hippel 2006).
-The accessibility of prompt adaptation suggests a path toward more broadly distributed productivity gains—provided that users are supported by scaffolds and interfaces that enable iterative refinement (Rogers 2003).
+Participants in our study, drawn from a general population rather than a specialized group of prompt engineers, improved performance through trial-and-error within a single session.
+This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (attewell1992technology, von2006democratizing).
+The relative accessibility of prompt adaptation highlights one potential pathway toward more broadly distributed productivity gains, provided that users are supported by scaffolds and interfaces that enable iterative refinement (rogers2003diffusion).
 At the same time, this accessibility introduces risk.
 Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve.
-This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (Henderson and Clark 1990).
+This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (henderson1990architectural).
 Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
-Finally, our results caution against automated prompting systems that intervene without aligning with user intent.
-In our experiment, DALL-E 3’s automatic prompt rewriting—designed to enhance usability—reduced realized performance gains by 58%.
-This does not imply that automated prompting is inherently counterproductive.
-In some settings, automation may simplify use or improve consistency.
-But when the goals of a rewriting system diverge from those of the user—as in our task, where general improvements (e.g., aesthetic quality or detail) were prioritized over faithful image replication—such interventions can interfere with the adaptive strategies users are developing.
-Similar challenges have been observed in other domains: for instance, in clinical documentation, automated prompt optimization improved consistency but still required manual revision to ensure quality and alignment (Yao et al.
-2024).
-These findings reinforce the broader insight—rooted in both human–AI collaboration and dynamic capabilities research—that adaptive user behavior is not merely a response to system performance but a critical input.
-Interventions that disrupt this adaptive process risk eroding the very dynamic complements that make advanced systems productive (Fügener et al.
-2022).
-While this study offers new insight into prompt adaptation, it has several limitations.
-First, our focus was restricted to a single transition (from DALL-E 2 to DALL-E 3) and one type of generative AI (text-to-image generation).
+Our study has several limitations.
+First, our experimental interface was intentionally memoryless: each prompt was processed independently, and participants were informed that prior interactions would not influence subsequent generations.
+This simplification was necessary to isolate prompt-level adaptation in a controlled setting, but it departs from the conversational context that characterizes many real-world AI systems; our estimates therefore characterize performance and adaptation conditional on a memoryless interaction structure.
+Second, our focus was restricted to a single transition (from DALL-E 2 to DALL-E 3) and one type of generative AI (text-to-image generation).
 Although our conceptual framework suggests these mechanisms may generalize to other domains, further research is needed to assess how these dynamics play out in text generation, programming, scientific research, and other high-stakes settings.
-Second, we observe short-run adaptation behavior in a controlled setting, whereas longer-term learning dynamics, organizational feedback structures, or team-based workflows may shape prompting strategies differently in real-world environments.
-Third, while our replay analysis isolates the effect of prompt adaptation, we do not identify which specific types of prompt modifications (e.g., lengthening, lexical substitution, structural rephrasing) have a causal effect on performance.
-Future work should investigate these finer-grained mechanisms to better inform prompting practices and interface design.
+Third, we observe short-run adaptation behavior in a controlled setting, whereas longer-term learning dynamics, organizational feedback structures, or team-based workflows may shape prompting strategies differently in real-world environments.
+Finally, while our replay analysis helps isolate the contribution of prompt adaptation relative to model improvements, it does not fully disentangle the causal effects of specific prompt modifications (e.g., lengthening, lexical substitution, structural rephrasing), particularly given the path-dependent nature of user adaptation.
 Building on these limitations, several promising research directions emerge.
-One direction is to examine prompt adaptation longitudinally—tracking how users develop durable heuristics that generalize across domains and respond to shifting model behavior over time.
+One important next step is to examine prompt adaptation in conversational, memory-enabled settings, where prior interactions persist and can be leveraged for cumulative learning by both users and models, and to study how this co-evolution impacts performance.
+Relatedly, future studies might explore prompt adaptation longitudinally, observing how users develop durable heuristics that generalize across domains and respond to shifting model behavior over time.
 Another line of inquiry involves organizational-level complements to prompting, such as shared prompt repositories, collaborative refinement practices, or analytics dashboards that surface effective patterns.
@@
 For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance.
-The near-equal contributions of model improvements and prompt adaptation suggest that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (Arthur 2009, Dosi 1982).
+In some task settings, prompt adaptation can account for as much as half of the realized performance gains from model upgrades, underscoring that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (arthur2009nature, dosi1982technological).
 As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
-Appendix A Experiment Design
-A.1 Task Design
+Appendix A Conceptual Framework and Derivations
+This appendix develops the formal derivations and proofs underlying the conceptual framework presented in the main text.
+Whereas the main text emphasizes intuition and empirical implications, the goal here is to make the underlying mechanics transparent.
+We first analyze a bounded quality function, where output quality is capped at 1, and then extend the analysis to unbounded tasks.
+The exposition below focuses on the bounded case; the unbounded case follows by analogy.
+A.1 Model Setup
+Consider a task where users attempt to replicate a target object (e.g., an image) using a generative model. Perfect replication yields a quality of 1. Define the quality function
+$Q(\theta,s,x)=1-\exp[-\theta sx],$
+where $\theta\in(0,1]$ denotes model capacity; $s\in(0,1]$ denotes user skill at writing prompts; and $x\geq 0$ denotes the user’s effort in generating prompts.
+This specification has two desirable properties:
+  1. 1.
+    Boundedness. Even for very large $x$, $Q<1$; quality cannot exceed perfection.
+  2. 2.
+    Diminishing returns. The marginal gain from effort, $Q_{x}=\theta se^{-\theta sx}$, declines as $x$ increases.
+Prompting entails a linear cost function
+$c(x)=kx,\qquad k>0,$
+interpreted as the time or cognitive cost per unit of effort. The user’s utility from a given combination of model capacity, skill, and effort is
+$U(\theta,s,x)=1-e^{-\theta sx}-kx.$
+A.2 User Optimization
+The user chooses effort $x\geq 0$ to maximize $U(\theta,s,x)$. The first-order condition is:
+$\theta se^{-\theta sx}=k\;\;\Rightarrow\;\;x^{*}(\theta,s)=\frac{1}{\theta s}\ln\!\Bigl(\frac{\theta s}{k}\Bigr).$
+An interior optimum requires $\theta s>k$, ensuring that $x^{(}\theta,s)>0$ and the logarithm is well defined.
+Intuitively, the user’s effective productivity $\theta s$ must exceed the marginal cost of effort $k$.
+This inequality is readily satisfied when prompting costs are small (i.e., $k\approx 0$).
+Substituting the optimal effort into the quality function yields the user’s maximum attainable quality:
+$Q^{*}(\theta,s)=1-\frac{k}{\theta s}.$ (6)
+This closed-form expression will serve as the basis for all subsequent comparative statics and distributional analysis.
+A.3 Comparative Statics
+The partial derivatives of $Q^{*}$ are straightforward:
+$\frac{\partial Q^{*}}{\partial\theta}=\frac{k}{\theta^{2}s}>0,\qquad\frac{\partial Q^{*}}{\partial s}=\frac{k}{\theta s^{2}}>0,\qquad\frac{\partial Q^{*}}{\partial k}=-\frac{1}{\theta s}<0.$
+Hence, output quality increases in both model capacity and user skill, and decreases with the cost of effort.
+Improvements in model quality or prompting skill therefore unambiguously raise output quality, whereas higher marginal costs of effort reduce it.
+A.4 Distributional Impact of Model Improvements (Overall Equalizing Effect)
+We now examine how increases in model capacity $\theta$ reshape the distribution of optimal quality $Q^{*}$ across users.
+Because $Q^{*}$ is strictly increasing in skill $s$, dispersion in $Q^{*}$ reflects variation in skill, since $Q^{*}$ is a monotone transformation of $s$ (and hence of $G_{s}$).
+Let $G_{Q^{*}}(\pi)$ denote the inverse CDF of $Q^{*}$—that is, the quality achieved by the user at the $\pi$-th percentile of the quality distribution—and let $G_{s}(\pi)$ denote the inverse CDF of the skill distribution.
+Then:
+$G_{Q^{*}}(\pi)=1-\frac{k}{\theta G_{s}(\pi)}.$
+Proposition 1 (Equalizing Effect of Model Improvements).
+For any strictly increasing skill distribution $G_{s}(\pi)$, we have
+$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}<0.$
+Proof.
+Differentiating once gives
+$\frac{\partial G_{Q^{*}}(\pi)}{\partial\theta}=\frac{k}{\theta^{2}G_{s}(\pi)}>0,$
+so all users benefit from higher $\theta$. Differentiating again with respect to $\pi$,
+$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}=-\frac{k}{\theta^{2}G_{s}^{2}(\pi)}\,G_{s}^{\prime}(\pi)<0,$
+because $G_{s}^{\prime}(\pi)>0$. ∎
+Model improvements therefore compress the distribution of output quality, reducing dispersion across users.
+Intuitively, greater model capacity raises quality throughout the distribution but shifts the lower portion upward more strongly in relative terms, since outputs at the top are already close to the performance ceiling.
+A.5 Decomposing Total Improvements in Quality
+The total change in optimal quality with respect to model capacity, $\theta$, can be expressed as a combination of a direct effect of the model itself and an indirect effect that operates through re-optimization of effort.
+To distinguish between these two direct and indirect channels, we introduce a counterfactual quality function that treats the model capacity determining effort and the model capacity entering the quality function as distinct parameters.
+$Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))=1-\exp[-\theta_{2}s\,x^{*}(\theta_{1},s)]=1-\Bigl(\frac{k}{\theta_{1}s}\Bigr)^{\!\theta_{2}/\theta_{1}},$
+which represents the quality achieved when a user employs the effort that would be optimal for model $\theta_{1}$ while using a model with capacity $\theta_{2}$.
+The model and prompting effects are obtained by differentiating this counterfactual quality function with respect to $\theta_{2}$ and $\theta_{1}$, respectively, and evaluating the results at $\theta_{1}=\theta_{2}=\theta$:
+$M(\theta,s)$ $=\frac{\partial Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))}{\partial\theta_{2}}=\frac{k}{\theta^{2}s}\ln\!\Bigl(\frac{\theta s}{k}\Bigr)$
+$P(\theta,s)$ $=\frac{\partial Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))}{\partial\theta_{1}}=\frac{k}{\theta^{2}s}\!\left[1-\ln\!\Bigl(\frac{\theta s}{k}\Bigr)\right].$
+$M(\theta,s)$ captures the direct mechanical improvement in quality when model capacity increases but effort is held fixed.
+In contrast, $P(\theta,s)$, captures the behavioral response—the gain that results from users adjusting their effort once the model improves.
+By construction,
+$\frac{dQ^{*}(\theta,s)}{d\theta}=M(\theta,s)+P(\theta,s).$
+A.6 Distributional Implications of the Model and Prompting Effects
+Although the decomposition in Appendix A.5 holds pointwise for each user, the distributional consequences of model and prompting effects require characterizing how these effects reshape the distribution of outcomes.
+Let $G_{Q^{cf}}(\pi)$ denote the inverse CDF of the counterfactual quality:
+$G_{Q^{cf}}(\pi)=1-\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)^{\!\frac{\theta_{2}}{\theta_{1}}}.$
+Proposition 2 (Model Effect: Equalizing Channel).
+Holding prompting effort fixed, an increase in model capacity compresses the distribution of output quality: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more.
+More formally,
+$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{2}\,\partial\pi}<0.$
+Proof.
+Differentiating with respect to $\theta_{2}$,
+$\frac{\partial G_{Q^{cf}}(\pi)}{\partial\theta_{2}}=-\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)^{\!\frac{\theta_{2}}{\theta_{1}}}\frac{1}{\theta_{1}}\ln\!\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr).$
+Differentiating again and setting $\theta_{1}=\theta_{2}=\theta$,
+$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{2}\,\partial\pi}=\frac{k}{\theta^{2}G_{s}^{2}(\pi)}G_{s}^{\prime}(\pi)\Bigl[\ln\!\Bigl(\frac{k}{\theta G_{s}(\pi)}\Bigr)+1\Bigr].$
+For sufficiently small $k$, the bracketed term is negative, implying that the entire expression is negative. ∎
+Proposition 3 (Prompting Effect: Expanding Channel).
+Allowing users to re-optimize their prompting effort expands the distribution of output quality: performance differences across users widen as the upper portion of the distribution shifts further upward.
+More formally,
+$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{1}\,\partial\pi}>0.$
+Proof.
+Differentiating with respect to $\theta_{1}$,
+$\frac{\partial G_{Q^{cf}}(\pi)}{\partial\theta_{1}}=\frac{\theta_{2}}{\theta_{1}^{2}}\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)^{\!\frac{\theta_{2}}{\theta_{1}}}\Bigl[\ln\!\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)+1\Bigr].$
+Differentiating again and setting $\theta_{1}=\theta_{2}=\theta$,
+$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{1}\,\partial\pi}=-\frac{1}{\theta}\Bigl(\frac{k}{\theta G_{s}(\pi)}\Bigr)\frac{G_{s}^{\prime}(\pi)}{G_{s}(\pi)}\Bigl[\ln\!\Bigl(\frac{k}{\theta G_{s}(\pi)}\Bigr)+2\Bigr].$
+For sufficiently small $k$, the term in brackets is negative, so the derivative is positive. ∎
+A.7 Summary of Distributional Impacts
+In this exponential-bounded specification, the distributional consequences of all effects are analytically unambiguous:
+Effect Cross-partial sign Distributional interpretation
+Model effect <0<0 Equalizing (Compression)
+Prompting effect >0>0 Expanding (Dispersion)
+Total effect <0<0 Equalizing (Compression)
+The overall equalizing pattern established in Section A.4 thus arises because the compressive model channel dominates the dispersive prompting channel in magnitude.
+Intuitively, as model capacity increases, gains in output quality are largest at the lower end of the distribution, where outcomes remain far from the quality ceiling.
+Improvements at the upper end are more limited, leading to an overall compression in the distribution of performance.
+Behavioral adaptation through prompting has the opposite tendency, stretching the upper tail, but not enough to offset the dominant equalizing shift.
+A.8 Unbounded Quality
+We now consider a concave but unbounded quality specification,
+$Q(\theta,s,x)=\log\!\big(1+(\theta+s)x\big)$
+with the same linear cost $c(x)=kx$.
+This formulation preserves monotonicity and concavity in $x$, $\theta$, and $s$, but relaxes the upper bound on attainable quality.
+The user’s utility function is therefore
+$U(\theta,s,x)=\log\!\big(1+(\theta+s)x\big)-kx$
+User optimization.
+The first-order condition is
+$\frac{\theta+s}{1+(\theta+s)x}=k$
+Assuming an interior solution ($\theta+s>k$), the optimal effort, quality, and utility are
+$x^{*}(\theta,s)$ $=\frac{1}{k}-\frac{1}{\theta+s}$
+$Q^{*}(\theta,s)$ $=\log\!\left(\frac{\theta+s}{k}\right)$
+$U^{*}(\theta,s)$ $=\log\!\left(\frac{\theta+s}{k}\right)-1+\frac{k}{\theta+s}$
+Comparative statics.
+Optimal quality increases with both model capacity and user skill, and decreases with prompting cost:
+$\frac{\partial Q^{*}}{\partial\theta}$ $=\frac{1}{\theta+s}>0\qquad\frac{\partial Q^{*}}{\partial s}=\frac{1}{\theta+s}>0\qquad\frac{\partial Q^{*}}{\partial k}=-\frac{1}{k}<0$
+Distributional Impact of Model Improvements (Overall Equalizing Effect)
+We now analyze how increases in model capacity $\theta$ reshape the distribution of optimal quality $Q^{*}$ across users.
+Let $G_{Q^{*}}(\pi)$ denote the inverse CDF of $Q^{*}$ and $G_{s}(\pi)$ the inverse CDF of skill.
+Then:
+$G_{Q^{*}}(\pi)=\log\!\left(\frac{\theta+G_{s}(\pi)}{k}\right)$
+Proposition 4 (Equalizing Effect of Model Improvements, Unbounded Case).
+For any strictly increasing skill distribution $G_{s}(\pi)$,
+$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}=-\frac{G_{s}^{\prime}(\pi)}{(\theta+G_{s}(\pi))^{2}}<0$
+Proof.
+Differentiating once gives
+$\frac{\partial G_{Q^{*}}(\pi)}{\partial\theta}=\frac{1}{\theta+G_{s}(\pi)}>0$
+so all users gain from higher $\theta$. Differentiating again with respect to $\pi$,
+$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}=-\frac{G_{s}^{\prime}(\pi)}{(\theta+G_{s}(\pi))^{2}}<0$
+Model improvements therefore compress the quality distribution, raising performance throughout the distribution but proportionally more at lower quantiles—mirroring the equalizing effect found in the bounded model.
+Counterfactual quality and decomposition.
+As in the bounded model, we decompose total quality improvements into direct (model) and indirect (prompting) components using a counterfactual formulation:
+$Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))=\log\!\Big(1+(\theta_{2}+s)\,x^{*}(\theta_{1},s)\Big)$
+where $\theta_{2}$ governs the model capacity in the quality function and $\theta_{1}$ determines the user’s optimal effort.
+Differentiating with respect to $\theta_{2}$ and $\theta_{1}$ and evaluating at $\theta_{1}=\theta_{2}=\theta$ gives
+$M(\theta,s)=\frac{(\theta+s)-k}{(\theta+s)^{2}},\qquad P(\theta,s)=\frac{k}{(\theta+s)^{2}}$
+so that the total effect of model improvement equals the sum of model and prompting effects: $\frac{dQ^{*}(\theta,s)}{d\theta}=M(\theta,s)+P(\theta,s)$ as before.
+Distributional implications.
+Let $G_{Q^{cf}}(\pi)$ denote the inverse CDF of counterfactual quality.
+It can be shown that $Q^{cf}$ is strictly increasing in $s$ if $k<\theta_{1}+s$, a condition needed for an interior solution.
+Thus:
+$G_{Q^{cf}}(\pi;\theta_{1},\theta_{2})=\log\!\Bigg(1+\frac{\theta_{2}+G_{s}(\pi)}{k}-\frac{\theta_{2}+G_{s}(\pi)}{\theta_{1}+G_{s}(\pi)}\Bigg)$
+Proposition 5 (Model Effect: Equalizing Channel, Unbounded Case).
+Holding prompting effort fixed, an increase in model capacity compresses the distribution of output quality: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more.
+Formally,
+$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{2}\,\partial\pi}=\frac{2k-\big(\theta+G_{s}(\pi)\big)}{(\theta+G_{s}(\pi))^{3}}G_{s}^{\prime}(\pi)<0$
+which is negative for sufficiently small $k$.
+Proposition 6 (Prompting Effect: Equalizing Channel, Unbounded Case).
+Allowing users to re-optimize their prompting effort compresses the distribution of output quality: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more.
+Formally,
+$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{1}\,\partial\pi}=-\frac{2k}{(\theta+G_{s}(\pi))^{3}}G_{s}^{\prime}(\pi)<0.$
+Appendix B Experiment Design Details
+In this appendix, we provide details on the design of both experiments reported in the main text.
+B.1 Randomization
+For both experiments, we randomized participants across two dimensions: the target task (target image for the image replication experiment and target brief for the logo generation experiment) and the text-to-image generative AI model that participants had access to.
+We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target item-model cells per experiment.
+We conducted a balance check after the conclusion of each experiment with a $\chi^{2}$ test across all cells.
+For the image replication experiment, we obtained $\chi^{2}=7.056,\;df=44,\;p>0.999$, thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups.
+For the logo generation experiment, we obtained $\chi^{2}=5.383,\;df=44,\;p>0.999$, thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups.
+In both experiments, participants were unaware of the underlying randomization.
+B.1.1 Balance Checks
+Table A.1 reports covariate balance checks for the image replication experiment across prior GPT usage, gender, and age.
+We find no statistically significant evidence of imbalance across any of these dimensions.
+Table A.2 reports the corresponding balance checks for the logo generation experiment.
+While one pairwise comparison yields a $p$-value below 0.05, this is not unexpected given that we conduct 18 balance tests across the two experiments.
+Importantly, the statistically significant comparison arises in a contrast involving the DALL-E 3 (Revised) arm.
+As described in the main text, this arm is included primarily as a secondary, diagnostic condition to assess the consequences of automated system-side prompt rewriting, and it is not central to the paper’s primary treatment contrasts or replay-based identification strategy.
+We do not rely on direct comparisons between DALL-E 3 (Revised) and DALL-E 3 (Verbatim) for our headline results, and inspection of the balance table indicates that the deviation is driven by the DALL-E 3 (Revised) group rather than broad differences across the core arms of interest.
+Finally, to assess whether this pattern could have been induced by sample restrictions or data processing, in Table A.3 we re-conduct the logo-generation balance checks on the full sample prior to applying any exclusions.
+The same low $p$-value persists, indicating that it is not created by filtering and is consistent with random variation in assignment rather than some systematic bias in our treatment of the data.
+Table A.1: Covariate Balance Across All Three Arms (Image Replication Experiment)
+Covariate / Level N(D2) Prop(D2) N(D3r) Prop(D3r) N(D3v) Prop(D3v) D2 vs. D3r D2 vs. D3v D3r vs. D3v
+p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df)
+GPT Usage p = 0.745, $\chi^{2}$ = 1.948 (4) p = 0.654, $\chi^{2}$ = 2.449 (4) p = 0.629, $\chi^{2}$ = 2.587 (4)
+None 215 0.329 186 0.305 186 0.296
+Little 190 0.291 178 0.292 195 0.311
+Some 179 0.274 173 0.284 170 0.271
+A lot 70 0.110 72 0.118 76 0.121
+NA/Missing 1 0.002 0 0.000 2 0.003
+Gender p = 0.727, $\chi^{2}$ = 1.309 (3) p = 0.319, $\chi^{2}$ = 3.517 (3) p = 0.231, $\chi^{2}$ = 4.297 (3)
+Female 354 0.541 332 0.545 341 0.544
+Male 279 0.427 261 0.428 271 0.432
+Other or prefer not to say 21 0.032 16 0.026 13 0.021
+NA/Missing 1 0.002 0 0.000 4 0.006
+Age Group p = 0.847, $\chi^{2}$ = 2.013 (5) p = 0.649, $\chi^{2}$ = 3.334 (5) p = 0.250, $\chi^{2}$ = 6.631 (5)
+[18,26) 112 0.171 109 0.179 99 0.158
+[26,32) 136 0.208 122 0.200 122 0.195
+[32,39) 139 0.213 132 0.217 147 0.235
+[39,49) 127 0.194 127 0.208 119 0.190
+[49,83] 140 0.214 119 0.195 138 0.220
+NA/Missing 1 0.002 0 0.000 4 0.006
+  • Sample sizes: DALL-E 2 (D2) = 655, DALL-E 3 Revised (D3r) = 609, DALL-E 3 Verbatim (D3v) = 629.
+    $N(\cdot)$ columns report the count of observations in each treatment arm belonging to that category, $\text{Prop}(\cdot)$ gives the proportion of respondents in that category for the arm.
+Chi-squared test columns show p-values and $\chi^{2}$ test statistics with degrees of freedom for each pairwise comparison.
+Table A.2: Covariate Balance Across All Three Arms (Logo Generation Experiment)
+Covariate / Level N(D2) Prop(D2) N(D3r) Prop(D3r) N(D3v) Prop(D3v) D2 vs. D3r D2 vs. D3v D3r vs. D3v
+p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df)
+GPT Usage p = 0.098, $\chi^{2}$ = 7.843 (4) p = 0.165, $\chi^{2}$ = 6.497 (4) p = 0.032, $\chi^{2}$ = 10.542 (4)
+None 67 0.107 42 0.069 70 0.113
+Little 117 0.186 123 0.201 100 0.162
+Some 205 0.327 227 0.371 234 0.379
+A lot 237 0.378 219 0.358 213 0.345
+NA/Missing 2 0.003 1 0.002 0 0.000
+Gender p = 0.901, $\chi^{2}$ = 0.581 (3) p = 0.563, $\chi^{2}$ = 2.046 (3) p = 0.417, $\chi^{2}$ = 2.838 (3)
+Female 323 0.516 326 0.533 319 0.520
+Male 297 0.474 277 0.453 293 0.480
+Other / Prefer not to say 6 0.010 7 0.011 5 0.008
+NA/Missing 2 0.003 2 0.003 0 0.000
+Age Group p = 0.465, $\chi^{2}$ = 4.610 (5) p = 0.501, $\chi^{2}$ = 4.348 (5) p = 0.392, $\chi^{2}$ = 5.203 (5)
+[18,29) 126 0.201 120 0.196 113 0.185
+[29,35) 103 0.164 126 0.206 118 0.193
+[35,43) 145 0.231 129 0.211 134 0.219
+[43,52) 123 0.196 105 0.172 129 0.211
+[52,87] 129 0.206 130 0.213 123 0.201
+NA/Missing 2 0.003 2 0.003 0 0.000
+  • Sample sizes: DALL-E 2 (D2) = 628, DALL-E 3 Revised (D3r) = 612, DALL-E 3 Verbatim (D3v) = 617.
+    $N(\cdot)$ columns report the count of observations in each treatment arm belonging to that category, $\text{Prop}(\cdot)$ gives the proportion of respondents in that category for the arm.
+Chi-squared test columns show p-values and $\chi^{2}$ test statistics with degrees of freedom for each pairwise comparison.
+Table A.3: Covariate Balance Across All Three Arms (Logo Generation Experiment, Pre-filtering; N = 1924)
+Covariate / Level N(D2) Prop(D2) N(D3r) Prop(D3r) N(D3v) Prop(D3v) D2 vs. D3r D2 vs. D3v D3r vs. D3v
+p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df)
+GPT Usage p = 0.134, $\chi^{2}$ = 7.044 (4) p = 0.267, $\chi^{2}$ = 5.203 (4) p = 0.034, $\chi^{2}$ = 10.441 (4)
+None 67 0.103 44 0.070 73 0.114
+Little 121 0.186 123 0.196 101 0.158
+Some 212 0.326 235 0.375 240 0.374
+A lot 248 0.382 223 0.356 227 0.354
+NA/Missing 3 0.005 2 0.003 5 0.008
+Gender p = 0.876, $\chi^{2}$ = 0.690 (3) p = 0.825, $\chi^{2}$ = 0.900 (3) p = 0.762, $\chi^{2}$ = 1.163 (3)
+Female 330 0.508 332 0.530 332 0.518
+Male 311 0.478 285 0.455 304 0.474
+Other / Prefer not to say 7 0.011 7 0.011 5 0.008
+NA/Missing 3 0.005 3 0.005 5 0.008
+Age Group p = 0.568, $\chi^{2}$ = 3.873 (5) p = 0.877, $\chi^{2}$ = 1.793 (5) p = 0.636, $\chi^{2}$ = 3.418 (5)
+[18,29) 130 0.200 122 0.195 118 0.184
+[29,35) 111 0.171 130 0.208 121 0.189
+[35,43) 147 0.226 130 0.208 140 0.218
+[43,52) 128 0.197 109 0.174 133 0.207
+[52,87] 132 0.203 133 0.212 129 0.201
+NA/Missing 3 0.005 3 0.005 5 0.008
+  • Sample sizes: DALL-E 2 (D2) = 651, DALL-E 3 Revised (D3r) = 627, DALL-E 3 Verbatim (D3v) = 646. Includes 7 users missing survey crosswalk.
+    $N(\cdot)$ columns report the count of observations in each treatment arm belonging to that category, $\text{Prop}(\cdot)$ gives the proportion of respondents in that category for the arm.
+Chi-squared test columns show p-values and $\chi^{2}$ test statistics with degrees of freedom for each pairwise comparison.
+B.2 Generative Models
+For both experiments, we randomly assigned participants to 1 of 3 generative models:
+  1. 1.
+    DALL-E 2
+  2. 2.
+    DALL-E 3 (Verbatim) (sometimes simply referred to as DALL-E 3)
+  3. 3.
+    DALL-E 3 (Revised)
+Both the “verbatim” and “revised” versions of the DALL-E 3 treatment utilize the same underlying image-generating model; the distinction lies in the pre-processing applied before submitting user prompts to OpenAI’s image-generating API.
+OpenAI’s DALL-E 3 system, by design, employs a GPT-4 model to rewrite user prompts, adding more detail before processing the modified prompt using the DALL-E 3 image-generating model.
+During our experiment, it was not possible to explicitly disable this prompt rewriting feature of the DALL-E 3 system.
+To manage this behavior, we defined two treatments utilizing the DALL-E 3 model.
+In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process.
+In the DALL-E 3 (Verbatim) treatment arm, we prepend text to the participant’s prompt instructing the GPT-4 model to refrain from modifying it before passing it forward to DALL-E 3.
+This prepended text was never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint.
+[^11] See here and here for online documentation.
+We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The exact text we prepended to participant prompts is copied below:
+“I NEED to test how the tool works with my prompt as it is written. DO NOT add any detail; just use it AS IS:”
+Prepending this text to participants’ prompts did reduce the rate at which OpenAI’s endpoint modified prompts, but compliance was not perfect.
+Thus, we view the “verbatim” treatment arm as more of an intent-to-treat intervention.
+The GPT model still modified 59% of participant prompts.
+The average token sort ratio (TSR) between the original prompt and the modified prompt was 77 for the DALL-E 3 (Verbatim) arm, compared to an average token sort ratio of 44 across the entire DALL-E 3 (Revised) treatment arm (a TSR of 100 denotes an exact string match).
+Conditional on any modification (any observations with TSR << 100), the average TSR between the original prompt and the modified prompt was 61 for the DALL-E 3 (Verbatim) arm, compared to an average TSR of 44 across the entire DALL-E 3 (Revised) treatment arm.
+B.3 Model Endpoints
+We used the following model endpoints and parameters to generate images from prompts:
+  1. 1.
+    OpenAI JavaScript API: For both the image replication and logo generation experiments, we used the image generation endpoint of the official OpenAI JavaScript library (release version 3.1.0) to generate images for user prompts in the experimental interface.
+For all treatment arms, we set the image size parameter to be 1024 x 1024 pixels.
+For the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we set the quality parameter to standard and the style parameter to natural.
+  2. 2.
+    Azure OpenAI Service: For the image replication experiment, we used the image generation endpoints in the Python implementation of Azure OpenAI Service to generate all replay images based off the original user prompts.
+For prompts replayed through the DALL-E 2 treatment arm, we deployed a set of DALL-E 2 models on Azure OpenAI Service and set the API version for each to the 2023-06-01-preview version.
+For prompts replayed through the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we deployed a set of DALL-E 3 models on Azure OpenAI Service and set the API version for each to the 2023-12-01-preview version.
+The parameter values for image size, and quality and style for the DALL-E 3 treatment arms, were set to the same values as above.
+  3. 3.
+    OpenAI Python API: For the logo generation experiment, we used the image generation endpoints in the official OpenAI Python library (release version 0.27.8) to generate all replay images based off the original user prompts.
+The parameter values for image size, and quality and style for the DALL-E 3 treatment arms, were set to the same values as above.
+B.4 Image Replication Experiment
+In this subsection, we describe design elements specific to the image replication experiment.
+B.4.1 Task Design
 Participants were asked to reproduce a single target image as closely as possible using a text-to-image generative AI model (i.e., DALL-E 2, DALL-E 3 without prompt revision, or DALL-E 3 with prompt revision, all developed by OpenAI).
@@
 In response to each submitted prompt, the model would generate an image, which was then displayed to the participant next to their assigned target image.
-Participants were instructed to make at least 10 attempts at trying to recreate the target image within a 25-minute window, with no upper limit on their number of attempts.
+Participants were instructed to make at least 10 attempts to recreate the target image within a 25-minute window, with no upper limit on their number of attempts.
 All interactions between participants and the generative AI models occurred on a custom-built online interface designed to resemble OpenAI’s ChatGPT interface but with some adjustments related to our task (e.g., displaying the target image and the total number of attempts so far to the user).
@@
 We placed the text box where participants were able to write and submit their prompts at the bottom of the interface.
-Prompts were limited to a maximum of 1,000 characters.
-Participants were informed that their interactions with their assigned model would be memory-less, i.e., the model retained no memory of previous prompts and only used the current prompt to generate each image.
+Prompts were limited to a maximum of 1,000 characters (approximately 200-250 words).
+Participants were informed that their interactions with their assigned model would be memoryless, i.e., the model retained no memory of previous prompts and only used the current prompt to generate each image.
 Before the task, participants were provided with written and video instructions on how to interact with our experiment interface.
@@
 We also inquired about their self-assessed occupational skills and how often they 1) engaged in creative writing, 2) wrote specific instructions, and 3) engaged in any sort of computer programming.
-Finally, we collected socio-demographic data, such as age, gender, and occupation.
-A.2 Randomization
-We randomized participants across two dimensions: the target image and the text-to-image generative AI model that participants had access to.
-We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells.
-We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test across all cells.
-With $\chi^{2}=7.056,\;df=44$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
-$H_{0}:p_{1}=p_{2}=\cdots=p_{45}$
-Participants were unaware of this randomization.
-A.2.1 Generative Models
-We randomly assigned participants to 1 of 3 generative models:
-  1. 1.
-    DALL-E 2, which is referred to at points in the main text as the inferior model.
-  2. 2.
-    DALL-E 3 (Verbatim), which is referred to at points in the main text as the superior model.
-  3. 3.
-    DALL-E 3 (Revised), which is referred to at points in the main text as DALL-E 3 with Revision
-Both the “verbatim” and “revised” versions of the DALL-E 3 treatment utilize the same underlying image-generating model; the distinction lies in the pre-processing applied before submitting user prompts to OpenAI’s image-generating API.
-OpenAI’s DALL-E 3 system, by design, employs a GPT-4 model to rewrite user prompts, adding more detail before processing the modified prompt using the DALL-E image-generating model.
-During our experiment, it was not possible to explicitly disable this prompt rewriting feature of the DALL-E 3 system.
-To manage this behavior, we defined two treatments utilizing the DALL-E 3 model.
-In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process.
-In the DALL-E 3 (Verbatim) treatment arm, we prepend a string instructing the GPT-4 model to not modify the participant’s prompt before passing it forward to DALL-E 3.
-This string is never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint.
-[^8] See here and here for online documentation.
-We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
-“I NEED to test how the tool works with my prompt as it is written. DO NOT add any detail; just use it AS IS:”
-Prepending this string to participants’ prompt did reduce the rate at which OpenAI’s endpoint modified prompts, but compliance was not perfect.
-Thus, we view the “verbatim” treatment arm as more of an intent-to-treat intervention.
-The GPT model still modified 59% of participant prompts.
-The average token sort ratio (TSR) between the original prompt and the modified prompt was 77 for the DALL-E 3 (Verbatim) arm, compared to an average token sort ratio of 44 across the entire DALL-E 3 (Revised) treatment arm (a TSR of 100 denotes an exact string match).
-Conditional on any modification (any observations with TSR $<$ 100), the average TSR between the original prompt and the modified prompt was 61 for the DALL-E 3 (Verbatim) arm, compared to an average TSR of 44 across the entire DALL-E 3 (Revised) treatment arm.
-A.2.2 Target Images
+Finally, we collected socio-demographic data, such as age and gender.
+B.4.2 Target Images
 We randomly assigned participants to 1 of 15 target images.
@@
 In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
-A.3 Subjects
-Our Prolific-recruited US sample (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once.
+B.4.3 Subjects
+Our Prolific-recruited US sample for the image replication task (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once.
 We also prevented users from completing the task on mobile devices or tablets.
 Data was collected between December 12, 2023 and December 19, 2023.
-Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section C.5.1).
-The median time to complete our entire task, including a demographic survey, was 22 minutes.
-Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$15 USD per hour.
+Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in terms of the image most similar to the target (this was operationalized with the Dreamsim score between each image and the targe, which we describe in section D.1).
+The median (mean) time to complete our entire task, including a demographic survey, was 23.5 (26.4) minutes.
+Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$13 USD per hour.
 We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task.
@@
 For subsequent checks, participants were required to retry until they demonstrated understanding.
-A.4 Model Endpoints
-We used the following model endpoints and parameters to generate images from prompts:
-  1. 1.
-    OpenAI API: We used the image generation endpoint of the official OpenAI Node.js library to generate images for user prompts during the experiment.
-For all treatment arms, we set the image size parameter to be 1024 x 1024 pixels.
-For the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we set the quality parameter to standard and the style parameter to natural.
-  2. 2.
-    Azure OpenAI Service: We used the image generation endpoints in the Python implementation of Azure OpenAI Service to generate all replay images based off user prompts collected during the experiment.
-For prompts replayed through the DALL-E 2 treatment arm, we deployed a set of DALL-E 2 models on Azure OpenAI Service and set the API version for each to the 2023-06-01-preview version.
-For DALL-E 2, we created replay images in batches of 5.
-For prompts replayed through the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we deployed a set of DALL-E 3 models on Azure OpenAI Service and set the API version for each to the 2023-12-01-preview version.
-The parameter values for image size, and quality and style for the DALL-E 3 treatment arms, were set to the same values as in the experiment.
-A.5 Image Similarity Metrics
-The primary outcome in our experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al.
-2021).
-CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together.
-By embedding both the target image and each generated image into this space and computing the cosine similarity between them, we obtain a quantitative measure of how closely the generated image matches the target along both visual and conceptual dimensions.
-Because the output of each generative model is stochastic, the same prompt can yield different images across attempts.
-To account for this variability, we generated 10 images for each prompt and computed their cosine similarity to the target image individually.
-We then averaged these 10 similarity scores to produce an expected quality score for each prompt—the primary outcome variable used in our analysis.
-As a robustness check, we replicated all analyses using DreamSim, a recently developed perceptual similarity metric that aligns more closely with human judgments.
-[^9] DreamSim (Fu et al.
-2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
-Our results are robust to using DreamSim in place of CLIP cosine similarity.
-Full DreamSim-based analyses can be found in Appendix F.
-The two measures were highly correlated, and our findings were consistent across both.
-A.6 Pre-registration
+B.4.4 Pre-registration
 This study was pre-registered.
@@
 The pre-registration document can be found at URL removed to maintain anonymity.
-Appendix B Example Prompts
-Table 2 presents the complete set of prompts and their corresponding generated images that were used in our analysis shown in Figure 1.
+B.5 Logo Generation Experiment
+B.5.1 Task Design
+The logo generation task design was identical in format to that of the image replication experiment.
+The only difference was that participants were instructed to make the best logo possible according to the instructions presented on the right-hand side of the interface.
+They were similarly required to make at least 10 attempts in 25 minutes and to answer the same set of post-task survey questions.
+B.5.2 Logo Instructions
+Participants were randomly assigned to one of fifteen logo-generation tasks.
+Each task described an organization that needed a logo and included the instruction: “Please create the best possible design based on the client’s description and listed requirements.” All briefs provided (i) a title, (ii) one sentence of background, (iii) required visual elements, and (iv) the intended context of use.
+They were written with three goals in mind: to capture a wide variety of realistic use cases, to include a few objective elements so independent evaluators could assess them consistently, and to allow for open-ended creative possibilities.
+In no particular order, the fifteen logo tasks were:
+1. Craft Pine Brewery Logo
+The owners of Green Pine Brewery are seeking a logo for their family-owned craft brewery located in Vermont.
+Required visual elements: pine tree, hops, barley.
+Context for use: beer labels, merchandise, digital marketing.
+2. Logo for Community Park Sign
+The City of Springfield Parks Department is seeking a sign for the new Elmwood Community Park, a welcoming outdoor family space.
+Required visual elements: trees, playground equipment, children playing.
+Context for use: park entrance, leaflets, community posters.
+3. Soccer Team Logo
+The coach of Riverdale FC is seeking a logo for their soccer team to build team spirit and local pride.
+Required visual elements: the city’s iconic bridge over the river, soccer ball.
+Context for use: uniforms, merchandise, stadium banners/signage.
+4. University Logo
+Prestwick University is seeking a logo emphasizing academic excellence. Required visual elements: open book, torch. Context for use: digital marketing, printed materials, university merchandise.
+5. Coffee Shop Logo
+Downtown Brew, a trendy independent coffee shop, is seeking a logo to attract people as a place to work.
+Required visual elements: coffee cup, city skyline.
+Context for use: cups, signage, digital marketing.
+6. Artisan Bakery Logo
+The Daily Loaf, an artisan bakery known for sourdough breads, wants a logo highlighting craftsmanship.
+Required visual elements: loaf of bread, wheat stalk.
+Context for use: packaging, storefront, digital marketing.
+7. Fitness Studio Logo
+Peak Performance Gym wants a logo to attract wealthy young professionals. Required visual elements: dumbbell, mountain peak. Context for use: gym signage, apparel, digital marketing.
+8. Music Festival Logo
+The organizers of the annual Summer Beats Festival seek a logo to celebrate electronic music and summer vibes.
+Required visual elements: sound waves, the sun.
+Context for use: posters, tickets, merchandise.
+9. Veterinary Clinic Logo
+The Happy Tails Vet Clinic seeks a friendly logo for their animal health practice. Required visual elements: dog, cat, heart. Context for use: clinic signage, uniforms, magazine ads/paper leaflets.
+10. Local Farmers Market Logo
+The Green Valley Weekend Farmers Market seeks a logo emphasizing fresh, local produce and artisan crafts.
+Required visual elements: basket of vegetables, shovel.
+Context for use: market banners, tote bags, digital marketing.
+11. Ice Cream Shop Logo
+Super Scoops, an ice cream parlor, desires a playful and attractive logo. Required visual elements: ice cream cone, tongue. Context for use: storefront, menus, ice cream packaging.
+12. Bookstore Logo
+Pages & Co., a cozy independent bookstore, seeks a logo to attract book lovers. Required visual elements: smiling person holding a book. Context for use: signage, bookmarks, stickers.
+13. Mountain Resort Logo
+The owners of Highland Retreat, a mountain resort, want a logo emphasizing outdoor adventure.
+Required visual elements: mountain range, cabin.
+Context for use: resort signage, merchandise, digital marketing.
+14. Tech Startup Logo
+MegaSurf is a tech startup developing a mobile app that connects surfers in a neighborhood.
+Required visual elements: surfboard, ocean wave.
+Context for use: app icon, digital marketing, business cards.
+15. Nonprofit Organization Logo
+The leaders of Happy Housing seek a compassionate-looking logo representing their mission to ensure all children have a home.
+Required visual elements: smiling child, a roof with a chimney.
+Context for use: fundraising leaflets, t-shirts, event signage.
+B.5.3 Subjects
+Our Prolific-recruited US sample for the logo generation task (N = 2,067) was limited to fluent English speakers, and we prevented participants from completing the task more than once.
+We also prevented users from completing the task on mobile devices or tablets.
+Data was collected between June 30, 2025 and July 1, 2025.
+Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in BT score of their best-performing image.
+The median (mean) time to complete our entire task, including a demographic survey, was 26.1 (29.5) minutes.
+Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$11.50 USD per hour.
+We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task.
+The onboarding process also included multiple attention checks; participants who failed the first check were immediately disqualified.
+For subsequent checks, participants were required to retry until they demonstrated understanding.
+B.5.4 Pre-registration
+This study was pre-registered.
+The pre-registration document included our hypotheses, planned analyses, and sample size justification.
+The pre-registration document can be found at URL removed to maintain anonymity.
+Appendix C Data
+This appendix provides detailed information about the data collection, measurement, and variables used in our analyses, expanding on the overview provided in the main text.
+C.1 Survey Data
+We collected additional participant information via a Qualtrics survey that included:
+Demographics.
+We collect the following information from each participant: Ethnicity, Gender, Age, Highest level of education attained (some high school, high school, some college, associate’s degree, bachelor’s degree, master’s degree, doctoral degree, professional degree, other), Years of work experience, Annual Income (0-\$25k, \$25.001k-\$50k, \$50.001k-75k, \$75.001k-\$100k, \$100.001k-\$150k, \$150k+), and elicitation of sets of O*NET job skills that participants used in their occupation (reading comprehension, active listening, writing, speaking, critical thinking, social perceptiveness, coordination, instructing, programming, judgment and decision making, systems evaluations, science, active learning, learning strategies, monitoring, complex problem analysis, technology design, troubleshooting, quality control analysis, systems analysis).
+Opinions and Skills.
+The opinions and skills we collect information on are: computer programming proficiency and usage frequency (self-reported), Structured and creative writing proficiency and usage frequency (self-reported), Generative AI tool proficiency and usage frequency (self-reported), Attitudes towards net social impact of Generative AI (self-reported), Advice for (hypothetical) future participants on how to perform well on the task.
+C.2 Prompt Data
+For each prompt, we recorded the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
+C.3 Image Data
+We collected three different sets of images:
+  1. 1.
+    The participant-facing images (image replication experiment and logo generation experiment): These images were created directly by the participants, generated by the model to which they were randomly assigned using the prompt they submitted.
+For the image replication experiment, these were generated from December 12 to December 19, 2023.
+For the logo generation experiment, these were generated from June 30 to July 1, 2025.
+  2. 2.
+    Post-hoc resampled images (image replication experiment only): For any given prompt, the output of the text-to-image model is stochastic.
+To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the image replication experiment concluded.
+We provide full details on this procedure in Section E.
+These images were generated from December 26, 2023 to January 27, 2024.
+These images are not used for analyses presented in the main text, but were used for other pre-registered analyses.
+These additional analyses are discussed in Section I.
+  3. 3.
+    Post-hoc replayed images (image replication experiment and logo generation experiment): For both experiments, to decompose our overall effects into model and prompting effects, we generated “counterfactual images” for each prompt written under the DALL-E 2 and DALL-E 3 (Verbatim) treatments.
+In other words, we submitted all prompts written under both the DALL-E 2 and DALL-E 3 (Verbatim) treatments to both the DALL-E 2 and DALL-E 3 (Verbatim) endpoints.
+    For the image replication experiment, similar to the resampling procedure outlined above, we generated 10 images per prompt per model: we generated a single replay for each prompt-model pair from March 16 to 18, 2024, and then, to increase power, generated the replications for these replay images from June 14 to 27, 2024.
+This replay process produced a total of 20 images per prompt—10 under the original model, 10 under the counterfactual model.
+We re-submitted prompts to their original model to account for potential model drift, as this exploratory analysis was conducted multiple months after our initial data collection.
+For consistency, this replay data is used throughout the main text of our paper.
+    For the logo generation experiment, we generated a single replay for each prompt-model pair.
+This took place from July 3 to 9, 2025.
+In this case, given the negligible gap in time between participant data collection and participant image replay, there was no need to account for potential model drift as we did in the image replication experiment.
+C.4 Image Replication Sample Construction
+The image replication experiment sample analyzed in the main text was constructed as follows.
+The initial “raw” dataset collected during the experiment is comprised of 24,672 rows of raw prompt data (one prompt per row) generated by 2,059 participants.
+We first removed rows with blank prompt entries, invalid prolific IDs, and unsuccessful attempts (logging errors).
+These exclusion criteria were pre-registered.
+This left us with 2,029 participants and 24,123 prompts.
+We next removed participants from our sample if they failed to submit at least 10 prompts.
+This exclusion criterion was pre-registered and also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” This left us with 1,914 participants and 23,566 prompts.
+Although participants were allowed to submit as many prompts as they desired in the 25-minute time span, we limited all analyses to each participant’s first 10 prompts—the minimum required to receive payment for the task.
+This exclusion criteria was not pre-registered, and is noted in the list of deviations from pre-registration in Section I.I.2.
+We restrict our analysis dataset in this way because participants who chose to submit more than 10 prompts may have been systematically different than those who did not.
+Excluding any prompt beyond the 10th attempt allows us to alleviate selection bias concerns.
+This left us with 1,914 participants and 19,140 prompts.
+We next removed participants who submitted the same prompt at least five times in a row at any point during the task.
+This exclusion criterion was pre-registered.
+To avoid reward hacking, we did not specify to participants the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants and 18,990 prompts.
+We then removed participants who failed to complete the Qualtrics survey.
+This exclusion criteria was pre-registered.
+This left us with 1,893 participants and 18,930 prompts.
+We also removed prompts from our dataset according to a number of post-hoc, non-pre-registered exclusion criteria to ensure data quality and avoid selection bias.
+These criteria include the following:
+  1. 1.
+    Prompts sometimes trigger errors in OpenAI’s safety system because they contain language that might be deemed unsafe under OpenAI’s policies.
+The specific language that triggers these errors is constantly changing and not available publicly.
+If a prompt triggered a safety error during the replication or replay process, we re-submitted the prompt up to 50 times or until the 10 original arm replications/replay samples had been collected.
+We removed prompts if they failed to generate 10 replications on the original model or 10 replay samples under the counterfactual model during the replication/replay process.
+This affected 343 prompts across all three treatment arms.
+  2. 2.
+    Due to rare database synchronization latency issues, some prompts were assigned duplicate attempt numbers by the MongoDB database that we used to collect our data.
+This data collection error led to issues in the data analysis process.
+Thus, we excluded prompts with duplicate attempt numbers.
+This affected 34 prompts across all three treatment arms, and 20 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms, approximately 0.1% of the original data.
+In total, these post-hoc, non-pre-registered exclusion criteria affected 370 prompts across all remaining participants. Our final sample thus included 1,893 participants and 18,560 prompts.
+C.5 Logo Generation Sample Construction
+The logo generation experiment sample analyzed in the main text was constructed as follows.
+The initial “raw” dataset collected during the experiment is comprised of 23,159 rows of raw prompt data (one prompt per row) generated by 2,067 participants.
+We first removed rows with blank prompt entries, invalid prolific IDs, and unsuccessful attempts (logging errors).
+These exclusion criteria were pre-registered.
+This left us with 2,011 participants and 22,721 prompts.
+We next removed participants from our sample if they failed to submit at least 10 prompts.
+This exclusion criterion was pre-registered and also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” This left us with 1,924 participants and 22,312 prompts.
+Although participants were allowed to submit as many prompts as they desired in the 25-minute time span, we limited all analyses to each participant’s first 10 prompts—the minimum required to receive payment for the task.
+This exclusion criteria was not pre-registered, and is noted in the list of deviations from pre-registration in Section I.I.2.
+We restrict our analysis dataset in this way because participants who chose to submit more than 10 prompts may have been systematically different than those who did not.
+Excluding any prompt beyond the 10th attempt allows us to alleviate selection bias concerns.
+This left us with 1,924 participants and 19,240 prompts.
+We next removed participants who submitted the same prompt at least five times in a row at any point during the task.
+This exclusion criterion was pre-registered.
+To avoid reward hacking, we did not specify to participants the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,892 participants and 18,920 prompts.
+We additionally removed a small number of participants and their prompts for whom we discovered post-hoc that the treatment condition to which they were exposed in the experimental interface did not match up with their assigned task label in the MongoDB database that we used to collect our data.
+This exclusion criterion was not pre-registered.
+This left us with 1,864 participants and 18,640 prompts.
+We then removed participants who failed to complete the Qualtrics survey.
+This exclusion criteria was pre-registered.
+This left us with 1,857 participants and 18,570 prompts.
+We also removed prompts from our dataset according to a number of post-hoc, non-pre-registered exclusion criteria to ensure data quality and avoid selection bias.
+These criteria include the following:
+  1. 1.
+    Prompts sometimes trigger errors in OpenAI’s safety system because they contain language that might be deemed unsafe under OpenAI’s policies.
+The specific language that triggers these errors is constantly changing and not available publicly.
+If a prompt triggered a safety error during the replay process, we re-submitted the prompt up to 3 times.
+We removed prompts if they failed the complete set of replays under the counterfactual model during the replay process.
+  2. 2.
+    Due to rare database synchronization latency issues, some prompts were assigned duplicate attempt numbers by the MongoDB database that we used to collect our data.
+This data collection error led to issues in the data analysis process.
+Thus, we excluded prompts with duplicate attempt numbers.
+In total, these post-hoc, non-pre-registered exclusion criteria affected 145 prompts across all remaining participants. Our final sample thus included 1,857 participants and 18,425 prompts.
+C.5.1 Removing “Off-Topic” Submissions (Image Replication Experiment Only)
+While analyzing our data, we found that our sample contained a number of “off-topic” prompts that did not seem related to the task.
+As a robustness check on our main results, we used the following process to systematically identify and remove “off-topic” prompts.
+First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model.
+We then calculated the mean embedding for each target image.
+Next, we calculated the Euclidean distance between each prompt’s embedding vector and the mean embedding vector for prompts corresponding to the focal prompt’s assigned target image.
+Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector.
+This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms.
+All of our main text results are robust to the exclusion of these “off-topic” prompts.
+Appendix D Dependent Variables
+D.1 Image Similarity
+We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (fu2023dreamsim).
+In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers.
+Our results are qualitatively and quantitatively similar using DreamSim instead.
+CLIP Embedding Cosine Similarity.
+To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (radford2021learning) using the version of CLIP available on Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image.
+Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text.
+We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
+DreamSim.
+DreamSim is an image similarity measure proposed recently by (fu2023dreamsim).
+The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity.
+Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score.
+After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.
+We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure.
+We present the results obtained when conducting our main text analyses using DreamSim in Section H.1.
+D.2 Bradley-Terry Strength
+To generate a Bradley-Terry strength value for each original and replayed image in the logo generation experiment, we first collected all images and enumerated all pairwise comparisons within each of the 15 tasks.
+We then randomly sampled, without replacement, 20% of the pairwise comparisons within each task to feed into our LLM-based evaluation approach, resulting in approximately 20 million pairwise comparisons.
+For the LLM-based evaluation, for each pairwise comparison, we called the response endpoint of the official OpenAI Python library (release version 1.97.0) with the following prompt, with the task brief and the two images appended after:
+“You are provided a logo design request from a client, as well as two images (Image A and Image B).
+Please evaluate both images carefully and determine which one better fulfills the client’s specific requirements.
+Please respond with only the letter label of the image, either ‘A’ or ‘B’.
+For all requests, we used the multi-modal gpt-4.1-nano model and set the temperature parameter to be 0.
+The images were first base64-encoded before being included in the request input.
+We further randomized the order in which the images were presented in the prompt to avoid the effects of prompt position bias.
+To convert the resulting pairwise preferences into continuous strength values, we used the Bradley–Terry model.
+In this model, each item $i$ is assumed to have an associated latent strength parameter $\beta_{i}\in\mathbb{R}$, and the probability that item $i$ is preferred to item $j$ is given by
+$\Pr(i\succ j)\;=\;\frac{\exp(\beta_{i})}{\exp(\beta_{i})+\exp(\beta_{j})}.$
+Our goal is to estimate the set of strength parameters $\{\beta_{i}\}$ for all images.
+For each task, once we collected the LLM-generated pairwise comparison outcomes and converted them into win–loss matrices, we estimated the $\beta_{i}$ parameters by fitting the Bradley–Terry model via maximum likelihood using a standard iterative procedure.
+D.3 Prompt Length
+We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt.
+In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9954$, 95% CI: [0.99528, 0.99560]).
+D.4 Embedding-based Prompt Similarity
+We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity.
+Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (neelakantan2022text).
+The two similarity measures are defined as follows:
+Successive similarity:
+The successive similarity ($ss$) is a measure of the similarity of a participant’s prompt to their immediately preceding prompt.
+We define the successive similarity of a prompt $p_{i,n}$ written by user $i$ to the their immediately preceding prompt $p_{i,n-1}$ as:
+[…]
+where $\mathbf{E(p_{i,n})}$ is the vector embedding representation of participant $i$’s $n^{\text{th}}$ prompt, $p_{i,n}$.
+This measure starts with participant $i$’s 2nd attempt, as the calculation requires a previous attempt.
+Aggregate similarity:
+The aggregate similarity ($as$) is a measure of how dispersed each user’s prompts are around their “average prompt” (calculated by taking the element-wise average of all prompt embeddings produced by the user).
+We define the aggregate similarity for the 10 prompts written by a given user as:
+$as_{i}=\frac{1}{10}\sum_{n=1}^{10}\lVert\mathbf{E(p_{i,n})}-\overline{\mathbf{E(p_{i,n})}}\rVert_{2}^{2},$ (8)
+where $\mathbf{E(p_{i,n})}$ is again the vector embedding representation of participant $i$’s $n^{\text{th}}$ prompt, $p_{i,n}$, and $\overline{\mathbf{E(p_{i,n})}}$ is the element-wise mean of all 10 of participant $i$’s prompts.
+D.5 Successive Prompt Token Sort Ratio
+Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$.
+TSR is a fuzzy string-matching technique (singla2012string) that provides a continuous measure of how similar two strings are.
+D.6 Successive Prompt ‘Contains Previous Prompt’ Dummy
+Starting with each participant’s second prompt, we record whether each prompt $p_{i,n}$ contains the immediately preceding prompt $p_{i,n-1}$ as an exact substring.
+D.7 Prompt Composition
+We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt.
+SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation posuniversaldep.
+These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs.
+The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
+D.8 Strategic Shifts
+In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting.
+In order to do so, we adapt a method proposed in (torricelli2023role) (because they are conducting research in a different context, torricelli2023role refer to these shifts as “topical transitions” as opposed to “strategic shifts”).
+To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
+$MCS_{t}=\frac{2}{P_{t}(P_{t}-1)}\sum_{a=1}^{P_{t}}\sum_{b=a+1}^{P_{t}}\textrm{CosineSim}(\mathbf{E(p_{a,t})},\mathbf{E(p_{b,t})}).$ (9)
+where $P_{t}$ is the total number of prompts submitted in response to a given target image, and $a$ and $b$ are indices representing individual prompts for that target.
+We then label any given prompt as a strategic shift ($SS$) if the cosine similarity of its embedding vector with that of the previous prompt is lower than this target-image-level mean:
+[…]
+It is worth noting that torricelli2023role uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift.
+We use the task-level mean because in our setting, it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
+Appendix E Methods
+This appendix provides additional methodological details to supplement the analyses presented in the main text.
+E.1 Stratification
+The results shown in the main text and in the supplementary information are typically stratified by reference image and iteration.
+In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level).
+The exact stratification for each finding is indicated in section F.
+To stratify our results, we calculate a weighted average across j=1,,Jj=1,...,J cells defined by our stratification variables:
+$\overline{Y}_{strat}=\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}$
+To calculate the variance (and standard error) of this sample mean, we apply the following:
+$\widehat{\mathrm{Var}}(\overline{Y}_{strat})=\widehat{\mathrm{Var}}\left(\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}\right)=\sum_{j=1}^{J}\left(\frac{N_{j}}{N}\right)^{2}\frac{s_{j}^{2}}{N_{j}}$
+$N_{j}$ is the population size of stratum $j$.
+NN is the total population size across all strata.
+$\bar{Y}_{j}$ is the sample mean for stratum $j$.
+$J$ is the total number of strata.
+$s_{j}$ is the sample standard deviation of stratum $j$.
+Therefore, $s_{j}^{2}$ is sample variance stratum $j$.
+E.2 Z-Scoring
+Our analysis found statistically significant differences in performance variability across the 15 target images used in our experiments, as discussed in Section I.
+The main text also demonstrated that performance increases across successive attempts.
+To ensure our results are not driven by this image-level or attempt-level variation, we replicated all analyses using the within-image-attempt Z-score of CLIP-cosine similarity for each image produced by participants.
+Formally, this is:
+$Z(\text{Sim}_{i,n,t})=\frac{\text{Sim}_{i,n,t}-\text{Mean}_{n,t}\left(\text{Sim}_{i,n,t}\right)}{\text{SD}_{n,t}\left(\text{Sim}_{i,n,t}\right)}$ (11)
+where $\textrm{Sim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$.
+The mean and standard deviation are computed for each image-attempt pair, but across both the DALL-E 2 and DALL-E 3 treatment arms.
+We also applied this rescaling to test the robustness of our DreamSim-based analyses.
+Nearly all robustness analyses reported here use the Z-score scaled measure of performance within each image-attempt set.
+The only exception is analyses that examine improvements and prompts across attempts (e.g., Figure 1B), where Z-scores are computed only within each target image and across all attempts for that image.
+We use the same Z-scoring technique for robustness checks of the logo generation task.
+Instead of $t$ indexing the target image in these cases, it indexes the brief/vignette the participant was assigned to.
+E.3 Accounting for Model Stochasticity (Image Replication Experiment Only)
+Generative AI models produce stochastic outputs in response to a given prompt.
+This stochasticity is controlled by a parameter called temperature, which could not be modified using the DALL-E API at the time of our experiment.
+To account for this model stochasticity, we generated 10 images for each prompt submitted by participants across all treatment arms.
+We then calculated the similarity between each replication and its corresponding target image, and computed an ”expected” CLIP cosine similarity and DreamSim score for each prompt by averaging across these samples.
+Using these replicated images, we also calculated the standard deviation of cosine similarity induced by this stochasticity.
+With the expected cosine similarity and its standard deviation per prompt, we computed a normalized Z-score for the observed image relative to its replication distribution.
+This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and is used for further analysis in section I.
+We generated these additional samples both for the original prompts on their assigned treatment arms and by replaying prompts on the counterfactual arms, as introduced in Figure 2 in the main text.
+Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling.
+As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts.
+This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the “replaying” procedure (Section C.3).
+Appendix F Main Text Analyses
+This section provides detailed methodological information about the analyses presented in the main text.
+F.1 Image Replication Task
+F.1.1 Task Performance and ATEs
+The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations).
+It displays the average cosine similarity score stratified by the reference image.
+A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms.
+This may be due to participants’ initial misunderstanding of the model’s “memoryless” nature.
+Participants failed to recognize that context from previous prompts was not carried over to new iterations.
+We observed numerous prompts in the second iteration across users that explicitly referenced the first prompt, a behavior that rarely occurred in subsequent attempts.
+Alternatively, it may also be indicative of other early exploration by participants.
+However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
+The bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane.
+[^12] In Section F, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified.
+To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
+$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (12)
+$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
+The coefficient estimates generated by this model are: $\hat{\beta_{1}}=0.0011\;(0.0003)$, $p=0.0004$; $\hat{\beta_{2}}=0.0120\;(0.0037)$, $p=0.0013$; $\hat{\beta_{3}}=0.0010\;(0.0004)$, $p=0.0227$.
+The overall ATEs that we report between different pairs of treatment arms (DALL-E 2, DALL-E 3, and DALL-E 3 with revisions) in the main text are estimated from a two-way fixed effect (iteration and target image) model per each pair.
+Standard errors are cluster robust at the participant level.
+F.1.2 Prompt Characteristics
+Figure 1C compares the prompt length and prompt similarity of the two models.
+To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section C.4.
+The prompt length is the average number of words per model and iteration stratified by the reference image.
+The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section C.D.4 for details on similarity calculations).
+The color scale in Figure 1C shows the stratified average similarity to the previous prompt across all users per each model.
+We find that DALL-E 3 users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
+Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the DALL-E 2 participants.
+For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section C.D.4, as the dependent variable.
+When we average across all participants by model, we find that the average distance of prompts written by DALL-E 3 participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$).
+Standard errors are cluster-robust at the participant level.
+F.1.3 ATE Decomposition
+Figure 2 in the main text decomposes the ATE into the model and prompting effects.
+This decomposition is conceptually similar to a simple mediation analysis, with an important difference being that we can observe counterfactual outcomes (e.g., prompting DALL-E 3 as if it is DALL-E 2).
+This is not typically the case in mediation analysis, and makes causal identification rely on fewer assumptions.
+To obtain counterfactual outcomes, we fed or “replayed” the participant prompts when interacting with one model (e.g., inferior) on another model (e.g., superior).
+The notation (prompt, model) specifies which treatment arm the prompts were written under and which model was used in the replay.
+For example, (2,3) indicates replaying prompts written under DALL-E 2 on DALL-E 3.
+To be clear, (2,2) and (3,3) correspond to the original observed treatment arms, while (2,3) and (3,2) are the counterfactual outcomes of interest.
+[^13] To avoid problems with model drift for the image replication experiment, we regenerated images for all four possible combinations at the same time and used these images for all analyses in the main text.
+The left-most point in Figure 2 corresponds to the average CLIP cosine similarity to the target image of (2,2).
+To make the interpretation of the results clearer, we have subtracted this quantity from all average quality scores and added a dashed line throughout.
+The second point from the left corresponds to average similarity to the target of (2,3), the third point from the left to (3,3), and the rightmost points to (3,2).
+All average similarity scores are stratified by iteration and reference image, and the standard errors are bootstrapped and cluster-robust at the participant level.
+The model effect, as shown by the red braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (2,2).
+In the terminology of mediation analysis, the model effect would be referred to as the direct effect.
+The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (3,3) relative to (2,3).
+In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect.
+We can also test the difference in average quality between (3,3) and (3,2), as well as the difference between (3,2) and (2,2).
+Both of these differences are visible in Figure 2; the second is small and not statistically significant.
+The point estimates for the total, model, and prompting effects, along with their significance values reported in the main text, are obtained from two-way fixed-effects regressions that compare average outcomes across the two relevant conditions.
+Specifically, each effect is estimated using a model with iteration and target-image fixed effects, where an indicator for the effect of interest serves as the main independent variable:
+$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{effect}+\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$ (13)
+where $\beta_{1}$ is the coefficient on the effect type in question (i.e., model or prompting).
+To estimate the different effects, we simply use the above model and filter the data as appropriate.
+For example, to estimate the ATE, the data contains all (2,2) and (3,3) scores, and in this case effect=1 for observations in (3,3) group.
+Similarly, to estimate the direct or model effect, the data contains all (2,2) and (2,3) scores and effect=1 for observations in (2,3) group.
+Finally, to estimate the indirect or prompting effect, the data contains all (2,3) and (3,3) scores and effect=1 for observations in (3,3) group.
+The standard errors for each estimated model are cluster robust at the participant level, and p-values are adjusted accordingly.
+F.1.4 Distributional Impacts
+Table 1 in the main text demonstrates how total, model, and prompting QTEs vary across different score quantile levels.
+The total effect compares the cosine similarity of outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 2.
+The model effect compares outputs from DALL-E 2 users replayed on DALL-E 3 against the same users replayed on DALL-E 2.
+The prompting effect compares outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 3.
+To assess user skill, users within each target image, iteration and replay scenario are divided into 50 equally sized brackets based on their performance, assigning each bracket a rank corresponding to performance percentiles.
+For instance, when evaluating DALL-E 2 users replayed on DALL-E 2, each user’s skill level is determined by their percentile rank among all DALL-E 2 users replayed on DALL-E 2.
+Conversely, when the same user is replayed on DALL-E 3 (to estimate the model effect or to serve as a baseline in estimating the prompting effect), their performance percentile is determined by their rank relative to all DALL-E 2 users replayed on DALL-E 3, per each iteration and target image.
+After assigning performance percentile, we estimate the following linear model with two-way fixed effects (iteration and target image) for each effect type (model, prompting, and total):
+$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\text{effect}+\beta_{2}\text{effect}\times\text{Performance Percentile}_{i,n,t}\,+$ (14)
+$\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$
+where $\beta_{1}$ represents the coefficient associated with the effect type being analyzed (model, prompting, or total), and $\beta_{2}$ captures the interaction between the effect type and user skill level.
+User skill is measured as the user’s rank (binned) within the same iteration, target image, and replay scenario.
+Standard errors are calculated using cluster-robust methods at the participant level.
+This methodology closely follows the procedure described earlier in section F.1.3, with the additional step of incorporating an interaction term to evaluate skill-level heterogeneity.
+F.2 Logo Generation Task
+F.2.1 Task Performance and ATEs: Logo Generation Task
+The analytic framework for the logo generation task mirrors that of the image replication task (Section F.1.1).
+As before, we estimate Equation 12, treating iteration as a numeric variable and clustering standard errors at the participant level.
+The dependent variable here, however, is the average Bradley–Terry ($\beta$) strength derived from the pairwise comparison framework described in Section D.2.
+Figure 3B reports the estimated performance trajectory across attempts.
+Consistent with the image replication task, we again observe a pronounced dip during the second iteration across all model conditions, potentially reflecting participants’ temporary misunderstanding of the system’s memoryless nature.
+As before, performance quickly rebounds from the third iteration onward once participants realize that previous prompts are not retained.
+The ATE estimates in the main text are from a two-way fixed effect model (iteration and logo taks) with cluster robust standard errors at the participant level.
+For example, the overall average treatment effect (ATE) between DALL-E 3 and DALL-E 2 participants corresponds to a $0.8075$ increase in Bradley–Terry $\beta$ (95% CI: [0.7120, 0.9030], $p<10^{-10}$), equivalent to roughly $0.57$ standard deviations of improvement.
+Similarly, the estimate comparing DALL-E 3 with revision and DALL-E 2 is 0.834 (95% CI: [0.7358, 0.9316], $p<10^{-10}$).
+Equation 12 accounts for variation in ATE over different attempts.
+The estimated coefficients from that model for the logo generation task are: $\hat{\beta_{1}}=0.00493;(0.00628)$, $p=0.432$; $\hat{\beta_{2}}=0.747;(0.0598)$, $p<10^{-10}$; $\hat{\beta_{3}}=0.0135;(0.00819)$, $p=0.099$.
+The above estimates suggest a similar pattern: the gap between DALL-E 3 and DALL-E 2 slightly increased over the attempts, although this divergence in the logo task is weaker and lacks sufficient statistical power.
+F.2.2 Prompt Characteristics: Logo Generation Task
+Prompt-level patterns in the logo generation task follow the same structure for image replication and the result is shown in Figure 3C.
+Participants assigned to DALL-E 3 again produced significantly longer prompts (approximately 24.7% longer on average) and exhibited higher successive prompt similarity relative to DALL-E 2 users.
+As in the image replication task, superior-model users produced prompts that were more similar to their previous attempts.
+On average, successive prompt cosine similarity was $\beta=0.03689$ higher for DALL-E3 users ($p=0.00392$).
+Prompt sequences were also more internally consistent for superior-model participants (see section D.4).
+Their prompts lay $\beta=0.03642$ closer to their own centroid in embedding space than those of DALL-E 2 users ($p<10^{-6}$), using cluster-robust standard errors.
+Using the same analysis approach as in the image replication experiment, we also find that the proportion of nouns and adjectives is statistically indistinguishable across model conditions (43% for DALL-E 3 vs.
+42% for DALL-E 2; $p=0.12$)
+F.2.3 ATE Decomposition: Logo Generation Task
+The decomposition procedure for the logo generation task is identical to that described in Section F.1.3.
+We again estimate Equation 13 with two-way fixed effects for iteration and target item, cluster-robust standard errors at the participant level, and the effect-type indicator defined in the same way.
+The only methodological difference is the outcome: instead of CLIP cosine similarity, we use the Bradley–Terry $\beta$ strength assigned to each generated logo.
+As before, replaying participant prompts across models yields estimates of the total, model, and prompting effects.
+F.2.4 Distributional Impacts: Logo Generation Task
+We replicate the analysis of Section F.1.4 to assess how total, model, and prompting effects vary across the performance distribution.
+The same two-way fixed effects model in Equation 14 is used, substituting the Bradley–Terry $\beta$ scores in place of cosine similarity outcome.
+Table 2 in the main text presents the corresponding regression coefficients.
+Appendix G Example Prompts
+G.1 Image Replication
+Table A.4 presents the complete set of prompts and their corresponding generated images that were used in our analysis shown in Figure 1.
 The table displays three example target images (woman with shopping basket, abstract painting with cross, and online shopping interface), with three different participant-generated versions of each.
@@
 The prompts are presented verbatim from participant submissions, preserving all original formatting and punctuation (or lack thereof).
-Table 2: Participant prompts for the images provided in Figure 1
+Table A.4: Participant prompts for the images provided in Figure 1
 Image Prompt
@@
 [Uncaptioned image] “at the store a man using a touch screen monitor to shop for a shirt and a woman walking by with a store cart full of gift”
-Appendix C Measurement and Variables
-This appendix provides detailed information about the data collection, measurement, and variables used in our analyses, expanding on the overview provided in the main text.
-C.1 Survey Data
-We collected additional participant information via a Qualtrics survey that included:
-  • Demographics: Ethnicity, Gender, Age, Highest level of education attained (some high school, high school, some college, associate’s degree, bachelor’s degree, master’s degree, doctoral degree, professional degree, other), Years of work experience, Annual Income (0-\$25k, \$25.001k-\$50k, \$50.001k-75k, \$75.001k-\$100k, \$100.001k-\$150k, \$150k+), and elicitation of sets of O*NET job skills that participants used in their occupation (reading comprehension, active listening, writing, speaking, critical thinking, social perceptiveness, coordination, instructing, programming, judgment and decision making, systems evaluations, science, active learning, learning strategies, monitoring, complex problem analysis, technology design, troubleshooting, quality control analysis, systems analysis).
-  • Opinions and Skills: Computer programming proficiency and usage frequency (self-reported), Structured and creative writing proficiency and usage frequency (self-reported), Generative AI tool proficiency and usage frequency (self-reported), Attitudes towards net social impact of Generative AI (self-reported), Advice for (hypothetical) future participants on how to perform well on the task.
-C.2 Prompt Data
-For each prompt, we recorded the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
-C.3 Image Data
-We collected three different sets of images:
-  1. 1.
-    The participant-facing images (OpenAI API endpoint): The image shown to the participant during the experiment, generated by the model they were assigned to using the prompt they submitted.
-These images were generated from December 12-19, 2023.
-  2. 2.
-    Post-hoc resampled images (Azure OpenAI endpoint): For any given prompt, the output of the text-to-image model is stochastic.
-To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the experiment concluded.
-We provide full details on this procedure in Section D.
-These images were generated from December 26, 2023 - January 27, 2024.
-These images are not used for analyses presented in the main text, but were used for other pre-registered analyses.
-These additional analyses are discussed in Section G.
-  3. 3.
-    Post-hoc replayed images (Azure OpenAI endpoint): To decompose our overall effects into model and prompting effects, we generated “counterfactual images” for each prompt written under the DALL-E 2 and DALL-E 3 (Verbatim) treatments.
-In other words, we submitted all prompts written under both the DALL-E 2 and DALL-E 3 (Verbatim) treatments to both the DALL-E 2 and DALL-E 3 (Verbatim) endpoints.
-Similarly to the resampling procedure outlined above, we generated 10 images per prompt per model: we generated a single replay for each prompt-model pair from March 16-18, 2024, and then, to increase power, generated the replications for these replay images from June 14-27, 2024.
-This replay process produced a total of 20 images per prompt—10 under the original model, 10 under the counterfactual model.
-We re-submitted prompts to their original model to account for potential model drift, as this exploratory analysis was conducted multiple months after our initial data collection.
-For consistency, this replay data is used throughout the main text of our paper.
-C.4 Sample Construction
-The sample analyzed in the main text was constructed as follows:
-  • The initial “raw” dataset collected during the experiment is comprised of 24,672 rows of raw prompt data (one prompt per row) generated by 2,059 participants.
-  • We first removed rows with blank prompt entries, invalid prolific IDs, and unsuccessful attempts (logging errors).
-These exclusion criteria were pre-registered.
-This left us with 2,029 participants and 24,123 prompts.
-  • We next removed participants from our sample if they failed to submit at least 10 prompts or if a participant submitted the same prompt at least five times in a row at any point during the task.
-Both of these exclusion criteria were pre-registered.
-These exclusion criteria were also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” To avoid reward hacking, we did not specify the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants.
-[…]
-  • We next removed participants who failed to complete the Qualtrics survey. This exclusion criteria was pre-registered. This left us with 1,893 participants and 18,930 prompts.
-  • We also removed prompts from our dataset according to a number of post-hoc, non-pre-registered exclusion criteria to ensure data quality and avoid selection bias.
-If a prompt had any of the following flags, it was removed from the sample:
-    • Prompts sometimes trigger errors in OpenAI’s safety system because they contain language that might be deemed unsafe under OpenAI’s policies.
-The specific language that triggers these errors is constantly changing and not available publicly.
-If a prompt triggered a safety error during the replication or replay process, we re-submitted the prompt up to 50 times or until the 10 original arm replications/replay samples had been collected.
-We removed prompts if they failed to generate 10 replications on the original model or 10 replay samples under the counterfactual model during the replication/replay process.
-This affected 305 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms.
-It did not affect any DALL-E 3 (Revised) prompts, as we did not conduct replay analysis with the prompts from this treatment arm.
-    • Due to rare latency issues, some prompts were assigned duplicate attempt numbers by the MongoDB database that we used to collect our data.
-This data collection error led to issues in the data analysis process.
-Thus, we excluded prompts with duplicate attempt numbers.
-This affected 34 prompts across all three treatment arms, and 20 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms, approximately 0.1% of the original data.
-  • Our final sample included 1,893 participants and 18,560 prompts.
-C.4.1 “Off-Topic” Robustness Check
-While analyzing our data, we found that our sample contained a number of “off-topic” prompts that did not seem related to the task.
-As a robustness check on our main results, we used the following process to systematically identify and remove “off-topic” prompts.
-First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model.
-We then calculated the mean embedding for each target image.
-Next, we calculated the Euclidean distance between each prompt’s embedding vector and the mean embedding vector for prompts corresponding to the focal prompt’s assigned target image.
-Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector.
-This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms.
-All of our main text results are robust to the exclusion of these “off-topic” prompts.
-C.5 Dependent Variables
-C.5.1 Image Similarity
-We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al.
-2023).
-In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers.
-Our results are qualitatively and quantitatively similar using DreamSim instead.
-  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al.
-2021) from Hugging Face (OpenAI 2024) for each participant-generated image and for each target image.
-Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text.
-We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
-  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al.
-2023).
-The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity.
-Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score.
-After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.
-We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure.
-We present the results obtained when conducting our main text analyses using DreamSim in Section F.F.1.
-C.5.2 Prompt Length
-We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt.
-In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9954$, 95% CI: [0.99528, 0.99560]).
-C.5.3 Embedding-based Prompt Similarity
-We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity.
-Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al.
-2022).
-The two similarity measures are defined as follows:
-  • Successive similarity: The successive similarity ($ss$) is a measure of the similarity of a participant’s prompt to their immediately preceding prompt.
-We define the successive similarity of a prompt $p_{i,n}$ written by user $i$ to the their immediately preceding prompt $p_{i,n-1}$ as:
-    $ss_{i,n,n-1}=\frac{\mathbf{E(p_{i,n})}\cdot\mathbf{E(p_{i,n-1})}}{||\mathbf{E(p_{i,n})}||\,||\mathbf{E(p_{i,n-1})}||},$ (6)
-[…]
-  • Aggregate similarity: The aggregate similarity ($as$) is a measure of how dispersed each user’s prompts are around their “average prompt” (calculated by taking the element-wise average of all prompt embeddings produced by the user).
-We define the aggregate similarity for the 10 prompts written by a given user as:
-    $as_{i}=\frac{1}{10}\sum_{n=1}^{10}\lVert\mathbf{E(p_{i,n})}-\overline{\mathbf{E(p_{i,n})}}\rVert_{2}^{2},$ (7)
-    where $\mathbf{E(p_{i,n})}$ is again the vector embedding representation of participant $i$’s $n^{\text{th}}$ prompt, $p_{i,n}$, and $\overline{\mathbf{E(p_{i,n})}}$ is the element-wise mean of all 10 of participant $i$’s prompts.
-C.5.4 Successive Prompt Token Sort Ratio
-Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$.
-TSR is a fuzzy string-matching technique (Singla and Garg 2012) that provides a continuous measure of how similar two strings are.
-C.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
-Starting with each participant’s second prompt, we record whether each prompt $p_{i,n}$ contains the immediately preceding prompt $p_{i,n-1}$ as an exact substring.
-C.5.6 Prompt Composition
-We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt.
-SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal Dependencies Project (2024).
-These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs.
-The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
-C.5.7 Strategic Shifts
-In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting.
-In order to do so, we adapt a method proposed in (Torricelli et al.
-2023) (because they are conducting research in a different context, (Torricelli et al.
-2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”).
-To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
-$MCS_{t}=\frac{2}{P_{t}(P_{t}-1)}\sum_{a=1}^{P_{t}}\sum_{b=a+1}^{P_{t}-1}\textrm{CosineSim}(\mathbf{E(p_{a,t})},\mathbf{E(p_{b,t})}).$ (8)
-where $P_{t}$ is the total number of prompts submitted in response to a given target image, and $a$ and $b$ are indices representing individual prompts for that target.
-We then label any given prompt as a strategic shift ($SS$) if the cosine similarity of its embedding vector with that of the previous prompt is lower than this target-image-level mean:
-$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}\big{(}\mathbf{E(p}_{i,n,t}),\mathbf{E(p}_{i-1,n,t})\big{)}\\ &\quad<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (9)
-It is worth noting that Torricelli et al.
-(2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift.
-We instead use the task-level mean because in our setting, as it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
-Appendix D Methods
-This appendix provides additional methodological details to supplement the analyses presented in the main text.
-D.1 Stratification
-The results shown in the main text and in the supplementary information are typically stratified by reference image and iteration.
-In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level).
-The exact stratification for each finding is indicated in section E.
-To stratify our results, we calculate a weighted average across $j=1,...,J$ cells defined by our stratification variables:
-$\overline{Y}_{strat}=\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}$
-To calculate the variance (and standard error) of this sample mean, we apply the following:
-$\widehat{\mathrm{Var}}(\overline{Y}_{strat})=\widehat{\mathrm{Var}}\left(\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}\right)=\sum_{j=1}^{J}\left(\frac{N_{j}}{N}\right)^{2}\frac{s_{j}^{2}}{N_{j}}$
-where:
-  • $N_{j}$ is the population size of stratum $j$.
-  • $N$ is the total population size across all strata.
-  • $\bar{Y}_{j}$ is the sample mean for stratum $j$.
-  • $J$ is the total number of strata.
-  • $s_{j}$ is the sample standard deviation of stratum $j$. Therefore, $s_{j}^{2}$ is sample variance stratum $j$.
-D.2 Z-Scoring
-Our analysis found statistically significant differences in performance variability across the 15 target images used in our experiments, as discussed in Section G.
-The main text also demonstrated that performance increases across successive attempts.
-To ensure our results are not driven by this image-level or attempt-level variation, we replicated all analyses using the within-image-attempt Z-score of CLIP-cosine similarity for each image produced by participants.
-Formally, this is:
-$Z(\text{Sim}_{i,n,t})=\frac{\text{Sim}_{i,n,t}-\text{Mean}_{n,t}\left(\text{Sim}_{i,n,t}\right)}{\text{SD}_{n,t}\left(\text{Sim}_{i,n,t}\right)}$ (10)
-where $\textrm{Sim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$.
-The mean and standard deviation are computed for each image-attempt pair, but across both the DALL-E 2 and DALL-E 3 treatment arms.
-We also applied this rescaling to test the robustness of our DreamSim-based analyses.
-Nearly all robustness analyses reported here use the Z-score scaled measure of performance within each image-attempt set.
-The only exception is analyses that examine improvements and prompts across attempts (e.g., Figure 1B), where Z-scores are computed only within each target image and across all attempts for that image.
-D.3 Accounting for Model Stochasticity
-Generative AI models produce stochastic outputs in response to a given prompt.
-This stochasticity is controlled by a parameter called temperature, which could not be modified using the DALL-E API at the time of our experiment.
-To account for this model stochasticity, we generated 10 images for each prompt submitted by participants across all treatment arms.
-We then calculated the similarity between each replication and its corresponding target image, and computed an ”expected” CLIP cosine similarity and DreamSim score for each prompt by averaging across these samples.
-Using these replicated images, we also calculated the standard deviation of cosine similarity induced by this stochasticity.
-With the expected cosine similarity and its standard deviation per prompt, we computed a normalized Z-score for the observed image relative to its replication distribution.
-This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and is used for further analysis in section G.
-We generated these additional samples both for the original prompts on their assigned treatment arms and by replaying prompts on the counterfactual arms, as introduced in Figure 2 in the main text.
-Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling.
-As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts.
-This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section C.C.3.3).
-Appendix E Main Text Analyses
-This section provides detailed methodological information about the analyses presented in the main text.
-E.1 Conceptual Framework
-This appendix provides detailed derivations and proofs for the conceptual framework presented in the main text.
-While the main text introduces the conceptual framework and key insights, here we describe the model in more detail and develop its mathematical foundations and results more thoroughly.
-We model a task where users attempt to replicate an object (such as an image) using a generative model.
-Perfect replication yields a maximum quality value of 1.
-The bounded quality function is defined as:
-[…]
-where $\theta\in(0,1]$ denotes the model capacity, $s\in(0,1]$ denotes the user’s skill at writing prompts, and $x\geq 0$ is the level of user’s effort in generating the appropriate prompts.
-This function satisfies two intuitive properties:
-  1. 1.
-    Boundedness: Even as $x$ becomes large, $Q$ remains below 1, reflecting that the user cannot exceed perfect replication.
-  2. 2.
-    Diminishing returns: As user effort ($x$), increases, its marginal impact on quality becomes smaller.
-We assume a linear cost function of prompting:
-$c(x)\;=\;k\,x,$
-where $k>0$ represents the marginal cost of user effort, for example in terms of time spent. The user’s utility given the quality and cost functions is
-$U(\theta,s,x)=1-e^{-\theta sx}-k\,x.$
-E.1.1 User Optimization
-The user chooses the effort $x\geq 0$ such that it maximizes $U(\theta,s,x)$.
-We assume the user’s solution is always interior with positive effort which requires $\frac{\theta s}{k}>1$.
-This can be justified since the time cost associated with the task is minimal ($k\approx 0)$.
-As such, optimal user effort is $x^{*}(\theta,s)$:
-$x^{*}(\theta,s)\;=\;\frac{1}{\theta s}\,\ln\Bigl{(}\frac{\theta s}{k}\Bigr{)}>0$
-Thus, the optimal quality the user achieves at their optimal effort level becomes:
-$Q^{*}=Q\big{(}\theta,s,x^{*}(\theta,s)\big{)}=1-\frac{k}{\theta s}>0$
-E.1.2 Comparative Statics
-We now analyze how quality, $Q^{*}$, varies with $\theta$, $s$, and $k$.
-Comparative statics are easily shown by taking the derivative of $Q^{*}$ with respect to the model capacity, user’s skill, and prompting cost.
-Model Capacity: As model capacity $\theta$ increases, optimal quality $Q^{*}$ increases.
-$\frac{\partial Q^{*}}{\partial\Theta}\Bigg{|}_{\Theta=\theta}=\frac{k}{\theta^{2}s}>0$ (11)
-User’s Skill: As the skill of the user $s$ increases, the optimal quality $Q^{*}$ also increases.
-$\frac{\partial Q^{*}}{\partial S}\Bigg{|}_{S=s}=\frac{k}{\theta s^{2}}>0$ (12)
-Prompting Cost: As the prompting cost $k$ increases, the optimal quality $Q^{*}$ decreases.
-$\frac{\partial Q^{*}}{\partial K}\Bigg{|}_{K=k}=-\frac{1}{\theta s}<0$ (13)
-In the three derivatives above, the capital letters $\Theta,S$ and $K$ correspond to variables with respect to which derivatives are taken, and small letters $\theta,s$ and $k$ are the values at which derivatives are evaluated.
-E.1.3 Model Improvement and Skill Heterogeneity
-We next show that improvements in model capacity reduce the performance gap between high and low prompting skills. This result is mainly due to the diminishing returns in output quality.
-Result. As model capacity $\theta$ improves, the effect of user skill $s$ on optimal quality $Q^{*}$ diminishes.
-Proof: This can be easily shown by taking the derivative of equation 12, which denotes the total effect of user skill, with respect to a model improvement ($\theta$).
-$\frac{\partial^{2}Q^{*}}{\partial S\partial\Theta}\Bigg{|}_{S=s,\Theta=\theta}=-\frac{k}{\theta^{2}s^{2}}<0$ (14)
-As before capital letters $\Theta,S$ correspond to variables that are evaluated at $\theta,s$. $\square$
-E.1.4 Decomposition into Model and Prompt Effects
-As discussed in the main text, we decompose the total quality improvement into two parts: a model effect (the improvement from upgrading the model but holding the prompting effort fixed) and a prompting effect (the improvement from re-optimizing the effort in response to the upgraded model).
-To perform this decomposition, we consider a “counterfactual” quality function that uses one model’s capacity but another model’s optimal prompt:
-$Q\bigl{(}\theta_{2},s,x^{*}(\theta_{1},s)\bigr{)}=1-\exp\big{(}-\theta_{2}s\,x^{*}(\theta_{1},s)\big{)}$ (15)
-which is the quality obtained by using the model $\theta_{2}$ while choosing the prompting effort $x^{*}(\theta_{1},s)$ that would have been optimal if the user actually had model $\theta_{1}$.
-Note that if $\theta_{1}=\theta_{2}$, then $x^{*}(\theta_{1},s)$ is indeed the true optimum for $\theta_{2}$, yielding $Q^{*}$.
-Given the formulation above, we can now express the model effect ($M$) and the prompting effect ($P$) by taking the derivative of the counterfactual quality in equation 15 with respect to the model capacity directly or indirectly through its effect on the prompt.
-The model effect is the derivative of 15 with respect to the direct effect of model capacity, while keeping the user’s effort fixed.
-$M$ $=\;\frac{\partial Q\bigl{(}\Theta_{2},s,x^{*}(\Theta_{1},s)\bigr{)}}{\partial\Theta_{2}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$ (16)
-$=\;s\cdot x^{*}(\Theta_{1},s)\cdot\exp\left(-\Theta_{2}s\cdot x^{*}(\Theta_{1},s)\right)\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$
-$=\;\frac{k}{\theta^{2}s}\cdot\ln\left(\frac{\theta s}{k}\right)$
-We can obtain the prompt effect by taking the derivative of 15 with respect to its indirect impact on the user’s effort, while keeping the direct impact of the model capacity fixed.
-$P$ $=\;\frac{\partial Q\bigl{(}\Theta_{2},s,x^{*}(\Theta_{1},s)\bigr{)}}{\partial\Theta_{1}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$ (17)
-$=\;\Theta_{2}s\cdot\exp\left(-\Theta_{2}s\cdot x^{*}(\Theta_{1},s)\right)\cdot\,\frac{\partial x^{*}(\Theta_{1},s)}{\partial\Theta_{1}}\Bigg{|}_{\Theta_{1}=\Theta_{2}=\theta}$
-$=\;\frac{k}{\theta^{2}s}\left(1-\ln\left(\frac{\theta s}{k}\right)\right)$
-We now show that the model effect mainly benefits low-skilled users in contrast to the prompting effect which mainly benefits the high-skilled users.
-Proposition 3.
-As user skill $s$ increases, the model effect becomes smaller, for sufficiently small prompting cost $k$.
-Proof.
-To examine how the model effect varies by user skill, we can take the derivative of the model effect M in equation 16 with respect to $s$.
-$\frac{\partial M}{\partial S}\Bigg{|}_{S=s}$ $=\frac{\partial}{\partial S}\frac{k}{\theta^{2}S}\cdot\ln\left(\frac{\theta S}{k}\right)\Bigg{|}_{S=s}$
-$=\frac{k}{\theta^{2}s^{2}}\left(1-\ln\left(\frac{\theta s}{k}\right)\right)<0$
-for sufficiently small $k$. ∎
-Intuitively when prompting costs are low, users can afford to put substantial effort into writing prompts (so $x^{*}$ is high).
-Once the effort is large enough, the marginal effect of the model on quality (i.e.
-the model effect) will be decreasing by user skill.
-This happens due to the diminishing returns in the quality function.
-Once a high-skill user is already achieving near-peak quality by putting in large effort, small increases in $\theta$ deliver small additional benefit.
-However, the same model improvement yields a much larger marginal impact for a low-skilled user who is putting the same level of effort.
-Proposition 4.
-As user skill $s$ improves, the prompting effect becomes larger, for sufficiently small cost $k$.
-Proof.
-Proof To examine how the prompt effect varies by user skill, we can again take the derivative of the prompting effect P in equation 17 with respect to $s$.
-$\frac{\partial P}{\partial S}\Bigg{|}_{S=s}$ $=\frac{\partial}{\partial S}\frac{k}{\theta^{2}S}\cdot\big{(}1-\ln(\frac{\theta S}{k})\big{)}\Bigg{|}_{S=s}$ (18)
-$=\frac{k}{\theta^{2}s^{2}}\left(\ln\left(\frac{\theta s}{k}\right)-2\right)>0$
-for sufficiently small $k$. ∎
-Comparing equations E.1.4 and 18, we conclude that the prompting effect is less sensitive to user skill than the model effect, as its marginal with respect to skill has a smaller magnitude.
-This makes the overall marginal effect of user skill negative, matching the earlier finding in equation 14.
-E.2 Task Performance and ATEs
-The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations).
-It displays the average cosine similarity score stratified by the reference image.
-A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms.
-This is likely due to participants’ initial misunderstanding of the model’s “memoryless” nature.
-Participants failed to recognize that context from previous prompts was not carried over to new iterations.
-We observed numerous prompts in the second iteration across users that explicitly referenced the first prompt, a behavior that rarely occurred in subsequent attempts.
-However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
-The bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane.
-[^10] In Section E, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified.
-To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
+G.2 Example Logo Generation Prompts
+Table A.5 presents the complete set of prompts and their corresponding generated images that were used in our analysis shown in Figure 3 for the logo generation experiment.
+It is the exact same format and presents the analogous information as Table A.4, but for the logos.
+Table A.5: Participant prompts for the images provided in Figure 3
+Image Prompt
+[Uncaptioned image] “Acting as a graphic designer design a logo for mega surf.
+They are a tech startup company developing a mobile app that will connect surfers in a neighborhood.
+The logo must include a surfboard and an ocean wave ,be very attractive and appeal to new and existing users.
+It will be used for the app icon, digital marketing and business cards.”
+[Uncaptioned image] “the words ‘MegaSurf’ are the main focus of the image.
+a Large simple lightorange surf board ontop of a light teal wave two people holding surf baords can be seen on the shore.
+the logo is done in a simple sillhoutte style’
+[Uncaptioned image] “Create an image from the sandy beach showing a surfer on a surfboard and an ocean wave.”
+[Uncaptioned image] “create a logo for Prestwick University that emphasizes academic excellence, and includes a torch with a red flame, an open book with writing in it, and the color blue”
+[Uncaptioned image] “draw me a simple cartoon drawing of an open book and a torch”
+[Uncaptioned image] “on the last post make the book open”
+[Uncaptioned image] “Create a unique design for a logo for an electronic music event, featuring a sun that emits sound waves instead of sun rays and musical notes.
+The name of the event is ”Summer Beats Festival”
+[Uncaptioned image] “A logo for Summer Beats Festival.
+A large sun is the main character and he is the DJ dressed up in rave gear ready to party.
+Music is bumping loudly from speakers and you can see the sound waves coming off the speakers.”
+[Uncaptioned image] “The sun holding a microphone. Sound waves coming out of the microphone.”
+Appendix H Robustness Checks
+This section provides additional analyses that confirm the robustness of our main findings using alternative metrics and statistical approaches.
+H.1 Image Replication Task: DreamSim-based Analysis
+As discussed in Section C, we repeated all main-text analyses using DreamSim as an alternative image similarity metric.
+The results, presented below, demonstrate that our findings are not dependent on the specific similarity measure used.
+H.1.1 Overall ATEs
+In terms of DreamSim, participants using DALL-E 3 produced images that were, on average, $z=0.238$ standard deviations (95% CI = [0.152, 0.324]) closer to the target image ($\Delta DreamSim=0.0306$, $p<10^{-7}$) than those produced by participants using DALL-E 2.
+H.1.2 Figure 1
+We reran the regression in Section F with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
 $Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (15)
 $+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
-The coefficient estimates generated by this model are:
-  • $\hat{\beta_{1}}=0.0011$, $\hat{SE}(\beta_{1})=0.0003$, $p=0.0004$
-  • $\hat{\beta_{2}}=0.0120$, $\hat{SE}(\beta_{2})=0.0037$, $p=0.0013$
-  • $\hat{\beta_{3}}=0.0010$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0227$
-The overall ATEs that we report between different pairs of treatment arms (DALL-E 2, DALL-E 3, and DALL-E 3 with revisions) in the main text are estimated from a two-way fixed effect (iteration and target image) model per each pair.
-Standard errors are cluster robust at the participant level.
-E.3 Prompt Characteristics
-Figure 1C compares the prompt length and prompt similarity of the two models.
-To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section C.C.4.
-The prompt length is the average number of words per model and iteration stratified by the reference image.
-The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section C.C.5.3 for details on similarity calculations).
-The color scale in Figure 1C shows the stratified average similarity to the previous prompt across all users per each model.
-We find that superior model users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
-Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants.
-For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section C.C.5.3, as the dependent variable.
-When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$).
-Standard errors are cluster-robust at the participant level.
-E.4 ATE Decomposition
-Figure 2 in the main text decomposes the ATE into the model and prompting effects.
-This decomposition is conceptually similar to a simple mediation analysis, with an important difference being that we can observe counterfactual outcomes (e.g., prompting the superior model as if it is the inferior model).
-This is not typically the case in mediation analysis, and makes causal identification rely on fewer assumptions.
-To obtain counterfactual outcomes, we fed or “replayed” the participant prompts when interacting with one model (e.g., inferior) on another model (e.g., superior).
-The notation (prompt, model) specifies which treatment arm the prompts were written under and which model was used in the replay.
-For example, (2,3) indicates replaying prompts written under DALL-E 2 on DALL-E 3.
-To be clear, (2,2) and (3,3) correspond to the original observed treatment arms, while (2,3) and (3,2) are the counterfactual outcomes of interest.
-[^11] To avoid problems with model drift, we regenerated images for all four possible combinations at the same time and used these images for all analyses in the main text.
-The left-most point in Figure 2 corresponds to the average CLIP cosine similarity to the target image of (2,2).
-To make the interpretation of the results clearer, we have subtracted this quantity from all average quality scores and added a dashed line throughout.
-The second point from the left corresponds to average similarity to the target of (2,3), the third point from the left to (3,3), and the rightmost points to (3,2).
-All average similarity scores are stratified by iteration and reference image, and the standard errors are bootstrapped and cluster-robust at the participant level.
-The model effect, as shown by the red braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (2,2).
-In the terminology of mediation analysis, the model effect would be referred to as the direct effect.
-The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (3,3) relative to (2,3).
-In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect.
-We can also test the difference in average quality between (3,3) and (3,2), as well as the difference between (3,2) and (2,2).
-Both of these differences are visible in Figure 2; the second is small and not statistically significant.
-The standard errors in Figure 2 correspond to the uncertainty around the estimated average score for each of the four replay conditions.
-These uncertainty estimates are insufficient for exact inference on the direct, indirect, and treatment effects.
-The statistics and significance values reported in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
-$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{effect}+\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$ (20)
-where $\beta_{1}$ is the coefficient on the effect type in question (i.e., model or prompting).
-To estimate the different effects, we simply use the above model and filter the data as appropriate.
-For example, to estimate the ATE, the data contains all (2,2) and (3,3) scores, and in this case effect=1 for observations in (3,3) group.
-Similarly, to estimate the direct or model effect, the data contains all (2,2) and (2,3) scores and effect=1 for observations in (2,3) group.
-Finally, to estimate the indirect or prompting effect, the data contains all (2,3) and (3,3) scores and effect=1 for observations in (3,3) group.
-The standard errors for each estimated model are cluster robust at the participant level, and p-values are adjusted accordingly.
-E.5 Heterogeneity by Skill
-Table 1 in the main text demonstrates how total, model, and prompting effects vary across different user skill levels.
-The total effect compares the Cosine similarity of outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 2.
-The model effect compares outputs from DALL-E 2 users replayed on DALL-E 3 against the same users replayed on DALL-E 2.
-The prompting effect compares outputs from DALL-E 3 users replayed on DALL-E 3 to those from DALL-E 2 users replayed on DALL-E 3.
-To assess user skill, users within each target image, iteration and replay scenario are divided into 50 equally sized brackets based on their performance, assigning each bracket a rank corresponding to performance deciles.
-For instance, when evaluating DALL-E 2 users replayed on DALL-E 2, each user’s skill level is determined by their percentile rank among all DALL-E 2 users replayed on DALL-E 2.
-Conversely, when the same user is replayed on DALL-E 3 (to estimate the model effect or to serve as a baseline in estimating the prompting effect), their performance decile is determined by their rank relative to all DALL-E 2 users replayed on DALL-E 3, per each iteration and target image.
-After assigning performance deciles, we estimate the following linear model with two-way fixed effects (iteration and target image) for each effect type (model, prompting, and total):
-$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\text{effect}+\beta_{2}\text{effect}\times\text{Performance Decile}_{i,n,t}\,+$ (21)
-$\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$
-where $\beta_{1}$ represents the coefficient associated with the effect type being analyzed (model, prompting, or total), and $\beta_{2}$ captures the interaction between the effect type and user skill level.
-User skill is measured as the user’s rank (binned) within the same iteration, target image, and replay scenario.
-Standard errors are calculated using cluster-robust methods at the participant level.
-This methodology closely follows the procedure described earlier in section E.4, with the additional step of incorporating an interaction term to evaluate skill-level heterogeneity.
-Appendix F Robustness Checks
-This section provides additional analyses that confirm the robustness of our main findings using alternative metrics and statistical approaches.
-F.1 DreamSim-based Analysis
-As discussed in Section C, we repeated all main-text analyses using DreamSim as an alternative image similarity metric.
-The results, presented below, demonstrate that our findings are not dependent on the specific similarity measure used.
-F.1.1 Overall ATEs
-In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.238$ standard deviations (95% CI = [0.152, 0.324]) closer to the target image ($\Delta DreamSim=0.0306$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
-F.1.2 Figure 1
-We reran the regression in Section E with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
-$Y_{i,n,t}=$ $\beta_{0}+\beta_{1}\,\text{iteration}+\beta_{2}\,\mathbb{I}[\text{dalleVersion = 3}]_{i}$ (22)
-$+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion =3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$
-The coefficient estimates generated by this analysis are:
-  • $\hat{\beta_{1}}=0.0034$, $\hat{SE}(\beta_{1})=0.0005$, $p=1.3\times 10^{-12}$
-  • $\hat{\beta_{2}}=0.0200$, $\hat{SE}(\beta_{2})=0.0061$, $p=0.0011$
-  • $\hat{\beta_{3}}=0.0024$, $\hat{SE}(\beta_{3})=0.0007$, $p=0.0009$
-F.1.3 Figure 2
+The coefficient estimates generated by this analysis are: $\hat{\beta_{1}}=0.0034;(0.0005)$, $p=1.3\times 10^{-12}$; $\hat{\beta_{2}}=0.0200;(0.0061)$, $p=0.0011$; $\hat{\beta_{3}}=0.0024;(0.0007)$, $p=0.0009$.
+H.1.3 Figure 2
 Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text.
 The model effect accounts for 54.4% of the ATE ($\Delta DreamSim=0.0166$, $p<10^{-7}$), whereas the prompting effect accounts for 45.4% of the ATE ($\Delta DreamSim=0.01390$, $p=0.014$).
-F.2 Z-score-based Analysis
-As we discuss in Section D.2, we repeated all main-text analyses using the within-image-attempt Z-score of CLIP-cosine similarity to account for variation between images and attempts.
-When comparing across attempts, the Z-score was computed within each image, as mentioned in Section D.2.
-The results remain consistent throughout, and in some cases, the differences between the superior and inferior model are even more pronounced than when using raw cosine similarity.
-F.2.1 Overall ATEs
-As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.19$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.19, 95% CI = [0.100, 0.271]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2 (the inferior model).
+H.2 Image Replication Task: Z-score-based Analysis
+As we discuss in Section E.2, we repeated all main-text analyses using the within-image-attempt Z-score of CLIP-cosine similarity to account for variation between images and attempts.
+When comparing across attempts, the Z-score was computed within each image, as mentioned in Section E.2.
+The results remain consistent throughout, and in some cases, the differences between DALL-E 3 and DALL-E 2 are even more pronounced than when using raw cosine similarity.
+H.2.1 Overall ATEs
+As mentioned in the main text, participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.19, 95% CI = [0.100, 0.271]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2.
 Standard errors are clustered at the participant level.
-F.2.2 Figure 1
-On average, participants using the superior model produced images that were $z=0.19$ standard deviations closer (the ATE) to the target image than those using the inferior model.
+H.2.2 Figure 1
+On average, participants using DALL-E 3 produced images that were $z=0.19$ standard deviations closer (the ATE) to the target image than those using DALL-E 2.
 Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
@@
 $+\beta_{3}\,\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\epsilon_{i,n}$
-  • $\hat{\beta_{1}}=0.0129$, $\hat{SE}(\beta_{1})=0.0038$, $p=0.0007$
-  • $\hat{\beta_{2}}=0.1250$, $\hat{SE}(\beta_{2})=0.0457$, $p=0.0064$
-  • $\hat{\beta_{3}}=0.0128$, $\hat{SE}(\beta_{3})=0.0053$, $p=0.015$
-F.2.3 Figure 2
+The coefficient estimates generated by this analysis are: $\hat{\beta_{1}}=0.0129;(0.0038)$, $p=0.0007$; $\hat{\beta_{2}}=0.1250;(0.0457)$, $p=0.0064$; $\hat{\beta_{3}}=0.0128;(0.0053)$, $p=0.015$.
+H.2.3 Figure 2
 When we decompose the ATE into the model effect ($z=0.0791$; $p=8.35\times 10^{-6}$) and prompting effect ($z=0.1046$; $p=0.016$), they account for 43% and 56% of the treatment effect, respectively.
-And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.033$; $p=0.45$).
+And when we replay DALL-E 3 prompts on DALL-E 2, there is not a statistically significant increase in performance relative to DALL-E 2 prompts played on DALL-E 2 (z=0.033z=-0.033; $p=0.45$).
 In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z-score of the cosine similarity.
-Appendix G Pre-registration
-Prior to conducting our experiment, we developed and pre-registered a comprehensive set of hypotheses and analyses.
+H.3 Logo Generation Task: Z-score-based Analysis
+The Z-score procedure for the logo generation task follows the same approach described in Section H.2 for the image replication task.
+As before, we compute within-logo Z-scores to normalize Bradley–Terry (BT) strengths, allowing us to remove baseline variation in difficulty across logo task.
+The only substantive difference is the outcome: here the dependent variable is the BT $\beta$ score described in Section D.2 rather than CLIP cosine similarity.
+H.3.1 Overall ATEs
+Applying the same Z-scoring procedure, we find that participants using DALL-E 3 produced logos that were, on average, $z=0.569$ standard deviations higher in BT strength than those produced by DALL-E 2 users (95% CI = [0.502, 0.636], $p<10^{-10}$).
+As in the unscaled BT analysis reported in the main text, these results indicate a substantial and statistically significant advantage for DALL-E 3.
+H.3.2 Figure 3
+Repeating the regression specification used in Section H.2.2, but substituting Z-scored BT strengths as the dependent variable, we again observe that the performance gap between DALL-E 3 and DALL-E 2 widens modestly across attempts, but we lack sufficient power to detect statistically significant effects.
+One potential reason is the dip in second–attempt performance, which flattens the across-iteration trend and reduces the estimated slope in a linear model.
+The estimated coefficients are: $\hat{\beta_{1}}=0.0043;(0.0043)$, $p=0.32$; $\hat{\beta_{2}}=0.5104;(0.0412)$, $p<2.2\times 10^{-16}$; $\hat{\beta_{3}}=0.0067;(0.0057)$, $p=0.238$.
+H.3.3 Figure 4
+The replay-based decomposition is unchanged conceptually from Section H.2.3.
+Using Z-scored BT strengths, we estimate the model effect ($z=0.522$; $p<10^{-15}$) and prompting effect ($z=0.046$; $p=0.167$) to constitute about 91.8% and 8.1% of the total treatment effect, respectively.
+The relative sizes of model and prompting effects using the z-scaled scores are similar to the unscaled analysis reported in the main text.
+The replay of DALL-E 3 prompts on DALL-E 2 shows no meaningful improvement (z=0.0028z=-0.0028, $p=0.934$).
+Appendix I Full pre-registered analyses (Image Replication Experiment)
+Prior to conducting the image replication experiment, we developed and pre-registered a comprehensive set of hypotheses and analyses.
 This pre-registration is deposited at OSF at the following URL: URL removed to maintain anonymity.
-While the main text focuses on a subset of our pre-registered analyses that constitute the most important and timely findings, this appendix provides a comprehensive overview of all analyses specified in our pre-registration.
-In our pre-registration, we outlined plans to conduct each analysis using six [^12] See section G.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
+While the main text of our paper focuses on a subset of our pre-registered analyses, this appendix provides a comprehensive overview of all analyses specified in our pre-registration.
+Note that for the logo generation task, we limited our pre-registered hypotheses to the analyses appearing in the main text of this paper.
+In our pre-registration, we outlined plans to conduct each analysis using eight [^14] See section I.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables.
 possible outcome variables, all representing different transformations of the same underlying data.
 These include CLIP embedding cosine similarity and DreamSim, each with three rescaling methods:
+  1. 1.
     No rescaling: The outcome variable used in its original form.
-  • Z-score rescaling: The outcome variable transformed into a Z-score following the procedure detailed in Section D.2.
-  • Percentile rank rescaling: The outcome variable converted to a percentile rank relative to all other prompts submitted for the same target image.
-While we are still in the process of completing all analyses with each of these outcome variables, our preliminary results suggest that our findings are robust across these different transformations.
-As additional analyses are completed, we will update this appendix.
-Full results, data, and replication code available online upon publication.
-G.1 Hypotheses and Results
+  2. 2.
+    Z-score rescaling: The outcome variable is transformed into a Z-score following the procedure detailed in Section E.2.
+  3. 3.
+    Percentile rank rescaling: The outcome variable is converted to a percentile rank relative to all other prompts submitted for the same target image.
+There is one additional rescaling method that we pre-registered, but do not conduct our analyses with this rescaling.
+The reason that we do not use it is detailed in the Section discussing deviations from the pre-registration (Section I.2).
+I.1 Hypotheses and Results
 Below, we present each hypothesis exactly as stated in our pre-registration document, followed by a summary of our findings.
 This comprehensive review demonstrates the robustness of the key results highlighted in the main text, while also providing additional insights into the relationship between model capabilities, user behavior, and task performance.
-H1
+H1.
 There are differences in prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) across demographic attributes and other observables, such as educational background and occupational skills.
-Analysis approach: For this hypothesis, we conducted multiple ANOVA tests, one for each demographic variable, with the relevant performance measure as the dependent variable and the model (DALL-E version) as a covariate.
+Analysis approach:
+For this hypothesis, we conducted multiple ANOVA tests, one for each demographic variable, with the relevant performance measure as the dependent variable and the model (DALL-E version) as a covariate.
 As a robustness check, we repeated these analyses using the non-parametric Kruskal-Wallis U test.
 To account for multiple comparisons, we applied the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
-Results: Our analyses revealed several consistent demographic patterns in prompt engineering ability across different outcome measures:
-  • CLIP embedding cosine similarity
-    • No rescaling: Using ANOVA, we found significant associations between performance and: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills (particularly critical thinking, active listening, and quality control).
+Our analyses revealed several consistent demographic patterns in prompt engineering ability across different outcome measures:
+CLIP no rescaling.
+Using ANOVA, we found significant associations between performance and: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills (particularly critical thinking, active listening, and quality control).
 In linear models examining directionality, we found that participants who reported critical thinking as a job skill and little usage of generative AI or imagery writing skill performed better, on average.
@@
 The more conservative Kruskal-Wallis tests identified fewer significant variables: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender, and age.
-    • Z-score rescaling: ANOVA tests revealed significant associations with similar demographic factors as the unscaled measure, plus additional significant associations with self-reported occupational skills in technology design, social perceptiveness, and troubleshooting.
+CLIP Z-score rescaling.
+ANOVA tests revealed significant associations with similar demographic factors as the unscaled measure, plus additional significant associations with self-reported occupational skills in technology design, social perceptiveness, and troubleshooting.
 Linear models showed that participants reporting critical thinking and troubleshooting as job skills, little generative AI use, little imagery writing skill, some programming skill, and some instruction writing skill performed better.
@@
 Kruskal-Wallis tests again identified a smaller set of significant predictors: self-reported programming frequency, generative AI outlook, programming skill, gender, and age.
-    • Percentile rank rescaling: ANOVA identified significant associations between performance and computer programming frequency, self-reported programming ability and frequency, generative AI outlook, age, gender, generative AI use, and critical thinking as an occupational skill.
+CLIP Percentile rank rescaling.
+ANOVA identified significant associations between performance and computer programming frequency, self-reported programming ability and frequency, generative AI outlook, age, gender, generative AI use, and critical thinking as an occupational skill.
 Linear models showed better performance among those reporting critical thinking as a job skill, little generative AI use, and some programming skill.
@@
 Kruskal-Wallis tests revealed the same set of significant predictors as with other scalings: programming frequency, generative AI outlook, programming skill, gender, and age.
-  • DreamSim
-    • No rescaling: ANOVA tests identified significant associations between performance and: programming frequency, self-reported programming ability, generative AI outlook, age, gender, generative AI use, imagery and instructional writing skills, and several occupational skills (critical thinking, learning strategies, technology design, and quality control).
+DreamSim no rescaling.
+ANOVA tests identified significant associations between performance and: programming frequency, self-reported programming ability, generative AI outlook, age, gender, generative AI use, imagery and instructional writing skills, and several occupational skills (critical thinking, learning strategies, technology design, and quality control).
 Linear models showed better performance among those reporting critical thinking and social perceptiveness as job skills, little generative AI use, and little imagery writing or programming skills.
@@
 Kruskal-Wallis tests identified significant relationships with: programming frequency, generative AI outlook, imagery writing skill, gender, age, and learning strategies as an occupational skill.
-    • Z-score rescaling: Using ANOVA, significant associations included all factors identified in the unscaled analysis, plus education and additional occupational skills (social perceptiveness).
+DreamSim Z-score rescaling.
+Using ANOVA, significant associations included all factors identified in the unscaled analysis, plus education and additional occupational skills (social perceptiveness).
 Linear model directional findings were consistent with the unscaled measure, with additional negative associations for graduate degrees and self-reported technology design, quality control, and learning strategies skills.
 Kruskal-Wallis tests showed significant relationships with: programming frequency, generative AI outlook, imagery writing skill, gender, age, and occupational skills in learning strategies, social perceptiveness, and technology design.
-    • Percentile rank rescaling: ANOVA identified significant associations with: programming frequency, programming ability, generative AI outlook, age, gender, generative AI use, imagery writing skill, education, and occupational skills in social perceptiveness, learning strategies, and technology design.
+DreamSim Percentile rank rescaling.
+ANOVA identified significant associations with: programming frequency, programming ability, generative AI outlook, age, gender, generative AI use, imagery writing skill, education, and occupational skills in social perceptiveness, learning strategies, and technology design.
 Linear models showed better performance among participants reporting social perceptiveness as a job skill, little generative AI use, and little imagery writing or programming skills.
@@
 Kruskal-Wallis tests revealed significant relationships with: programming frequency, generative AI use, generative AI outlook, imagery writing skill, gender, age, and occupational skills in learning strategies, social perceptiveness, and technology design.
+Interpretation.
 Across these analyses, several consistent patterns emerge.
@@
 These patterns were generally consistent across different outcome measures and scaling approaches, suggesting robust demographic differences in prompt engineering ability.
-H2
+H2.
 There are observable differences in the prompting techniques of successful prompt engineers and unsuccessful prompt engineers.
 Such prompting techniques might include the use of longer prompts, the use of structured prompting techniques, and/or specific patterns in the way that the participant iterates on their prompts over time.
-Analysis approach: We investigated this hypothesis by examining the relationship between exploration/exploitation strategies in the prompting space and performance outcomes.
+Analysis approach.
+We investigated this hypothesis by examining the relationship between exploration/exploitation strategies in the prompting space and performance outcomes.
 We characterized participants as more ”exploitative” if they wrote prompts that were similar to their previous prompts, and more ”exploratory” if their prompts exhibited greater deviation from previous attempts.
-We operationalized these concepts using several metrics:
-  1. 1.
-    Measures positively associated with exploitation:
-    • Average token sort ratio compared to the previous prompt
-    • Average cosine similarity between the embeddings of consecutive prompts
-    • Fraction of times a prompt contains the previous prompt as an exact substring
-  2. 2.
-    Measures negatively associated with exploitation:
-    • Variance of the prompt embedding
-    • Number of topical transitions
+We operationalized these concepts using several metrics.
+The measures positively associated with exploitation are: average token sort ratio compared to the previous prompt, average cosine similarity between the embeddings of consecutive prompts, and the fraction of times a prompt contains the previous prompt as an exact substring.
+The measures negatively associated with exploitation are: variance of the prompt embedding and the number of topical transitions
 We also measured each participant’s average prompt length.
@@
 We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket.
-Our estimates adjusted for other covariates by matching user-iteration observations on target image, DALL-E model, iteration number, and exact quality of the previous attempt (Sävje et al.
-2021).
-Results: Our analyses revealed consistent patterns across different outcome measures:
-  • CLIP embedding cosine similarity
-    • No rescaling: At the user level, we found strong and statistically significant associations between performance and prompting strategies, even after Benjamini-Hochberg adjustment for multiple testing.
+Our estimates adjusted for other covariates by matching user-iteration observations on target image, DALL-E model, iteration number, and exact quality of the previous attempt (savje2021generalized).
+Our analyses revealed consistent patterns across different outcome measures:
+CLIP No rescaling.
+At the user level, we found strong and statistically significant associations between performance and prompting strategies, even after Benjamini-Hochberg adjustment for multiple testing.
 Specifically, token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were all positively associated with performance.
@@
 Similarly, prompts that included the previous prompt showed higher performance, with larger improvements as previous performance increased.
-    • Z-score rescaling: Results were consistent with those for unscaled cosine similarity.
+CLIP Z-score rescaling.
+Results were consistent with those for unscaled cosine similarity.
 The primary difference was that the non-linear relationship between performance and continuous measures of exploitation (TSR ratio and cosine similarity with previous prompt) became more pronounced for the bottom two brackets of previous performance.
 For low-performing prompts in the previous attempt, moderate exploitation levels yielded optimal performance, with significant performance deterioration at both high and low exploitation levels.
-    • Percentile rank rescaling: We found the same general patterns described above, with the non-linearity in exploration/exploitation versus performance in the bottom two brackets of previous performance even stronger than observed with Z-score rescaled cosine similarity.
-  • DreamSim
-    • No rescaling: Results were consistent with those described for unscaled cosine similarity.
-    • Z-score rescaling: Results matched those described for Z-score cosine similarity, with particularly stark non-linearity between performance and continuous measures of exploitation at the user-prompt level for the bottom two brackets of previous performance.
-    • Percentile rank rescaling: Results were consistent with those described for percentile rank rescaled cosine similarity.
+CLIP Percentile rank rescaling.
+We found the same general patterns described above, with the non-linearity in exploration/exploitation versus performance in the bottom two brackets of previous performance even stronger than observed with Z-score rescaled cosine similarity.
+Dream Sim No rescaling.
+Results were consistent with those described for unscaled cosine similarity.
+DreamSim Z-score rescaling.
+Results matched those described for Z-score cosine similarity, with particularly stark non-linearity between performance and continuous measures of exploitation at the user-prompt level for the bottom two brackets of previous performance.
+DreamSim Percentile rank rescaling.
+Results were consistent with those described for percentile rank rescaled cosine similarity.
+Interpretation.
 These findings suggest a nuanced relationship between exploration/exploitation strategies and performance.
@@
 This pattern was consistent across different outcome measures and scaling approaches.
-H3
+H3.
 There are differences in prompt engineering techniques (as measured through metrics such as prompt length and iteration-to-iteration token sort ratio) across demographic attributes and other observables, such as educational background and occupational skills.
-Analysis approach: To test this hypothesis, we estimated linear models with each demographic trait as the independent variable, treatment arm fixed effects as controls, and various prompting behaviors (described in Section C.C.5) as the dependent variables.
+Empirical approach:
+To test this hypothesis, we estimated linear models with each demographic trait as the independent variable, treatment arm fixed effects as controls, and various prompting behaviors (described in Section C.D) as the dependent variables.
 To account for multiple testing, we adjusted p-values across all models using the Benjamini-Hochberg procedure.
-Results: Our analysis revealed several significant demographic differences in prompting strategies, with consistent patterns emerging across different measures of prompt similarity and evolution.
+Our analysis revealed several significant demographic differences in prompting strategies, with consistent patterns emerging across different measures of prompt similarity and evolution.
 The detailed findings for each dependent variable are summarized below:
-  • Prompt embedding variance: We did not find any significant differences across demographic traits in the overall variance of prompt embeddings, suggesting that the breadth of prompt space exploration was relatively consistent across different demographic groups.
-  • Strategic Shifts: We observed statistically significant differences based on age, programming frequency, and instructional writing frequency.
+Prompt embedding variance.
+We did not find any significant differences across demographic traits in the overall variance of prompt embeddings, suggesting that the breadth of prompt space exploration was relatively consistent across different demographic groups.
+Strategic shifts.
+We observed statistically significant differences based on age, programming frequency, and instructional writing frequency.
 Specifically, younger participants, those with low programming frequency, and those with some instructional writing experience demonstrated fewer topical transitions across their prompts, indicating a more consistent approach to prompt development.
-  • Successive Prompt Token Sort Ratio: Significant differences emerged by age, education, programming frequency, and imagery/instructional writing frequencies.
+Successive prompt token sort ratio.
+Significant differences emerged by age, education, programming frequency, and imagery/instructional writing frequencies.
 Older users, those with post-graduate degrees, those with high programming frequency, and those with high writing frequency (both for precise instructions and imagery) wrote prompts that were less similar to their previous attempts as measured by token sort ratio.
 This suggests these groups took a more exploratory approach to prompt iteration.
-  • Successive similarity: We found significant differences across numerous demographic factors: age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency, and certain occupational skills.
+Successive similarity.
+We found significant differences across numerous demographic factors: age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency, and certain occupational skills.
 On average, older users, males, those with post-graduate degrees, frequent programmers, self-reported skilled programmers, those with positive outlooks toward generative AI, frequent generative AI users, those who frequently write instructions or imagery, and those reporting critical thinking and social perceptiveness as occupational skills wrote prompts with lower cosine similarity to their previous attempts.
 This consistent pattern across multiple demographic factors suggests robust differences in exploration-exploitation tendencies.
-  • Successive Prompt ’Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook toward generative AI, and imagery writing frequency.
+Successive prompt ’contains previous prompt’ dummy.
+We found significant differences by age, gender, outlook toward generative AI, and imagery writing frequency.
 Older users and those with neutral outlooks toward generative AI were less likely to write prompts that contained their previous prompt as an exact substring.
 In contrast, males and those with some imagery writing skills were more likely to build directly upon their previous prompts by including them in subsequent attempts.
+Interpretation.
 These findings reveal a nuanced picture of demographic differences in prompt engineering approaches.
@@
 This apparent contradiction suggests that the relationship between background, prompting strategy, and performance is complex and potentially mediated by other factors not captured in these analyses.
-H4
+H4.
 Insofar as the output returned by a generative AI model in response to a prompt is stochastic, the subsequent prompting strategies and prompting outcomes of participants that get lower-than-expected, higher-than-expected, or approximately expected outputs in response to their first prompt are different.
-Analysis approach: To test this hypothesis, we examined how the random variation in image quality resulting from model stochasticity affects subsequent user behavior.
-We first calculated a Z-score for each participant’s generated image relative to the expected distribution for that prompt, following the procedure outlined in Section C.C.3 (with computational details in Section D).
+Empirical approach:
+To test this hypothesis, we examined how the random variation in image quality resulting from model stochasticity affects subsequent user behavior.
+We first calculated a Z-score for each participant’s generated image relative to the expected distribution for that prompt, following the procedure outlined in Section C.C.3 (with computational details in Section E).
 This Z-score quantifies how much better or worse the actually generated image was compared to what would be expected for that prompt on average.
@@
 Additionally, we conducted analyses at the user level to determine whether the Z-score of the first prompt affects average performance across all subsequent attempts.
-Results: Our analyses revealed several consistent patterns across different outcome measures:
-  • CLIP embedding cosine similarity
-    • No rescaling: We found statistically significant evidence that higher Z-scores for images in previous prompts were associated with increased cosine similarity in subsequent attempts.
+Our analyses revealed several consistent patterns across different outcome measures:
+CLIP No rescaling.
+We found statistically significant evidence that higher Z-scores for images in previous prompts were associated with increased cosine similarity in subsequent attempts.
 When comparing performance across the trichotomous Z-score variable, the difference between the top bracket (”higher-than-expected”) and bottom bracket (”lower-than-expected”) was statistically significant, with the top bracket showing better performance in the next attempt.
@@
 Linear models at the user level did not reveal statistically significant relationships between the first prompt’s Z-score and average performance in subsequent attempts.
-    • Z-score rescaling: Results were similar to those for unscaled cosine similarity.
+CLIP Z-score rescaling.
+Results were similar to those for unscaled cosine similarity.
 We found a statistically significant difference between the top and bottom Z-score brackets, with higher performance in the next attempt for the top bracket.
@@
 Unlike with the unscaled measure, we found no statistically significant effects at the user level, either when comparing Z-score brackets or when using linear models.
-    • Percentile rank rescaling: Results were consistent with those for Z-score rescaled cosine similarity.
-  • DreamSim
-    • No rescaling: Results showed similar patterns to CLIP embedding cosine similarity.
+CLIP Percentile rank rescaling:
+Results were consistent with those for Z-score rescaled cosine similarity.
+DreamSim No rescaling.
+Results showed similar patterns to CLIP embedding cosine similarity.
 Higher Z-scores in previous prompts were associated with better performance in subsequent attempts.
@@
 At the user level, we found no statistically significant relationships between the first prompt’s Z-score and performance in subsequent attempts.
-    • Z-score rescaling: Results closely matched those for unscaled DreamSim scores, with statistically significant differences between the top bracket and both the bottom and middle brackets, but no significant difference between the bottom and middle brackets.
+DreamSim Z-score rescaling.
+Results closely matched those for unscaled DreamSim scores, with statistically significant differences between the top bracket and both the bottom and middle brackets, but no significant difference between the bottom and middle brackets.
 No significant effects emerged from linear models or from user-level analyses.
-    • Percentile rank rescaling: Results aligned with those for Z-score rescaled DreamSim, with one exception: at the user-attempt level, we observed a statistically significant difference only between the top and bottom brackets, not between the top and middle brackets.
-  • Prompting Behaviors
-    • Prompt Length: We found statistically significant evidence that higher Z-scores led to longer subsequent prompts.
+DreamSim Percentile rank rescaling.
+Results aligned with those for Z-score rescaled DreamSim, with one exception: at the user-attempt level, we observed a statistically significant difference only between the top and bottom brackets, not between the top and middle brackets.
+Prompt Length.
+We found statistically significant evidence that higher Z-scores led to longer subsequent prompts.
 When using the trichotomous variable, the difference between the top and bottom Z-score brackets was significant, with longer prompts on average in the top bracket.
@@
 At the user level, the Z-score of the first prompt did not significantly affect the length of subsequent prompts.
-    • Successive Similarity: Higher Z-scores were associated with increased similarity between consecutive prompts.
+Successive Similarity.
+Higher Z-scores were associated with increased similarity between consecutive prompts.
 The top Z-score bracket showed significantly higher prompt similarity compared to both the bottom and middle brackets, though no significant difference emerged between the bottom and middle brackets.
@@
 No significant effects were found at the user level.
-    • Successive Prompt Token Sort Ratio: Results mirrored those for successive cosine similarity, with higher Z-scores associated with increased token sort ratios between successive prompts.
+Successive Prompt Token Sort Ratio.
+Results mirrored those for successive cosine similarity, with higher Z-scores associated with increased token sort ratios between successive prompts.
 This pattern was consistent across both trichotomous variable comparisons and linear models.
 No significant effects emerged at the user level.
-    • Successive Prompt ’Contains Previous Prompt’ Dummy: Higher Z-scores significantly increased the likelihood that subsequent prompts would contain the previous prompt as an exact substring.
+Successive prompt ’contains previous prompt’ dummy.
+Higher Z-scores significantly increased the likelihood that subsequent prompts would contain the previous prompt as an exact substring.
 The difference between top and bottom Z-score brackets was significant, though differences involving the middle bracket were not.
@@
 No significant effects were found at the user level.
+Interpretation
 These findings reveal a consistent pattern: when the stochastic nature of generative AI produces unexpectedly high-quality outputs for a given prompt, users tend to build more directly upon that prompt in their subsequent attempts—writing longer prompts that are more similar to and often directly incorporate the previous prompt.
@@
 The tendency to build upon prompts that produce better-than-expected results represents an intuitive but potentially suboptimal learning strategy, as it may attribute too much importance to random variations rather than focusing on the underlying quality of the prompt itself.
-H5
+H5.
 Average prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) and prompting strategies will depend on the capacity of the model that participants are interacting with.
-Analysis approach: To test this hypothesis, we conducted pairwise comparisons across the three treatment arms using two-sample t-tests to evaluate differences in both performance outcomes and prompting behaviors.
+Empirical approach:
+To test this hypothesis, we conducted pairwise comparisons across the three treatment arms using two-sample t-tests to evaluate differences in both performance outcomes and prompting behaviors.
 For robustness, we also performed ANOVA tests to examine whether any significant differences existed across all three treatment arms simultaneously.
@@
 Results: Our analyses revealed several key differences in both performance and prompting behavior across the different model conditions:
-  • CLIP embedding cosine similarity
-    • Z-score rescaling: We found statistically significant evidence for performance differences across treatment arms when examining participants’ first attempts, average performance across all attempts, and best attempts (all measured using taskwide Z-scores rather than task-iteration Z-scores).
+CLIP Z-score rescaling.
+We found statistically significant evidence for performance differences across treatment arms when examining participants’ first attempts, average performance across all attempts, and best attempts (all measured using taskwide Z-scores rather than task-iteration Z-scores).
 For first attempt performance, DALL-E 3 (Verbatim) participants significantly outperformed both DALL-E 3 (Revised) and DALL-E 2 participants, while no statistically significant difference emerged between DALL-E 3 (Revised) and DALL-E 2 participants.
@@
 ANOVA tests for all three performance measures (first, average, and best attempts) confirmed significant effects of model assignment.
-    • Percentile rank rescaling: Results using percentile rank rescaling mirrored those found with Z-score rescaling.
+CLIP Percentile rank rescaling.
+Results using percentile rank rescaling mirrored those found with Z-score rescaling.
 DALL-E 3 (Verbatim) participants outperformed both other conditions on first attempt performance and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2 for these metrics.
@@
 ANOVA tests confirmed significant effects of model assignment for all three performance measures.
-  • DreamSim
-    • Z-score rescaling: Results using DreamSim with Z-score rescaling showed the same pattern as CLIP embedding cosine similarity.
+DreamSim Z-score rescaling.
+Results using DreamSim with Z-score rescaling showed the same pattern as CLIP embedding cosine similarity.
 DALL-E 3 (Verbatim) participants significantly outperformed both other conditions on first attempt and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2.
@@
 ANOVA tests confirmed significant effects across all performance measures.
-    • Percentile rank rescaling: Results with percentile rank rescaling of DreamSim scores were consistent with the Z-score findings.
+DreamSim Percentile rank rescaling.
+Results with percentile rank rescaling of DreamSim scores were consistent with the Z-score findings.
 DALL-E 3 (Verbatim) participants outperformed both other conditions on first and average performance, with no differences between DALL-E 3 (Revised) and DALL-E 2.
@@
 ANOVA tests confirmed significant effects across all performance measures.
-  • Prompting Behaviors
-    • Mean prompt Length: We found statistically significant differences in prompt length across treatment arms.
+Mean prompt Length.
+We found statistically significant differences in prompt length across treatment arms.
 Participants using DALL-E 2 wrote significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups.
@@
 ANOVA results confirmed a significant effect of model version on prompt length.
-    • Aggregate Similarity: We found no statistically significant differences in the overall variability of prompts across the three treatment arms.
+Aggregate Similarity.
+We found no statistically significant differences in the overall variability of prompts across the three treatment arms.
 ANOVA results confirmed no significant effect of model version on this measure, suggesting that participants explored similar breadths of prompt space regardless of the model they were using.
-    • Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences across treatment arms.
+These results contrast with our findings from the main analysis, where prompts written by DALL-E 3 participants were, on average, more consistent than those written by DALL-E 2 participants.
+This difference reflects statistical power and specification: the main analysis uses a higher-powered model with task and attempt fixed effects, whereas the aggregate-similarity analysis here relies on a simpler cross-sectional comparison.
+Successive Similarity.
+Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences across treatment arms.
 ANOVA results confirmed no significant effect of model version on successive prompt similarity, indicating that the tendency to build upon or deviate from previous prompts was similar across models.
-    • Successive Prompt Token Sort Ratio: No statistically significant differences emerged across treatment arms.
-ANOVA results confirmed no significant effect of model version on this measure of textual similarity between successive prompts.
-    • Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences across treatment arms.
+Successive prompt token sort ratio.
+No statistically significant differences emerged across treatment arms. ANOVA results confirmed no significant effect of model version on this measure of textual similarity between successive prompts.
+Successive prompt ‘contains previous prompt’ dummy:
+Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences across treatment arms.
 ANOVA results confirmed no significant effect of model version on this measure.
+Interpretation.
 These findings paint an interesting picture of how model capacity influences both performance and prompting behavior.
@@
 However, we found no significant differences across models in measures of prompt similarity or evolution, suggesting that while participants provided more detail when using more capable models, their overall strategies for iterating on prompts remained relatively consistent regardless of model assignment.
-H6
+H6.
 Variability in participants’ ability to prompt engineer effectively and prompting strategies will depend on the capacity of the model that participants are interacting with.
-Analysis approach: To test this hypothesis, we conducted two types of analyses.
+Empirical approach:
+To test this hypothesis, we conducted two types of analyses.
 First, we performed F-tests comparing the variance of participant performance and prompting behaviors between all three pairs of model conditions (DALL-E 2 vs.
@@
 For example, positive effects for low quantiles and negative/null effects for high quantiles would indicate inequality reduction.
-F-test Results (with BH Adjusted p-values):
-  • DALL-E 3 (revised) vs. DALL-E 2
+F-tests.
+We first present results (with BH Adjusted p-values) for DALL-E 3 (revised) vs. DALL-E 2:
+  1. 1.
     Mean prompt length (words): DALL-E 3 Revised showed significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
+  2. 2.
     Prompt embedding variance: DALL-E 3 Revised also showed significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
+  3. 3.
     Cosine similarity with target image, Z-Score: DALL-E 3 Revised showed significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
+  4. 4.
     Failed to reject null of no differences in variance: Mean raw DreamSim score vs.
@@
 target image (p = 0.8301), mean proportion of prompts containing previous prompt (p = 0.5461), Z-score of DreamSim score (p = 0.4016), number of topical transitions (cosine similarity) (p = 0.3505), and mean percentile rank of CosineSim (p = 0.0795).
-  • DALL-E 3 (verbatim) vs. DALL-E 3 (revised)
+We next present results (with BH Adjusted p-values) for DALL-E 3 (verbatim) vs. DALL-E 3 (revised):
+  1. 1.
     Mean percentile rank of CosineSim: DALL-E 3 (verbatim) showed significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
+  2. 2.
     Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also showed significantly less variance (ratio 0.7352, $p=0.0009$).
+  3. 3.
     Mean percentile rank of DreamSim: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7888, $p=0.0141$).
+  4. 4.
     Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7961, $p=0.0159$).
+  5. 5.
     DreamSim with target image, Z-Score: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.8195, $p=0.0377$).
+  6. 6.
     Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) showed significantly more variance (ratio 1.2303, $p=0.0301$).
+  7. 7.
     Failed to reject null of no differences in variance: Mean raw DreamSim score vs.
@@
 target image (p = 0.5916), prompt embedding variance (p = 0.9771), mean cosine similarity to previous prompt (p = 0.9035), mean prompt length (words) (p = 0.6244), number of topical transitions (cosine similarity) (p = 0.5916), and mean token sort ratio with previous prompt (p = 0.5206).
-  • DALL-E 2 vs. DALL-E 3 (verbatim)
+Finally, the F-tests for DALL-E 2 vs. DALL-E 3 (verbatim):
+  1. 1.
     Mean prompt length (words): DALL-E 3 (verbatim) showed significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
+  2. 2.
     Prompt embedding variance: DALL-E 3 (verbatim) also showed significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
+  3. 3.
     Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7501, $p=0.0015$).
+  4. 4.
     Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7904, $p=0.0141$).
+  5. 5.
     Mean percentile rank of DreamSim: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.8005, $p=0.0159$).
+  6. 6.
     Failed to reject null of no differences in variance: Mean percentile rank of CosineSim (p = 0.1842), Z-score of DreamSim score (p = 0.3294), Mean Raw CosineSim (p = 0.8151), Z-score of CosineSim score with target image (p = 0.9906), mean token sort ratio with previous prompt (p = 0.8748), mean cosine similarity to previous prompt (p = 0.807), number of topical transitions (token sort ratio) (p = 0.3505), and number of topical transitions (cosine similarity) (p = 0.085).
-QTE Highlighted Results: Through our quantile treatment effects analysis and visual inspection of QTE plots, we found:
-  • Evidence of dispersion reduction: Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALL-E 3 Revised vs.
+QTE Highlighted Results.
+Through our quantile treatment effects analysis and visual inspection of QTE plots, we found clear evidence of dispersion reduction.
+The Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALL-E 3 Revised vs.
 DALL-E 2 comparison showed patterns consistent with inequality reduction, with larger positive effects at lower quantiles.
-  • Evidence of dispersion increase: Mean prompt length for all three pairwise model comparisons (DALL-E 3 Verbatim vs.
+We also found evidence of dispersion increase: Mean prompt length for all three pairwise model comparisons (DALL-E 3 Verbatim vs.
 DALL-E 2, DALL-E 3 Revised vs.
@@
 As participants repeatedly try to complete a task with a given model, the quality of their attempts will increase, and the extent to which the quality increases varies as a function of model capacity.
-Analysis approach: To evaluate how performance improves across successive attempts and whether this improvement varies by model, we conducted stratified two-sample tests comparing participants’ initial scores with their best scores.
+Empirical approach:
+To evaluate how performance improves across successive attempts and whether this improvement varies by model, we conducted stratified two-sample tests comparing participants’ initial scores with their best scores.
 These comparisons were performed both within each treatment arm (DALL-E 2, DALL-E 3 Verbatim, and DALL-E 3 Revised) and overall across all participants.
 This approach allowed us to determine both whether participants generally improved with practice and whether the rate of improvement differed across models with varying capabilities.
-Results: Our analyses revealed consistent performance improvements across all treatment arms and outcome measures:
-  • CLIP embedding cosine similarity
-    • No rescaling: We found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
+Our analyses revealed consistent performance improvements across all treatment arms and outcome measures.
+CLIP No rescaling.
+We found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
 This indicates that participants were able to improve their performance through iterative attempts regardless of which model they were using.
-    • Z-score rescaling: Results mirrored those of the unscaled analysis, with statistically significant improvements from initial to best performance observed both within each treatment arm and overall.
-    • Percentile rank rescaling: Consistent with other scaling approaches, we observed statistically significant improvements from initial to best performance both within each treatment arm and overall.
-  • DreamSim
-    • No rescaling: DreamSim results aligned with CLIP similarity findings, showing statistically significant improvements from initial to best performance both within each treatment arm and overall.
-    • Z-score rescaling: We again found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
-    • Percentile rank rescaling: Results were consistent with other analyses, showing statistically significant improvements from initial to best performance across all conditions.
+CLIP Z-score rescaling.
+Results mirrored those of the unscaled analysis, with statistically significant improvements from initial to best performance observed both within each treatment arm and overall.
+CLIP Percentile rank rescaling.
+Consistent with other scaling approaches, we observed statistically significant improvements from initial to best performance both within each treatment arm and overall.
+DreamSim No rescaling.
+DreamSim results aligned with CLIP similarity findings, showing statistically significant improvements from initial to best performance both within each treatment arm and overall.
+DreamSim Z-score rescaling.
+We again found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
+DreamSim Percentile rank rescaling.
+Results were consistent with other analyses, showing statistically significant improvements from initial to best performance across all conditions.
+Interpretation.
 These findings demonstrate that participants consistently improved their performance through iterative attempts across all model conditions and performance measures.
@@
 The extent to which participants can recreate images using models such as DALL-E 2/3 will vary across images.
-Analysis approach: To evaluate this hypothesis, we employed two complementary approaches.
+Empirical approach.
+To evaluate this hypothesis, we employed two complementary approaches.
 First, we used GPT-4V (a multimodal generative AI model capable of processing both text and images) to generate optimal prompts for each target image.
@@
 This two-pronged approach allowed us to assess image difficulty from both AI and human perspectives.
-Results:
-  • CLIP embedding cosine similarity: When ranking target images by GPT-4V’s ability to generate effective prompts, we found substantial variation across images:
+CLIP Results.
+When ranking target images by GPT-4V’s ability to generate effective prompts, we found substantial variation across images:
   1. 1.
@@
 The highest average cosine similarity (easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, while the lowest score (hardest image to replicate) was $CoSim=0.669$ for Business Image #4.
-  • DreamSim: When using DreamSim as our similarity metric (inverted as 1-DreamSim to create a similarity rather than distance measure), GPT-4V’s prompt generation performance ranked as follows:
+DreamSim Results.
+When using DreamSim as our similarity metric (inverted as 1-DreamSim to create a similarity rather than distance measure), GPT-4V’s prompt generation performance ranked as follows:
   1. 1.
@@
 The highest 1-DreamSim score (easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, while the lowest score (hardest image to replicate) was $\tilde{D}=0.356$ for Photography Image #4.
+Interpretation
 These results clearly demonstrate substantial variation in image replicability across our target set.
@@
 The pre-registered exploratory analyses were described as follows:
-”We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate.
+“We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate.
 We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
@@
 This approach allowed us to isolate how much improvement came from the model’s enhanced capabilities versus users adapting their prompting strategies to take advantage of those capabilities.
-G.2 Deviations From Pre-registration
+I.2 Deviations From Pre-registration
 We report below all deviations from our pre-registered analysis plan.
 These deviations primarily resulted from statistical or methodological considerations that became apparent during data analysis, rather than from substantive changes to our research questions or hypotheses.
-  • Statistical tests: Our pre-registration specified t-tests and Mann-Whitney U tests for many hypotheses.
+Statistical tests.
+Our pre-registration specified t-tests and Mann-Whitney U tests for many hypotheses.
 However, this approach proved inappropriate for variables with multiple categories (which characterized most demographic traits).
 We therefore employed ANOVA and Kruskal-Wallis tests instead, which provide equivalent information for multi-category variables.
-  • Z-score computation: Our pre-registration was insufficiently precise about computing Z-score performance measures.
-As detailed in Section D.2, we calculated Z-scores within image-attempt pairs to adjust for variations across target images and attempts, except when comparing performance across attempts, where Z-scores were computed within target images only.
+Z-score computation.
+Our pre-registration was insufficiently precise about computing Z-score performance measures.
+As detailed in Section E.2, we calculated Z-scores within image-attempt pairs to adjust for variations across target images and attempts, except when comparing performance across attempts, where Z-scores were computed within target images only.
 The pre-registration had anticipated computing Z-scores only within target images for all analyses.
-  • Additional demographic variables: We inadvertently omitted Education and Generative AI outlook from our pre-registered list of demographics for measuring task performance heterogeneity.
+Additional demographic variables.
+We inadvertently omitted Education and Generative AI outlook from our pre-registered list of demographics for measuring task performance heterogeneity.
 Given their theoretical importance, we included these variables in our analyses despite this accidental omission.
-  • Additional prompt exclusion criteria: Beyond our pre-registered exclusion criteria, we removed additional prompts that did not appear to constitute ”good-faith efforts” based on their text content (see Section C.4 for details).
+Additional prompt exclusion criteria.
+Beyond our pre-registered exclusion criteria, we removed additional prompts that did not appear to constitute ”good-faith efforts” based on their text content (see Section C.4 for details).
 Robustness checks confirmed that our results remain consistent when including these prompts.
-  • T-tests vs.
-Z-tests: For testing H5, we used t-tests rather than z-tests because they were easier to implement.
-As these tests are asymptotically equivalent, this change does not meaningfully affect our results.
-  • Z-score outlier handling: When testing H4, we observed that Z-scores did not follow a normal distribution and included extreme values (ranging from -21.9 to 7.1).
-To prevent these outliers from disproportionately influencing our linear models, we excluded observations with absolute Z-scores greater than 3 (2.65
-  • GPT-4V rescaling omission: Our pre-registration included a rescaled version of performance metrics based on GPT-4V prompt quality.
+T-tests vs. Z-tests.
+For testing H5, we used t-tests rather than z-tests because they were easier to implement. As these tests are asymptotically equivalent, this change does not meaningfully affect our results.
+Z-score outlier handling.
+When testing H4, we observed that Z-scores did not follow a normal distribution and included extreme values (ranging from -21.9 to 7.1).
+To prevent these outliers from disproportionately influencing our linear models, we excluded observations with absolute Z-scores greater than 3 (2.65% of user-attempt observations).
+Robustness checks including these outliers yielded nearly identical results, with minor differences in significance levels for some comparisons.
+GPT-4V rescaling omission.
+Our pre-registration included a rescaled version of performance metrics based on GPT-4V prompt quality.
 However, this rescaling would have amounted to subtracting a constant from each unscaled score, and since all our analyses adjusted for target image (either through post-stratification or as a covariate), results would have been identical to those using unscaled measures.
 We therefore omitted these redundant robustness checks.
-G.3 OSF Pre-registration
-A complete copy of the pre-registration deposited on OSF on December 11, 2023 will be included beginning on the next page in the final manuscript, but has been removed here to maintain anonymity.

### v7

@@
 1 Introduction
-Generative AI is being integrated into work practices across the economy (zhang2023generative, bright2024generative), yielding notable productivity gains in tasks as diverse as software development, writing, and scientific research (brynjolfsson2023generative, dell2023navigating, Noy2023, peng2023impact, Yu2024AIImpact).
-Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning2024Automated), including tasks as complex as chemical research and proving mathematical theorems (RomeraParedes2024Math, Boiko2023AutonomousResearch).
+Generative AI is being integrated into work practices across the economy (Zhang and Kamel Boulos 2023, Bright et al.
+2024), yielding notable productivity gains in tasks as diverse as software development, writing, and scientific research (Brynjolfsson et al.
+2023, Dell’Acqua et al.
+2023, Noy and Zhang 2023, Peng et al.
+2023, Yu 2024).
+Recent research points to even greater potential ahead, demonstrating advances in automating core scientific processes (Manning et al.
+2024), including tasks as complex as chemical research and proving mathematical theorems (Romera-Paredes et al.
+2024, Boiko et al.
+2023).
 The adoption of generative AI is also occurring at an unprecedented pace, with recent research showing that approximately 28% of U.S.
-workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (bright2024generative, bick2024rapid).
+workers are already using generative AI in their jobs—a rate that significantly outpaces early adoption of personal computers and internet technology at comparable points in their diffusion (Bright et al.
+2024, Bick et al.
+2024).
 As with many other general-purpose technologies, the effectiveness of generative AI depends not only on the technology itself, but on users’ ability to craft inputs that produce high-quality results.
@@
 These prompts can range from simple commands (e.g., “write a short story about a robot”) to highly detailed specifications tailored to particular outputs (e.g., a series of paragraphs instructing an AI system to implement a complete piece of software).
-In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (brynjolfsson2000beyond).
+In this way, prompting serves as a complementary skill—one that, like spreadsheet modeling in the early PC era, can determine the productivity impact of a given tool (Brynjolfsson and Hitt 2000).
 Prompting has quickly become an area of active research and practice.
-Scholars have developed taxonomies of prompt engineering techniques (oppenlaender2023), documented recurring patterns in prompt construction (schulhoff2024prompt), and examined how developers embed prompts into software systems (liang2024prompts).
-Other studies have explored prompting strategies for specific applications, including image generation (donyehiya2023, xie2023) and clinical documentation (yao2024clinicians).
+Scholars have developed taxonomies of prompt engineering techniques (Oppenlaender 2023), documented recurring patterns in prompt construction (Schulhoff et al.
+2024), and examined how developers embed prompts into software systems (Liang et al.
+2024).
+Other studies have explored prompting strategies for specific applications, including image generation (Don-Yehiya et al.
+2023, Xie et al.
+2023) and clinical documentation (Yao et al.
+2024).
 In parallel, practitioners have built prompt libraries, shared tutorials, and developed tools to support prompt design.
@@
 Many prompt libraries and tutorials present effective prompts as reusable artifacts.
-But prompts that work well with one model version may underperform or break entirely with the next (liang2024prompts, meincke2025prompting).
+But prompts that work well with one model version may underperform or break entirely with the next (Liang et al.
+2024, Meincke et al.
+2025).
 While recent research increasingly views prompting as an adaptive process, empirical evidence remains limited on how these strategies evolve—both as users refine prompts for a single model and as they adjust to model updates—and on how these changes ultimately affect performance.
@@
 Others are creative, where goals are open-ended, evaluation is subjective, and improvement depends more on exploration and curation than on precision (e.g., make a beautiful picture).
-Recent work helps to establish these regimes: vafa2025s investigate the steerability of models, i.e., how effectively users can reach a specified target, whereas orwig2024language and zhou2024generative describe generative AI–assisted creation as a co-creative process of idea generation, filtering, and aesthetic evaluation.
+Recent work helps to establish these regimes: Vafa et al.
+(2025) investigate the steerability of models, i.e., how effectively users can reach a specified target, whereas Orwig et al.
+(2024) and Zhou and Lee (2024) describe generative AI–assisted creation as a co-creative process of idea generation, filtering, and aesthetic evaluation.
 In organizational settings, steering tasks (or bounded tasks, herein) dominate operational workflows such as classification, brand-compliant content generation, or reproducing structured materials where quality can be defined ex ante.
@@
 In terms of related literature and additional theory, our research builds on work in information systems, emphasizing the importance of dynamic, user-driven complements to digital technologies.
-Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (bharadwaj2000resource, joshi2010changing, teece1997dynamic).
-Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (jasperson2005comprehensive).
-Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (fugener2022cognitive).
-And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (bohm2023co-evolution).
-We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (brynjolfsson1993productivity, brynjolfsson2000beyond, brynjolfsson2021productivity, david1990dynamo).
+Research on IT-enabled dynamic capabilities has shown that the value of new systems depends not only on technical infrastructure but on organizations’ ability to reconfigure routines and user behaviors in response to ongoing change (Bharadwaj 2000, Joshi et al.
+2010, Teece et al.
+1997).
+Related work on post-adoptive IT use has demonstrated that users often engage only superficially with new systems and that meaningful performance gains tend to emerge only when users experiment with and refine their interaction strategies over time (Jasperson et al.
+2005).
+Recent research on human-AI collaboration further underscores that interface design and task structure shape the degree to which users can learn from and adapt to model behavior (Fügener et al.
+2022).
+And the concept of co-evolution has been introduced to describe how humans and generative AI systems jointly adapt over time, forming interdependent capabilities that neither could realize alone (Böhm and Schedlberger 2023).
+We also engage with work on general-purpose technologies, which has long emphasized that the productivity gains from technical advances depend on the development of new human and organizational complements (Brynjolfsson 1993, Brynjolfsson and Hitt 2000, Brynjolfsson et al.
+2021, David 1990).
 We conceptualize prompt adaptation as a dynamic complement that co-evolves with model capability, emerges through use rather than formal training, and whose importance depends on task structure and system design.
@@
 Each unit of effort incurs a linear cost $c(x)=kx$ for some $k>0$, reflecting the total cognitive and temporal costs of refining one’s prompts.
-We assume $k$ is small ($k\!\approx\!0$), consistent with the low marginal cost of iterative prompting and using generative AI models more generally (shahidi2025coasean).
+We assume $k$ is small ($k\!\approx\!0$), consistent with the low marginal cost of iterative prompting and using generative AI models more generally (Shahidi et al.
+2025).
 We begin with a bounded quality function,
@@
 3.2.1 Outcome Definition: Expected CLIP Embedding Cosine Similarity
-The primary outcome in the image replication experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (radford2021learning).
+The primary outcome in the image replication experiment is the similarity between each participant-generated image and the assigned target image, measured using the cosine similarity of CLIP embeddings (Radford et al.
+2021).
 CLIP (Contrastive Language–Image Pretraining) is a neural network trained to jointly embed images and text into a shared latent space, such that semantically or visually similar items lie close together.
@@
 As a robustness check, we replicated all analyses using DreamSim, a recently developed alternative to cosine similarity on CLIP embeddings that is based on perceptual similarity and aligns more closely with human judgments.
-[^8] DreamSim (fu2023dreamsim) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
+[^8] DreamSim (Fu et al.
+2023) is designed to better capture human perceptions of image similarity than traditional embedding-based methods.
 Our results are robust to using DreamSim in place of CLIP cosine similarity.
@@
 Because each logo-generation task was open-ended and lacked an objective reference image, we could not assess output quality using embedding-based similarity metrics (e.g., cosine similarity of CLIP representations).
-Instead, we employed the Bradley–Terry (BT) model (bradley1952rank), a standard approach for inferring latent quality scores when only relative judgments are available (zermelo1929berechnung, ford1957solution).
+Instead, we employed the Bradley–Terry (BT) model (Bradley and Terry 1952), a standard approach for inferring latent quality scores when only relative judgments are available (Zermelo 1929, Ford 1957).
 At a high level, the BT model is a probabilistic model that estimates a strength parameter (which we refer to hereafter as “BT $\beta$”) for each item based on the outcomes of many pairwise comparisons.
@@
 These cross-model evaluations provide empirical analogues of $Q(\theta_{2},s,x^{*}(\theta_{1},s))$ and $Q(\theta_{1},s,x^{*}(\theta_{2},s))$, respectively, and together with the original prompt–model pairings, allow us to implement the decomposition into model and prompting effects shown in Equation 5.
-In the image replication experiment, replay was conducted several weeks after the initial data collection, so we additionally re-generated outputs for the original prompt–model combinations contemporaneously with replay to guard against model drift (chen2023gptchanging).
+In the image replication experiment, replay was conducted several weeks after the initial data collection, so we additionally re-generated outputs for the original prompt–model combinations contemporaneously with replay to guard against model drift (Chen et al.
+2023).
 In the logo generation experiment, replay occurred shortly after data collection, so this additional step was unnecessary.
@@
 but poorly aligned system prompts, especially when hidden, can degrade performance and/or limit users’ ability to discover or exploit emergent but useful ways of interacting with a generative AI system.
-Such complexities are highlighted by related work, such as yao2024clinicians, which studies automated prompt optimization in the context of clinical documentation.
-More generally, our results both align with and help clarify themes in the literature on human–AI collaboration and the economics of technology, particularly the role of dynamic complements such as prompt adaptation that co-evolve with technical change (bohm2023co-evolution).
+Such complexities are highlighted by related work, such as Yao et al.
+(2024), which studies automated prompt optimization in the context of clinical documentation.
+More generally, our results both align with and help clarify themes in the literature on human–AI collaboration and the economics of technology, particularly the role of dynamic complements such as prompt adaptation that co-evolve with technical change (Böhm and Schedlberger 2023).
 As generative AI models advance, often substantially from month to month, organizations that fail to adapt their prompting strategies in task settings where precision and steerability matter may forgo a meaningful share of the economic value these upgrades make possible.
 This co-evolutionary pattern is not unique to AI.
-It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (brynjolfsson1993productivity, david1990dynamo).
+It echoes dynamics observed in earlier general-purpose technologies, where technical improvements often yielded modest returns until complementary skills and practices evolved to match (Brynjolfsson 1993, David 1990).
 However, the pace of generative AI advancement introduces a distinctive challenge: the window for adaptation is far shorter.
-Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (brynjolfsson2000beyond, brynjolfsson2021productivity).
+Organizations that treat prompting as a one-time investment rather than an ongoing capability risk failing to capture the full value of model upgrades—echoing the productivity paradox observed when complementary assets lag behind technological potential (Brynjolfsson and Hitt 2000, Brynjolfsson et al.
+2021).
 Our findings complement this stream of work by offering direct empirical evidence that a specific dynamic complement—user adaptation, specifically through prompt refinement—can account for a substantial share of realized performance gains from model upgrades in task settings where precision and steerability matter.
-Seen from a broader organizational perspective, these findings also speak to long-standing work on IT-enabled dynamic capabilities (bharadwaj2000resource, joshi2010changing, teece1997dynamic) and post-adoptive IT use (jasperson2005comprehensive).
+Seen from a broader organizational perspective, these findings also speak to long-standing work on IT-enabled dynamic capabilities (Bharadwaj 2000, Joshi et al.
+2010, Teece et al.
+1997) and post-adoptive IT use (Jasperson et al.
+2005).
 Prompt adaptation, as we observe it, is not grounded in formal training or specialized skill.
 Participants in our study, drawn from a general population rather than a specialized group of prompt engineers, improved performance through trial-and-error within a single session.
-This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (attewell1992technology, von2006democratizing).
-The relative accessibility of prompt adaptation highlights one potential pathway toward more broadly distributed productivity gains, provided that users are supported by scaffolds and interfaces that enable iterative refinement (rogers2003diffusion).
+This contrasts with earlier IT transitions, where complementary capabilities often required extensive training (Attewell 1992, Von Hippel 2006).
+The relative accessibility of prompt adaptation highlights one potential pathway toward more broadly distributed productivity gains, provided that users are supported by scaffolds and interfaces that enable iterative refinement (Rogers 2003).
 At the same time, this accessibility introduces risk.
 Over-optimizing for a particular model version may reduce users’ ability to adapt as systems evolve.
-This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (henderson1990architectural).
+This behavioral lock-in resembles challenges in architectural innovation, where tightly coupled routines inhibit flexibility when key system components change (Henderson and Clark 1990).
 Supporting long-term adaptation may require not just prompt training, but workflows and learning mechanisms that encourage ongoing experimentation.
@@
 For researchers and practitioners interested in the economics of AI, our findings reinforce the idea that complements play a central role in shaping performance.
-In some task settings, prompt adaptation can account for as much as half of the realized performance gains from model upgrades, underscoring that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (arthur2009nature, dosi1982technological).
+In some task settings, prompt adaptation can account for as much as half of the realized performance gains from model upgrades, underscoring that skill development and technological evolution must be treated as interdependent elements of innovation trajectories (Arthur 2009, Dosi 1982).
 As generative AI continues to advance, organizations that invest in adaptive, not static, complementary skills will be better positioned to realize its full value.
@@
 D.1 Image Similarity
-We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (fu2023dreamsim).
+We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al.
+2023).
 In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers.
@@
 CLIP Embedding Cosine Similarity.
-To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (radford2021learning) using the version of CLIP available on Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image.
+To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al.
+2021) using the version of CLIP available on Hugging Face (OpenAI 2024) for each participant-generated image and for each target image.
 Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text.
@@
 DreamSim.
-DreamSim is an image similarity measure proposed recently by (fu2023dreamsim).
+DreamSim is an image similarity measure proposed recently by (Fu et al.
+2023).
 The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity.
@@
 We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity.
-Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (neelakantan2022text).
+Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al.
+2022).
 The two similarity measures are defined as follows:
@@
 Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$.
-TSR is a fuzzy string-matching technique (singla2012string) that provides a continuous measure of how similar two strings are.
+TSR is a fuzzy string-matching technique (Singla and Garg 2012) that provides a continuous measure of how similar two strings are.
 D.6 Successive Prompt ‘Contains Previous Prompt’ Dummy
@@
 We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt.
-SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation posuniversaldep.
+SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal Dependencies Project (2024).
 These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs.
@@
 In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting.
-In order to do so, we adapt a method proposed in (torricelli2023role) (because they are conducting research in a different context, torricelli2023role refer to these shifts as “topical transitions” as opposed to “strategic shifts”).
+In order to do so, we adapt a method proposed in (Torricelli et al.
+2023) (because they are conducting research in a different context, Torricelli et al.
+(2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”).
 To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
@@
 […]
-It is worth noting that torricelli2023role uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift.
+It is worth noting that Torricelli et al.
+(2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift.
 We use the task-level mean because in our setting, it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
@@
 We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket.
-Our estimates adjusted for other covariates by matching user-iteration observations on target image, DALL-E model, iteration number, and exact quality of the previous attempt (savje2021generalized).
+Our estimates adjusted for other covariates by matching user-iteration observations on target image, DALL-E model, iteration number, and exact quality of the previous attempt (Sävje et al.
+2021).
 Our analyses revealed consistent patterns across different outcome measures: